### v3

# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

Abstract
In this paper, we extend two classes of functions classically involved in asymptotic stability analyses for studying a maximization problem on the reachable values of a discrete-time dynamical system. This particular maximization problem is called a peaks computation problem. The problem is to find a couple composed of an initial state and a time which maximizes a given function over states. The paper focuses on the time component of the optimal solution which is an integer as the time is discrete. We develop a method to provide an upper bound of this integer from a formula which requires a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. A first result proves that the formula provides, in theory, the optimal integer. However, in practice, the computation cannot be so precise. Then, we develop two alternative methods. The first is based on discontinuous and non strictly increasing/decreasing $\mathcal{KL}$-like functions named $\mathcal{KL}_{\rm gen}$ functions. We prove that the existence of a $\mathcal{KL}_{\rm gen}$ upper bound is equivalent to the existence of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of the strictly increasing continuous function from a $\mathcal{KL}_{\rm gen}$ function needs an extension of the famous Sontag’s lemma. Finally, we construct a new type of Lyapunov functions well designed for our peaks computation problem. These functions are called Opt-Lyapunov functions. Opt-Lyapunov functions are well designed since we establish an equivalence theorem between the existence of an Opt-Lyapunov function and of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of a Opt-Lyapunov function from a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence is insipred by the Yoshizawa construction of Lyapunov functions.
Keywords: Discrete-time, Peaks Computation Problem, Supremum of real sequences, Lyapunov functions, $\mathcal{KL}$ functions.
1 Introduction
In this paper, we study the problem of peaks computation for discrete-time dynamical systems in finite dimensional state spaces. A peaks computation problem can be viewed as a particular maximization problem for which the constraints set is the reachable values set of a discrete-time dynamical system. The objective function of this optimization problem is a function over the states. This particular maximization problem can model various situations in engineering, economics, physics, biology… For example, in population dynamics, for a predator-prey system, one may be interested in the maximal size of each species even if the system asymptotically oscillates between several equilibra. This problem boils down to maximize separetely each coordinate of the state variable over the system modeling the population dynamics.
Besides its importance to analyze critical concrete situations, the literature about peaks analysis problem in the community of dynamical systems analysis is pretty thin. The peaks computation problem appears in a discrete-time setting for linear systems and where the objective function is the Euclidean norm [APS18]. The problem also appears for the analysis of stable linear systems with general quadratic objective functions [Adj21, Adj23]. A slighty different version of the problem is also encountered in [AG24]. The difference comes from the fact that the state variable is, moreover, constrained to stay in a some given set. The problem is studied in a continuous-time setting and in, algebraically, a polynomial environment in [MHS20]. Note that in [MHS20], the authors address the problem of verifying safety properties (see e.g. [BYG17] for verification problems on discrete-time systems) for dynamical systems and formulate the safety verification problem as a peaks computation problem. Safety verification problems aim to prove that the trajectories of the system are contained in some safe set. A safety verification problem can be translated into a peaks computation problem when the safe set is a sublevel of a sufficiently regular function. In this case, the objective function of the peaks computation problem is the function associated with the sublevel set. The safety analysis from the peaks computation point-of-view is also proposed for switched system [Adj17] and for program analysis in [AGM15]. For safety analysis of dynamical systems, one can think that a reachability analysis (see e.g. [RKML06] and reference therein) equivalent to a feasibility analysis in our context is sufficient to prove some properties on the dynamical system. However, maximality in our context is a synonym for criticality. The optimal value in this situation can be viewed as the value which penalizes the most the system with respect to some criteria and defines the extreme condition where the system could enter into.
In [MHS20] the authors are interested in computing an upper bound of the peaks optimal value for the continuous case. The techniques used is based on Sums-Of-Squares (SOS) and Linear Matrix Inequalities (LMI) formulations (the interested reader can consult [Las15] for those subjects). The same goal motivates the authors of [APS18] and, as they limit their study to linear system and norm objective functions, they directly use semidefinite programming (see [BTN01] for further details on the subject). The reference [AG24] cannot be fairly compared due to the presence of the additional constraint. Moreover, again, the studied systems are linear or piecewise linear and the objective functions are linear. The goal of the current paper is to provide the theoretical bases to solve exactly a peaks computation problem in a general way. The developments made in this paper can be viewed as a theoretical generalization of works done in [Adj21, Adj23]
In this paper, we propose to study peaks computation problems from classical tools of dynamical systems stability theory (e.g. [Ela05, Chap. 4] for discrete-time). This approach comes from a simple observation: the existence of peaks depends on the asymptotic behaviour of the system. The importance of stability analysis is time-tested and a stability analysis is a prerequisite to the implementation of numerical methods. Whereas general stability properties can be difficult to obtain, asymptotic stability properties can be easily presented and understood from $\mathcal{KL}$ bounds (e.g. [ATP10] or [NT04]) i.e. the norms of the states are globally bounded from above by a $\mathcal{KL}$ function. However, proving that a candidate $\mathcal{KL}$ function is an upper bound over the norms of state has to be done for all integers which can be difficult. On the other hand, Lyapunov functions are functional certificates for the system stability. Some results guarantee the existence of Lyapunov functions: a system is globally asymptotic stable if and only if there exists a Lyapunov function (see  e.g. [Kha02]). Being a Lyapunov function is a timeless property and only depends on dynamics. In few advantageous situations, Lyapunov functions can even be computed using numerical and optimization solvers (see the survey [GH15]).
Formally, the data of a peaks computation problem is decomposed into an autonomous discrete-time dynamical system and an objective function $\varphi$ over states. The dynamical system consists of a set of initial conditions $X^{\mathrm{i}n}$ and the dynamics of the system i.e. a function $T$ which maps a state to a new state. The associated peaks computation problem aims in maximizing $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{i}n}$ and $k$ is a natural integer. The optimal solution of a peaks computation problem is then a couple $(x_{\rm opt},k_{\rm opt})$ where $x_{\rm opt}\in X^{\mathrm{in}}$ and $k_{\rm opt}$ is a natural integer. Note that when an integer $k$ is fixed, maximizing $\varphi(T^{k}(x))$ over $x\in X^{\mathrm{in}}$ is a standard maximization problem. The main difficulty of peaks computation problem is to compute an optimal $k_{\rm opt}$ i.e. an integer $n$ for which the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is the greatest. Our approach is to find an upper bound $K$ of the (greatest) integer maximizer. We impose to the upper bound $K$ to satisfy the following condition : for all $n$ greater than $K$, the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is smaller than the supremum of $\varphi(T^{K}(x))$. Once this integer $K$ computed, to solve the peaks computation problem can be done by solving a finite number of maximization problems. We can extract the optimal solution $(x_{\rm opt},k_{\rm opt})$ from one of those maximization problems. Thus we reduce the peaks computation problem to find the smallest possible integer $K$ satisfying the previous condition. We, then, provide a general formula to compute such integer. This formula relies on well-chosen pairs of a strictly increasing and continuous function on $[0,1]$, $h$, and a scalar $\beta\in(0,1)$. In this case, well-chosen means that the pair $(h,\beta)$ forms a sequence $(h(\beta^{k}))_{k}$ such that $(h(\beta^{k}))_{k}$ is greater than the sequence of the suprema of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$. A well-chosen pair $(h,\beta)$ is said to be useful for our problem if $h(0)$ is strictly smaller than a supremum of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$. We prove that there exists an optimal useful pair, in the sense of that the pair permits to find the greatest optimal integer. In practice, this pair is not computable. Therefore, we propose two other methods to construct useful pairs. First, we extend the class of $\mathcal{KL}$ functions to the class of functions which are possibly discontinuous, unsigned (on the state variable but nonnegative on the time variable), increasing on states and decreasing on time. We call this class of functions $\mathcal{KL}_{\rm gen}$. As for classical $\mathcal{KL}$ functions, we use $\mathcal{KL}_{\rm gen}$ functions as upper bounds on $\varphi(T^{k}(x))$ (for the classical case with $\mathcal{K}\mathcal{L}$, $\varphi$ is a norm). We establish an equivalence theorem where we prove that a useful pair $(h,\beta)$ can be constructed from a $\mathcal{KL}_{\rm gen}$ upper bound function and we can construct $\mathcal{KL}_{\rm gen}$ upper bounds from useful pairs $(h,\beta)$. The construction of useful pairs $(h,\beta)$ from $\mathcal{KL}_{\rm gen}$ upper bound functions relies on an extension of Sontag’s result [Son98, Prop. 7] for $\mathcal{KL}_{\rm gen}$. Finally, we propose a new concept of Lyapunov functions adapted for the peaks computation problem namely the Opt-Lyapunov functions. The word adapted means the existence of an Opt-Lyapunov function is equivalent to the existence of a useful pair $(h,\beta)$. Once again, the proofs of each implication are constructive. The construction of an Opt-Lyapunov function from a useful pair $(h,\beta)$ is inspired by Yoshizawa construction of Lyapunov functions [Yos66]. The main difference between classical Lyapunov functions and this new concept is that we allow Opt-Lyapunov functions to have several zeros and to take infinity values. Those two new possibities make irrelevant the use of classical comparison functions (see e.g. [Kel14] for a comprehensive survey). In our context, when we have access to classical Lyapunov functions, we can link continuous $\varphi$ with classical Lyapunov functions through comparison functions. Without this link, we need to develop a concept of weaker relations between $\varphi$ and an Opt-Lyapunov function. In this paper, we introduce the notion of certificate of compatibility to overcome the loss of the link.
In summary, the main contributions of the paper are the following:
  • we propose a novel approach of the peaks computation problem by bounding the number of static optimization problem to solve to obtain the optimal value;
  • we obtain a general upper bound method for bounded from above sequence based on homeomorphic deformations of geometric sequences;
  • we generalize the concept of $\mathcal{K}\mathcal{L}$ functions to $\mathcal{KL}_{\rm gen}$ adapted to peaks computation problems for which we can formulate peaks computation problems as a $\mathcal{KL}_{\rm gen}$ upper bounds existence ;
  • we introduce a new concept of Lyapunov functions well designed for peaks computation problem for which the comparison functions approach is replaced by a compatibility notion.
The paper is organized as follows. Section 2 presents the peaks computation problem addressed in this paper and the main difficulty related to the search of an integer as an optimal solution of the peaks computations problem. Section 3 recalls some basic results about the supremum of sequences and how it is related to their asymptotic behaviour. In Section 3, we also find the fundamental result about the existence of a useful pair of a strictly increasing and continuous function on $[0,1]$ and a scalar in $(0,1)$. This pair furnishes, in theory, the exact integer optimal solution. However, this pair is difficult to compute in practice. Then, in Section 4, after developing a simple result using the classical $\mathcal{K}\mathcal{L}$ class, we introduce the new class $\mathcal{KL}_{\rm gen}$. We prove that the existence of $\mathcal{KL}_{\rm gen}$ upper bounds of $\varphi(T^{k}(x))$ is equivalent to the existence of a useful pair. In Section 5, we present the concept of Opt-Lyapunov functions well designed for our maximization problem. The classical comparison functions associated to classical Lyapunov functions are replaced by a new notion called compatibility. We also prove that the existence of a compatible Opt-Lyapunov function is equivalent to the existence of a useful pair. In Section 6, we summarize the results obtained highlighting the concrete constructions between useful pairs of strictly increasing and continuous functions on $[0,1]$ and convergent geometric sequences, $\mathcal{KL}_{\rm gen}$ class and Opt-Lyapunov functions.
Notations: $\mathbb{R}$ stands for the set of reals, $\mathbb{R}_{+}$ the set of nonnegative reals, $\mathbb{R}^{*}$ the set of nonzero reals and $\mathbb{R}_{+}^{*}$ the set of strictly positive reals. The vector space of vectors of $d$ reals is denoted by $\mathbb{R}^{d}$. The set of natural integers is denoted by $\mathbb{N}$ whereas $\mathbb{N}^{*}$ denotes the set of nonzero natural integers.
2 Problem Statement And Motivated Example
2.1 Problem Statement
Let $X^{\mathrm{in}}$ be a subset of $\mathbb{R}^{d}$ and $T$ a nonzero self-map on $\mathbb{R}^{d}$. Let us consider the associated discrete-time dynamical system defined as follows:
$x_{0}\in X^{\mathrm{in}},\ x_{k+1}=T(x_{k})\enspace.$
Let us consider $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$. We are interested in the reachable values that maximizes $\varphi$. More precisily, writing $T^{k}$ for the $k$-fold composition of $T$ or the identity when $k=0$, we are interested in solving the following maximization problem called a peaks computation problem:
Problem 1 (Peaks Computation Problem).
Maximize $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
Without loss of generality, we can suppose that $\varphi(0)=0$; otherwise, we change $\varphi$ by $\varphi-\varphi(0)$ and then we add $\varphi(0)$ at the end of the optimization procedure.
To solve Problem 1 means that we have to compute the optimal value $\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}$ and, if there exists, to characterize and to compute a couple of a vector $x\in X^{\mathrm{in}}$ and an integer $k$ which maximizes the value $\varphi(T^{k}(x))$. We can solve Problem 1 sequentially by solving a sufficient number of static maximization problems :
Problem Pk (Static Problems).
Maximize $\varphi(T^{k}(x))$ under the constraint $x\in X^{\mathrm{in}}$.
Some hypotheses can be made to ensure that Problem Pk has a finite optimal values and an optimal solution. For example, if $X^{\mathrm{in}}$ is compact and the function $\varphi\circ T^{k}$ is upper semicontinuous, Problem Pk has an optimal solution. About computations, some special instances of Problem Pk can be solved numerically more or less efficiently. For example, when $\varphi\circ T^{k}$ is concave and $X^{\mathrm{in}}$ is convex, the problem is said to be convex and numerous methods exist (see e.g. [BTN01, Ber15]), or when $\varphi\circ T^{k}$ is polynomial and $X^{\mathrm{in}}$ a basic semialgebraic set (see e.g. [AL11, Las15]). The main issue remains to guess the interesting $k$, the integer for which the optimal value of Problem Pk is the greatest. A direct computation of such $k$ seems to be very difficult. Hence, the method we use in this paper is to solve instances of Problem Pk for $k=0,\ldots,K$ where $K$ has to determined. The integer $K$ has the particularity of ensuring that the optimal value of Problem 1 corresponds to the optimal value of some Problem Pk where $k$ is smaller than $K$.
In this paper, we will adopt the following notations:
$\nu:=(\nu_{k})_{k\in\mathbb{N}}\text{ where }\forall\,k\in\mathbb{N},\ \nu_{k}:=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\text{ and }\nu_{\rm opt}:=\sup\{\nu_{k}:k\in\mathbb{N}\}$ (1)
Sometimes we will also use the terminlogy the supremum of $\nu$ to speak about $\nu_{\rm opt}$. The sequence $\nu$ corresponds to the sequence of optimal values of the maximization problems depicted at Problem Pk and $\nu_{\rm opt}$ is the optimal value of Problem 1. Following those notations, the main problem of the paper can be formulated mathematically as follows:
Problem 2 (Safe Number of Problem $P_{k}$ to Solve).
Find $K\in\mathbb{N}$ such that $\nu_{\rm opt}=\max\{\nu_{k}:k=0,\ldots,K\}$.
In this paper, we provide simple characterizations for which a solution $K$ of Problem 2 exists. Then we construct a formula based on pairs composed of a strictly increasing continuous functions on $[0,1]$ and a scalar in $(0,1)$. Finally, we propose constructions of those strictly increasing continuous functions on [0,1] based on extensions of classical tools on stability theory for discrete-time dynamical systems such as $\mathcal{K}\mathcal{L}$ class of functions and Lyapunov functions.
The developments made in the paper require real-valued sequences; hence the only assumption made in this paper is the following :
Assumption 1.
For all $k\in\mathbb{N},\ \nu_{k}=\sup\left\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\right\}$ is finite.
A simple situation where Assumption 1 holds when $X^{\mathrm{in}}$ is a bounded set, $T$ maps from bounded sets to bounded sets and $\varphi$ is upper semicontinuous.
2.2 An unstable linear system and an objective function with finite optimal value
Throughout the paper, we illustrate our method and the computations on the example presented in this subsection. The peaks computation problem can be solved more simply and we compare the results obtained with our methodology with the one obtained analytically.
Let us consider the following dynamics on $\mathbb{R}^{2}$:
$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}.$
We parameterize the initial conditions set by a scalar $\mu\in(0,1/3]$ and we define:
$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}.$
Finally, the objective function is parameterized by a positive scalar $p$:
$\mathbb{R}^{2}\ni(x,y):\mapsto\varphi(x,y):=y^{2}-x^{2}+px$
The associated peaks computation problem is the following:
$\begin{array}[]{rc}\displaystyle{\operatorname*{Max}_{(x,y),n}}&\varphi(A^{n}(x,y)^{\intercal})\\ s.t.&\left\{\begin{array}[]{l}(x,y)\in X^{\mathrm{in}}\\ n\in\mathbb{N}\end{array}\right.\end{array}$
The system is linear and the objective function is quadratic. However, we remark that $A$ admits $1/2$ and $3/2$ as eigenvalues. Therefore, $A$ is unstable and for some $x$, $\|A^{n}x\|$ goes to +$\infty$+ ∞ as $n$ goes to +$\infty$+ ∞. Therefore, we cannot use the method developed in [Adj23]. Indeed, the system does not admit a quadratic Lyapunov function. We also note that $\varphi$ is not concave nor convex.
As the matrix $A$ is diagonalizable, the powers of it are easy to compute and for all $n\in\mathbb{N}$ we have :
$A^{n}=\dfrac{1}{2^{n+1}}\begin{pmatrix}3^{n}+1&2(3^{n}-1)\\ \frac{3^{n}-1}{2}&3^{n}+1\end{pmatrix}\text{ and }\forall\,y\in\mathbb{R},\ A^{n}\begin{pmatrix}2y\\ y\end{pmatrix}=\left(\frac{3}{2}\right)^{n}\begin{pmatrix}2y\\ y\end{pmatrix}$
Moreover, the initial conditions set is actually a segment each element of it can be written as $(2y,y)^{\intercal}$ with $y\in[\mu,1/2]$. Using this, we have for all $y\in[\mu,1/2]$, for all $n\in\mathbb{N}$:
$\varphi_{n}(y):=\varphi(A^{n}(2y,y)^{\intercal})=-3\left(\dfrac{3}{2}\right)^{2n}y^{2}+2p\left(\dfrac{3}{2}\right)^{n}y$
Let $n\in\mathbb{N}$. The function $y\mapsto\varphi_{n}(y)$ is concave and infinitely derivable. It admits an unique maximizer on $\mathbb{R}$ which is $y_{n}^{*}:=\left(\frac{2}{3}\right)^{n}\frac{p}{3}$. The optimal value of $\varphi_{n}$ on $\mathbb{R}$ is $p^{2}/3$. The function $\varphi_{n}$ is strictly increasing before $y_{n}^{*}$ and strictly decreasing after. As we maximize on $[\mu,1/2]$, we have to check whether $y_{n}^{*}$ belongs to this interval. To avoid the case where $\frac{1}{2}$ is not used as a maximizer we assume that $p>3/2$. In this case, $y_{0}^{*}>1/2$.
As the sequence $(y_{n}^{*})_{n\in\mathbb{N}}$ strictly decreases to zero, we can define two integers:
$\underline{n}:=\lfloor\ln(3/2p)/\ln(2/3)\rfloor+1\text{ and }\overline{n}:=\lfloor\ln(3\mu/p)/\ln(2/3)\rfloor+1$
The integer $\underline{n}$ corresponds to the smallest integer such that $y_{n}^{*}<1/2$ whereas $\overline{n}$ is the smallest integer such that $y_{n}^{*}<\mu$. Moreover, as we assume that $\mu\in(0,1/3]$, we have $(\ln(3\mu/p)-\ln(3/2p))/\ln(2/3)=-(\ln(\mu)+\ln(2))\ln(3/2)\geq-(\ln(1/3)+\ln(2))/\ln(3/2)=1$. We conclude that $\overline{n}>\underline{n}$, as for all $x,y\in\mathbb{R}_{+}$ such that $x-y\geq 1$, we have $\lfloor x\rfloor\geq\lfloor y\rfloor+1$.
Finally using $\underline{n}$ and $\overline{n}$ , we can define the sequence of interest $(\nu_{n})_{n\in\mathbb{N}}$ and for all $n\in\mathbb{N}$:
$\nu_{n}=\left\{\begin{array}[]{lr}\varphi_{n}(1/2)&\text{ if }n\in[0,\underline{n}-1]\cap\mathbb{N}\\ \dfrac{p^{2}}{3}&\text{ if }n\in[\underline{n},\overline{n}-1]\cap\mathbb{N}\\ \varphi_{n}(\mu)&\text{ if }n\geq\overline{n}\end{array}\right.$ (2)
From now on, when we refer to this example, we will assume that $\mu\in(0,1/3]$ and $p\geq 3/2$.
3 Suprema and Rank Maximizer of General Sequences
3.1 Notations and Basic Results
We introduce the set of real sequences which are bounded from above:
$\Lambda:=\{u=(u_{0},u_{1},\ldots)\in\mathbb{R}^{\mathbb{N}}:\sup_{n\in\mathbb{N}}u_{n}<+\infty\}.$
For an element of $\Lambda$, we are interested in computing the supremum of its terms.
Proposition 1.
$u\in\Lambda\iff\displaystyle{\limsup_{n\to+\infty}u_{n}\in\mathbb{R}\cup\{-\infty\}}$.
For $u\in\Lambda$, we are also looking for the set of maximizers i.e. $\operatorname{Argmax}(u):=\{k\in\mathbb{N}:u_{k}=\displaystyle{\sup_{n\in\mathbb{N}}u_{n}}\}$. For computations purpose, we need to characterize $\operatorname{Argmax}(u)$. Consequently, we introduce, for $u\in\Lambda$, the two sets of ranks:
$\Delta_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}\right\}\ \text{ and }\ \Delta^{s}_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}>\sup_{j>k}u_{j}\right\}$
It is easy to see that $\Delta_{u}\subseteq\Delta^{s}_{u}$. Moreover, if $k$ belongs to $\Delta_{u}$ (resp. $\Delta^{s}_{u}$), then any integer greater than $k$ belongs to $\Delta_{u}$ (resp. $\Delta^{s}_{u}$). Besides, we will need the set of ranks for which the associated term is greater than the limsup (resp. strictly greater):
${\mathrm{R}}_{u}:=\left\{k\in\mathbb{N}:u_{k}\geq\limsup_{n\to+\infty}u_{n}\right\}\ \text{ and }\ {\mathrm{R}}_{u}^{s}:=\left\{k\in\mathbb{N}:u_{k}>\limsup_{n\to+\infty}u_{n}\right\}$
It is obvious that ${\mathrm{R}}_{u}^{s}\subseteq{\mathrm{R}}_{u}$. Note that even if $u\in\Lambda$, ${\mathrm{R}}_{u}$ and hence ${\mathrm{R}}_{u}^{s}$ can be empty. To construct the characterizations, we need to consider the smallest elements of $\Delta_{u}$, $\Delta^{s}_{u}$, ${\mathrm{R}}_{u}$ and ${\mathrm{R}}_{u}^{s}$.
$\mathbf{K}_{u}=\inf\Delta_{u};\ \qquad\mathbf{K}^{s}_{u}=\inf\Delta^{s}_{u};\ \qquad k_{u}=\inf{\mathrm{R}}_{u}\quad\text{ and }\ k^{s}_{u}=\inf{\mathrm{R}}_{u}^{s}$
We recall that the infimum of the empty set is equal to +$\infty$+ ∞. Hence, $\mathbf{K}_{u}<+\infty$ (resp. $\mathbf{K}^{s}_{u}<+\infty$) if and only if $\Delta_{u}\neq\emptyset$ (resp. $\Delta^{s}_{u}\neq\emptyset$); and $k_{u}<+\infty$ (resp. $k^{s}_{u}<+\infty$) if and only if ${\mathrm{R}}_{u}\neq\emptyset$ (resp. ${\mathrm{R}}_{u}^{s}\neq\emptyset$). Furthermore, if the sets of integers defined above are nonempty, they contain their infimum i.e. $\mathbf{K}_{u}\in\Delta_{u}$, $\mathbf{K}^{s}_{u}\in\Delta^{s}_{u}$, $k_{u}\in{\mathrm{R}}_{u}$ and $k^{s}_{u}\in{\mathrm{R}}_{u}^{s}$ when those integers are finite.
Proposition 2.
If $u\notin\Lambda$ then $\Delta_{u}=\emptyset$ and ${\mathrm{R}}_{u}=\emptyset$. Obviously, it also implies that $\Delta^{s}_{u}=\emptyset$ and ${\mathrm{R}}_{u}^{s}=\emptyset$.
Proposition 3.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    $\Delta^{s}_{u}\neq\emptyset\iff{\mathrm{R}}_{u}^{s}\neq\emptyset$;
  2. 2.
    $\Delta_{u}\neq\emptyset\iff{\mathrm{R}}_{u}\neq\emptyset$;
  3. 3.
    $\Delta^{s}_{u}=\emptyset\iff\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$.
Proposition 4.
The following inequalities hold:
$k_{u}\leq k^{s}_{u};\qquad k_{u}\leq\mathbf{K}_{u};\qquad k^{s}_{u}\leq\mathbf{K}^{s}_{u}\quad\text{ and }\mathbf{K}_{u}\leq\mathbf{K}^{s}_{u}\enspace.$
Moreover, if $k^{s}_{u}<+\infty$, then $k^{s}_{u}\leq\mathbf{K}_{u}$.
Proposition 5.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    $\operatorname{Argmax}(u)\subseteq\Delta_{u}$;
  2. 2.
    $\operatorname{Argmax}(u)\neq\emptyset\iff\Delta_{u}\neq\emptyset$;
  3. 3.
    $\mathbf{K}_{u}=\inf\operatorname{Argmax}(u)$;
  4. 4.
    $\operatorname{Argmax}(u)=\emptyset$ or $\operatorname{Argmax}(u)$ nonempty and unbounded $\implies\sup_{n\in\mathbb{N}}u_{n}=\limsup_{k\to+\infty}u_{k}$;
  5. 5.
    If $\Delta^{s}_{u}\neq\emptyset$ then $\mathbf{K}^{s}_{u}=\max\operatorname{Argmax}(u)$.
Note that $\operatorname{Argmax}(u)$ can be nonempty and bounded when $\Delta^{s}_{u}$ is empty. In this case, by the third statement of Prop 3, we have $\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$. However, the sequence finitely touches $\limsup_{n\to+\infty}u_{n}$ before approaching, up to a subsequence, $\limsup_{n\to+\infty}u_{n}$ from below (strictly).
Example 1.
Let us consider the peaks computation problem depicted at Subsection 2.2. The sequence $(\nu_{k})_{k\in\mathbb{N}}$ presented at Eq. (2) has a greatest maximizer as $\limsup_{k\to+\infty}\nu_{k}=\lim_{k\to+\infty}\varphi_{k}(\mu)=-\infty$. The set $\operatorname{Argmax}(u)$ would not be bounded from above if $\mu=0$ and thus $\overline{n}=+\infty$. Indeed, in this case, we would get $\limsup_{k\to+\infty}\nu_{k}=p^{2}/3=\sup_{k\in\mathbb{N}}\nu_{k}$. ∎
Proposition 6.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    Assume that ${\mathrm{R}}_{u}\neq\emptyset$. Let $k\geq k_{u}$ such that $\sup_{k_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
  2. 2.
    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $k\geq k^{s}_{u}$ such that $\sup_{k^{s}_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
  3. 3.
    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. For all $\mathbf{K}_{u}\leq k$,
    $\sup_{0\leq j\leq k}u_{j}=\sup_{k^{s}_{u}\leq j\leq k}u_{j}=\sup_{k_{u}\leq j\leq k}u_{j}=\sup_{j\in\mathbb{N}}u_{j}\enspace.$
In summary, for $u\in\Lambda$, to compute $\sup_{k\in\mathbb{N}}u_{k}$, we need to study first the emptiness of ${\mathrm{R}}_{u}^{s}$. Indeed, if ${\mathrm{R}}_{u}^{s}$ is empty, we know that (the third statement of Prop. 3) $\sup_{n\in\mathbb{N}}u_{n}$ is equal to $\limsup_{n}u_{n}$. If ${\mathrm{R}}_{u}^{s}$ is not empty, then we have to compute $\mathbf{K}_{u}$ to find $\sup_{k\in\mathbb{N}}u_{k}$ and a maximizer (the second statement of Prop. 5). However, to identify $\mathbf{K}_{u}$, we need knowledge on the past and the future of the sequence. The good point is that any over-approximation $k$ of $\mathbf{K}_{u}$ permits to know $\sup_{k\in\mathbb{N}}u_{k}$ by computing $\sup_{0\leq j\leq k}u_{j}$ (the third statement of Prop 6).
3.2 A Formula Based on Pairs of Strictly Increasing Continuous Functions and Convergent Geometric Sequences
Let us write $\mathbb{I}\left(\mathbb{R}\right)$ the set of all intervals (convex sets) on $\mathbb{R}$ (open, closed, or semi-open). For a function $f$ from $\mathbb{R}$ to $\mathbb{R}$, we define the set of intervals:
$\mathrm{SIC}(f)=\{I\in\mathbb{I}\left(\mathbb{R}\right):f:I\mapsto f(I)\text{ is strictly increasing and continuous }\}.$
Recall that a strictly increasing function $f$ is continuous on an interval $[a,b]$ if and only if $f$ is surjective from $[a;b]$ to $[f(a);f(b)]$. Hence, if $I$ belongs to $\mathrm{SIC}(f)$, then $f$ is bijective from $I$ to $f(I)$ and we can define its inverse on $I$ i.e. a function $g:f(I)\mapsto I$ such that $f\circ g=g\circ f=\mathrm{Id}$ where $\mathrm{Id}:x\mapsto x$ (viewed as a function from $I$ to itself). For $I\in\mathrm{SIC}(f)$, we denote by ${{f}_{\left(I\right)}^{-1}}$ the inverse of $f$ on $I$. Note that ${{f}_{\left(I\right)}^{-1}}$ is also strictly increasing and continuous (surjective) from $f(I)$ to $I$. Finally, we pay a particular attention to functions strictly increasing and continuous on $[0,1]$ and we write :
$\Omega([0,1]):=\{h:\mathbb{R}\mapsto\mathbb{R}:[0,1]\in\mathrm{SIC}(h)\}.$
The functions of $\Omega([0,1])$ are part of the upper bound on the sequence of interest $u$. To complete the upper bound, we need a convergent geometric sequence. We introduce, for $u\in\mathbb{R}^{\mathbb{N}}$ the following sets:
$\mathbb{\Gamma}(u):=\{(h,\beta)\in\Omega([0,1])\times(0,1):u_{k}\leq h(\beta^{k}),\ \forall\,k\in\mathbb{N}\}$
and
$\mathbb{\Gamma}_{\rm fun}(u):=\{h\in\Omega([0,1]):\exists\,\beta\in(0,1)\text{s.t. }(h,\beta)\in\mathbb{\Gamma}(u)\}\enspace.$
Finally, we introduce for $u\in\mathbb{R}^{\mathbb{N}}$ and $h\in\mathbb{\Gamma}_{\rm fun}(u)$:
$\mathbb{\Gamma}_{\rm sc}(u,h):=\{\beta\in(0,1):(h,\beta)\in\mathbb{\Gamma}(u)\}\enspace.$
[…]
Proposition 7.
Let $u\in\Lambda$. Then :
  1. 1.
    For all $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we have $\limsup_{k\to+\infty}u_{k}\leq h(0)$;
  2. 2.
    For all $k\in\mathbb{N}^{*}$, $u_{k}<h(1)$ and $u_{0}\leq h(1)$.
In Prop. 7, the first statement only requires $h$ to be upper semicontinuous. However, a strictly increasing function $f$ from $I$ to $f(I)$ is surjective if and only if it is continuous on $I$.
Corollary 1.
Let $u\in\Lambda$. If there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ such that $u_{0}=h(1)$ then $u_{0}=\max_{k\in\mathbb{N}}u_{k}$.
It is easy to find a pair of a strictly increasing continuous function and a convergent geometric sequence to construct an upper bound of the sequence $u$ when the supremum of $u$ is finite.
Proposition 8.
$u\in\Lambda$ if and only if $\mathbb{\Gamma}(u)\neq\emptyset$.
Proof.
Suppose $u\in\Lambda$. By definition, $\sup_{n\in\mathbb{N}}u_{n}\in\mathbb{R}$ and the function $h:x\mapsto x+\sup_{n\in\mathbb{N}}u_{n}$ belongs to $\Omega([0,1])$ and for all $\beta\in(0,1)$ and all $k\in\mathbb{N}$, $u_{k}\leq\beta^{k}+\sup_{n\in\mathbb{N}}u_{n}$ and $\mathbb{\Gamma}(u)\neq\emptyset$. Now, let $u\in\mathbb{R}^{\mathbb{N}}$ and suppose that $\mathbb{\Gamma}(u)\neq\emptyset$. Let $(h,\beta)\in\mathbb{\Gamma}(u)$. From Prop. 7 and as $h$ is finite-valued then $\limsup_{k}u_{k}\leq h(0)\in\mathbb{R}$. From Prop. 1, this implies that $u\in\Lambda$. ∎
Some upper bounds $h(\beta^{k})$ appearing in $\mathbb{\Gamma}(u)$ cannot be used in our formula. It depends on how $h(0)$ is placed with respect to the terms of $u$.
Definition 1 (Useful Strictly Increasing Continuous Functions).
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. A function $h\in\mathbb{\Gamma}_{\rm fun}(u)$ is said to be useful for $u$ if the set :
$\mathcal{S}(u,h):=\{k\in\mathbb{N}:u_{k}>h(0)\}$ (3)
is nonempty.
By extension, a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ is useful for $u$ if $h$ is useful for $u$.
From the first statement of Prop. 7, if $h$ is useful for $u$, there exists $k\in\mathbb{N}$ such that $u_{k}>h(0)\geq\limsup_{n\to+\infty}u_{n}$ i.e. ${\mathrm{R}}_{u}^{s}\neq\emptyset$. We will see that the contrary also holds for i.e. if ${\mathrm{R}}_{u}^{s}\neq\emptyset$, then there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$.
In Prop. 9, we present some theoretical properties on $\mathbb{\Gamma}_{\rm fun}(u)$ and $\mathbb{\Gamma}_{\rm sc}(u,h)$.
Proposition 9.
Let $u\in\Lambda$.
  1. 1.
    Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$. Let us define :
    $\underline{\beta}_{u,h}:=\left\{\begin{array}[]{lr}\displaystyle{\sup_{k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}}\exp\left(\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{k}\right)}&\text{ if }\mathcal{S}(u,h)\cap\mathbb{N}^{*}\neq\emptyset\\ \\ 0&\text{ otherwise}\end{array}\right.;$
    Then $\mathbb{\Gamma}_{\rm sc}(u,h)$ is the interval $(0,1)$ if $\underline{\beta}_{u,h}=0$ and $[\underline{\beta}_{u,h},1)$ if $\underline{\beta}_{u,h}>0$.
  2. 2.
    For all finite family families $\{(h_{1},\beta_{1}),(h_{2},\beta_{2}),\ldots,(h_{n},\beta_{n})\}\subset\mathbb{\Gamma}(u)$:
    $\max_{i=1,\ldots,n}\beta_{i}\in\bigcap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i});$
  3. 3.
    The set $\mathbb{\Gamma}_{\rm fun}(u)$ is a convex sublattice (for the classical order, $\min$ and $\max$ on functions) subset of $\Omega([0,1])$ i.e.
    • (lattice) for all finite families $\{h_{1},h_{2},\ldots,h_{n}\}\subset\mathbb{\Gamma}_{\rm fun}(u)$, $\min_{i=1,\ldots,n}h_{i}\in\mathbb{\Gamma}_{\rm fun}(u)$ and $\max_{i=1,\ldots,n}h_{i}\in\mathbb{\Gamma}_{\rm fun}(u)$;
    • (convex) for all $t\in[0,1]$ and $h,g\in\mathbb{\Gamma}_{\rm fun}(u)$, $th+(1-t)g\in\mathbb{\Gamma}_{\rm fun}(u)$.
  4. 4.
    For all finite families $\{h_{1},h_{2},\ldots,h_{n}\}\subset\mathbb{\Gamma}_{\rm fun}(u)$, we have actually:
    $\mathbb{\Gamma}_{\rm sc}(u,\min_{i=1,\ldots,n}h_{i})=\bigcap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})\subseteq\mathbb{\Gamma}_{\rm sc}(u,\max_{i=1,\ldots,n}h_{i})$
Proof.
1. Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$. As $h$ is strictly increasing on $[0,1]$, it follows that if $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ it also holds for all $\beta^{\prime}\in[\beta,1)$. Now, it suffices to prove that $\underline{\beta}_{u,h}$ is the infimum of $\mathbb{\Gamma}_{\rm sc}(u,h)$ and it belongs to $\mathbb{\Gamma}_{\rm sc}(u,h)$ when it is strictly positive. First, we suppose that $\underline{\beta}_{u,h}=0$. By definition, this is the same as for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h(0)$. As $h$ is strictly increasing, we have for all $\beta\in(0,1)$, for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h(0)<h(\beta^{k})$. We have also $u_{0}\leq h(1)=h(\beta^{0})$ for all $\beta\in(0,1)$ as $h\in\mathbb{\Gamma}_{\rm fun}(u)$. Then we have $\mathbb{\Gamma}_{\rm sc}(u,h)=(0,1)$ and $\underline{\beta}_{u,h}=0=\inf\mathbb{\Gamma}_{\rm sc}(u,h)$. Now, suppose that $\underline{\beta}_{u,h}>0$. If $k=0$, $u_{0}\leq h(\underline{\beta}_{u,h}^{0})$ as $h\in\mathbb{\Gamma}_{\rm fun}(u)$. If $k\notin\mathcal{S}(u,h)$ and $k\neq 0$ then $u_{k}\leq h(0)<h(\underline{\beta}_{u,h}^{k})$. Let $k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}$, $\exp(\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))/k)\leq\underline{\beta}_{u,h}$ and thus ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\leq\underline{\beta}_{u,h}^{k}$ and then $u_{k}=h(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))\leq h(\underline{\beta}_{u,h}^{k})$. Let consider $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$. We have for all $k\in\mathbb{N}$, $u_{k}\leq h(\beta^{k})$. Let $k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}$. By simple calculus, we have $\exp(\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))/k)\leq\beta$. Taking the supremum over $\mathcal{S}(u,h)\cap\mathbb{N}^{*}$, we obtain $\underline{\beta}_{u,h}\leq\beta<1$. We have proved that $\underline{\beta}_{u,h}\in(0,1)$, for all $k\in\mathbb{N}$, $u_{k}\leq h(\underline{\beta}_{u,h}^{k})$ and for all $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$, $\underline{\beta}_{u,h}\leq\beta$ We conclude that $\underline{\beta}_{u,h}\in\mathbb{\Gamma}_{\rm sc}(u,h)$ and $\underline{\beta}_{u,h}=\inf\mathbb{\Gamma}_{\rm sc}(u,h)$.
2. The result follows from the fact that all elements of $\mathbb{\Gamma}_{\rm fun}(u)$ and for all $k\in\mathbb{N}^{*}$, the functions $x\mapsto x^{k}$ are increasing on the positive reals set.
3. (Lattice) First, the set $\Omega([0,1])$ is a lattice as finitely $\min$ or $\max$ of strictly increasing continuous functions are still strictly increasing and continuous. Now, it suffices to prove the results for the case $n=2$. Let $h_{1},h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$. There exist $\beta_{1},\beta_{2}\in(0,1)$ such that for all $k\in\mathbb{N}$, $u_{k}\leq h_{1}(\beta_{1}^{k})$ and $u_{k}\leq h_{2}(\beta_{2}^{k})$. Let us write $\beta=\max\{\beta_{1},\beta_{2}\}$. Then for all $k\in\mathbb{N}$, $u_{k}\leq\min\{h_{1}(\beta_{1}^{k}),h_{2}(\beta_{2}^{k})\}\leq\min\{h_{1}(\beta^{k}),h_{2}(\beta^{k}\}=\min\{h_{1},h_{2}\}(\beta^{k})$. The same techniques can be used for the maximum.
(Convexity) Let $t\in[0,1]$, $h_{1},h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$. There exist $\beta_{1},\beta_{2}\in(0,1)$ such that for all $k\in\mathbb{N}$, $u_{k}\leq h_{1}(\beta_{1}^{k})$ and $u_{k}\leq h_{2}(\beta_{2}^{k})$. Then $u_{k}=tu_{k}+(1-t)u_{k}\leq th_{1}(\beta_{1}^{k})+(1-t)h_{2}(\beta_{2}^{k})\leq th_{1}(\max\{\beta_{1},\beta_{2}\}^{k})+(1-t)h_{2}(\max\{\beta_{1},\beta_{2}\}^{k})=(th_{1}+(1-t)h_{2})(\max\{\beta_{1},\beta_{2}\}^{k})$. We conclude that $th_{1}+(1-t)h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$.
4. Let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,\min_{i=1,\ldots,n}h_{i})$. Then $u_{k}\leq\min_{i=1,\ldots,n}h_{i}(\beta^{k})\leq h_{i}(\beta^{k})$ for all $k\in\mathbb{N}$. Hence, $\beta\in\cap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})$. Now, let $\beta\in\cap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})$. For all $i=1,\ldots,n$, for all $k\in\mathbb{N}$, $u_{k}\leq h_{i}(\beta^{k})$. Taking the minimum over $i$ leads to the conclusion. The inclusion on the right-hand side follows readily from the definition of the finite maximum of functions. ∎
We present the main result of this section. It asserts that we can find an optimal useful affine function for $u$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
Theorem 1 (Existence of a Useful Optimal Affine Function).
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exist $a>0$, $b\in(0,1)$ and $c\in\mathbb{R}$ such that:
  1. 1.
    $\{k\in\mathbb{N}:u_{k}>c\}\neq\emptyset$;
  2. 2.
    $u_{k}\leq ab^{k}+c$ for all $k\in\mathbb{N}$;
  3. 3.
    $u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c$.
Proof.
If ${\mathrm{R}}_{u}^{s}$ is nonempty then $u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}>\limsup_{k\to\infty}u_{k}\in\mathbb{R}\cup\{-\infty\}$ and for all $t\in(\limsup_{k\to\infty}u_{k},\max_{k\in\mathbb{N}}u_{k})$ there exists $k\in\mathbb{N}$, such that $u_{k}>t$. So let $c$ be in $(\limsup_{k\to\infty}u_{k},\sup_{k\in\mathbb{N}}u_{k})$. Now as $c>\limsup_{k\to+\infty}u_{k}$, there exists $N_{c}\in\mathbb{N}$ s.t. $\sup_{k\geq N_{c}}x_{k}\leq c$. This is the same as for all $k\geq N_{c}$, $u_{k}\leq c$. We conclude that for all $k\geq N_{c}$, $u_{k}\leq c+ab^{k}$ for all $a>0$ and all $b\in(0,1)$. Note that as $c<\max_{k\in\mathbb{N}}u_{k}$ and $\mathbf{K}^{s}_{u}$ is the greatest element of the argmax of $u$ then $N_{c}$ is strictly greater than $\mathbf{K}^{s}_{u}$. Now let introduce the set :
$S_{c}:=\{(a,b)\in\mathbb{R}_{+}^{*}\times(0,1):u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}=ab^{\mathbf{K}^{s}_{u}}+c\}.$
It is easy to see that $S_{c}$ is nonempty, indeed, for all $b$ in $(0,1)$, $(\frac{u_{\mathbf{K}^{s}_{u}}-c}{b^{\mathbf{K}^{s}_{u}}},b)\in S_{c}$. Moreover, if $\mathbf{K}^{s}_{u}>0$, for all $(a,b)\in S_{c}$, for all $0<k<\mathbf{K}^{s}_{u}$, $u_{k}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c<ab^{k}+c$ as $b\in(0,1)$. We also have $u_{0}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c\leq a+c$ as $b\in(0,1)$. Hence, the result $u_{k}\leq ab^{k}+c$ holds on the sets $[0,\mathbf{K}^{s}_{u}]\cap\mathbb{N}$ and $[N_{c},+\infty)\cap\mathbb{N}$ with any element $(a,b)\in S_{c}$. It suffices to find $(a,b)\in S_{c}$, such that $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]\cap\mathbb{N}$. Recall that we have $\mathbf{K}^{s}_{u}+1\leq N_{c}$.
Let us write $(-\infty,0)\ni\gamma:=\max_{k\in(\mathbf{K}^{s}_{u},N_{c}]}(u_{\mathbf{K}^{s}_{u}}-u_{k})/(\mathbf{K}^{s}_{u}-k)$. This cannot be zero since $\mathbf{K}^{s}_{u}$ is the maximal integer $k$ such that $u_{k}=\max_{n\in\mathbb{N}}u_{n}$. Now, let $(a,b)\in\mathbb{R}_{+}\times(0,1)$ and define $f_{a,b}:\mathbb{R}\ni y\mapsto ab^{y}+c$. The function $f_{a,b}$ is derivable and strictly convex and $f^{\prime}_{a,b}(y)=ab^{y}\ln(b)$ for all $y\in\mathbb{R}$. If $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in S_{c}$ then, by convexity of $f_{a,b}$, the tangent of $f$ at $y=\mathbf{K}^{s}_{u}$ satisfies for all $k\in\mathbb{N}$, $f(k)\geq f^{\prime}(\mathbf{K}^{s}_{u})(k-\mathbf{K}^{s}_{u})+f(\mathbf{K}^{s}_{u})$ which is the same as $ab^{k}+c\geq\gamma(k-\mathbf{K}^{s}_{u})+u_{\mathbf{K}^{s}_{u}}$. By definition of $\gamma$, we have for all $k\in(\mathbf{K}^{s}_{u},N_{c})$, $(\mathbf{K}^{s}_{u}-k)\gamma\leq u_{\mathbf{K}^{s}_{u}}-u_{k}$ and so $u_{k}\leq u_{\mathbf{K}^{s}_{u}}+(k-\mathbf{K}^{s}_{u})\gamma$. We conclude that if $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in S_{c}$, $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]$. Finally, we are looking for $(a,b)\in\mathbb{R}_{+}\times(0,1)$ such that :
$f_{a,b}(\mathbf{K}^{s}_{u})=u_{\mathbf{K}^{s}_{u}}\text{ i.e. }(a,b)\in S_{c}\text{ and }f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$
From $(a,b)\in S_{c}$, we get $ab^{\mathbf{K}^{s}_{u}}=u_{\mathbf{K}^{s}_{u}}-c$. Hence, $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ if and only if $\ln(b)=\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c}<0$. So $b=\exp(\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c})\in(0,1)$ and $a=(u_{\mathbf{K}^{s}_{u}}-c)\exp(-\frac{\mathbf{K}^{s}_{u}\gamma}{u_{\mathbf{K}^{s}_{u}}-c})>0$. ∎
Th 1 brings deeper details compared with Prop. 8. The existence of a useful and optimal function is important for the construction and the optimality of the proposed formula. Moreover, the proof of Th 1 exhibits the parameters of the affine function to choose. Unfortunately (but naturally) all the parameters depend on $\mathbf{K}^{s}_{u}$ and $u_{\mathbf{K}^{s}_{u}}$ what is exactly what we want to compute.
Corollary 2.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exists a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ useful for $u$ such that $u_{\mathbf{K}^{s}_{u}}=h(\beta^{\mathbf{K}^{s}_{u}})$.
Proposition 10.
Let $u\in\Lambda$. Then ${\mathrm{R}}_{u}^{s}\neq\emptyset$ if and only if there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$.
Proof.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. From Th 1, there exists a function $h$ of the form $x\mapsto ax+c$ which satisfies $\mathcal{S}(u,h)\neq\emptyset$. For the reverse implication, let assume ${\mathrm{R}}_{u}^{s}=\emptyset$. From Prop. 7, for all $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we have $\mathcal{S}(u,h)\subseteq{\mathrm{R}}_{u}^{s}$. This proves that the set $\mathcal{S}(u,h)$ is empty for all $h\in\mathbb{\Gamma}_{\rm fun}(u)$. ∎
Example 2.
Let us consider the peaks computation problem mentioned at Subsection 2.2. We now construct an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for the sequence $\nu$ presented at Eq. (2).
Let $n_{0}$ be the smallest integer $n$ such that $\varphi_{n}(\mu)<0$. With simple computations, we get $n_{0}=\lfloor\ln(2p/(3\mu))/\ln(3/2)\rfloor+1$. As $\ln(2p/(3\mu))/\ln(3/2)-\ln(3\mu/p)/\ln(2/3)=(\ln(2p)-\ln(p))/\ln(3/2)$, we have $n^{0}\geq\overline{n}$. For all $n\geq\overline{n}$, $\nu_{n}=\varphi_{n}(\mu)$ and thus for all $n\geq n_{0}$, $\nu_{n}<0$.
Following the proof of Th 1, we can impose to $h$ to be linear i.e. $h:x\mapsto ax$. Indeed, the value $c$ of Th 1 can be chosen to be 0 as we have $\lim_{k\to+\infty}\nu_{k}=-\infty$ and $p^{2}/3>0$. In this case, if $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$, $h$ is useful for $\nu$. A simple upper bound over $(\nu_{n})_{n}$ is furnished from any couple $(a,\beta)$ such that:
$a>0,\ \beta\in(0,1)\text{ and }a\beta^{n_{0}}\geq p^{2}/3$
For example, we can choose $a\beta^{n_{0}}=p^{2}/3$ for which $a=2p^{2}/3$ and $\beta=2^{-1/n_{0}}$ is a solution. ∎
Th. 1 provides a theoretical result and to have in hand an optimal function $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$ is very unlikely to happen. However, if we have any function $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we are able to produce a safe upper bound of $\mathbf{K}^{s}_{u}$ and then a solution of Problem 2. To compute it, we define, for $u\in\mathbb{R}^{\mathbb{N}}$, we introduce the function $\mathfrak{F}_{u}$, from ×$\mathbb{N}\times\Omega([0,1])\times(0,1)\mapsto\mathbb{R}\cup\{+\infty\}$ defined for all $k\in\mathbb{N}$, and $h\in\Omega([0,1])$ and $\beta\in(0,1)$ by:
$\mathfrak{F}_{u}(k,h,\beta):=\left\{\begin{array}[]{lr}\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{\ln(\beta)}&\text{ if }(h,\beta)\in\mathbb{\Gamma}(u)\text{ and }k\in\mathcal{S}(u,h)\\ +\infty&\text{ otherwise}\end{array}\right.$ (4)
We will show that $\mathfrak{F}_{u}$ has the desired properties to solve Problem 2.
Proposition 11.
Let us take a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ useful for $u$ and let $k\in\mathcal{S}(u,h)$. The following statements hold:
  1. 1.
    $\mathfrak{F}_{u}(k,h,\beta)$ is well-defined and strictly positive if $u_{0}<h(1)$ and null if $k=0$ and $u_{0}=h(1)$;
  2. 2.
    $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1=\min\{j\in\mathbb{N}:h(\beta^{j})<u_{k}\}=\min\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$;
  3. 3.
    $\mathfrak{F}_{u}(k,h,\beta)\geq k$;
  4. 4.
    For all $j\in\mathbb{N}$, $j>\mathfrak{F}_{u}(k,h,\beta)$, $u_{j}<u_{k}$;
  5. 5.
    Let $j\in\mathcal{S}(u,h)$. If $u_{j}\leq u_{k}$ then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
Proof.
1. Let $(h,\beta)\in\mathbb{\Gamma}(u)$. As $h\in\mathbb{\Gamma}_{\rm fun}(u)$ and we only consider $u_{k}>h(0)$, we have ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})>0$. We also have by definition of $h$, if $k\in\mathbb{N}^{*}$, $u_{k}<h(1)$ meaning that ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\in(0,1)$. If $u_{0}<h(1)$, the strict inequality is also valid at $k=0$. By definition, $\beta\in(0,1)$. We can apply the natural logarithm to ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})$ and $\beta$; those values are both strictly negative and then the quotient is positive. If we have $u_{0}=h(1)$ and $k=0$, $\mathfrak{F}_{u}(k,h,\beta)=0$.
2. As $h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and $k\in\mathcal{S}(u,h)$, there exists $j\in\mathbb{N}$ such that $h(\beta^{j})<u_{k}$. Then the sets $\{j\in\mathbb{N}:h(\beta^{j})<u_{k}\}$ and $\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$ are equal and nonempty. Let $j\in\mathbb{N}$ s.t. $\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})$. Applying the natural logarithm and dividing by $\ln(\beta)<0$ leads to $j>\mathfrak{F}_{u}(k,h,\beta)$ and as $j\in\mathbb{N}$ it implies that $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$. By definition of integer parts and as $\beta\in(0,1)$, $\ln(\beta)(\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1)<\mathfrak{F}_{u}(k,h,\beta)\ln(\beta)=\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))$. Applying the exponential leads to $\beta^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})$ and $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ belongs to $\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$ and is its minimum integer.
3. The pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(u)$ and thus for all $k\in\mathbb{N}$, $u_{k}\leq h(\beta^{k})$ which is the same as ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\leq\beta^{k}$. Applying the natural logarithm and dividing by $\ln(\beta)$ leads to the result.
4. Let $j>\mathfrak{F}_{u}(k,h,\beta)$ then $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ and, as $h$ is increasing, $h(\beta^{j})\leq h(\beta^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1})<u_{k}$ from the second statement. Moreover, by definition of $(h,\beta)$, we have $u_{j}\leq h(\beta^{j})$. It follows that $u_{j}<u_{k}$.
[…]
We relate some structural properties of the sets $\mathbb{\Gamma}_{\rm fun}(u)$ and $\mathbb{\Gamma}_{\rm sc}(u,h)$ depicted at Prop. 9 with the computations of $\mathfrak{F}_{u}(k,h,\beta)$.
Proposition 12.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. The following properties hold:
  1. 1.
    Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$ and $k\in\mathcal{S}(u,h)$. Let $\beta,\beta^{\prime}\in\mathbb{\Gamma}_{\rm sc}(u,h)$ such that $\beta\leq\beta^{\prime}$. Then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(k,h,\beta^{\prime})$.
  2. 2.
    Let $g,h\in\mathbb{\Gamma}_{\rm fun}(u)$. Let $k\in\mathcal{S}(u,\min\{g,h\})$. Let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,\min\{g,h\})$. Then $\mathfrak{F}_{u}(k,\min\{g,h\},\beta)\leq\min\{\mathfrak{F}_{u}(k,g,\beta),\mathfrak{F}_{u}(k,h,\beta)\}$.
Proof.
1. The function $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is increasing as it is the composition of (×$(\times\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})))\circ{\cdot}^{-1}\circ\ln$ where $\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))$ is negative.
2. First from Prop 9, $f:=\min\{g,h\}\in\mathbb{\Gamma}_{\rm fun}(u)$. Second, for all $x\in[0,1]$, $f(x)\leq g(x)$ and we can apply ${{f}_{\left([0,1]\right)}^{-1}}$ which is strictly increasing on $[0,1]$ and $x\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(g(x))$. Thus, we write $y=g(x)$ and thus $x=\displaystyle{{g}_{\left([0,1]\right)}^{-1}}(y)$. We conclude that ${{g}_{\left([0,1]\right)}^{-1}}(y)\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(y)$. Using the same technique, we can prove that ${{h}_{\left([0,1]\right)}^{-1}}(y)\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(y)$. The result follows from the fact that $\ln$ is increasing and $\ln(\beta)$ is negative. ∎
The first statement of Prop. 12 has to be read as follows : we have to choose the smallest possible $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ to reduce the approximation gap. The second statement has a similar meaning in a functional sense : if we have several functions in $\mathbb{\Gamma}_{\rm fun}(u)$ we have to consider the minimum functions to reduce the overapproximation gap.
Proposition 13.
Let $u\in\Lambda$ and $(h,\beta)\in\mathbb{\Gamma}(u)$. Then for all $j\in\mathbb{N}$, $\mathbf{K}^{s}_{u}\leq\lfloor\mathfrak{F}_{u}(j,h,\beta)\rfloor$. Moreover, if $h$ is useful for $u$, we have:
$\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)\enspace.$
Proof.
By definition of $\mathfrak{F}_{u}$, if $u_{j}\leq h(0)$, the inequality holds. Now let $j\in\mathcal{S}(u,h)$, we have $u_{j}\leq\sup_{k\in\mathbb{N}}u_{k}=u_{\mathbf{K}^{s}_{u}}$. Hence, from the fifth statement of Prop. 11, $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$. We conclude from the third statement of Prop. 11 that $\mathbf{K}^{s}_{u}\leq\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor$ and $\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)$ whenever the set $\mathcal{S}(u,h)$ is nonempty. Indeed, if $\mathcal{S}(u,h)$ is nonempty, then $u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}$ is also strictly greater than $h(0)$. ∎
The result of Prop. 13 is not surprising but can be directly used in a very particular context. Suppose that we have an explicit closed formula for $u_{k}$ but not derivable. The function $h$ can regularize and strictly convexify the closed formula. We thus obtain with some calculus the exact value of $\mathbf{K}^{s}_{\nu}$ as the integer part (or integer part +1) of the zero of the derivative of $s\to\mathfrak{F}_{u}(s,h,\beta)$ (extended in the reals).
We can go further than the results of Prop 13 to get a solution of Problem 2. Th. 2 constitutes the next step of Prop 13.
Theorem 2 (Solution of Problem 2).
Let $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$ be useful for $u$ and let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ and $k\in\mathcal{S}(u,h)$. Then $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor$ is a solution of Problem 2 that is :
$\max_{n\in\mathbb{N}}u_{n}=\max\{u_{k}:k=0,\ldots,\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor\}.$
Proof.
It suffices to combine Prop. 6 and Prop. 13. ∎
Th. 2 shows how the function depicted at Eq. (4) is used to solve Problem 2. It suffices to construct an upper bound of the sequence of interest of the form of $(h(\beta^{k}))_{k\in\mathbb{N}}$ where $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$; moreover, $h(0)$ has to be strictly smaller than some terms of the sequence of interest. In theory, when the limsup of the sequence of interest is not the supremum of the sequence, such couples exist and one of them is optimal (Th. 1) and $h$ can even be chosen as an affine function.
Example 3.
To give a practical view of the function provided at Eq. (4), we evaluate $\mathfrak{F}_{\nu}(k,h,\beta)$ for the sequence $\nu$ described at Eq (2) of Subsection 2.2. The pairs $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ considered are the ones proposed at Example 2.
We evaluate the function for different values of $p$ and $\mu$ to two choices of pairs $(h,\beta)$. In Example 2, we proposed linear functions $h$ characterized by a scalar $a$. We recall that the pair are assumed to satisfy the conditions: $a>0$, $\beta\in(0,1)$ and $a\beta^{n_{0}}\geq p^{2}/3$ where $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\mu)<0$. First, as the function $\mathfrak{F}_{u}$ depends on the values of the term $\nu_{k}$, we show the seven first $\nu_{0},\ldots,\nu_{6}$ (rounded to the nearest hundredth) for certain values of $p$ and $\mu$ at Table 1.
$\nu_{k}$
k 0 1 2 3 4 5 6 $\mathbf{K}^{s}_{\nu}$ : $\nu_{\rm opt}$
$p=30$ $\mu=\frac{1}{3}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 8 : 300
$\mu=\frac{1}{10}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 11 : 300
$\mu=\frac{1}{1000}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 22 : 300
$p=9$ $\mu=\frac{1}{3}$ 8.25 11.81 16.45 21.83 26.34 27. 25.09 5 : 27
$\mu=\frac{1}{10}$ 8.25 11.81 16.45 21.83 26.34 27. 27. 8 : 27
$\mu=\frac{1}{1000}$ 8.25 11.81 16.45 21.83 26.34 27. 27. 19 : 27
$p=3$ $\mu=\frac{1}{3}$ 2.25 2.81 3 2.95 1.58 -4.03 -20.47 2 : 3
$\mu=\frac{1}{10}$ 2.25 2.81 3 3. 3. 3. 2.94 5 : 3
$\mu=\frac{1}{1000}$ 2.25 2.81 3. 3. 3. 3. 3. 17 : 3
Table 1: Few values of the terms $\nu_{k}$ defined at Eq. (2) depending on the problem parameters $p$ and $\mu$.
In Table 1, we observe that the terms $\nu_{k}$ strictly increases until the optimal values (the smallest integer maximizer $\mathbf{K}_{\nu}$) and strictly decreases after (after the greatest integer maximizer $\mathbf{K}^{s}_{\nu}$). This observation will be formally proved and exploited later on. Another point related to this observation is that the supremum of $\nu_{k}$ for $k\notin\operatorname{Argmax}(\nu)$ seems to be either $\nu_{\underline{n}-1}$ or $\nu_{\overline{n}}$. For example, for $p=3$ and $\mu=1/3$, $\nu_{3}>\nu_{1}$ whereas for $p=9$ and $\mu=1/3$, $\nu_{4}>\nu_{5}$. Actually, we will see that the rank preceding $\underline{n}$ is not the one to consider to compute $\nu_{k}$ for $k\notin\operatorname{Argmax}(\nu)$.
Now, we evaluate the function $\mathfrak{F}_{\nu}$ for our example. To simplify, we write $\mathfrak{F}_{\nu}(k,a,\beta)$ instead of $\mathfrak{F}_{\nu}(k,h,\beta)$. We choose two pairs: first, $a=p^{2}$ and $\beta=(2/3)^{1/n_{0}}$ and second $a=2p^{2}/3$ and $\beta=(1/2)^{n_{0}}$. The values of $\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ for these two choices are provided at Table 2.
$\lfloor\mathfrak{F}_{\nu}(k,p^{2},(2/3)^{1/n_{0}})\rfloor$ $\lfloor\mathfrak{F}_{\nu}(k,p^{2}/3,(1/2)^{1/n_{0}})\rfloor$
k $\underline{n}$ $\overline{n}$ $n_{0}$ 0 1 2 3 4 $\mathbf{K}^{s}_{\nu}$ 0 1 2 3 4 $\mathbf{K}^{s}_{\nu}$
$p$ $\mu=\frac{1}{3}$ 8 9 10 84 74 65 56 47 27 43 37 32 26 21 9
= $\mu=\frac{1}{10}$ 8 12 13 97 84 72 61 50 35 56 49 42 35 28 12
30 $\mu=\frac{1}{1000}$ 8 23 24 202 179 156 131 113 65 104 91 77 64 52 23
$p$ $\mu=\frac{1}{3}$ 5 6 7 39 33 27 22 19 18 18 15 12 9 7 7
= $\mu=\frac{1}{10}$ 5 9 10 56 47 39 32 27 27 27 21 17 13 10 9
9 $\mu=\frac{1}{1000}$ 5 20 21 118 99 82 67 58 56 56 46 36 27 21 20
$p$ $\mu=\frac{1}{3}$ 2 3 4 13 11 10 $10^{*}$ $17^{*}$ 10 5 4 3 $4^{*}$ $7^{*}$ 3
= $\mu=\frac{1}{10}$ 2 4 5 23 20 18 18 18 18 9 7 7 7 7 7
3 $\mu=\frac{1}{1000}$ 2 18 18 61 51 48 48 48 48 25 19 18 18 18 18
Table 2: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h,\beta)\rfloor$ described at Eq. (4) for the sequence $\nu$ of Eq. (2) and $(h,\beta)\in\mathbb{\Gamma}(\nu)$ proposed at Ex. 2
The numbers of Table 2 at columns $\lfloor\mathfrak{F}_{\nu}(k,p^{2},(2/3)^{1/n_{0}})\rfloor$ and $\lfloor\mathfrak{F}_{\nu}(k,2p^{2}/3,(1/2)^{1/n_{0}})\rfloor$ represent safe upper bounds on the number of terms of the sequence $\nu$ of Eq. (2) to compare to obtain the supremum of the terms $\nu_{k}$. From the analysis made at Subsection 2.2, $\mathbf{K}^{s}_{\nu}$ is well identified and is equal to $\overline{n}-1$. We observe that all numbers of the concerned columns are greater than $\overline{n}-1$ as stated at Prop. 13. For both choices of $a$ and $\beta$, the best upper bound is furnished by the column $\lfloor\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},a,\beta)$ as asserted at Prop. 13. Note that four integers are followed by an asterisk. In practice, those numbers are useless since not computed. Indeed, applying the fifth statement of Prop. 11, it suffices to compute a new stopping integer at $k$ when we have $\nu_{j}<\nu_{k}$ for all $j<k$. Otherwise, the stopping integer is greater than at least one previous computed integer. Moreover, for the case $a=2p^{2}/3$ and $\beta=(1/2)^{1/n_{0}}$, following Th 2, the terms $\nu_{k}$, for $k>3$, are not computed as $\mathfrak{F}_{\nu}(3,2p^{2}/3,(1/2)^{1/n_{0}})=3$ and we only have to compare $\nu_{0},\nu_{1},\nu_{2}$ and $\nu_{3}$ to obtain the maximal values of the terms $\nu_{k}$.
Finally, Table 2 highlights the importance of the choice of the pair $(h,\beta)\in\mathbb{\Gamma}(u)$ to compute a less conservative upper bound of $\mathbf{K}^{s}_{u}$. For general peaks computation problems of the form Problem 1, those upper bounds serve to fix the number of instances of Problem Pk to solve to compute the optimal value of the peaks computation problem. As Problem Pk are constrained optimization problems, a too conservative upper bound can dragstically increase the computational time. ∎
Theorem 3 (Formula Optimality).
For all $u\in\mathbb{R}^{\mathbb{N}}$, we have the following formula:
$\mathbf{K}^{s}_{u}=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{\ln(\beta)}$
and the $\inf$ can be replaced by $\min$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
Proof.
Let us define
$F:=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta).$
If $u\notin\Lambda$, then from Prop. 8, $\mathbb{\Gamma}(u)=\emptyset$ and thus $F=+\infty$. Moreover, from Prop. 2, $\mathbf{K}^{s}_{u}=+\infty$.
Suppose that $u\in\Lambda$. From Prop. 8, $\mathbb{\Gamma}(u)\neq\emptyset$. Assume that $\Delta^{s}_{u}$ is empty. This is the same as $\mathbf{K}^{s}_{u}=+\infty$. From Prop. 3, ${\mathrm{R}}_{u}^{s}=\emptyset$ and from Prop. 10, $\mathcal{S}(u,h)=\emptyset$ which leads to $F=+\infty$. Assume that $\Delta^{s}_{u}$ is nonempty and thus ${\mathrm{R}}_{u}^{s}\neq\emptyset$. From Prop. 13, $\mathbf{K}^{s}_{u}\leq F$. Moreover, from Corollary 2, there exists $(h,\beta)\in\mathbb{\Gamma}(u)$ such that $u_{\mathbf{K}^{s}_{u}}=h(\beta^{\mathbf{K}^{s}_{u}})>h(0)$ and $F<+\infty$. Taking the inverse of $h$ on $[0,1]$ and the natural logarithm, we get $\mathbf{K}^{s}_{u}=\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$. By definition of $F$, $F\leq\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)=\mathbf{K}^{s}_{u}$. We conclude that $\mathbf{K}^{s}_{u}=F$. ∎
In general, we cannot rely on pre-analysis of the system as it is done in Subsection 2.2 to solve a peaks computation problem. Thanks to Th. 2, we reduce a peaks computation problem to the construction of a pair $(h,\beta)$ useful for $\nu$. In consequence, we are faced to the following problem on the sequence $\nu$ defined at Eq (1).
Problem 3.
Construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$.
Next sections of the paper are focused on classical tools used in stability theory of dynamical systems such as $\mathcal{K}\mathcal{L}$-stability and Lyapunov functions. We show how these tools can be adapted to solve Problem 3.
4 Extended $\mathcal{K}\mathcal{L}$ Theory for Discrete-Time Systems Maximization
In this section, we come back to the initial problem that is a peaks computation problem i.e. the maximization of a function over the reachable values of a discrete-time dynamical system. We recall that the data are the following:
  • an initial conditions set $X^{\mathrm{in}}\subseteq\mathbb{R}^{d}$;
  • a nonzero map $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$;
  • a function $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$ such that $\varphi(0)=0$.
We recall that we want compute $\nu_{\rm opt}=\sup\{\nu_{k}:k\in\mathbb{N}\}$ where $\nu_{k}=\sup\{\varphi(T^{k}(x):x\in X^{\mathrm{in}}\}$. To use the results of Section 3 with the sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$, we suppose that Assumption 1 holds.
We propose to simplify the notation of the maximal value of an objective function $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ on $X^{\mathrm{in}}$, we introduce the following new notation:
$\overline{f_{X^{\mathrm{in}}}}:=\sup\{f(x):x\in X^{\mathrm{in}}\}$
and the set of functions which are bounded from above on $X^{\mathrm{in}}$:
$\operatorname{Fin}\left(X^{\mathrm{in}}\right):=\{f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}:\overline{f_{X^{\mathrm{in}}}}<+\infty\}$
With these notations, $\nu_{k}=\overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}$ and by following Assumption 1, we have for all $k\in\mathbb{N}$, $\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$. Following Subsection 3.2, we have to study, first, the limsup of the sequence $\nu$ and then construct, when the limsup is finite, an element $(h,\beta)$ of $\mathbb{\Gamma}(\nu)$ i.e. $h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and for all $k\in\mathbb{N}$, $h(\beta^{k})$ is greater than $\nu_{k}$. We also have to guarantee that if ${\mathrm{R}}_{\nu}^{s}$ is nonempty, $h$ has to be useful for $\nu$ i.e. $\mathcal{S}(\nu,h)$ is nonempty. Without this latter condition, the function $h$ is so useless as $\mathfrak{F}_{u}(k,h,\beta)=+\infty$. In this section, we will revisit $\mathcal{KL}$ class of functions to study the limsup of $\nu$ and to construct our necessary pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$.
4.1 Using the classical $\mathcal{KL}$ class of functions
To study global asymptotic stability of discrete-time systems can be done by means of comparison functions (see e.g. [GGLW14]). For a comprehensive literature about comparison functions, the interested reader can refer to  [Kel14]. Comparison functions can be used for our problem to study the limsup of the sequence $\nu$.
We start by recalling some necessary definitions about comparison functions.
Definition 2 ($\mathcal{K}$/$\mathcal{K}_{\infty}$ functions classes).
A function $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}$ if $\alpha$ is strictly increasing continuous and $\alpha(0)=0$. If moreover, $\lim_{x\to+\infty}\alpha(x)=+\infty$, $\alpha$ is said to belong to $\mathcal{K}_{\infty}$.
Definition 3 ($\mathcal{L}$ functions class).
A function $\sigma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to $\mathcal{L}$ if it is continuous, strictly decreasing and $\lim_{x\to+\infty}\sigma(x)=0$.
Definition 4 ($\mathcal{K}\mathcal{L}$ functions class).
A function $\gamma:\mathbb{R}_{+}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}\mathcal{L}$ if for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ belongs to $\mathcal{K}$ and if for all $s\in\mathbb{R}_{+}$, $t\mapsto\gamma(s,t)$ belongs to $\mathcal{L}$.
The class $\mathcal{K}$ is included in $\Omega([0,1])$ but $\mathcal{K}$ functions are strictly positive except at 0 where they are null. Hence, $h\in\mathcal{K}\cap\mathbb{\Gamma}_{\rm fun}(\nu)$ is useful for $\nu$ if at least one term $\nu_{k}$ is strictly positive. Thus, in this subsection, Assumption 2 must hold.
Assumption 2.
There exists $k\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))>0$.
We recall the result obtained by Sontag.
Lemma 1 (Proposition 7 of [Son98]).
For all $\gamma\in\mathcal{K}\mathcal{L}$, there exists $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $s,t\in\mathbb{R}_{+}$, $\gamma(s,t)\leq\theta_{1}(\theta_{2}(s)e^{-t})$.
Theorem 4.
Assume that there exist a function $\gamma\in\mathcal{K}\mathcal{L}$ and a function $\psi:\mathbb{R}^{d}\mapsto\mathbb{R}_{+}$ such that $\overline{\psi_{X^{\mathrm{in}}}}>0$. We also assume that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$:
$\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\enspace.$
Then:
  1. 1.
    $\limsup_{n\to+\infty}\nu_{n}\leq 0$ and if, moreover, Assumption 2 holds then ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$;
  2. 2.
    there exists $h\in\mathcal{K}_{\infty}$ s.t. $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ and, if, moreover, Assumption 2 holds then $h$ is useful for $\nu$.
Proof.
[…]
2. This a direct consequence of the Sontag’s result (Lemma 1 here). Indeed, there exist $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\leq\theta_{1}(\theta_{2}(\psi(x))e^{-k})$. Then $\nu_{k}\leq h(e^{-1}):=\sup_{x\in X^{\mathrm{in}}}\theta_{1}(\theta_{2}(\psi(x)e^{-k}))$. It is easy to see, as $\theta_{1}$ and $\theta_{2}$ are increasing and continuous, that $h$ can be also written $y\mapsto\theta_{1}(\theta_{2}(\overline{\psi_{X^{\mathrm{in}}}}y))$. The function $h$ is thus the compositions of $\theta_{1}$, $\theta_{2}$ and $y\mapsto\overline{\psi_{X^{\mathrm{in}}}}y$ which belong to $\mathcal{K}_{\infty}$ (the third as $\overline{\psi_{X^{\mathrm{in}}}}>0$) and thus belongs to $\mathcal{K}_{\infty}$. Finally, under Assumption 2, there exists $k\in\mathbb{N}$, such that $\nu_{k}>0=h(0)$. ∎
Usually i.e. when $\varphi=\|\cdot\|$, the inequality $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)$ (for $\psi=\|\cdot\|$) is true for all $x\in\mathbb{R}^{d}$ and all $k\in\mathbb{N}$ if and only if the system $x_{k+1}=T(x_{k}),x_{0}\in\mathbb{R}^{d}$ is said to be uniformly global asymptotic stability (UGAS) (see for example [NT04, Prop. 1] or [ATP10, Th. 2]. This implies that $x_{k}$ tends to 0 as $k$ tends to +$\infty$+ ∞. Thus, with a general function $\varphi$, we obtain a one-sided information on the asymptotic behaviour of the dynamics. We could obtain the other side applying the theorem for $-\varphi$.
4.2 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
In this paper, we are not interested in stability of the dynamical system, we can relax the strict increase (or decrease), the continuity and the positivity (on the state variable) conditions required for $\mathcal{KL}$ functions.
Definition 5 (The Class $\mathcal{KL}_{\rm gen}$).
A function $\gamma:\mathbb{R}\times\mathbb{R}_{+}\mapsto\mathbb{R}$ belongs to $\mathcal{KL}_{\rm gen}$ if for all $s\in\mathbb{R}$, $t\mapsto\gamma(s,t)$ is increasing and for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ is decreasing.
Definition 6 ($\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds).
A pair $(\gamma,\theta)\in\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ if and only if :
$\forall\,k\in\mathbb{N},\ \forall\,x\in X^{\mathrm{in}},\ \varphi(T^{k}(x))\leq\gamma(\theta(x),k)$
Moreover, $(\gamma,\theta)$ is said to useful for $\nu$ such that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if the set:
$\left\{k\in{\mathrm{R}}_{\nu}^{s}:\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}\right\}$
is nonempty.
Following the notations introduced earlier, we have the equivalence:
$\nu\in\Lambda\iff\sup_{k\in\mathbb{N}}\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$
From Prop. 1 we have the equivalence, we also have:
$\nu\in\Lambda\iff\limsup_{k\to+\infty}\nu_{k}=\limsup_{k\to+\infty}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))=\limsup_{k\to+\infty}\ \overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}\in\mathbb{R}\cup\{-\infty\}$
We can establish a first result which states that the existence of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds is equivalent to the fact that the sequence $\nu$ is bounded from above. This is an adaptation of classical results dealing with stability and $\mathcal{KL}$ upper bounds.
Theorem 5 ($\mathbb{\Gamma}(\nu)$ to $\mathcal{KL}_{\rm gen}$).
The sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ belongs to $\Lambda$ if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. If, moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ then $(\gamma,\theta)$ can be chosen to be useful for $\nu$.
Proof.
Suppose that there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. As $s\mapsto\gamma(s,t)$ is increasing for all $t\geq 0$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\gamma(\theta(x),k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ and thus $\nu_{k}=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ for all $k\in\mathbb{N}$. Now, taking the limsup on both sides leads to $\limsup_{k\mapsto+\infty}\nu_{k}\leq\inf_{k\in\mathbb{N}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},0)<+\infty$ as $t\mapsto\gamma(s,t)$ is decreasing for all $s\in\mathbb{R}$.
Suppose that $\nu=(\nu_{k})_{k\in\mathbb{N}}\in\Lambda$. From Prop. 8, $\Gamma(\nu)$ is nonempty and let $(h,\beta)$ be in $\Gamma(\nu)$. Then, we define for all $s\in\mathbb{R}$ and $t\in\mathbb{R}_{+}$, $\gamma(s,t):=\min\{s,h(\beta^{t})\}$ and for all $x\in\mathbb{R}^{d}$, $\theta(x):=\sup_{k\in\mathbb{N}}\varphi(T^{k}(x))$. It is clear that $\gamma\in\mathcal{KL}_{\rm gen}$ as $\beta\in(0,1)$. Now as $\nu\in\Lambda$, $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\nu_{n}\in\mathbb{R}$. Moreover, we have $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{x\in X^{\mathrm{in}}}\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))=\overline{\theta_{X^{\mathrm{in}}}}$. By definition of $\theta$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\theta(x)$. By definition of $(h,\beta)$, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq h(\beta^{k})$. This is the same as for all $x\in X^{\mathrm{in}}$ and for all $k\in X^{\mathrm{in}}$, $\varphi(T^{k}(x))\leq h(\beta^{k})$. We conclude that $\varphi(T^{k}(x))\leq\min\{\theta(x),h(\beta^{k})\}$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$. Note that for a function $\gamma$ constructed as below, for all $s\in\mathbb{R}$, $\inf_{t\geq 0}\gamma(s,t)=\inf_{t\geq 0}\min\{s,h(\beta^{t})\}=\min\{s,h(0)\}$. This follows from the continuity of $h$, the fact $h$ is strictly increasing and $\beta\in(0,1)$.
If moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then from Lemma 10, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$. We could define $\gamma$ using this function $h$. Now, let $k\in\mathcal{S}(\nu,h)$. As $\nu_{k}\leq\nu_{\rm opt}=\overline{\theta_{X^{\mathrm{in}}}}$, we conclude that $\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)=\min\{\overline{\theta_{X^{\mathrm{in}}}},h(0)\}=h(0)$ which ends the proof. ∎
Example 4.
Let us come back to the analysis of the peaks computation problem of Subsection 2.2. We propose to compute a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for the triplet $(X^{\mathrm{in}},T,\varphi)$ defined at Subsection 2.2.
Following the proof of Th 5, one can define the $\mathcal{KL}_{\rm gen}$ function as $(s,t)\mapsto\min\{s,h(\beta^{t})\}$ where $(h,\beta)$ is a pair proposed at Example 2 and $\theta\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ as $\mathbb{R}^{2}\ni x\mapsto\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))$.
Another possible pair consists in the $\mathcal{KL}_{\rm gen}$ function: ×$\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\min\{s,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$ and the constant function in $\operatorname{Fin}\left(X^{\mathrm{in}}\right)$: $\mathbb{R}^{2}\ni x\mapsto p^{2}/3$. We recall that $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\alpha)<0$. We can easily check that $\varphi(T^{k}(x))\leq\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(k+1))^{-1}\}$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Moreover, as $\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}=0$, there exists $k\in\mathbb{N}$ such that $\nu_{k}>\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$.
In the proof of Th. 5, we exploit Prop. 8 and Prop. 10 which link the nonemptiness of the sets $\mathbb{\Gamma}(u)$, $\mathcal{S}(u,h)$ and ${\mathrm{R}}_{u}^{s}$. We can reformulate Th 5 as Corollary 3.
Corollary 3.
The set $\mathbb{\Gamma}(\nu)$ is nonempty if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. If, moreover, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ then there exists a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$.
In Th 5 and in Corollary 3, we construct a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ from a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$. Now, the question is to know whether is it possible to construct a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. This is exactly what was done at Th 4 from the Sontag’s result. Thus, we provide a positive answer. The result is based on an extension of the Sontag’s result for the class $\mathcal{KL}_{\rm gen}$. This permits to fix the value $\beta\in(0,1)$ to only consider problem on functions in $\Omega([0,1])$.
Theorem 6 ($\mathcal{KL}_{\rm gen}$ to $\mathbb{\Gamma}(\nu)$).
Suppose that there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. Then there exists $h\in\Omega([0,1])$ such that $(h,e^{-1})\in\Gamma(\nu)$. If moreover, the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ is useful for $\nu$ then $h$ can be chosen to be useful for $\nu$.
Lemma 2.
Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a decreasing function and let $m\geq\inf_{\mathbb{R}_{+}}f$ ($m$ can be equal to $\infty$- ∞ if $\inf_{\mathbb{R}_{+}}f$ is). There exists a function $g:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that:
  1. 1.
    $f\leq g$;
  2. 2.
    $\inf_{\mathbb{R}_{+}}g=m$.
Proof.
If $m\geq f(0)$, the function $x\mapsto g(x)=m+1/(x+1)$ satisfies the above properties. Note that this case includes the case where $f$ is constant on $\mathbb{R}_{+}$.
Now assume that $f(0)>m\geq\inf_{\mathbb{R}_{+}}f$. We denote for all couple of pairs $\{(x,y),(z,t)\}$ where $x,y,z,t\in\mathbb{R}$ by $\mathrm{Aff}\{(x,y),(z,t)\}$ the unique affine function $\ell:\mathbb{R}\mapsto\mathbb{R}$ such that $\ell(x)=y$ and $\ell(z)=t$. We define the functions $h_{0}$ and $h_{n}$ for all $n\in\mathbb{N}^{*}$ as follows :
$F_{0}=f(0)\mathbb{1}_{[0,1]}\text{ and }F_{n}=f(n-1)\mathbb{1}_{\left[n;n+\frac{1}{2}\right]}+\mathrm{Aff}\left\{\left(n+\frac{1}{2};f(n-1)\right);\left(n+1;f(n)\right)\right\}\mathbb{1}_{\left[n+\frac{1}{2};n+1\right]}$
Finally, we define, for all $x\in\mathbb{R}_{+}$, the function $F$:
$F(x)=\sum_{n\in\mathbb{N}}F_{n}(x)+\dfrac{1}{x+1}\text{ and note that}\sum_{n\in\mathbb{N}}F_{n}(x)=\left\{\begin{array}[]{lr}f(0)&\text{ if }x\in[0,1]\\ f(\lfloor x\rfloor-1)&\text{ if }x\in[\lfloor x\rfloor,\lfloor x\rfloor+1/2]\\ \ell\left(x\right)&\text{ if }x\in(\lfloor x\rfloor+1/2,\lfloor x\rfloor+1)\end{array}\right.$
where $\ell=\mathrm{Aff}\left\{\left(\lfloor x\rfloor+1/2;f(\lfloor x\rfloor-1)\right);\left(\lfloor x\rfloor+1;f(\lfloor x\rfloor)\right)\right\}$.
The function $\sum_{n\in\mathbb{N}}F_{n}$ is clearly continuous and decreasing (as $f$ does) on $\mathbb{R}_{+}$. The function $\sum_{n\in\mathbb{N}}F_{n}$ is not strictly decreasing but adding $x\mapsto(x+1)^{-1}$ makes $F$ to be strictly decreasing and continuous. We also have $f\leq\sum_{n\in\mathbb{N}}F_{n}$ as $f$ is decreasing. As $x\mapsto(x+1)^{-1}$ is strictly positive we have $f\leq F$. Finally, we have $\inf_{\mathbb{R}_{+}}F=\inf_{\mathbb{R}_{+}}f$. Indeed, for all $n\in\mathbb{N}^{*}$ and all $x\in[n,n+1)$, $F_{n}\leq f(n-1)$, it implies that $\inf_{\mathbb{R}_{+}}F=\inf_{x\in\mathbb{R}_{+}}\inf_{n\in\mathbb{N}^{*}}F_{n}(x)\leq\inf_{n\in\mathbb{N}^{*}}f(n-1)=\inf_{\mathbb{R}_{+}}f$ (as $f$ decreases). The reverse inequality holds as $f\leq F$.
The function $F$ does not necessarily satisfy $\inf_{\mathbb{R}_{+}}h=m$. Hence, we define for all $x\in\mathbb{R}_{+}$:
$g(x)=\left\{\begin{array}[]{lr}\max\{F(x),(x+1)^{-1}+m\}&\text{ if }m>\inf_{\mathbb{R}_{+}}f\\ F(x)&\text{ if }m=\inf_{\mathbb{R}_{+}}f\end{array}\right.$
Suppose that $m>\inf_{\mathbb{R}_{+}}f$. The function $g$ is clearly strictly decreasing and continuous and satisfies $f\leq g$ as $F$ does. As $m>\inf_{\mathbb{R}_{+}}f=\inf_{\mathbb{R}_{+}}F$, there exists $x_{0}$ such that $F(x)<m$ for all $x\geq x_{0}$ and then $g(x)=(x+1)^{-1}+m$ for all $x\geq x_{0}$. This implies that $\inf_{\mathbb{R}_{+}}g=m$. ∎
Lemma 3 (Sontag’s extension).
Let $\gamma\in\mathcal{KL}_{\rm gen}$. Let $s\in\mathbb{R}$ and let $m\geq\inf_{t\geq 0}\gamma(s,t)$ be finite. Then there exists $h\in\Omega([0,1])$ such that for all $t\geq 0$, $\gamma(s,t)\leq h(e^{-t})$ and $h(0)=m$.
Proof.
By definition, $t\mapsto\gamma(s,t)$ is decreasing and applying Lemma 2, there exists a function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that $\gamma(s,t)\leq\sigma(t)$ for all $t\geq 0$ and $\inf_{\mathbb{R}_{+}}\sigma=m$. Let us define for all $r\in(0,1]$, $h(r)=\sigma(-\ln r)$ and $\lim_{x\to 0^{+}}h(x)=\lim_{y\to+\infty}\sigma(y)=m$. Then defining $h(0)=m$ makes $h$ continuous at 0. Finally, $h$ belongs to $\Omega([0,1])$. ∎
Proof of Th. 6.
By assumption, there exist a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. This implies that $\varphi(T^{k}(x))\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$. By Lemma 3, there exists $h\in\Omega([0,1])$ such that $\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq h(e^{-k})$ for all $k\in\mathbb{N}$. This leads to $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$. If there exists $k\in\mathbb{N}$ such that $\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}$, we can choose, as it is stated in Lemma 3, $h$ such that $h(0)=m\in(\inf_{t\in\mathbb{R}}\gamma(\overline{\psi},t),\nu_{k})$. ∎
We extend the class of $\mathcal{K}\mathcal{L}$ functions which classically serve for dynamical systems stability. In our context, we have a similar use and our $\mathcal{KL}_{\rm gen}$ class constitutes the main part of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds of $\varphi(T^{k}(x))$. We prove that such upper bounds exist if and only if $\nu\in\Lambda$. Moreover, if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, we are able to construct a function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ and such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
Unfortunately, in practice, to prove that a pair $(\gamma,\theta)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ seems to be as difficult as proving that a pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(\nu)$. We thus propose a new type of Lyapunov functions for which the conditions does not involve the whole trajectories of the analyzed dynamical system.
5 Lyapunov type Functions for Discrete-time Systems Maximization
Classically, Lyapunov functions are used as certificates of stability (see  [Ela05] or [KP01]). They bring a constructive approach to prove the stability of the dynamical system through converse Lyapunov theorems which prove that stability notions are equivalent to the existence of Lyapunov functions [Kha02]. For some linear, polynomial or piecewise polynomial systems, Lyapunov functions can be computed from numerical optimization solvers based on linear or semidefinite programming (for example see [GH15] and references therein).
Classical quadratic Lyapunov functions solve our problem when the dynamical system is stable and affine and the objective function is quadratic [Adj21]. In Subsection 5.1, we generalize this approach theoretically. Then in Subsection 5.2, we define a more general class of Lyapunov functions. The existence of those functions does not ensure the stability of the system but just the finiteness of the limsup of the sequence $\nu$ defined at Eq. (1). Last but not least, we etablish an equivalence theorem between the existence of those new Lyapunov functions and the existence of an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
5.1 Classical Lyapunov Functions Theory
5.1.1 Definition and Useful Facts
[…]
Definition 7 (Classical Lyapunov Functions).
A function $V:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is said to a Lyapunov function if and only if :
  1. 1.
    There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that:
    $\alpha_{1}\circ\|\cdot\|\leq V\leq\alpha_{2}\circ\|\cdot\|$
  2. 2.
    There exists $\lambda\in(0,1)$ such that
    $V\circ T\leq\lambda V$
[…]
  1. 1.
    there exists $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite111a function $g:X\mapsto\mathbb{R}_{+}$ where $X\subset\mathbb{R}^{d}$ and $0\in X$ is said to be positive definite (on $X$) if $g(0)=0$ and $g(x)>0$ for all $x\in X\backslash\{0\}$ and continuous such that $V\circ T\leq V-\alpha\circ\|\cdot\|$;
  2. 2.
    there exists $\rho:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite, continuous and such that $\mathrm{Id}-\rho$ is positive definite satisfying $V\circ T\leq\rho\circ V$.
We will propose a proof for this equivalence later on the new class of Lyapunov functions used for the maximization problem.
First, we remark that from the first statement of Def. 7, $V$ is positive definite. If we require $V$ to be continuous and such that $\lim V(x)=+\infty$ as $\|x\|\to+\infty$ (see e.g. [Adj24]), then the first statement of Def. 7 is actually equivalent to $V$ is postive definite. Second, if there exists a classical Lyapunov function for $T$, we $T$ must be null at zero [^2] Actually, if $T(e)=e$ for a nonzero vector, we could replace $V$ by $V(\cdot-e)$ and the lower and the upper bounds on $V(\cdot-e)$ by respectively $\alpha_{1}(\|\cdot-e\|)$ and $\alpha_{2}(\|\cdot-e\|)$. Finally, from the second statement of Def 7, we can define a kind of norm operator relative to a Lyapunov function. We have, for all Lyapunov function $V$ associated to the dynamic $T$, for all $k\in\mathbb{N}$:
$N_{T^{k}}(V):=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{d}\\ x\neq 0\end{subarray}}\dfrac{V(T^{k}(x))}{V(x)}\in(0,1)$ (5)
We will simply write $N_{T}(V)$ when $k=1$. By definition, we have for all $x\in\mathbb{R}^{d}$, $V(T(x))\leq N_{T}(V)V(x)$. Using this and the recurrence principle, we get, for all $k\in\mathbb{N}$, $N_{T^{k}}(V)\leq(N_{T}(V))^{k}$.
5.1.2 A Direct Classical Lyapunov Theorem
Let $f$ be a function from $\mathbb{R}^{d}$ to $\mathbb{R}$. We recall that the function from $\mathbb{R}_{+}$ to $\mathbb{R}$:
$\alpha^{f}:s\mapsto\sup_{\|y\|\leq s}f(y)$ (6)
is continuous when $f$ is continuous (see e.g. [Kha02] or extensions in [Adj24]). Obviously, the function $\alpha^{f}$ is increasing. Moreover, if $f(0)=0$ then $\alpha^{f}$ is nonnegative and satisfies $\alpha^{f}(0)=0$. Last but not least, we have for all $x\in\mathbb{R}^{d}$, $f(x)\leq\alpha^{f}(\|x\|)$.
Theorem 7 (Classical Lyapunov to $\mathbb{\Gamma}(\nu)$).
Assume that $\varphi$ is continuous and there exists a Lyapunov function for $T$. Then there exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $h(0)=0$. Thus, if Assumption 2 holds, $h$ is useful for $\nu$.
Proof.
If there exists a Lyapunov function $V$ for $T$, there exists $\alpha_{1}\in\mathcal{K}_{\infty}$ such that $\alpha_{1}\circ\|\cdot\|\leq V$. Recall that we have assumed that $\varphi(0)=0$. With the continuity assumption of $\varphi$, we conclude that $\alpha^{\varphi}$ (defined at Eq. (6)) is nonnegative, increasing, continuous and is equal to 0 at 0. Two small issues : $\alpha^{\varphi}$ is not guaranteed to be strictly increasing and surjective from $\mathbb{R}_{+}$ to $\mathbb{R}_{+}$ (the limit of $\alpha^{\varphi}$ at +$\infty$+ ∞ can be finite). We thus add to $\alpha^{\varphi}$ a function $\psi\in\mathcal{K}_{\infty}$ and we have for $y\in\mathbb{R}^{d}$, $\varphi(y)\leq(\alpha^{\varphi}+\psi)(\|y\|)$ and $\alpha:=\alpha^{\varphi}+\psi$ belongs to $\mathcal{K}_{\infty}$.
From $\alpha_{1}\circ\|\cdot\|\leq V$ and $\alpha_{1}\in\mathcal{K}_{\infty}$, we get $\varphi\leq\alpha\circ\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\circ V$ as $\alpha$ is increasing. Moreover, we have, for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $V(T^{k}(x))\leq N_{T^{k}}(V)V(x)\leq(N_{T}(V))^{k}V(x)\leq(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}$. As $\alpha$ and ${{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}$ are increasing, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\left(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}\right)\right)$. We thus define $h:x\mapsto\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}(x\overline{V_{X^{\mathrm{in}}}})\right)$ and $\beta=N_{T}(V)$. We have $h(0)=0$ and we conclude that $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$ when Assumption 2 holds. ∎
The goal of extending the concept of Lyapunov functions is to weaken the hypothesis of Th 7 by removing the continuity and the positivity assumptions on $\varphi$. The extension also allows to prove the equivalence between the existence of useful pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and the existence of this new type of Lyapunov functions.
5.2 Opt-Lyapunov Functions for Discrete-time Systems Maximization
5.2.1 Definitions and Useful Facts
We recall that the domain of function $g:\mathbb{R}^{d}\mapsto(-\infty,+\infty]$ is the set of $x\in\mathbb{R}^{d}$ such that $f(x)<+\infty$. We denote by $\operatorname{dom}(f)$ the domain of a extended real valued function.
Definition 8 (Opt-Lyapunov functions).
A function $V$ is said to be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ if and only if:
  1. 1.
    $V:\mathbb{R}^{d}\mapsto[0,+\infty]$;
  2. 2.
    $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$;
  3. 3.
    $V\circ T\leq\lambda V$ for some $\lambda\in(0,1)$.
Example 5.
Let us consider the dynamical system of the example of Subsection 2.2. Recall that $\mu\in(0,1/3)$ and the set of initial conditions is :
$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}$
and the dynamics is the following:
$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}$
The function defined, for all $x=(x_{1},x_{2})^{\intercal}\in\mathbb{R}^{2}$ as follows:
$V(x)=\left\{\begin{array}[]{cr}\dfrac{2\mu^{2}}{x_{1}x_{2}}&\text{ if }x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}\\ \\ 0&\text{ otherwise}\end{array}\right.$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Let us prove that it is true. Let $x\in\mathbb{R}^{2}$. If $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, then $x=(2,1)^{\intercal}y$ with $y\in\mathbb{R}^{*}$. First, the coordinates of $x$ has the same sign and so $V$ is nonnegative. The fact that $V(x)$ is nonnegative when for $x\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $x=0$ is obvious. Second, let $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ and let $y\neq 0$ such that $x=(2,1)^{\intercal}y$. Then, $V(T(x))=V((3y,3y/2)^{\intercal})=(4\mu^{2})/(9y^{2})$ and $V(x)=\mu^{2}/y^{2}$ then $V(T(x))/V(x)=(4y^{2})/(9y^{2})=4/9$. Any value in $[4/9,1)$ can be chosen. Let us take $5/9$. As we have $T(z)\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ if and only if $z\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, we conclude that for all $z\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $z=0$, we have $V(T(z))=V(z)=0$. We conclude that $V(T(z))\leq 5V(z)/9$ for all $z\in\mathbb{R}^{2}$. Finally, since $\mu>0$, we have $\max_{x\in X^{\mathrm{in}}}V(x)=\max_{y\in[\mu,1/2]}(2\mu^{2})/(2y^{2})^{-1}=(2\mu^{2})(2\mu^{2})^{-1}=1$.
Proposition 14.
If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ then :
  1. 1.
    for all $k\in\mathbb{N}$, $T^{k}(X^{\mathrm{in}})\subseteq V^{-1}([0,1])\subseteq\operatorname{dom}(V)$;
  2. 2.
    for all fixed points $x$ of $T$, $V(x)\in\{0,+\infty\}$.
Proof.
1. From the second statement for all $x\in X^{\mathrm{in}}$, $V(x)\in[0,1]$. From the third statement and the fact that $V$ is nonnegative, we have, for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $0\leq V(T^{k}(x))\leq V(x)\leq 1$.
2. Let $x$ be a fixed point of $T$. If $V(x)=+\infty$, there is nothing to prove. Now, suppose that $V(x)$ is finite. From the third statement, $0\leq V(T(x))=V(x)\leq\lambda V(x)$ for some $\lambda\in(0,1)$. Hence, $V(x)(1-\lambda)=0$ and we conclude that $V(x)=0$. ∎
Corollary 4.
If $X^{\mathrm{in}}$ is included in the fixed point set of $T$ then a function $V:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that $V\circ T\leq\lambda V$ for all some $\lambda\in(0,1)$ cannot satisfy $\overline{V_{X^{\mathrm{in}}}}>0$.
Corollary 4 extends the fact that if a classical Lyapunov function (as defined in Def 7) exists for a self-map $T$, $T$ cannot have non-zero fixed point. If an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ exists, $X^{\mathrm{in}}$ contains a vector $x$ such that $x\neq f(x)$.
Prop. 15 shows that the set of Opt-Lyapunov functions contains the classical Lyapunov functions.
Proposition 15 (Classical is Opt).
If $V$ is a classical Lyapunov function for $T$ and $X^{\mathrm{in}}$ is not reduced to $\{0\}$, then $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Actually, we can replace equivalently the second statement of Def. 8 by $\overline{V_{X^{\mathrm{in}}}}$ is finite and strictly positive. Indeed, with this condition, we recover an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ with $\overline{W_{X^{\mathrm{in}}}}\in(0,1]$ by taking $W:=V/\overline{V}$.
Proposition 16.
The existence of a function $W:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that :
  1. 1.
    $\overline{W_{X^{\mathrm{in}}}}$ is strictly positive and finite ;
  2. 2.
    $W\circ T\leq\rho\circ W$ for a function $\rho:[0,+\infty]\to[0,+\infty]$ such that:
    • $\rho(x)=0\iff x=0$ and $\rho(x)=+\infty\iff x=+\infty$;
    • $\rho_{\mid\mathbb{R}_{+}}\in\mathcal{K}$;
    • $\rho(x)<x$ for all $x\in\mathbb{R}_{+}^{*}$.
is equivalent to the existence of a Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Lemma 4.
Let $\gamma$ be strictly greater than 1. Let $f\in\mathcal{K}_{\infty}$ verifying for all $x>0$, $f(x)>x$. Then, there exists a function $g\in\mathcal{K}$ which satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
Lemma 4 is a slighty modified version of [Kel14, Lemma 25] for which $f-\mathrm{Id}$ is supposed to be in $\mathcal{K}_{\infty}$. To be self-contained, we propose in Appendix a proof inspired by the one provided in [Kel14].
Lemma 5.
Let $\lambda\in(0,1)$. Let $f\in\mathcal{K}$ verifying for all $x>0$, $f(x)<x$. Then, there exists a function $g\in\mathcal{K}$ satisfying $g(f(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
Again a proof of Lemma 5 is given in [Kel14]. A proof is provided in Appendix for the convenience of the reader.
Proof of Prop. 16.
Suppose that such $W$ and $\rho$ exist. Let $\lambda\in(0,1)$. From Lemma 5, there exists $g$ such that $g(\rho(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$. By extension, we define $g(+\infty)=+\infty$. Then as $g$ is increasing, we have $g(W(T(x)))\leq g(\rho(W(x))=\lambda g(W(x))$ for all $x\in\mathbb{R}^{d}$. Now, as $g$ is strictly increasing there exists some $x\in X^{\mathrm{in}}$ such that $g(W(x))$ is strictly positive and $\overline{g\circ W_{X^{\mathrm{in}}}}$ is strictly positive and finite. Finally, $(g\circ W)/\overline{g\circ W_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
The reverse implication is obvious with $\rho(x)=\lambda x$ for all $x\in\mathbb{R}_{+}$. ∎
5.2.2 Operator Ratios From Proper Positive Semidefinite Functions
We extend the concept of operator ratios (defined at Eq. (5)) for which the arguments are positive semidefinite functions rather than positive definite functions. The arguments can also be functions which take infinite values. So, let us denote by $\mathcal{PSD}$ the set of nonzero proper positive semidefinite functions from $\mathbb{R}^{d}$ to $[0,+\infty]$ i.e. :
$\mathcal{PSD}:=\{f:\mathbb{R}^{d}\mapsto[0,+\infty]:\exists\,x\in\mathbb{R}^{d},\ f(x)\in\mathbb{R}_{+}^{*}\}$ (7)
We are able to extend the operator ratio function to $\mathcal{PSD}$ functions.
Definition 9 (Operator Ratios PSD Based).
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $k\in\mathbb{N}$. We associate to $U$ and $k$ the operator ratio function $\overline{N_{U^{k}}}:\mathcal{PSD}\mapsto[0,+\infty]$ defined as follows:
$\overline{N_{U^{k}}}(P):=\sup_{P(x)\in\mathbb{R}_{+}^{*}}\dfrac{P(U^{k}(x))}{P(x)}$
As for $N_{U^{k}}$, we will simply write $\overline{N_{U}}$ for $k=1$.
Recall that $\mathrm{Id}$ denotes the identity function $\mathbb{R}^{d}\ni x:\mapsto x$. It is straightforward to see that, if $P\in\mathcal{PSD}$, $\overline{N_{\mathrm{Id}}}(P)=1$.
Example 6.
We proved at Example 5 that:
$\overline{N_{T}}(V)=\sup_{\{x\in\mathbb{R}^{2}:V(x)>0\}}\dfrac{V(T(x))}{V(x)}=\max_{x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}}\dfrac{V(T(x))}{V(x)}=4/9.$
The operator function is well-defined but the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$ can fail to be true as we can see at Example 7.
Example 7 (A counterexample to the power commutation).
Let us consider the function $U:x\mapsto-x$ and $P:x\mapsto\max\{x,0\}$. We have $\overline{N_{U}}(P)=\sup_{P(x)>0}P(U(x))/P(x)=\sup_{x>0}P(-x)/x=\sup_{x>0}0/x=0$ whereas $\overline{N_{U^{2}}}(P)=\sup_{x>0}P(x)/x=\sup_{x>0}x/x=1$. Finally, we have $\overline{N_{U^{2}}}(P)>\overline{N_{U}}(P)^{2}$.
To obtain the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$, we have to consider a smaller class of operators $P$ and we introduce for a self-map $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, the set $\mathcal{D}\left(U\right)$ of operators defined as follows:
$\mathcal{D}\left(U\right):=\{P\in\mathcal{PSD}:\forall\,x\in\mathbb{R}^{d},\ P(U(x))=0\implies P(U^{2}(x))=0\}$ (8)
We remark that $P\in\mathcal{D}\left(U\right)$ if and only if $P(U^{2}(x))\in(0,+\infty]\implies P(U(x))\in(0,+\infty]$. Moreover, if $P\in\mathcal{D}\left(U\right)$, this leads to for all $k\geq 1$, $P(U^{k+1}(x))\in(0,+\infty]\implies P(U^{k}(x))\in(0,+\infty]$ and we also have if $P(U^{k}(x))\in(0,+\infty]$ for some $k\geq 1$ then $P(U^{j}(x))\in(0,+\infty]$ for all $1\leq j\leq k$.
Example 8 ($P$ not in $\mathcal{D}\left(U\right)$).
Let consider $U(x)=-x$ and $P(x)=\max\{x,0\}$. Let $x$ be a strictly positive real. Then, $P(U(x))=P(-x)=0$. However, $P(U^{2}(x))=P(x)=x>0$. So $P\notin\mathcal{D}\left(U\right)$.
Proposition 17.
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $P\in\mathcal{D}\left(U\right)$. Then, for all $k\in\mathbb{N}^{*}$, $\overline{N_{U^{k}}}(P)\leq\left(\overline{N_{U}}(P)\right)^{k}$ and if $\overline{N_{U}}(P)$ is finite this inequality holds for $k=0$.
Proof.
For $k=0$, the result holds as $\overline{N_{\mathrm{Id}}}(P)=1$ if there exists $x\in\mathbb{R}^{d}$ such that $P(x)>0$ and 0 otherwise. If $\overline{N_{U}}(P)$ is finite $(\overline{N_{U}}(P))^{0}$ is defined and equal to 1.
Now let $k\in\mathbb{N}^{*}$. If $\overline{N_{U}}(P)=+\infty$, the inequality is obviously true. Suppose now that $\overline{N_{U}}(P)<+\infty$. From the defintion of $\overline{N_{U}}(P)$, we have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j-1}(x))\in\mathbb{R}_{+}^{*}$, $P(U^{j}(x))\leq\overline{N_{U}}(P)P(U^{j-1}(x))$. From the fact that $P\in\mathcal{D}\left(U\right)$, we also have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j}(x))=0$, $P(U^{j+1}(x))\leq\overline{N_{U}}(P)P(U^{j}(x))$ as it implies that $P(U^{j+1}(x))=0$. Finally, we have, for all $x\in\mathbb{R}^{d}$ such that $P(x)>0$, $P(U^{k}(x))\leq\overline{N_{U}}(P)P(U^{k-1}(x))$. By dividing by $P(x)$ and taking the supremum, we obtain $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)\overline{N_{U^{k-1}}}(P)$ and we conclude by recurrence. ∎
Proposition 18.
If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, then $V\in\mathcal{D}\left(U\right)$ and $\overline{N_{T}}(V)\in[0,1)$.
Proof.
First $V^{-1}(\mathbb{R}_{+}^{*})$ is nonempty from the second statement of Def. 8. The third statement implies that $V(T^{2}(x))\leq\lambda V(T(x))$ for all $x\in\mathbb{R}^{d}$ for some $\lambda\in(0,1)$ and thus $V(T^{2}(x))>0$ implies that $V(T(x))>0$. We conclude that $V\in\mathcal{D}\left(T\right)$. The third statement also implies that $V(T(x))\leq\lambda V(x)$ for all $x\in\operatorname{dom}(V)$ for some $\lambda\in(0,1)$ and thus $N_{T}(V)\leq\lambda<1$. It can be equal to 0 if and only if for all $x$ such that $V(x)>0$, $V(T(x))=0$. ∎
To prove that a function $V\in\mathcal{PSD}$ is an Opt-Lyapunov for $X^{\mathrm{in}},T)$, we could use any value greater $\overline{N_{T}}(V)$. The computation of $\overline{N_{T}}(V)$ ensures to take the best value possible. It is important in the context of using the function $\mathfrak{F}_{u}(k,h,\beta)$ with an Opt-Lyapunov function since $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is decreasing (Prop. 12).
Proposition 19.
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. Suppose that there exists a classical Lyapunov function $V$ for $U$. Then, $V\in\mathcal{D}\left(U\right)$ and $\overline{N_{U}}(V)=N_{U}(V)$.
Proof.
If there exists a Lyapunov function for $U$, then $U(0)=0$. From the double inequalities $\alpha\circ\|\cdot\|\leq V\leq\beta\circ\|\cdot\|$ and thus $V(x)=0$ iff $x=0$. Thus if $V(U^{2}(x))>0$ then $U^{2}(x)\neq 0$ and thus $U(x)\neq 0$ hence $V(U(x))>0$ and $V\in\mathcal{D}\left(U\right)$. The equality between $\overline{N_{U}}(V)$ and $N_{U}(V)$ follows readily from the fact that $\{y\in\mathbb{R}^{d}:V(y)>0\}=\mathbb{R}^{d}\backslash\{0\}$. ∎
5.2.3 $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility and $(X^{\mathrm{in}},T,\varphi)$-compatibility
In the proof of Th. 7, the construction of $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ from a classical Lyapunov function $V$ is based on the existence of a function $\alpha\in\mathcal{K}_{\infty}$ such that $\varphi\leq\alpha\circ V$. The existence of this latter comes from the fact that in Th. 7, $\varphi$ is continuous and, by definition of $V$, $\alpha^{\prime}\circ\|\cdot\|\leq V$ for some $\alpha^{\prime}\in\mathcal{K}_{\infty}$. However, when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$ is not necessarily positive definite and we cannot guarantee the existence of $\alpha\in\mathcal{K}_{\infty}$ such that $\alpha\circ\varphi\leq V$. We have to use a different class of function to find a similar relation between $\varphi$ and $V$. First, we propose to restrict the interval where a function $\alpha$ should be continuous and strictly increasing. Indeed, the analysis only concerns the values $\{\varphi(T^{k}(x)),k\in\mathbb{N},x\in X^{\mathrm{in}}\}$ and $\{V(x),x\in X^{\mathrm{in}}\}$. Hence, we introduce the following interval :
$I_{X^{\mathrm{in}},T}^{\varphi}:=\overline{\operatorname{conv}}(\{\varphi(T^{k}(x)):k\in\mathbb{N},\ x\in X^{\mathrm{in}}\})\in\mathbb{I}\left(\mathbb{R}\right)$
Proposition 20.
The following assertions hold:
  1. 1.
    For all $k\in\mathbb{N}$, $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{n\in\mathbb{N}}\nu_{n}$ and $\liminf_{n\in\mathbb{N}}\nu_{n}$ also belong to $I_{X^{\mathrm{in}},T}^{\varphi}$.
  2. 2.
    $I_{X^{\mathrm{in}},T}^{\varphi}=\left[\inf_{k\in\mathbb{N}}\nu_{k},\sup_{k\in\mathbb{N}}\nu_{k}\right]\cap\mathbb{R}$
Proof.
Let $k\in\mathbb{N}$. There exists for all $n\in\mathbb{N}$, $x_{n}\in X^{\mathrm{in}}$ such that $\nu_{k}-1/n\leq\varphi(T^{k}(x_{n}))\leq\nu_{k}$ and thus the sequence $(\varphi(T^{k}(x_{n})))_{n\in\mathbb{N}}$ converges to $\nu_{k}$ and thus $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{k\in\mathbb{N}}\nu_{k}$ and $\liminf_{k\in\mathbb{N}}\nu_{k}$ are limits of subsequences of $(\nu_{k})_{k\in\mathbb{N}}$ and they belongs to $I_{X^{\mathrm{in}},T}^{\varphi}$. From Prop. 1, if $\limsup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\liminf_{k\in\mathbb{N}}\nu_{k}$) is infinite, the value $\sup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\inf_{k\in\mathbb{N}}\nu_{k}$ is equal to +$\infty$+ ∞ (resp. $\infty$- ∞).
Now, if one of the value $\inf_{k\in\mathbb{N}}\nu_{k}$ or $\sup_{k\in\mathbb{N}}\nu_{k}$ is not finite, the infinity passes to the related bound of $I_{X^{\mathrm{in}},T}^{\varphi}$. Now if those values are both finite, we can extract sequences in $I_{X^{\mathrm{in}},T}^{\varphi}$ which converge respectively to $\inf_{m\in\mathbb{N}}\nu_{m}$ and $\sup_{m\in\mathbb{N}}\nu_{m}$. For example, we can define $\varphi(T^{k_{n}}(x_{k_{n}}))\subset I_{X^{\mathrm{in}},T}^{\varphi}$ such that for all $n\in\mathbb{N}$, $\inf_{m\in\mathbb{N}}\nu_{m}-1/n\leq\nu_{k_{n}}-1/n\leq\varphi(T^{k_{n}}(x_{k_{n}}))\leq\nu_{k_{n}}\leq\inf_{m\in\mathbb{N}}\nu_{m}+1/n$. This sequence converge to $\inf_{m\in\mathbb{N}}\nu_{m}$. ∎
Let $V$ be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. To pass from $\alpha\circ\varphi\leq V$ to $\varphi\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\alpha\circ V$, $\alpha$ has to be surjective and strictly increasing as a function of $J\to\{V(x),x\in X^{\mathrm{in}}\}\subseteq[0,1]$. From the earlier discussion, $J$ must also contain $I_{X^{\mathrm{in}},T}^{\varphi}$. Therefore, we define the following set of functions:
$\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}:=\{\alpha:\mathbb{R}\mapsto\mathbb{R}:\exists\,I\in\mathbb{I}\left(\mathbb{R}\right)\text{ s.t. }\alpha(I)=[0,1]\text{ and}\operatorname{conv}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)\in\mathrm{SIC}(\alpha)\}$
Proposition 21.
Let $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$. There is only one interval $I$ such that $\alpha(I)=[0,1]$ and $I\in\mathrm{SIC}(\alpha)$.
Proof.
Let us consider $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and $I$ such that $\alpha(I)=[0,1]$. As $\alpha$ is continuous and strictly increasing on $I$, ${{\alpha}_{\left(I\right)}^{-1}}$ exists and is also continuous and strictly increasing from $[0,1]\to I$. Hence, $I$ is compact. Moreover, such interval $I$ is unique. Otherwise, consider $I,J\in\mathrm{SIC}(\alpha)$ for which we have $I\neq J$ and $\alpha(I)=\alpha(J)=[0,1]$. Suppose there exists $x\in I$ such that $x\leq\min J$. We have $\alpha(x)<\alpha(\min J)=0$ but $x\in I$ and $\alpha(x)\geq\alpha(\min I)=0$. ∎
We present some useful properties of functions belonging to $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and their relation with $\nu$.
Proposition 22.
For all $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that there exists an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$, satisfying $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, we have $\alpha(\limsup_{k\to+\infty}\nu_{k})\leq 0$.
Proof.
Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. By assumption, $\alpha(\varphi(T^{k}(x)))\leq V(\varphi(T^{k}(x)))\leq\lambda^{k}V(x)$. Taking the sup over $X^{\mathrm{in}}$ and from the increasing condition and the continuity of $\alpha$, $\alpha(\nu_{k})\leq\lambda^{k}\overline{V_{X^{\mathrm{in}}}}$. Taking the limsup, we get $\limsup_{k\mapsto+\infty}\alpha(\nu_{k})=\alpha(\limsup_{k\mapsto+\infty}\nu_{k})\leq 0$. ∎
Corollary 5.
Suppose there exist $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ $V$ such that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Then:
  1. 1.
    $\nu\in\Lambda$;
  2. 2.
    If ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $\alpha(\varphi(T^{k}(x)))\leq 0$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$.
Proof.
The first statement follows readily from Prop. 22. The second statement follows from Prop. 3 and Prop. 22, since we have, if ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$. ∎
The set $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ defines the class of functions that we must consider to obtain relations $\alpha\circ\varphi\leq V$ when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Another important consideration to take into account is usefulness of functions $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ based on $\alpha$. An Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is always nonnegative and if $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}$, the inequality is obvious. Then $\alpha(\nu_{k})$ must be strictly positive for some $k\in\mathbb{N}$. By Prop. 22 and Corollary 5, we have automatically $k\in{\mathrm{R}}_{\nu}^{s}$. We propose a special definition for those particular functions $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.
Definition 10 ($(X^{\mathrm{in}},T,\varphi)$-Certificate of compatibility).
Let $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$. A function $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that $\alpha(\nu_{k})>0$ for some $k\in{\mathrm{R}}_{\nu}^{s}$ and $\alpha(\varphi(T^{k}(x)))\leq g(T^{k}(x))$ for all $k\in\mathbb{N}$ and all $x\in X^{\mathrm{in}}$ is called a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$.
The problem is to decide which conditions on $X^{\mathrm{in}}$, $T$, $\varphi$ and $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$ ensure the existence of $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$. Therefore, we introduce the concept of $(X^{\mathrm{in}},T,\varphi)$-compatibility.
Definition 11 ($(X^{\mathrm{in}},T,\varphi)$-compatibility).
A function $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$ is said to be $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$ if and only if there exist $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{k}-\eta$, we have $g(T^{j}(x))>\varepsilon$.
We will say that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if there exists some $k\in{\mathrm{R}}_{\nu}^{s}$ at which $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
To be $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$ implies that if $T^{k}(x)$ is $\eta$-optimal for $\varphi$ then $g(T^{k}(x))$ is strictly positive. We extend the strict positivity of $g$ to all orbits of $T$ starting in $X^{\mathrm{in}}$ greater than $\nu_{k}-\eta$.
Proposition 23.
If $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$, it is $(X^{\mathrm{in}},T,\varphi)$-compatible at every $j\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{j}\geq\nu_{k}$.
Proof.
Suppose that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in\mathbb{N}$. Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{k}-\eta$, we have $g(T^{j}(x))>\varepsilon$. Let $j\in\mathbb{N}$ such that $\nu_{j}\geq\nu_{k}$. Let $x\in X^{\mathrm{in}}$, and $l\in\mathbb{N}$ such that $\varphi(T^{l}(x))>\nu_{j}-\eta$. This implies that $\varphi(T^{l}(x))>\nu_{j}-\eta\geq\nu_{k}-\eta$ and thus $g(T^{l}(x))>\varepsilon$. ∎
The existence of $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$ is actually equivalent to the $(X^{\mathrm{in}},T,\varphi)$-compatibility of $g$.
Theorem 8.
A function $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if and only if there exists $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$.
Proof.
Suppose that there exists $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that $\alpha(\nu_{k})>0$ for some $k\in{\mathrm{R}}_{\nu}^{s}$ and $\alpha(T^{k}(x))\leq g(T^{k}(x))$ for all $k\in\mathbb{N}$ all $x\in X^{\mathrm{in}}$. Let $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{k})>0$. From Prop. 23, we can suppose that $\nu_{k}>\inf_{n\in\mathbb{N}}\nu_{n}$ and so there exists $\delta>0$ such that $(\nu_{k}-\delta,\nu_{k})\subset I_{X^{\mathrm{in}},T}^{\varphi}$. As $\alpha$ is continuous and strictly increasing on $I_{X^{\mathrm{in}},T}^{\varphi}$, for $\varepsilon:=\alpha(\nu_{k})/2$, there exists $\eta\in(0,\delta]$, such that $0<\nu_{k}-s\leq\eta$ implies that $\varepsilon<\alpha(s)<\alpha(\nu_{k})+\varepsilon$. Now let us consider $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{j}(x))>\nu_{k}-\eta$. We have either $\nu_{k}>\varphi(T^{j}(x))$ or $\varphi(T^{j}(x))\geq\nu_{k}$. In the first case, we have, by hypothesis, $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))>\varepsilon$. In the second case, as $\alpha$ is increasing and $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))\geq\alpha(\nu_{k})>\varepsilon$.
Now, assume that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at some $k\in{\mathrm{R}}_{\nu}^{s}$. Then, from Prop.23, $g$ is also $(X^{\mathrm{in}},T,\varphi)$-compatible at $k=\mathbf{K}_{\nu}$. Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{\rm opt}-\eta$, we have $g(T^{j}(x))>\varepsilon$. Then we can define : $\alpha:s\mapsto s-\nu_{\rm opt}+\min\{\eta,\varepsilon\}$. Let $x\in X^{\mathrm{in}}$ and $j\in\mathbb{N}$. Then either $\varphi(T^{j}(x))\leq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$ or $\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}$. For the first case, $\alpha(\varphi(T^{j}(x)))\leq 0\leq g(T^{j}(x)))$. For the second, $0<\alpha(\varphi(T^{j}(x)))\leq\min\{\eta,\varepsilon\}$. By assumption on $g$, $g(T^{j}(x))>\varepsilon$. Note that the set $\{(x,j)\in X^{\mathrm{in}}\times\mathbb{N}:\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty. Finally, as $\alpha$ is strictly increasing, for all $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$, we have $\alpha(\nu_{k})>0$. Note that the set $\{k\in\mathbb{N}:\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty as $[\mathbf{K}_{\nu},\mathbf{K}^{s}_{\nu}]\cap\mathbb{N}$ is included in it. ∎
Example 9.
Let consider the Opt-Lyapunov function $V$ of Example 5. We prove that $V$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
We define for all $s\in\mathbb{R}$, $\alpha(s):=s-p^{2}/3+\min\{\zeta,\varepsilon\}$ where $0<\zeta<p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ and $\varepsilon:=\min\{\varepsilon_{k},k\in\operatorname{Argmax}{\nu}\}$ where $\varepsilon_{k}=\inf\{V(T^{k}(x)):x\in X^{\mathrm{in}}\}$. The scalar $\varepsilon$ is strictly positive as the function $[\mu,1/2]\ni y\mapsto V(T^{k}((2,1)y)$ is continuous and strictly positive for all $k\in\mathbb{N}$ and $\operatorname{Argmax}(\nu)$ is a finite set of integers. The function $\alpha$ is clearly strictly increasing and continuous on $\mathbb{R}$.
We have for all $j\in\mathbb{N}$ and for all $x\in X^{\mathrm{in}}$, $\alpha(\varphi(T^{j}(x)))=\varphi(T^{j}(x))-p^{2}/3+\min\{\zeta,\varepsilon\}<\varphi(T^{j}(x))-p^{2}/3+p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$. Then $\alpha(\varphi(T^{j}(x)))$ is strictly negative for all $j\notin\operatorname{Argmax}(\nu)$. Now let $j\in\operatorname{Argmax}(\nu)$. Suppose that $\varphi(T^{j}(x))>p^{2}/3-\min\{\zeta,\varepsilon\}$. As we have $\varphi(T^{j}(x))\leq p^{2}/3$, we get $\alpha(\varphi(T^{j}(x)))\leq\varepsilon\leq V(T^{k}(x))$ for all $k\in\operatorname{Argmax}(\nu)$ and all $x\in X^{\mathrm{in}}$. This leads to $\alpha(\varphi(T^{j}(x)))\leq V(T^{j}(x))$. On the other hand, $\alpha(\varphi(T^{j}(x)))\leq 0$ for all $j\in\operatorname{Argmax}(\nu)$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))\leq p^{2}/3-\min\{\eta,\varepsilon\}$. We have proved that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$. Moreover, as the set of $\{x\in X^{\mathrm{in}}:\varphi(T^{j}(x))>p^{2}/3-\min\{\eta,\varepsilon\}\}$ is nonempty for all $j\in\operatorname{Argmax}(\nu)$, we conclude that for all $j\in\operatorname{Argmax}(\nu)$, there exists $x\in X^{\mathrm{in}}$ such that $0<\alpha(\varphi(T^{j}(x)))$ and as $\alpha$ is increasing $0<\alpha(\nu_{j})$ for all $j\in\operatorname{Argmax}(\nu)$ and $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$ making $V$ $(X^{\mathrm{in}},T,\varphi)$-compatible.
Actually, by simple computations, we have $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$. One can compute $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ as follows. Let $n\in\mathbb{N}$. We have, for all $y\in\mathbb{R}_{+}^{*}$, $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$. Thus $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$ is strictly negative if and only if $n\geq n_{y}^{*}:=\max\{\lfloor\ln(4p/(15y))/\ln(3/2)\rfloor+1,0\}$. Then for all $y\in\mathbb{R}_{+}^{*}$, $\mathbb{N}\ni n\mapsto\varphi_{n}(y)$ is maximized at $n_{y}^{*}$. We deduce that $n_{1/2}^{*}=\lfloor\ln(8p/15)/\ln(3/2)\rfloor+1$ and, as $\mu\in(0,1/3)$, $n_{\mu}^{*}=\lfloor\ln(4p/(15\mu))/\ln(3/2)\rfloor+1$. This implies that $n_{1/2}^{*}\leq\underline{n}-1$ and thus we have $\nu_{n}\leq\nu_{n_{1/2}^{*}}$ for all $n\leq\underline{n}-1$. This also implies that $n_{\mu}^{*}<\overline{n}$ and so $\nu_{n}<\nu_{\overline{n}}$ for all $n>\overline{n}$. Finally, $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}=\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\}$.
5.2.4 Direct and Converse Opt-Lyapunov Theorems
We present both direct and converse Opt-Lyapunov theorems using the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion. More precisely, the direct version aims to prove that if a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function exists (and such that $\overline{N_{T}}(V)>0$), then we can construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$. The converse version exhibits a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function when a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$. Moreover, we are able to produce a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for this Opt-Lyapunov function.
We remark that from Prop. 14, if $V$ is an Opt-Lyapunov for $(X^{\mathrm{in}},T)$ then $\{T^{k}(x):k\in\mathbb{N},x\in X^{\mathrm{in}}\}\subseteq\operatorname{dom}(V)$.
Proposition 24.
Assume that there exists a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov $V$ such that $N_{V}(T)=0$. Then $\mathbf{K}^{s}_{\nu}=0$ and $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$.
Proof.
If $N_{V}(T)=0$ then for all $x\in\mathbb{R}^{d}$ such that $V(x)>0$, $V(T(x))=0$. This also implies that $V(T^{k}(x))=0$ for all $k\in\mathbb{N}^{*}$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. Thus, for all $k\in\mathbb{N}^{*}$, $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))=0$ and as $\alpha$ is increasing and continuous, $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}^{*}$. As for some $j\in{\mathrm{R}}_{\nu}^{s}$, $\alpha(\nu_{j})>0$, this leads to $j=0$ and as $\alpha$ is strictly increasing and continuous on $I_{X^{\mathrm{in}},T}^{\varphi}$ then for all $k\in\mathbb{N}^{*}$, $\nu_{0}>\nu_{k}$ and $\mathbf{K}^{s}_{\nu}=0$. This also implies that $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$. ∎
Theorem 9 (Direct Opt-Lyapunov Theorem).
Assume that there exists a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov $V$ such that $N_{V}(T)>0$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$. Let $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)$ where $\alpha(I)=[0,1]$. We define the function $h:x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$. Then $(h,N_{V}(T))\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$.
Proof.
The function ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and continuous. Thus, we have, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(V(T^{k}(x))$. By Def. 9 and Prop. 17, $V(T^{k}(x))\leq(N_{V}(T))^{k}V(x)\leq N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing, we conclude that $\nu_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}\right)$. Note that $x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$ is strictly increasing and continuous on $J\supset[0,1]$ as ${{\alpha}_{\left(J\right)}^{-1}}$ is and $\overline{V_{X^{\mathrm{in}}}}$ is not zero. Moreover, by Prop. 18 and hypothesis, $N_{V}(T)\in(0,1)$. As $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, there exists $j\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{j})>0$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and $\alpha(I)\ni 0$, $\nu_{j}>\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(0)=\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(0\times\overline{V_{X^{\mathrm{in}}}}\right)=h(0)$. ∎
Example 10.
Following Th. 9, from the Opt-Lyapunov function $V$ defined at Example 5 and the ($X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, $\alpha$ defined at Example 9, the function $h:[0,1]\ni s\mapsto\alpha_{\mathbb{R}}^{-1}(s\overline{V_{X^{\mathrm{in}}}})$ belongs to $\mathbb{\Gamma}_{\rm fun}(\nu)$ and $\overline{N_{T}}(V)=4/9\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$. Then $h_{V}(s)=s+p^{2}/3-\min\{\zeta,\varepsilon\}$ with $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$ and $\zeta\in(0,p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})$ where $n_{1/2}^{*}$ is the maximizer for the function $n\mapsto\varphi_{n}(1/2)$.
We draw a table as we done at Example 3 to compare the direct (from the function $h$ of Example 3) and the Opt-Lyapunov approaches. We have some flexibility about $\zeta$ so we propose two values for a comparison purpose. Hence, we compare the computations for $h_{\zeta_{1}}$ associated to $\zeta_{1}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/2$ and $h_{\zeta_{2}}$ associated to $\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$. The evaluations for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},4/9)\rfloor$ and for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},4/9)\rfloor$ are given at Table 3.
$p=30$ $p=9$ $p=3$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mathbf{K}^{s}_{\nu}$ 8 11 22 5 8 19 2 3 17
$\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ 9 12 23 7 9 20 3 7 18
$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},\frac{4}{9})\rfloor$ 8 14 36 5 11 33 4 8 31
$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},\frac{4}{9})\rfloor$ 8 14 36 9 11 33 12 12 31
Table 3: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},\beta)\rfloor$ described at Eq. (4) for the sequence $\nu$ of Eq. (2) and two pairs $(h_{\zeta_{1}},4/9)$ and $(h_{\zeta_{2}},4/9)$ of Example 10.
The evaluations of $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},4/9)\rfloor$ from the Opt-Lyapunov function proposed at Example 5 and associated $(X^{\mathrm{in}},T,\varphi)$-certificates of compatibility provide worst upper bounds when $\mu$ is small (equal to $1/10$ or $1/1000$) but the optimal value for $\mu=1/3$. A small value for $\zeta$ ($\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$) makes decrease the precision for $\mu=1/3$. We recall that the y-intercept of the affine function $h_{\zeta}$ is $p^{2}/3-\min\{\varepsilon,\eta\}$ where $\varepsilon$ is a very small value when $\mu$ is small. Hence, when $\mu=1/10$ or $\mu=1/1000$, $\zeta_{1}$ or $\zeta_{2}$ are still greater than $\varepsilon$.
Note that the values $h_{\zeta_{i}}^{-1}(\nu_{k})$ are strictly positive if and only if for $k\in\operatorname{Argmax}(\nu)$. This is sufficient to compute an upper bound of $\mathbf{K}^{s}_{\nu}$. ∎
We establish a ”converse Lyapunov theorem”. The construction of the Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is inspired by the one proposed by Yoshizawa [Yos66].
Theorem 10 (Converse Opt-Lyapunov Theorem).
Let $(h,\beta)\in\mathbb{\Gamma}(\nu)$ be useful for $\nu$. Let us define $s\mapsto\omega(s):=\min\left\{h(1),\max\left\{s,h(0)\right\}\right\}$. Then, the function $V\in\mathcal{PSD}$ defined as follows:
$x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega\left(\varphi(T^{k}(x))\right)\right)$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Moreover, the function $\hat{h}$ defined as follows:
$\hat{h}:\mathbb{R}\mapsto\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in(h(0),h(1)]\end{array}\right.$
is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$ making the function $V$ $(X^{\mathrm{in}},T,\varphi)$-compatible.
Proof.
First, we prove that the function $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Let us prove that $V$ takes its values in $[0,+\infty]$. Let $x\in\mathbb{R}^{d}$. As we have $h(1)\geq\omega(s)\geq h(0)$, we have, as ${{h}_{\left([0,1]\right)}^{-1}}$ is increasing, $1\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\omega(\varphi(T^{k}(x))))\geq 0$ for all $k\in\mathbb{N}$. We conclude, as $\beta\in(0,1)$ that $V(x)\geq 0$ and $V(x)=+\infty$ for all $x\in\mathbb{R}^{d}$ such that $\varphi(T^{k}(x))\geq h(1)$ for all $k$ sufficiently large.
By definition of $\mathbb{\Gamma}(\nu)$, we have for all $x\in X^{\mathrm{in}}$ and all $n\in\mathbb{N}$, $\varphi(T^{n}(x))\leq h(\beta^{n})<h(1)$ and so $V(x)\leq\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(\beta^{k}))=\sup_{k\in\mathbb{N}}\beta^{-k}\beta^{k}=1$. This implies that $\overline{V_{X^{\mathrm{in}}}}\in[0,1]$. By assumption, $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and $\mathcal{S}(\nu,h)$ is nonempty. Let us take $k\in\mathbb{N}$ such that $h(1)>\nu_{k}>h(0)$. So there exists $\varepsilon>0$ such that $\nu_{k}-\varepsilon>h(0)$. For this $\varepsilon$, there exists $x\in X^{\mathrm{in}}$ such that $\nu_{k}\geq\varphi(T^{k}(x))\geq\nu_{k}-\varepsilon$. Finally, for this $x$, $V(x)\geq\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\varphi(T^{k}(x))>\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$ as ${{h}_{\left([0,1]\right)}^{-1}}$ strictly increases. We conclude that $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$. Finally, we have to prove that $V\circ T\leq\beta V$. Let $y\in\mathbb{R}^{d}$, if $V(y)=+\infty$, the result holds. Now suppose that $y\in\operatorname{dom}V$, we have:
$\begin{array}[]{rl}V(T(y))=&\displaystyle{\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{k+1}(y))\right)}\\ =&\displaystyle{\sup_{j\in\mathbb{N}^{*}}\beta^{-j+1}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ =&\displaystyle{\beta\sup_{j\in\mathbb{N}^{*}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ \leq&\displaystyle{\beta\sup_{j\in\mathbb{N}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)=\beta V(y)}\end{array}$
We conclude that $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
[…]
6 A Technical Summary
We have proposed a function $\mathfrak{F}_{\nu}$ at Eq. (4) whose the integer part of values provide solutions for Problem 2. This function is based on couples $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$. Without this condition, $(h,\beta)$ is useless as $\mathfrak{F}_{\nu}$ is equal to +$\infty$+ ∞. When the supremum of $(\nu_{k})_{k}$ is strictly greater than its limit superior, a useful couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ exists (Th. 1) and $h$ can be chosen as an affine function. Furthermore, we can compute the greatest maximizer of $(\nu_{k})_{k}$ from this couple (Th. 3) and thus we can solve Problem 1. Unfortunately, to directly compute a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is as difficult as solving Problem 1 and we have to construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from alternative methods. We propose alternatives coming from classical tools from stability theory of difference equations.
The inequality $\varphi(T^{k}(x))\leq h(\beta^{k})$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ can be viewed as a kind of $\mathcal{KL}$ inequality that appears in asymptotic stability studies. Indeed, $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$ making the sequence $(h(\beta^{k}))_{k}$ strictly decreasing and convergent to $h(0)$. From this observation, we propose a more general $\mathcal{K}\mathcal{L}$ class of functions called $\mathcal{KL}_{\rm gen}$ where the positivity with respect to the state variable and the continuity assumptions are removed (compared with the $\mathcal{K}\mathcal{L}$ class). We have proved (Th. 5) that a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ exists if and only if $\nu_{\rm opt}$ is finite. Then, we extend Sontag’s result to the $\mathcal{KL}_{\rm gen}$ class (Lemma 2 and Lemma 3) to construct $h\in\Omega([0,1])$, useful for $\nu$, from a $\mathcal{KL}_{\rm gen}$ function and such that $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ (Th. 6). One of the advantage of the Sontag’s result extension is that the scalar part of elements of $\mathbb{\Gamma}(\nu)$ can be fixed and it remains to construct $h$ from the $\mathcal{KL}_{\rm gen}$ certificate. In practice, to decide whether a function $(\gamma,\operatorname{Fin}\left(X^{\mathrm{in}}\right))$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$ seems difficult to make.
One constructive alternative of $\mathcal{K}\mathcal{L}$ certificates consists in Lyapunov functions. Classical Lyapunov functions satisfy pure functional inequations and are not defined from the orbits of the dynamical system. One of the most interesting theoretical result for Lyapunov functions is the necessity of their existence for the stability (namely converse Lyapunov theorems). For our specific problem, we propose a new type of Lyapunov function called Opt-Lyapunov functions for which we obtain both direct (Th. 9) and converse theorem (Th 10). The main difference between classical and Opt-Lyapunov functions is that Opt-Lyapunov functions are nonnegative but not definite (many zeros) and can take infinite values (outside reachable values taken by the dynamical system). The fact that Opt-Lyapunov functions are not definite (and not continuous) makes classical comparison functions impossible to use. In the classical case, comparison functions serve as link between our objective function $\varphi$ and the classical Lyapunov function. This link (and comparison functions) is replaced by the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion (Th. 8). Theorem 11 sums up the equivalences proved in this paper relatively to the peaks computation problem.
Theorem 11 (Summary).
The following statements are equivalent:
  1. 1.
    The set ${\mathrm{R}}_{\nu}^{s}$ is nonempty;
  2. 2.
    There $(h,\beta)\in\mathbb{\Gamma}(\nu)$ for which $\mathcal{S}(\nu,h)$ is nonempty;
  3. 3.
    There exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)$ is finite and $\nu_{\rm opt}=\max_{k\in\mathbb{N}}\{\nu_{k}:k=0,\ldots,\mathbf{K}^{s}_{\nu}\}$;
  4. 4.
    There exists $\theta\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ and $\gamma\in\mathcal{KL}_{\rm gen}$ such that $\varphi(T^{k}(x))\leq\gamma(\theta(x),k)$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$ and $\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{n}$ for some $n\in\mathbb{N}$;
  5. 5.
    There exists an Opt-Lyapunov $(X^{\mathrm{in}},T,\varphi)$ compatible function.
The proof of Theorem is done by sketching the functions constructed in the proofs of earlier theorems. Figure 1 provides a graphical view of different results involved to prove Th. 11. In Figure 1, the notation $\mathcal{OL}_{\rm comp}$ stands for the set of $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov functions.
[…]
Figure 1: The scheme showing the mappings between $\mathcal{KL}_{\rm gen}$, $\mathcal{OL}_{\rm comp}$ and $\mathbb{\Gamma}(\nu)$
As $V\in\mathcal{OL}_{\rm comp}$ is $(X^{\mathrm{in}},T,\varphi)$-compatible, we associate with them a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. For a such a function $V$, we denote by $\alpha_{V}$ the chosen $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
We now clarify the circled numbers that appear in Figure 1.
[…]
    $\begin{array}[]{ccc}\mathbb{\Gamma}(\nu)&\longrightarrow&\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)\\ (h,\beta)&\longmapsto&\left(\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\displaystyle{\min\{s,h(\beta^{t})\}},\displaystyle{\mathbb{R}^{d}\ni x\mapsto\sup_{k\in\mathbb{N}}\varphi(T^{k}(x))}\right)\end{array}$
[…]
    $\begin{array}[]{ccc}\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)&\longrightarrow&\mathbb{\Gamma}(\nu)\\ (\gamma,\theta)&\longmapsto&(h,e^{-1})\end{array}$
    The function $h$ cannot be explicitly formulated and is constructed following the proofs of Lemma 2 and Lemma 3.
[…]
    $\begin{array}[]{ccccc}\mathcal{OL}_{\rm comp}&\longrightarrow&\mathcal{OL}_{\rm comp}\times\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}&\longrightarrow&\mathbb{\Gamma}(\nu)\\ V&\longmapsto&(V,\alpha_{V})&\longmapsto&(h_{V},\overline{N_{T}}(V))\end{array}$
    where $h_{V}$ is defined as follows:
    $[0,1]\ni s\mapsto h_{V}(s)=\displaystyle{{\alpha_{V}}_{\left(J\right)}^{-1}}(s\overline{V_{X^{\mathrm{in}}}})$
    where $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)\in\mathrm{SIC}(\alpha_{V})$ where $\alpha_{V}(I)=[0,1]$.
[…]
    $\begin{array}[]{ccc}\mathbb{\Gamma}(\nu)&\longrightarrow&\mathcal{OL}_{\rm comp}\times\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}\\ (h,\beta)&\longmapsto&(V,\alpha_{V})\end{array}$
    where:
    • $\bullet$
      ${\mathbb{R}^{d}\ni x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{k}(x)))\right)}$;
    • $\bullet$
      ${\mathbb{R}\ni s\mapsto\omega(s)=\min\{h(1),\max\{s,h(0)\}\}}$;
    • $\bullet$
      $(-\infty,h(1)]\ni s\mapsto\alpha_{V}(s)=\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in[h(0),h(1)]\end{array}\right.$
Appendix
Proof of Lemma 4.
Lemma 4.
Let $\gamma$ be strictly greater than 1. Let $f\in\mathcal{K}_{\infty}$ verifying for all $x>0$, $f(x)>x$. Then, there exists a function $g\in\mathcal{K}$ which satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
Proof.
As we have $f(x)>x$ for all $x>0$, we have $f(1)>1$ and thus the interval $[1,f(1))$ is nonempty. We define $h:[1,f(1))$ continuous, strictly increasing whose values are strictly positive such that $h(1)=1$ and $\lim_{x\mapsto f(1)^{-}}h(x)=\gamma$. We could take for example $h$ the affine map linking $(1,1)$ to $(f(1),\gamma)$. Now, as the sequence $(f^{k}(1))_{k}$ is strictly increasing and unbounded (if $M=\sup_{k\in\mathbb{N}}f^{k}(1)>0$ is finite then as $f$ is increasing and continuous $f(M)=M$ but $M<f(M)$ by assumption on $f$). As $f$ is invertible on $\mathbb{R}_{+}$ and we have $f^{-1}(x)<x$ for all $x>0$, the sequence $(f^{-k}(1))_{k}$ is strictly decreasing and positive and converges to 0 as $k$ tends to +$\infty$+ ∞. Hence $\mathbb{R}_{+}^{*}$ is union of disjoint intervals $[f^{k}(1),f^{k+1}(1))$ where $k\in\mathbb{Z}$. Then for all $x>0$, there exists a unique $k\in\mathbb{Z}$ such that $x\in[f^{k}(1),f^{k+1}(1))$. We define a function $g$ on $\mathbb{R}_{+}$ as follows $g(0)=0$ and for all $x>0$, $g(x):=\gamma^{k}h(f^{-k}(x))$ where $k$ is the unique integer such that $x\in[f^{k}(1),f^{k+1}(1))$. As $g$ is equal to $h$ on $[1,f(1)]$, then $g$ is strictly increasing, continuous and strictly positive on $[1,f(1)]$. Then, on the open interval $(f^{k}(1),f^{k+1}(1))$ for $k\neq 0$, $g$ is the composition $t\mapsto\gamma^{k}t$, $h$ and $f^{-k}$ and thus is strictly increasing, continuous and strictly positive on $(f^{k}(1),f^{k+1}(1))$ for all $k\in\mathbb{Z}$. Now, for all $k\neq 0$, $\lim_{y\to{f^{k}(1)}^{-}}g(y)=\lim_{y\to{f^{k}(1)}^{-}}\gamma^{k-1}h(f^{-k+1}(y))=\lim_{z=f^{-k}(y)\to 1^{-}}\gamma^{k-1}h(f(z))=\gamma^{k-1+1}=\gamma^{k}$. We also have $\lim_{y\to{f^{k}(1)}^{+}}g(y)=\lim_{y\to{f^{k}(1)}^{+}}\gamma^{k}h(f^{-k}(y))=\lim_{z\to 1^{-}}\gamma^{k}h(z)=\gamma^{k}h(1)=\gamma^{k}$. We conclude that $g$ is continuous on $\mathbb{R}_{+}^{*}$. Now, we prove that $g$ is continuous at 0 (at right). Let us consider a sequence $(x_{n})_{n\in\mathbb{N}}$ of positive reals which converges to 0. Let $n\in\mathbb{N}$. There exists a unique $k_{n}\in\mathbb{Z}$ such that $f^{k_{n}}(1)\leq x_{n}\leq f^{k_{n}+1}(1)$. We get $1\leq f^{-k_{n}}(x_{n})\leq f(1)$ and as $h$ is increasing on $[1,f(1))$ and continuous (at left) at $f(1)$, $1=h(1)\leq h(f^{-k_{n}}(x_{n}))\leq h(f(1))=\gamma$. Finally, $\gamma^{k_{n}}\leq g(x_{n})=\gamma^{k_{n}}h(f^{-k_{n}}(x_{n}))\leq\gamma^{k_{n}+1}$. As $(x_{n})_{n\in\mathbb{N}}$ tends to 0, the sequence $(k_{n})_{n\in\mathbb{N}}$ tends to $\infty$- ∞ and we conclude that $(g(x_{n}))_{n\in\mathbb{N}}$ tends to $0=g(0)$ and $g$ is continuous at 0. We have already proved that $g$ is strictly increasing on each interval $[f^{k}(1),f^{k+1}(1))$. It is easy to see that it is also the case when we pick two reals $x,y$ in two different intervals : $x<y$ implies $g(x)<g(y)$.
Finally, we have to prove that $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$. As $f(0)=0$, this holds for $x=0$. Now let us take $x>0$. This real $x$ belongs to some interval $[f^{k}(1),f^{k+1}(1))$. Thus, $f(x)\in[f^{k+1}(1),f^{k+2}(1))$ and $g(f(x))=\gamma^{k+1}h(f^{-k-1}(f(x))=\gamma\gamma^{k}h(f^{-k}(x))=\gamma g(x)$. This ends the proof. ∎
Proof of Lemma 5
Lemma 5.
Let $\lambda\in(0,1)$. Let $f\in\mathcal{K}$ verifying for all $x>0$, $f(x)<x$. Then, there exists a function $g\in\mathcal{K}$ satisfying $g(f(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
Proof.
The function $f^{-1}$ is continuous, strictly increasing and definite and satisfies $f^{-1}(x)>x$ for all $x>0$. Then using Lemma 4, there exists $g\in\mathcal{K}$ such that $g(f^{-1}(x))=\lambda^{-1}g(x)$ for all $x\in\mathbb{R}$. Then for all $y\in\mathbb{R}_{+}$, there exists a unique $x\in\mathbb{R}_{+}$ such that $y=f^{-1}(x)$ or equivalently $x=f(y)$. Then, we have $g(y)=\lambda^{-1}g(f(y))$ and we conclude that $g(f(y))=\lambda g(y)$ for all $y\in\mathbb{R}_{+}$. ∎

### v4

@@
-Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis
+Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peak Analysis
 Abstract
-In this paper, we extend two classes of functions classically involved in asymptotic stability analyses for studying a maximization problem on the reachable values of a discrete-time dynamical system.
-This particular maximization problem is called a peaks computation problem.
-The problem is to find a couple composed of an initial state and a time which maximizes a given function over states.
-The paper focuses on the time component of the optimal solution which is an integer as the time is discrete.
-We develop a method to provide an upper bound of this integer from a formula which requires a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence.
-A first result proves that the formula provides, in theory, the optimal integer.
-However, in practice, the computation cannot be so precise.
-Then, we develop two alternative methods.
-The first is based on discontinuous and non strictly increasing/decreasing $\mathcal{KL}$-like functions named $\mathcal{KL}_{\rm gen}$ functions.
+In this paper, we extend two classes of functions that are classically involved in asymptotic stability analyses.The purpose of this extension is to study a maximization problem on the reachable values of a discrete-time dynamical system.
+This particular maximization problem is called a peak computation problem.
+The problem consists in finding a couple composed of an initial state and a time that maximizes a given function over states.
+The paper focuses on the time component of the optimal solution, which is an integer as the time is discrete.
+We apply a method developed in previous papers to compute an upper bound of the greatest index maximizer of a real sequence.
+This previous method is based on a formula that requires a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence that provides an upper bound of the analyzed sequence.
+This pair is proven to exist.
+However, in practice, the computation cannot be done in a general setting.
+In this paper, we developed two alternative methods.
+The first is based on discontinuous and non-strictly increasing/decreasing $\mathcal{KL}$-like functions called $\mathcal{KL}_{\rm gen}$ functions.
 We prove that the existence of a $\mathcal{KL}_{\rm gen}$ upper bound is equivalent to the existence of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence.
-The construction of the strictly increasing continuous function from a $\mathcal{KL}_{\rm gen}$ function needs an extension of the famous Sontag’s lemma.
-Finally, we construct a new type of Lyapunov functions well designed for our peaks computation problem.
+The construction of a strictly increasing continuous function from a $\mathcal{KL}_{\rm gen}$ function requires an extension of the famous Sontag’s lemma.
+Finally, we construct a new type of Lyapunov function that is well designed for our peak computation problem.
 These functions are called Opt-Lyapunov functions.
-Opt-Lyapunov functions are well designed since we establish an equivalence theorem between the existence of an Opt-Lyapunov function and of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence.
-The construction of a Opt-Lyapunov function from a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence is insipred by the Yoshizawa construction of Lyapunov functions.
-Keywords: Discrete-time, Peaks Computation Problem, Supremum of real sequences, Lyapunov functions, $\mathcal{KL}$ functions.
+Opt-Lyapunov functions are suitable as we establish an equivalence theorem between the existence of an Opt-Lyapunov function and that of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence.
+The construction of an Opt-Lyapunov function from a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence is inspired by the Yoshizawa construction of Lyapunov functions.
+Keywords: Discrete-time, Peak Computation Problems, Supremum of real sequences, Lyapunov functions, $\mathcal{KL}$ functions.
 1 Introduction
-In this paper, we study the problem of peaks computation for discrete-time dynamical systems in finite dimensional state spaces.
-A peaks computation problem can be viewed as a particular maximization problem for which the constraints set is the reachable values set of a discrete-time dynamical system.
-The objective function of this optimization problem is a function over the states.
-This particular maximization problem can model various situations in engineering, economics, physics, biology… For example, in population dynamics, for a predator-prey system, one may be interested in the maximal size of each species even if the system asymptotically oscillates between several equilibra.
-This problem boils down to maximize separetely each coordinate of the state variable over the system modeling the population dynamics.
-Besides its importance to analyze critical concrete situations, the literature about peaks analysis problem in the community of dynamical systems analysis is pretty thin.
-The peaks computation problem appears in a discrete-time setting for linear systems and where the objective function is the Euclidean norm [APS18].
-The problem also appears for the analysis of stable linear systems with general quadratic objective functions [Adj21, Adj23].
-A slighty different version of the problem is also encountered in [AG24].
-The difference comes from the fact that the state variable is, moreover, constrained to stay in a some given set.
-The problem is studied in a continuous-time setting and in, algebraically, a polynomial environment in [MHS20].
-Note that in [MHS20], the authors address the problem of verifying safety properties (see e.g.
-[BYG17] for verification problems on discrete-time systems) for dynamical systems and formulate the safety verification problem as a peaks computation problem.
-Safety verification problems aim to prove that the trajectories of the system are contained in some safe set.
-A safety verification problem can be translated into a peaks computation problem when the safe set is a sublevel of a sufficiently regular function.
-In this case, the objective function of the peaks computation problem is the function associated with the sublevel set.
-The safety analysis from the peaks computation point-of-view is also proposed for switched system [Adj17] and for program analysis in [AGM15].
-For safety analysis of dynamical systems, one can think that a reachability analysis (see e.g.
-[RKML06] and reference therein) equivalent to a feasibility analysis in our context is sufficient to prove some properties on the dynamical system.
-However, maximality in our context is a synonym for criticality.
-The optimal value in this situation can be viewed as the value which penalizes the most the system with respect to some criteria and defines the extreme condition where the system could enter into.
-In [MHS20] the authors are interested in computing an upper bound of the peaks optimal value for the continuous case.
-The techniques used is based on Sums-Of-Squares (SOS) and Linear Matrix Inequalities (LMI) formulations (the interested reader can consult [Las15] for those subjects).
-The same goal motivates the authors of [APS18] and, as they limit their study to linear system and norm objective functions, they directly use semidefinite programming (see [BTN01] for further details on the subject).
-The reference [AG24] cannot be fairly compared due to the presence of the additional constraint.
-Moreover, again, the studied systems are linear or piecewise linear and the objective functions are linear.
-The goal of the current paper is to provide the theoretical bases to solve exactly a peaks computation problem in a general way.
-The developments made in this paper can be viewed as a theoretical generalization of works done in [Adj21, Adj23]
-In this paper, we propose to study peaks computation problems from classical tools of dynamical systems stability theory (e.g.
-[Ela05, Chap.
-4] for discrete-time).
-This approach comes from a simple observation: the existence of peaks depends on the asymptotic behaviour of the system.
-The importance of stability analysis is time-tested and a stability analysis is a prerequisite to the implementation of numerical methods.
-Whereas general stability properties can be difficult to obtain, asymptotic stability properties can be easily presented and understood from $\mathcal{KL}$ bounds (e.g.
-[ATP10] or [NT04]) i.e.
-the norms of the states are globally bounded from above by a $\mathcal{KL}$ function.
-However, proving that a candidate $\mathcal{KL}$ function is an upper bound over the norms of state has to be done for all integers which can be difficult.
-On the other hand, Lyapunov functions are functional certificates for the system stability.
-Some results guarantee the existence of Lyapunov functions: a system is globally asymptotic stable if and only if there exists a Lyapunov function (see  e.g.
-[Kha02]).
-Being a Lyapunov function is a timeless property and only depends on dynamics.
-In few advantageous situations, Lyapunov functions can even be computed using numerical and optimization solvers (see the survey [GH15]).
-Formally, the data of a peaks computation problem is decomposed into an autonomous discrete-time dynamical system and an objective function $\varphi$ over states.
-The dynamical system consists of a set of initial conditions $X^{\mathrm{i}n}$ and the dynamics of the system i.e.
-a function $T$ which maps a state to a new state.
-The associated peaks computation problem aims in maximizing $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{i}n}$ and $k$ is a natural integer.
-The optimal solution of a peaks computation problem is then a couple $(x_{\rm opt},k_{\rm opt})$ where $x_{\rm opt}\in X^{\mathrm{in}}$ and $k_{\rm opt}$ is a natural integer.
+In this paper, we study the problem of peak computation for discrete-time dynamical systems in finite-dimensional state spaces.
+A peak computation problem can be viewed as a particular maximization problem for which the constraint set is the reachable value set of a discrete-time dynamical system.
+The objective function of this optimization problem is a function of the states.
+This particular maximization problem can be used to model various situations in engineering, economics, physics, and biology… For example, in population dynamics, for a predator-prey system, one may be interested in the maximal size of each species, even if the system asymptotically oscillates between several equilibria.
+This problem boils down to maximizing each coordinate of the state variable separately over the system modeling the population dynamics.
+Besides its importance in analyzing critical concrete situations, the literature on peak analysis problems in the community of dynamical systems analysis is quite limited.
+The peak computation problem appears in a discrete-time setting for linear systems, where the objective function is the Euclidean norm [APS18].
+This problem also appears in the analysis of stable linear systems with general quadratic objective functions [Adj21, Adj25c].
+A slightly different version of the problem was encountered in [AG24].
+The difference arises from the fact that the state variable is constrained to remain in a given set.
+The problem was studied in a continuous-time setting and algebraically in a polynomial environment in [MHS20].
+In  [MHS20], the authors address the problem of verifying safety properties (see e.g.,  [BYG17] for verification problems on discrete-time systems) for dynamical systems and formulate the safety verification problem as a peak computation problem.
+Safety verification problems aim to prove that the trajectories of a system are contained in a safe set.
+A safety verification problem can be translated into a peak computation problem when the safe set is a sublevel of sufficiently regular function.
+In this case, the objective function of the peak computation problem is associated with the sublevel set.
+A safety analysis from the peak computation point of view was also proposed for switched systems  [Adj17] and for program analysis in [AGM15].
+For the safety analysis of dynamical systems, one can think that a reachability analysis (see e.g.,  [RKML06] and references therein) equivalent to a feasibility analysis in our context is sufficient to prove some properties of the dynamical system.
+However, maximality in our context is synonymous with criticality.
+The optimal value in this situation can be viewed as the value which penalizes the most the system with respect to some criteria and defines the extreme condition in which the system can enter.
+In [MHS20] the authors are interested in computing an upper bound of the peak optimal value for the continuous case.
+The techniques used are based on sum-of-squares (SOS) and linear matrix inequality (LMI) formulations (the interested reader can consult [Las15] for those subjects).
+The same goal motivated the authors of [APS18].
+They limited their study to linear systems and norm objective functions and directly used semidefinite programming (see [BTN01] for further details on the subject).
+Reference  [AG24] cannot be fairly compared because of the additional constraint.
+Moreover, the studied systems are linear or piecewise linear, and the objective functions are linear.
+The goal of the current paper is to provide the theoretical bases to solve exactly a peak computation problem in a general way.
+The developments made in this paper can be viewed as a theoretical generalization of works done in [Adj21, Adj25c, Adj25a] and a specialization of [Adj25b] to specific sequences of optimal values arising in peak analysis computations.
+In this paper, we propose to study peak computation problems from extensions of classical tools of dynamical systems stability theory (for example,  [Ela05, Chap.
+4] for discrete time).
+This approach is based on a simple observation: the existence of peaks depends on the asymptotic behavior of the system.
+The importance of stability analysis is time-tested, and stability analysis is a prerequisite for the implementation of numerical methods.
+Although general stability properties can be difficult to obtain, asymptotic stability properties can be easily presented and understood from $\mathcal{KL}$ bounds (e.g.,  [ATP10] or [NT04]), that is, the norms of the states are globally bounded from above by a $\mathcal{KL}$ function.
+However, proving that a candidate $\mathcal{KL}$ function is an upper bound over the norms of the state must be done for all integers, which can be difficult.
+In contrast, Lyapunov functions are functional certificates for system stability.
+Some results guarantee the existence of Lyapunov functions: a system is globally asymptotically stable if and only if there exists a Lyapunov function (see, e.g., [Kha02]).
+Being a Lyapunov function is a timeless property that depends only on the dynamics.
+In a few advantageous situations, Lyapunov functions can be computed using numerical and optimization solvers (see the survey [GH15]).
+Formally, the data of a peak computation problem are decomposed into an autonomous discrete-time dynamical system and an objective function $\varphi$ over states.
+The dynamical system consists of a set of initial conditions $X^{\mathrm{in}}$ and the dynamics of the system, that is, a function $T$ that maps a state to a new state.
+The associated peak computation problem aims to maximize $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{in}}$ and $k$ is a natural integer.
+An optimal solution of a peak computation problem is then a couple $(x_{\rm opt},k_{\rm opt})$ where $x_{\rm opt}\in X^{\mathrm{in}}$ and $k_{\rm opt}$ is a natural integer.
 Note that when an integer $k$ is fixed, maximizing $\varphi(T^{k}(x))$ over $x\in X^{\mathrm{in}}$ is a standard maximization problem.
-The main difficulty of peaks computation problem is to compute an optimal $k_{\rm opt}$ i.e.
-an integer $n$ for which the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is the greatest.
-Our approach is to find an upper bound $K$ of the (greatest) integer maximizer.
-We impose to the upper bound $K$ to satisfy the following condition : for all $n$ greater than $K$, the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is smaller than the supremum of $\varphi(T^{K}(x))$.
-Once this integer $K$ computed, to solve the peaks computation problem can be done by solving a finite number of maximization problems.
-We can extract the optimal solution $(x_{\rm opt},k_{\rm opt})$ from one of those maximization problems.
-Thus we reduce the peaks computation problem to find the smallest possible integer $K$ satisfying the previous condition.
-We, then, provide a general formula to compute such integer.
-This formula relies on well-chosen pairs of a strictly increasing and continuous function on $[0,1]$, $h$, and a scalar $\beta\in(0,1)$.
-In this case, well-chosen means that the pair $(h,\beta)$ forms a sequence $(h(\beta^{k}))_{k}$ such that $(h(\beta^{k}))_{k}$ is greater than the sequence of the suprema of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$.
-A well-chosen pair $(h,\beta)$ is said to be useful for our problem if $h(0)$ is strictly smaller than a supremum of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$.
-We prove that there exists an optimal useful pair, in the sense of that the pair permits to find the greatest optimal integer.
+The main difficulty of the peak computation problem is to compute an optimal $k_{\rm opt}$, that is, an integer $n$ for which the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is the greatest.
+The optimal value of a peak computation problem can be interpreted as the sequence of the optimal value of standard maximization problems where the integer $k$ is fixed.
+From this interpretation, the integer element $k_{\rm opt}$ of an optimal solution becomes an index for which the associated term maximizes the sequence.
+Therefore, we can apply the method developed in [Adj25b].
+It consists in computing an upper bound of the (greatest) index maximizer of a real sequence that is bounded from above.
+The computation of this upper bound is based on a formula involving a pair $(h,\beta)$ where $h$ is a strictly increasing continuous function on $[0,1]$ and $\beta$ is a scalar belonging to $(0,1)$.
+The pair $(h,\beta)$ is also constrained to form a sequence $(h(\beta^{k}))_{k\in\mathbb{N}}$ greater than the sequence of the suprema of $\varphi\circ T^{k}$ over $X^{\mathrm{in}}$.
+To be applicable, the pair $(h,\beta)$ has to be useful, that is, $h(0)$ must be strictly smaller than, at least, a supremum of $\varphi\circ T^{k}$ over $X^{\mathrm{in}}$.
+In [Adj25b], we proved that such a pair $(h,\beta)$ exists and can be even chosen to be optimal, meaning that the pair permits finding the greatest optimal integer.
 In practice, this pair is not computable.
-Therefore, we propose two other methods to construct useful pairs.
-First, we extend the class of $\mathcal{KL}$ functions to the class of functions which are possibly discontinuous, unsigned (on the state variable but nonnegative on the time variable), increasing on states and decreasing on time.
+Therefore, we propose two additional methods for constructing useful pairs.
+First, we extend the class of $\mathcal{KL}$ functions to the class of functions that are possibly discontinuous, unsigned (on the state variable but nonnegative on the time variable), increasing on states, and decreasing on time.
 We call this class of functions $\mathcal{KL}_{\rm gen}$.
 As for classical $\mathcal{KL}$ functions, we use $\mathcal{KL}_{\rm gen}$ functions as upper bounds on $\varphi(T^{k}(x))$ (for the classical case with $\mathcal{K}\mathcal{L}$, $\varphi$ is a norm).
-We establish an equivalence theorem where we prove that a useful pair $(h,\beta)$ can be constructed from a $\mathcal{KL}_{\rm gen}$ upper bound function and we can construct $\mathcal{KL}_{\rm gen}$ upper bounds from useful pairs $(h,\beta)$.
+We establish an equivalence theorem, where we prove that a useful pair $(h,\beta)$ can be constructed from a $\mathcal{KL}_{\rm gen}$ upper bound function, and we can construct $\mathcal{KL}_{\rm gen}$ upper bounds from useful pairs $(h,\beta)$.
 The construction of useful pairs $(h,\beta)$ from $\mathcal{KL}_{\rm gen}$ upper bound functions relies on an extension of Sontag’s result [Son98, Prop.
 7] for $\mathcal{KL}_{\rm gen}$.
-Finally, we propose a new concept of Lyapunov functions adapted for the peaks computation problem namely the Opt-Lyapunov functions.
-The word adapted means the existence of an Opt-Lyapunov function is equivalent to the existence of a useful pair $(h,\beta)$.
-Once again, the proofs of each implication are constructive.
-The construction of an Opt-Lyapunov function from a useful pair $(h,\beta)$ is inspired by Yoshizawa construction of Lyapunov functions [Yos66].
-The main difference between classical Lyapunov functions and this new concept is that we allow Opt-Lyapunov functions to have several zeros and to take infinity values.
-Those two new possibities make irrelevant the use of classical comparison functions (see e.g.
-[Kel14] for a comprehensive survey).
-In our context, when we have access to classical Lyapunov functions, we can link continuous $\varphi$ with classical Lyapunov functions through comparison functions.
-Without this link, we need to develop a concept of weaker relations between $\varphi$ and an Opt-Lyapunov function.
-In this paper, we introduce the notion of certificate of compatibility to overcome the loss of the link.
-In summary, the main contributions of the paper are the following:
-  • we propose a novel approach of the peaks computation problem by bounding the number of static optimization problem to solve to obtain the optimal value;
-  • we obtain a general upper bound method for bounded from above sequence based on homeomorphic deformations of geometric sequences;
-  • we generalize the concept of $\mathcal{K}\mathcal{L}$ functions to $\mathcal{KL}_{\rm gen}$ adapted to peaks computation problems for which we can formulate peaks computation problems as a $\mathcal{KL}_{\rm gen}$ upper bounds existence ;
+Finally, we propose a new concept of Lyapunov functions adapted for the peaks computation problem, namely, the Opt-Lyapunov functions.
+The word adapted means that the existence of an Opt-Lyapunov function is equivalent to the existence of a useful pair $(h,\beta)$.
+Once again, the proofs of all implications are constructive.
+The construction of an Opt-Lyapunov function from a useful pair $(h,\beta)$ is inspired by the Yoshizawa construction of Lyapunov functions [Yos66].
+The main difference between classical Lyapunov functions and this new concept is that we allow Opt-Lyapunov functions to have several zeros and take infinite values.
+These two new possibilities make the use of classical comparison functions irrelevant (see e.g.,  [Kel14] for a comprehensive survey).
+In our context, when we have access to classical Lyapunov functions, we can link, when $\varphi$ is continuous, $\varphi$ with classical Lyapunov functions using comparison functions.
+Without this link, we must develop a concept of weaker relations between $\varphi$ and an Opt-Lyapunov function.
+In this paper, we introduce the notion of a certificate of compatibility to overcome the loss of links.
+In summary, this study makes two main contributions.
+  • We generalize the concept of $\mathcal{K}\mathcal{L}$ functions to $\mathcal{KL}_{\rm gen}$ adapted to peaks computation problems for which we can formulate peaks computation problems as a $\mathcal{KL}_{\rm gen}$ upper bounds existence.
   • We introduce a new concept of Lyapunov functions well designed for peaks computation problem for which the comparison functions approach is replaced by a compatibility notion.
 The paper is organized as follows.
-Section 2 presents the peaks computation problem addressed in this paper and the main difficulty related to the search of an integer as an optimal solution of the peaks computations problem.
-Section 3 recalls some basic results about the supremum of sequences and how it is related to their asymptotic behaviour.
-In Section 3, we also find the fundamental result about the existence of a useful pair of a strictly increasing and continuous function on $[0,1]$ and a scalar in $(0,1)$.
-This pair furnishes, in theory, the exact integer optimal solution.
-However, this pair is difficult to compute in practice.
-Then, in Section 4, after developing a simple result using the classical $\mathcal{K}\mathcal{L}$ class, we introduce the new class $\mathcal{KL}_{\rm gen}$.
-We prove that the existence of $\mathcal{KL}_{\rm gen}$ upper bounds of $\varphi(T^{k}(x))$ is equivalent to the existence of a useful pair.
-In Section 5, we present the concept of Opt-Lyapunov functions well designed for our maximization problem.
-The classical comparison functions associated to classical Lyapunov functions are replaced by a new notion called compatibility.
+Section 2 presents the peak computation problem addressed in this paper and the main difficulty related to the search for an integer as an optimal solution of the problem.
+In Section 3, we introduce an illustrative example to which we apply the novel techniques in the following sections.
+This example is the simplest version of the Leslie model for studying the dynamics of a population decomposed into age classes.
+Then, in Section 4, we introduce the new class $\mathcal{KL}_{\rm gen}$.
+We prove that the existence of $\mathcal{KL}_{\rm gen}$ upper bounds on $\varphi(T^{k}(x))$ is equivalent to the existence of a useful pair.
+In Section 5, we present the concept of Opt-Lyapunov functions well-designed for our maximization problem.
+The classical comparison functions associated with classical Lyapunov functions are replaced by a new notion called compatibility.
 We also prove that the existence of a compatible Opt-Lyapunov function is equivalent to the existence of a useful pair.
-In Section 6, we summarize the results obtained highlighting the concrete constructions between useful pairs of strictly increasing and continuous functions on $[0,1]$ and convergent geometric sequences, $\mathcal{KL}_{\rm gen}$ class and Opt-Lyapunov functions.
+In Section 6, we summarize the results obtained, highlighting the concrete constructions between useful pairs of strictly increasing and continuous functions on $[0,1]$, convergent geometric sequences, $\mathcal{KL}_{\rm gen}$ class, and Opt-Lyapunov functions.
 Notations: $\mathbb{R}$ stands for the set of reals, $\mathbb{R}_{+}$ the set of nonnegative reals, $\mathbb{R}^{*}$ the set of nonzero reals and $\mathbb{R}_{+}^{*}$ the set of strictly positive reals.
@@
 2 Problem Statement And Motivated Example
-2.1 Problem Statement
-Let $X^{\mathrm{in}}$ be a subset of $\mathbb{R}^{d}$ and $T$ a nonzero self-map on $\mathbb{R}^{d}$. Let us consider the associated discrete-time dynamical system defined as follows:
-$x_{0}\in X^{\mathrm{in}},\ x_{k+1}=T(x_{k})\enspace.$
-Let us consider $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$.
-We are interested in the reachable values that maximizes $\varphi$.
-More precisily, writing $T^{k}$ for the $k$-fold composition of $T$ or the identity when $k=0$, we are interested in solving the following maximization problem called a peaks computation problem:
-Problem 1 (Peaks Computation Problem).
-Maximize $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
-Without loss of generality, we can suppose that $\varphi(0)=0$; otherwise, we change $\varphi$ by $\varphi-\varphi(0)$ and then we add $\varphi(0)$ at the end of the optimization procedure.
-To solve Problem 1 means that we have to compute the optimal value $\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}$ and, if there exists, to characterize and to compute a couple of a vector $x\in X^{\mathrm{in}}$ and an integer $k$ which maximizes the value $\varphi(T^{k}(x))$.
-We can solve Problem 1 sequentially by solving a sufficient number of static maximization problems :
-Problem Pk (Static Problems).
-Maximize $\varphi(T^{k}(x))$ under the constraint $x\in X^{\mathrm{in}}$.
-Some hypotheses can be made to ensure that Problem Pk has a finite optimal values and an optimal solution.
-For example, if $X^{\mathrm{in}}$ is compact and the function $\varphi\circ T^{k}$ is upper semicontinuous, Problem Pk has an optimal solution.
-About computations, some special instances of Problem Pk can be solved numerically more or less efficiently.
-For example, when $\varphi\circ T^{k}$ is concave and $X^{\mathrm{in}}$ is convex, the problem is said to be convex and numerous methods exist (see e.g.
-[BTN01, Ber15]), or when $\varphi\circ T^{k}$ is polynomial and $X^{\mathrm{in}}$ a basic semialgebraic set (see e.g.
-[AL11, Las15]).
-The main issue remains to guess the interesting $k$, the integer for which the optimal value of Problem Pk is the greatest.
-A direct computation of such $k$ seems to be very difficult.
-Hence, the method we use in this paper is to solve instances of Problem Pk for $k=0,\ldots,K$ where $K$ has to determined.
-The integer $K$ has the particularity of ensuring that the optimal value of Problem 1 corresponds to the optimal value of some Problem Pk where $k$ is smaller than $K$.
-In this paper, we will adopt the following notations:
-$\nu:=(\nu_{k})_{k\in\mathbb{N}}\text{ where }\forall\,k\in\mathbb{N},\ \nu_{k}:=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\text{ and }\nu_{\rm opt}:=\sup\{\nu_{k}:k\in\mathbb{N}\}$ (1)
-Sometimes we will also use the terminlogy the supremum of $\nu$ to speak about $\nu_{\rm opt}$.
-The sequence $\nu$ corresponds to the sequence of optimal values of the maximization problems depicted at Problem Pk and $\nu_{\rm opt}$ is the optimal value of Problem 1.
-Following those notations, the main problem of the paper can be formulated mathematically as follows:
-Problem 2 (Safe Number of Problem $P_{k}$ to Solve).
-Find $K\in\mathbb{N}$ such that $\nu_{\rm opt}=\max\{\nu_{k}:k=0,\ldots,K\}$.
-In this paper, we provide simple characterizations for which a solution $K$ of Problem 2 exists.
-Then we construct a formula based on pairs composed of a strictly increasing continuous functions on $[0,1]$ and a scalar in $(0,1)$.
-Finally, we propose constructions of those strictly increasing continuous functions on [0,1] based on extensions of classical tools on stability theory for discrete-time dynamical systems such as $\mathcal{K}\mathcal{L}$ class of functions and Lyapunov functions.
-The developments made in the paper require real-valued sequences; hence the only assumption made in this paper is the following :
+2.1 The motivation : a discrete-time peak computation problem
+2.1.1 Definition of the problem
+Let us consider a discrete-time dynamical system $(X^{\mathrm{in}},T)$ on $\mathbb{R}^{d}$ where
+  • $X^{\mathrm{in}}$ is the nonempty subset of $\mathbb{R}^{d}$ of initial conditions;
+  • $T$ is a nonzero self-map on $\mathbb{R}^{d}$ that updates the state-variable.
+The system is autonomous and evolves without control, disturbances, or external inputs.
+In this paper, we are interested in the reachable values that maximize a given function.
+Then, given $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$, we define the associated discrete-time peak computation problem as follows:
+$\begin{array}[]{lcl}\displaystyle{\operatorname*{Max}_{x,k}}&&\displaystyle{\varphi(T^{k}(x))}\\ &\text{ s. t.}&\displaystyle{x\in X^{\mathrm{in}},k\in\mathbb{N}}\end{array}$ (DPCP)
+where $T^{k}$ is the $k$-fold composition of $T$ if $k$ is not null, and the identity if $k=0$.
+Like any optimization problem, to solve a problem of the form (DPCP), we must compute
+  • its optimal value, i.e., $\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}$;
+  • its optimal solutions, i.e., couples $(\overline{x},\overline{k})\in X^{\mathrm{in}}\times\mathbb{N}$ such that
+    $\varphi(T^{\overline{k}}(\overline{x}))=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}.$
+An optimal solution $(x,k)$ of Problem (DPCP) is thus composed of an initial condition $x$ leading to a peak and its associated date $k$ of peak-realization.
+Without loss of generality, we can suppose that $\varphi(0)=0$; otherwise, we change $\varphi$ for $\varphi-\varphi(0)$ and then we add $\varphi(0)$ at the end of the optimization procedure.
+This is formulated in Assumption 1.
+Assumption 1.
+$\varphi(0)=0$.
+2.1.2 A problem on integers
+The optimal value of Problem (DPCP) can be described using a sequence of optimal values. Indeed, we can introduce a static classical optimization problem in which integer $k$ is fixed as follows:
+$\begin{array}[]{lcl}\displaystyle{\operatorname*{Max}_{x}}&&\displaystyle{\varphi(T^{k}(x))}\\ &\text{ s. t.}&\displaystyle{x\in X^{\mathrm{in}}}\end{array}$ ($P_{k}$)
+Then, we define the sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ for all $k\in\mathbb{N}$ by:
+$\nu_{k}:=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}$ (1)
+and its supremum:
+$\nu_{\rm opt}:=\sup\{\nu_{k}:k\in\mathbb{N}\}.$ (2)
+Naturally, if the optimal value of Problem  ($P_{k}$) is infinite, which is the same as $\nu_{k}=+\infty$ for some $k\in\mathbb{N}$, then the optimal value of Problem (DPCP) is infinite.
+We assume that the following holds:
 Assumption 2.
 For all $k\in\mathbb{N},\ \nu_{k}=\sup\left\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\right\}$ is finite.
-A simple situation where Assumption 1 holds when $X^{\mathrm{in}}$ is a bounded set, $T$ maps from bounded sets to bounded sets and $\varphi$ is upper semicontinuous.
-2.2 An unstable linear system and an objective function with finite optimal value
-Throughout the paper, we illustrate our method and the computations on the example presented in this subsection.
-The peaks computation problem can be solved more simply and we compare the results obtained with our methodology with the one obtained analytically.
-Let us consider the following dynamics on $\mathbb{R}^{2}$:
-$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}.$
-We parameterize the initial conditions set by a scalar $\mu\in(0,1/3]$ and we define:
-$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}.$
-Finally, the objective function is parameterized by a positive scalar $p$:
-$\mathbb{R}^{2}\ni(x,y):\mapsto\varphi(x,y):=y^{2}-x^{2}+px$
-The associated peaks computation problem is the following:
-$\begin{array}[]{rc}\displaystyle{\operatorname*{Max}_{(x,y),n}}&\varphi(A^{n}(x,y)^{\intercal})\\ s.t.&\left\{\begin{array}[]{l}(x,y)\in X^{\mathrm{in}}\\ n\in\mathbb{N}\end{array}\right.\end{array}$
-The system is linear and the objective function is quadratic.
-However, we remark that $A$ admits $1/2$ and $3/2$ as eigenvalues.
-Therefore, $A$ is unstable and for some $x$, $\|A^{n}x\|$ goes to +$\infty$+ ∞ as $n$ goes to +$\infty$+ ∞.
-Therefore, we cannot use the method developed in [Adj23].
-Indeed, the system does not admit a quadratic Lyapunov function.
-We also note that $\varphi$ is not concave nor convex.
-As the matrix $A$ is diagonalizable, the powers of it are easy to compute and for all $n\in\mathbb{N}$ we have :
-$A^{n}=\dfrac{1}{2^{n+1}}\begin{pmatrix}3^{n}+1&2(3^{n}-1)\\ \frac{3^{n}-1}{2}&3^{n}+1\end{pmatrix}\text{ and }\forall\,y\in\mathbb{R},\ A^{n}\begin{pmatrix}2y\\ y\end{pmatrix}=\left(\frac{3}{2}\right)^{n}\begin{pmatrix}2y\\ y\end{pmatrix}$
-Moreover, the initial conditions set is actually a segment each element of it can be written as $(2y,y)^{\intercal}$ with $y\in[\mu,1/2]$.
-Using this, we have for all $y\in[\mu,1/2]$, for all $n\in\mathbb{N}$:
-$\varphi_{n}(y):=\varphi(A^{n}(2y,y)^{\intercal})=-3\left(\dfrac{3}{2}\right)^{2n}y^{2}+2p\left(\dfrac{3}{2}\right)^{n}y$
-Let $n\in\mathbb{N}$.
-The function $y\mapsto\varphi_{n}(y)$ is concave and infinitely derivable.
-It admits an unique maximizer on $\mathbb{R}$ which is $y_{n}^{*}:=\left(\frac{2}{3}\right)^{n}\frac{p}{3}$.
-The optimal value of $\varphi_{n}$ on $\mathbb{R}$ is $p^{2}/3$.
-The function $\varphi_{n}$ is strictly increasing before $y_{n}^{*}$ and strictly decreasing after.
-As we maximize on $[\mu,1/2]$, we have to check whether $y_{n}^{*}$ belongs to this interval.
-To avoid the case where $\frac{1}{2}$ is not used as a maximizer we assume that $p>3/2$.
-In this case, $y_{0}^{*}>1/2$.
-As the sequence $(y_{n}^{*})_{n\in\mathbb{N}}$ strictly decreases to zero, we can define two integers:
-$\underline{n}:=\lfloor\ln(3/2p)/\ln(2/3)\rfloor+1\text{ and }\overline{n}:=\lfloor\ln(3\mu/p)/\ln(2/3)\rfloor+1$
-The integer $\underline{n}$ corresponds to the smallest integer such that $y_{n}^{*}<1/2$ whereas $\overline{n}$ is the smallest integer such that $y_{n}^{*}<\mu$.
-Moreover, as we assume that $\mu\in(0,1/3]$, we have $(\ln(3\mu/p)-\ln(3/2p))/\ln(2/3)=-(\ln(\mu)+\ln(2))\ln(3/2)\geq-(\ln(1/3)+\ln(2))/\ln(3/2)=1$.
-We conclude that $\overline{n}>\underline{n}$, as for all $x,y\in\mathbb{R}_{+}$ such that $x-y\geq 1$, we have $\lfloor x\rfloor\geq\lfloor y\rfloor+1$.
-Finally using $\underline{n}$ and $\overline{n}$ , we can define the sequence of interest $(\nu_{n})_{n\in\mathbb{N}}$ and for all $n\in\mathbb{N}$:
-$\nu_{n}=\left\{\begin{array}[]{lr}\varphi_{n}(1/2)&\text{ if }n\in[0,\underline{n}-1]\cap\mathbb{N}\\ \dfrac{p^{2}}{3}&\text{ if }n\in[\underline{n},\overline{n}-1]\cap\mathbb{N}\\ \varphi_{n}(\mu)&\text{ if }n\geq\overline{n}\end{array}\right.$ (2)
-From now on, when we refer to this example, we will assume that $\mu\in(0,1/3]$ and $p\geq 3/2$.
-3 Suprema and Rank Maximizer of General Sequences
-3.1 Notations and Basic Results
-We introduce the set of real sequences which are bounded from above:
-$\Lambda:=\{u=(u_{0},u_{1},\ldots)\in\mathbb{R}^{\mathbb{N}}:\sup_{n\in\mathbb{N}}u_{n}<+\infty\}.$
-For an element of $\Lambda$, we are interested in computing the supremum of its terms.
+A simple situation where Assumption 2 holds is when $X^{\mathrm{in}}$ is a bounded set, $T$ maps bounded sets to bounded sets, and $\varphi$ is upper semicontinuous.
+The existence of an optimal solution for Problem ($P_{k}$) is actually not mandatory to ensure Assumption 2.
+From these notations, $\nu_{\rm opt}$ is equal to the optimal value of Problem (DPCP), whereas $\nu_{k}$ is the optimal value of static optimization Problem  ($P_{k}$).
+Solving Problem (DPCP) involves computing $\nu_{\rm opt}$, an integer $\overline{k}$ such that $\nu_{\overline{k}}=\nu_{\rm opt}$ and an optimal solution $x\in X^{\mathrm{in}}$ of Problem ($P_{k}$).
+In this paper, we focus on the characterization and computation of the integer $\overline{k}$.
+Static problems of the form ($P_{k}$) are standard problems in optimization theory.
+The computation of an optimal solution for these problems relies on an optimization solver when a suitable solver is available.
+In theory, some special instances of Problem ($P_{k}$) can be solved numerically, more or less efficiently.
+For example, when $\varphi\circ T^{k}$ is concave and $X^{\mathrm{in}}$ is convex, numerous methods exist (see e.g.,  [BTN01, Ber15]), or when $\varphi\circ T^{k}$ is polynomial and $X^{\mathrm{in}}$ is a basic semialgebraic set (see e.g.,  [AL11, Las15]).
+In practice, the difficulty of resolving Problem ($P_{k}$) precedes the optimization procedure and remains in the writing of the optimization problem itself.
+Indeed, obtaining a closed form for $T^{k}$ or at least for $T^{k}(X^{\mathrm{in}})$ can be an ambitious task.
+For example, for logistic sequences, closed forms are only known for particular cases (see, e.g.,[GRR10]).
+Some specific difference equations or systems require intensive efforts to obtain closed forms for $T^{k}$ (see, e.g., [Els12, HR18, Ste14] and references therein).
+Similarly, obtaining an exact description of $T^{k}(X^{\mathrm{in}})$ when $X^{\mathrm{in}}$ is an infinite set incurs a high computational cost.
+For example, in [MHL15], the authors approximate the polynomial image of a basic semialgebraic set obtained from the $L^{1}$ limit of a sequence of semidefinite programs.
+In some specific cases, $T^{k}$ can be described explicitly.
+When $T$ is linear or conjugate to a linear map (i.e., $T=B^{-1}\circ A\circ B$ for a bijection $B$ and a linear map $A$), we can compute $T^{k}$ using an eigendecomposition if $T$ is diagonalizable (or in Jordan form).
+When $T$ is linear, $T^{k}(X^{\mathrm{in}})$ is always a polytope if $X^{\mathrm{in}}$ is so.
+Hence, when $T$ is linear and $X^{\mathrm{in}}$ is a polytope, Problem ($P_{k}$) becomes a linear program.
+While the theory developed in this paper is independent of the practical constructions of $T^{k}$ or $T^{k}(X^{\mathrm{in}})$, the computational illustration proposed in this paper is limited to an unstable linear system combined with a nonlinear objective function $\varphi$.
+A more general practical approach will be considered in future works.
+The approach developed in [Adj25b] consists in computing an upper bound of the greatest integer maximizer of Problem (DPCP).
+In [Adj25a], we employed the notion stopping integer, which is an integer $K$ such that
+$\nu_{\rm opt}=\max\{\nu_{k}:k=0,\ldots,K\}\text{ and }\sup_{k>K}\nu_{k}<\nu_{\rm opt}.$ (3)
+We define the set of stopping integers and the smallest stopping integer as follows:
+$\Delta^{s}_{\nu}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}\nu_{j}>\sup_{j>k}\nu_{j}\right\}\text{ and }\mathbf{K}^{s}_{u}:=\inf\Delta^{s}_{\nu}.$
+The set of stopping integers can be empty, and we can characterize its nonemptiness from the analysis of the limit superior of $\nu$ and thus introduce the set:
+${\mathrm{R}}_{\nu}^{s}:=\left\{k\in\mathbb{N}:\nu_{k}>\limsup_{n\to+\infty}\nu_{n}\right\}.$
+Moreover, we define the set of integer maximizers of Problem (DPCP) as
+$\operatorname{Argmax}(\nu):=\{k\in\mathbb{N}:\nu_{k}=\nu_{\rm opt}\}.$
+Recall that $\nu_{\rm opt}$ is finite if and only if the limit superior of $\nu$ is finite or equal to -$\infty$.
+This basic analysis can be completed using the sets $\Delta^{s}_{\nu}$ and ${\mathrm{R}}_{\nu}^{s}$.
 Proposition 1.
-$u\in\Lambda\iff\displaystyle{\limsup_{n\to+\infty}u_{n}\in\mathbb{R}\cup\{-\infty\}}$.
-For $u\in\Lambda$, we are also looking for the set of maximizers i.e.
-$\operatorname{Argmax}(u):=\{k\in\mathbb{N}:u_{k}=\displaystyle{\sup_{n\in\mathbb{N}}u_{n}}\}$.
-For computations purpose, we need to characterize $\operatorname{Argmax}(u)$.
-Consequently, we introduce, for $u\in\Lambda$, the two sets of ranks:
-$\Delta_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}\right\}\ \text{ and }\ \Delta^{s}_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}>\sup_{j>k}u_{j}\right\}$
-It is easy to see that $\Delta_{u}\subseteq\Delta^{s}_{u}$.
-Moreover, if $k$ belongs to $\Delta_{u}$ (resp.
-$\Delta^{s}_{u}$), then any integer greater than $k$ belongs to $\Delta_{u}$ (resp.
-$\Delta^{s}_{u}$).
-Besides, we will need the set of ranks for which the associated term is greater than the limsup (resp.
-strictly greater):
-${\mathrm{R}}_{u}:=\left\{k\in\mathbb{N}:u_{k}\geq\limsup_{n\to+\infty}u_{n}\right\}\ \text{ and }\ {\mathrm{R}}_{u}^{s}:=\left\{k\in\mathbb{N}:u_{k}>\limsup_{n\to+\infty}u_{n}\right\}$
-It is obvious that ${\mathrm{R}}_{u}^{s}\subseteq{\mathrm{R}}_{u}$.
-Note that even if $u\in\Lambda$, ${\mathrm{R}}_{u}$ and hence ${\mathrm{R}}_{u}^{s}$ can be empty.
-To construct the characterizations, we need to consider the smallest elements of $\Delta_{u}$, $\Delta^{s}_{u}$, ${\mathrm{R}}_{u}$ and ${\mathrm{R}}_{u}^{s}$.
-$\mathbf{K}_{u}=\inf\Delta_{u};\ \qquad\mathbf{K}^{s}_{u}=\inf\Delta^{s}_{u};\ \qquad k_{u}=\inf{\mathrm{R}}_{u}\quad\text{ and }\ k^{s}_{u}=\inf{\mathrm{R}}_{u}^{s}$
-We recall that the infimum of the empty set is equal to +$\infty$+ ∞.
-Hence, $\mathbf{K}_{u}<+\infty$ (resp.
-$\mathbf{K}^{s}_{u}<+\infty$) if and only if $\Delta_{u}\neq\emptyset$ (resp.
-$\Delta^{s}_{u}\neq\emptyset$); and $k_{u}<+\infty$ (resp.
-$k^{s}_{u}<+\infty$) if and only if ${\mathrm{R}}_{u}\neq\emptyset$ (resp.
-${\mathrm{R}}_{u}^{s}\neq\emptyset$).
-Furthermore, if the sets of integers defined above are nonempty, they contain their infimum i.e.
-$\mathbf{K}_{u}\in\Delta_{u}$, $\mathbf{K}^{s}_{u}\in\Delta^{s}_{u}$, $k_{u}\in{\mathrm{R}}_{u}$ and $k^{s}_{u}\in{\mathrm{R}}_{u}^{s}$ when those integers are finite.
-Proposition 2.
-If $u\notin\Lambda$ then $\Delta_{u}=\emptyset$ and ${\mathrm{R}}_{u}=\emptyset$. Obviously, it also implies that $\Delta^{s}_{u}=\emptyset$ and ${\mathrm{R}}_{u}^{s}=\emptyset$.
-Proposition 3.
-Let $u\in\Lambda$. The following assertions hold:
+If $\nu_{\rm opt}<+\infty$. The following assertions are made:
   1. 1.
-    $\Delta^{s}_{u}\neq\emptyset\iff{\mathrm{R}}_{u}^{s}\neq\emptyset$;
+    $\Delta^{s}_{\nu}\neq\emptyset\iff{\mathrm{R}}_{\nu}^{s}\neq\emptyset$;
   2. 2.
-    $\Delta_{u}\neq\emptyset\iff{\mathrm{R}}_{u}\neq\emptyset$;
+    $\Delta^{s}_{\nu}=\emptyset\iff\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$;
   3. 3.
-    $\Delta^{s}_{u}=\emptyset\iff\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$.
-Proposition 4.
-The following inequalities hold:
-$k_{u}\leq k^{s}_{u};\qquad k_{u}\leq\mathbf{K}_{u};\qquad k^{s}_{u}\leq\mathbf{K}^{s}_{u}\quad\text{ and }\mathbf{K}_{u}\leq\mathbf{K}^{s}_{u}\enspace.$
-Moreover, if $k^{s}_{u}<+\infty$, then $k^{s}_{u}\leq\mathbf{K}_{u}$.
-Proposition 5.
-Let $u\in\Lambda$. The following assertions hold:
+    If $\Delta^{s}_{\nu}\neq\emptyset$ then $\mathbf{K}^{s}_{\nu}=\max\operatorname{Argmax}(\nu)$.
+One way to compute an element of ${\mathrm{R}}_{\nu}^{s}$ is, given a term $\nu_{k}$, to look for an integer $K\in\mathbb{N}$ such that
+$\nu_{j}<\nu_{k},\text{ for all }j\geq K.$
+Such integers $K$ cannot be smaller than $\mathbf{K}^{s}_{\nu}$ and are thus stopping integers.
+The problem is to convert the term $\nu_{k}$ into a stopping integer.
+The most natural sequences allowing this conversion (which is a bijection) between terms and indices are the geometric sequences.
+However, assuming that $\nu$ is a geometric sequence is extremely restrictive.
+Instead of directly using geometric sequences, we use bijective deformations of positive convergent geometric sequences as the upper bounds of $\nu$.
+Let $\mathbb{I}\left(\mathbb{R}\right)$ be the set of all intervals on $\mathbb{R}$ with nonempty interiors.
+We define, for a function $f$ on reals, the set of intervals where $f$ is strictly increasing and continuous that is,
+[…]
+If $I\in\mathrm{SIC}(f)$, $f$ is bijective from $I$ to $f(I)$ and we denote by ${{f}_{\left(I\right)}^{-1}}$ the inverse of $f$ on $I$.
+Note that ${{f}_{\left(I\right)}^{-1}}$ is also strictly increasing and continuous from $f(I)$ to $I$.
+Finally, we pay particular attention to functions that are strictly increasing and continuous on $[0,1]$ and we write
+$\Omega([0,1]):=\{h:\mathbb{R}\mapsto\mathbb{R}:[0,1]\in\mathrm{SIC}(h)\}.$
+A convergent geometric sequence is required to complete the upper bound on $\nu$ and we introduce
+$\mathbb{\Gamma}(\nu):=\{(h,\beta)\in\Omega([0,1])\times(0,1):\nu_{k}\leq h(\beta^{k}),\ \forall\,k\in\mathbb{N}\}$
+and
+$\mathbb{\Gamma}_{\rm fun}(\nu):=\{h\in\Omega([0,1]):\exists\,\beta\in(0,1)\text{ s.t. }(h,\beta)\in\mathbb{\Gamma}(\nu)\}.$
+Finally, we introduce for $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$:
+$\mathbb{\Gamma}_{\rm sc}(\nu,h):=\{\beta\in(0,1):(h,\beta)\in\mathbb{\Gamma}(\nu)\}.$
+From $(h,\beta)\in\mathbb{\Gamma}(\nu)$, we can define
+$\mathfrak{F}_{\nu}(k,h,\beta):=\dfrac{\ln\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\nu_{k})}{\ln(\beta)}\text{ if }\nu_{k}>h(0)$ (4)
+for which we have
+$j\geq\lfloor\mathfrak{F}_{\nu}(k,h,\beta)\rfloor\implies\nu_{j}<\nu_{k}$
+The point requiring special consideration is the strict positivity of ${{h}_{\left([0,1]\right)}^{-1}}(\nu_{k})$ or, equivalently, the fact that $\nu_{k}>h(0)$.
+This depends on the choice of $h$ and we define the notion of usefulness.
+Definition 1 (Useful Strictly Increasing Continuous Functions).
+Suppose that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$. A function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ is said to be useful for $\nu$ if the set
+$\mathcal{S}(\nu,h):=\{k\in\mathbb{N}:\nu_{k}>h(0)\}$ (5)
+is nonempty.
+By extension, a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$ if $h$ is useful for $\nu$.
+Remark 1.
+Note that as $h\in\mathbb{\Gamma}(\nu)$ is continuous, we have $\limsup_{k\to+\infty}\nu_{k}\leq h(0)$ and thus $\mathcal{S}(\nu,h)\subset{\mathrm{R}}_{\nu}^{s}$.
+One of the main results of [Adj25b] is the existence of a useful element in $\mathbb{\Gamma}(\nu)$ when ${\mathrm{R}}_{\nu}^{s}$ is nonempty.
+Theorem 1.
+The following statements are true.
   1. 1.
-    $\operatorname{Argmax}(u)\subseteq\Delta_{u}$;
+    $\nu_{\rm opt}<+\infty$ if and only if $\mathbb{\Gamma}(\nu)\neq\emptyset$;
   2. 2.
-    $\operatorname{Argmax}(u)\neq\emptyset\iff\Delta_{u}\neq\emptyset$;
-  3. 3.
-    $\mathbf{K}_{u}=\inf\operatorname{Argmax}(u)$;
-  4. 4.
-    $\operatorname{Argmax}(u)=\emptyset$ or $\operatorname{Argmax}(u)$ nonempty and unbounded $\implies\sup_{n\in\mathbb{N}}u_{n}=\limsup_{k\to+\infty}u_{k}$;
-  5. 5.
-    If $\Delta^{s}_{u}\neq\emptyset$ then $\mathbf{K}^{s}_{u}=\max\operatorname{Argmax}(u)$.
-Note that $\operatorname{Argmax}(u)$ can be nonempty and bounded when $\Delta^{s}_{u}$ is empty.
-In this case, by the third statement of Prop 3, we have $\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$.
-However, the sequence finitely touches $\limsup_{n\to+\infty}u_{n}$ before approaching, up to a subsequence, $\limsup_{n\to+\infty}u_{n}$ from below (strictly).
-Example 1.
-Let us consider the peaks computation problem depicted at Subsection 2.2.
-The sequence $(\nu_{k})_{k\in\mathbb{N}}$ presented at Eq.
-(2) has a greatest maximizer as $\limsup_{k\to+\infty}\nu_{k}=\lim_{k\to+\infty}\varphi_{k}(\mu)=-\infty$.
-The set $\operatorname{Argmax}(u)$ would not be bounded from above if $\mu=0$ and thus $\overline{n}=+\infty$.
-Indeed, in this case, we would get $\limsup_{k\to+\infty}\nu_{k}=p^{2}/3=\sup_{k\in\mathbb{N}}\nu_{k}$.
-∎
-Proposition 6.
-Let $u\in\Lambda$. The following assertions hold:
+    ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if there exists $(h,\beta)\in\mathbb{\Gamma}_{\rm fun}(\nu)$ that is useful for $u$.
+2.1.3 Problem statement
+Even if we have proved that $\mathbb{\Gamma}(\nu)$ is nonempty, the proof furnished in [Adj25b] exhibits an affine function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ for which the slope and y-intercept depend on $\mathbf{K}^{s}_{\nu}$ and $\nu_{\rm opt}$.
+This leads to considering alternative approaches to construct an element in $\mathbb{\Gamma}(\nu)$.
+Problem 1.
+Construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$.
+In [Adj25a], we proposed to build $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a $\mathcal{K}\mathcal{L}$ stability certificate or a Lyapunov function.
+Problems of the form (DPCP) can have a finite optimal value and are such that ${\mathrm{R}}_{\nu}^{s}$ is nonempty without admitting a classical $\mathcal{K}\mathcal{L}$ stability certificate or a classical Lyapunov function.
+This is the case for the example developed in Section 3.
+To go further Problem 1, we are interested in a full characterization of the elements of $\mathbb{\Gamma}(\nu)$ that subsumes the previous results based on classical stability tools.
+Thus, the main contribution of this paper is the characterization of the elements of $\mathbb{\Gamma}(\nu)$ from an extension of $\mathcal{K}\mathcal{L}$ and the Lyapunov functions.
+We will prove that solving Problem 1 is equivalent to constructing a weaker version of $\mathcal{K}\mathcal{L}$ certificate and a weaker version of the Lyapunov functions.
+We will prove that we can construct an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from any extended $\mathcal{K}\mathcal{L}$ certificate and that an extended $\mathcal{K}\mathcal{L}$ certificate can be constructed from any $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
+The same results hold for our extended class of Lyapunov functions.
+2.2 Classical $\mathcal{K}\mathcal{L}$ and Lyapunov approaches
+The use of $\mathcal{K}\mathcal{L}$ and Lyapunov functions outside the scope of stability theory seems to be atypical and we bring some intuitions about the use in the context of solving Problem 1.
+In fact, classical $\mathcal{K}\mathcal{L}$ and Lyapunov functions furnish more or less naturally a function in $\Omega([0,1])$ and $\beta\in(0,1)$.
+To obtain an element of $\mathbb{\Gamma}(\nu)$ from a $\mathcal{K}\mathcal{L}$ certificate or a Lyapunov function requires a restrictive inequality or the continuity of $\varphi$.
+From the positivity of $\mathcal{K}\mathcal{L}$ and Lyapunov functions, usefulness holds if and only if the sequence $\nu$ has a positive term.
+All the details on the approach can be found in [Adj25a].
+2.2.1 The classical $\mathcal{K}\mathcal{L}$ approach
+First, recall that a $\mathcal{K}\mathcal{L}$ function $\gamma:\mathbb{R}_{+}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ is such that
   1. 1.
-    Assume that ${\mathrm{R}}_{u}\neq\emptyset$. Let $k\geq k_{u}$ such that $\sup_{k_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
+    For all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)\in\mathcal{K}_{\infty}$ i.e.
+the set of strictly increasing continuous functions $\mathbb{R}_{+}\to\mathbb{R}_{+}$ that are zero at zero and whose the limit at $\infty$ is $\infty$;
   2. 2.
-    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $k\geq k^{s}_{u}$ such that $\sup_{k^{s}_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
-  3. 3.
-    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. For all $\mathbf{K}_{u}\leq k$,
-    $\sup_{0\leq j\leq k}u_{j}=\sup_{k^{s}_{u}\leq j\leq k}u_{j}=\sup_{k_{u}\leq j\leq k}u_{j}=\sup_{j\in\mathbb{N}}u_{j}\enspace.$
-In summary, for $u\in\Lambda$, to compute $\sup_{k\in\mathbb{N}}u_{k}$, we need to study first the emptiness of ${\mathrm{R}}_{u}^{s}$.
-Indeed, if ${\mathrm{R}}_{u}^{s}$ is empty, we know that (the third statement of Prop.
-3) $\sup_{n\in\mathbb{N}}u_{n}$ is equal to $\limsup_{n}u_{n}$.
-If ${\mathrm{R}}_{u}^{s}$ is not empty, then we have to compute $\mathbf{K}_{u}$ to find $\sup_{k\in\mathbb{N}}u_{k}$ and a maximizer (the second statement of Prop.
-5).
-However, to identify $\mathbf{K}_{u}$, we need knowledge on the past and the future of the sequence.
-The good point is that any over-approximation $k$ of $\mathbf{K}_{u}$ permits to know $\sup_{k\in\mathbb{N}}u_{k}$ by computing $\sup_{0\leq j\leq k}u_{j}$ (the third statement of Prop 6).
-3.2 A Formula Based on Pairs of Strictly Increasing Continuous Functions and Convergent Geometric Sequences
-Let us write $\mathbb{I}\left(\mathbb{R}\right)$ the set of all intervals (convex sets) on $\mathbb{R}$ (open, closed, or semi-open).
-For a function $f$ from $\mathbb{R}$ to $\mathbb{R}$, we define the set of intervals:
-$\mathrm{SIC}(f)=\{I\in\mathbb{I}\left(\mathbb{R}\right):f:I\mapsto f(I)\text{ is strictly increasing and continuous }\}.$
-Recall that a strictly increasing function $f$ is continuous on an interval $[a,b]$ if and only if $f$ is surjective from $[a;b]$ to $[f(a);f(b)]$.
-Hence, if $I$ belongs to $\mathrm{SIC}(f)$, then $f$ is bijective from $I$ to $f(I)$ and we can define its inverse on $I$ i.e.
-a function $g:f(I)\mapsto I$ such that $f\circ g=g\circ f=\mathrm{Id}$ where $\mathrm{Id}:x\mapsto x$ (viewed as a function from $I$ to itself).
-For $I\in\mathrm{SIC}(f)$, we denote by ${{f}_{\left(I\right)}^{-1}}$ the inverse of $f$ on $I$.
-Note that ${{f}_{\left(I\right)}^{-1}}$ is also strictly increasing and continuous (surjective) from $f(I)$ to $I$.
-Finally, we pay a particular attention to functions strictly increasing and continuous on $[0,1]$ and we write :
-$\Omega([0,1]):=\{h:\mathbb{R}\mapsto\mathbb{R}:[0,1]\in\mathrm{SIC}(h)\}.$
-The functions of $\Omega([0,1])$ are part of the upper bound on the sequence of interest $u$.
-To complete the upper bound, we need a convergent geometric sequence.
-We introduce, for $u\in\mathbb{R}^{\mathbb{N}}$ the following sets:
-$\mathbb{\Gamma}(u):=\{(h,\beta)\in\Omega([0,1])\times(0,1):u_{k}\leq h(\beta^{k}),\ \forall\,k\in\mathbb{N}\}$
+    For all $s\in\mathbb{R}_{+}$, $t\mapsto\gamma(s,t)\in\mathcal{L}$ i.e. the set of decreasing continuous functions $\mathbb{R}_{+}\to\mathbb{R}_{+}$ whose the limit at $\infty$ is 0;
+The dynamical system $(X^{\mathrm{in}},T)$ admits a $\mathcal{K}\mathcal{L}$ certificate if and only if there exists $\gamma\in\mathcal{K}\mathcal{L}$ such that for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$
+$\|T^{k}(x)\|\leq\gamma(\|x\|,k)$ (6)
+If the inequality (6) holds, the famous result of Sontag [Son98, Proposition 7] directly provides a potential pair $(h,\beta)$.
+Recall that the result asserts that for any $\gamma\in\mathcal{K}\mathcal{L}$ there exists $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ verifying $\gamma(s,t)\leq\theta_{1}(\theta(s)e^{-t})$ for all $(s,t)\in\mathbb{R}_{+}\times\mathbb{R}_{+}$.
+Then, if $X^{\mathrm{in}}$ is bounded, we define
+$h:s\to\theta_{1}(\theta_{2}(s\sup_{x\in X^{\mathrm{in}}}\|x\|))\in\Omega([0,1]).$ (7)
+The difficulty is to exploit the inequality (6) to obtain $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$.
+This requires an additional assumption on $\varphi$.
+For example, if $\varphi$ is continuous, we obtain the desired upper bound from the famous Hahn/Khalil result [Hah67] and [Kha02, Lemma 4.3/Appendix C.4].
+This result asserts that for any positive definite222a function is positive definite if it is strictly positive except at 0 where it is null continuous coercive333a function $f$ is said to be coercive (or radially unbounded) if $\lim_{\|x\|\to+\infty}f(x)=+\infty$ function $f$ there exist $\overline{\alpha_{1}},\overline{\alpha_{2}}\in\mathcal{K}_{\infty}$ satisfying:
+$\overline{\alpha_{1}}\circ\|\cdot\|\leq f\leq\overline{\alpha_{2}}\circ\|\cdot\|.$ (8)
+From Assumption 1 $\max\{\varphi,\|\cdot\|\}$ is positive definite, coercive, and continuous if $\varphi$ is. Hence
+$(\alpha_{\varphi}\circ h,e^{-1})\in\mathbb{\Gamma}(\nu)\text{ where }\alpha_{\varphi}\in\mathcal{K}_{\infty}\text{ is such that }\varphi\leq\max\{\varphi,\|\cdot\|\}\leq\alpha_{\varphi}\circ\|\cdot\|.$ (9)
+We can relax the continuity assumption on $\varphi$ and instead imposing that, for some $\psi:\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\sup_{x\in X^{\mathrm{in}}}\psi(x)<+\infty$
+$\varphi(T^{k}(x))\leq\gamma(\psi(x),k),\text{ for all }x\in X^{\mathrm{in}}\text{ and }k\in\mathbb{N}.$ (10)
+This consists in replacing in the inequality (6) the norm on the left-hand side by $\varphi$ and the norm on the right-hand side by a nonnegative function $\psi$ with a strictly positive supremum on $X^{\mathrm{in}}$.
+In this case
+$(s\mapsto\theta_{1}(\theta_{2}(s\sup_{x\in X^{\mathrm{in}}}\psi(x))),e^{-1})\in\mathbb{\Gamma}(\nu).$ (11)
+Finally, as stated above, usefulness in the context of $\mathcal{K}\mathcal{L}$ certificate is equivalent to the existence of a strictly positive term for $\nu$.
+2.2.2 The Lyapunov approach
+In [Adj25c], we consider Lyapunov functions to solve Problem (DPCP).
+In [Adj25c], the dynamical system was a stable affine system, and the objective function $\varphi$ was a quadratic function.
+Recently, in [Adj25a], we generalized the approach for all discrete-time systems that admit a Lyapunov function.
+Recall that a function $V:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is a (global) Lyapunov function if and only if
+$\alpha_{1}\circ\|\cdot\|\leq V\leq\alpha_{2}\circ\|\cdot\|\text{ for some }\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ (12)
 and
-$\mathbb{\Gamma}_{\rm fun}(u):=\{h\in\Omega([0,1]):\exists\,\beta\in(0,1)\text{s.t. }(h,\beta)\in\mathbb{\Gamma}(u)\}\enspace.$
-Finally, we introduce for $u\in\mathbb{R}^{\mathbb{N}}$ and $h\in\mathbb{\Gamma}_{\rm fun}(u)$:
-$\mathbb{\Gamma}_{\rm sc}(u,h):=\{\beta\in(0,1):(h,\beta)\in\mathbb{\Gamma}(u)\}\enspace.$
+$V\circ T\leq\lambda V\text{ and for some }\lambda\in(0,1)$ (13)
+Note that (12) implies that $V$ is positive definite.
+Inequality (13) implies that
+$V(T^{k}(x))\leq\lambda^{k}V(x)\text{ for all }x\in X^{\mathrm{in}}.$ (14)
+If $X^{\mathrm{in}}$ is bounded and not reduced to $\{0\}$, then $\sup_{x\in X^{\mathrm{in}}}V(x)$ is finite and strictly positive. Hence, similarly to the $\mathcal{K}\mathcal{L}$ approach, we define
+$h:s\mapsto s\sup_{x\in X^{\mathrm{in}}}V(x)\in\Omega([0,1]).$
+To use $h$ as an element of $\mathbb{\Gamma}_{\rm fun}(\nu)$, we need to make additional assumptions. We can simply impose that $\varphi\leq V$ and from (14)
+$(h,\lambda)\in\mathbb{\Gamma}(\nu).$ (15)
+If $\varphi$ is continuous, the comparison between $\varphi$ and $V$ relies on the Hahn-Khalil result described in Eq.
+khalil.
+If $\varphi$ is continuous then, using the same notation as in the $\mathcal{K}\mathcal{L}$ approach
+$(\alpha_{\varphi}\circ\alpha_{1}^{-1}\circ h,\lambda)\in\mathbb{\Gamma}(\nu)$ (16)
+where $\alpha_{1}$ is the $\mathcal{K}_{\infty}$ function appearing in (12).
+In  [Adj25a], we slightly generalized the Lyapunov approach to allow local Lyapunov functions.
+Local means that the inequalities depicted in (12) and (13) hold only on an (forward) invariant set for the dynamics $T$.
+Moreover, for $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ and $k\in\mathcal{S}(\nu,h)$, the function $\beta\mapsto\mathfrak{F}_{\nu}(k,h,\beta)$ defined in Equation (4) increases.
+Therefore, we should consider the smallest element of $\beta\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
+For a Lyapunov function $V$ for $T$, we should consider the operator ratio
+$N_{T}(V):=\sup_{x\neq 0}\dfrac{V(T(x))}{V(x)}$ (17)
+as the scalar $\lambda$ in Equations (15) and (16). The study of the suitable properties of operator ratios in our context was developed in [Adj25a].
+2.3 Problem statement continues
+In Figure 1, we insist on the fact that the existence of a $\mathcal{K}\mathcal{L}$ certificate or a Lyapunov function combined with an additional assumption on $\varphi$ only implies the existence of a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
+In this paper, we exhibit two proper classes of functions ensuring direct and converse theorems for the condition ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$.
+The first class extends the $\mathcal{K}\mathcal{L}$ class of functions.
+We obtain the equivalence between the existence of a weaker $\mathcal{K}\mathcal{L}$ certificate and $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
+The second class extends the set of Lyapunov functions for which the existence of a weaker Lyapunov function is equivalent to the existence of $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
+The constructed pair $(h,\beta)$ is useful when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$.
+Lyapunov or $\mathcal{K}\mathcal{L}$ certificate + $\varphi$ smaller than the stability certificate or $\varphi$ continuous $\exists(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ [Adj25b, Th.2][Adj25a, Th.4,Th.5,Th.6,Th.7] ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ weaker Lyapunov functions / weaker $\mathcal{K}\mathcal{L}$ functions in this paper
+Figure 1: Previous results obtained using $\mathcal{K}\mathcal{L}$ and Lyapunov functions and the results developed in this paper.
+3 A running example from Leslie population evolution models
+In this paper, we illustrate the new tools that we develop using a running example.
+This example comes from population ecology and is a model for studying the evolution of a population according to some period.
+The model studied here is called a Leslie model and was introduced in 40s [Les45, Les48].
+3.1 Introduction
+Leslie population models analyze the growth of the population of a species structured in age classes.
+A class represents a stage of development of the species (juvenile, sub-adult, adult, etc.).
+From this structuration, the population is modeled as a vector of the sizes of the population of each age class.
+The evolution of the population is computed from a matrix called the Leslie matrix.
+The vector of the population sizes at a given year (or another time step) is the image of the vector of the population sizes of the previous year by the Leslie matrix.
+The coefficients of the first row are the fertility rates associated with each age class.
+Hence, those coefficients compute the number of individuals in the first age class that can be generated from each age class.
+The next rows attribute the survival rates to each age class.
+This rate aims to compute the number of individuals that pass from the current age class to the next.
+Hence, the survival rate belongs to $[0,1]$.
+We also note that this survival rate is the unique nonzero coefficient of the row.
+Usually, the questions raised about Leslie population models are about the asymptotic behavior of the population, the maximal growth rates of the population of each class, and the asymptotic proportions of the population of each age class with respect to the total population.
+The question concerning the proportions can be addressed in a different way: what is the proportion of the juvenile (the first class) with respect to the others?
+In this paper, we are interested in the maximum proportion.
+To illustrate our approach in the simplest way, we consider two age-class Leslie population models.
+This means that we only consider the juvenile and adult populations, and we are interested in the maximum proportion of juveniles over adults.
+3.2 Peak computation problem description
+We consider a matrix $L$ parameterized by scalars $a,b$ and $c$, defined as follows:
+$L=\begin{pmatrix}a&b\\ c&0\end{pmatrix}$
+The scalar $a$ represents the number of juveniles that can be generated per period by one juvenile, whereas $b$ represents the number of juveniles that can be generated per period by one adult.
+The scalar $c$ measures the proportion of juveniles that become adults at the end of the period.
+A value of 0 means that adults disappear after the period considered.
+We suppose that one juvenile and one adult can produce strictly more than one juvenile.
+Moreover, we recall that $c$ encodes the percentage of juveniles that become adults and thus cannot exceed one.
+We suppose that some juvenile individuals survive, and so $c$ cannot be null.
+More formally, we assume that
+$a\geq 1,\ b>1\text{ and }c\in(0,1)$ (18)
+To deal with several possible initial configurations simultaneously, we define
+$X^{\mathrm{in}}_{\ell}:=[l_{1},u_{1}]\times[l_{2},u_{2}]\text{ where }0<l_{1}<u_{1}\text{ and }0<l_{2}<u_{2}.$
+In the sequel, we use the following notation for the vertices of $X^{\mathrm{in}}$:
+$x^{ll}:=(l_{1},l_{2})^{\intercal},\ x^{ul}:=(u_{1},l_{2})^{\intercal},\ x^{lu}:=(l_{1},u_{2})^{\intercal}\text{ and }x^{uu}:=(u_{1},u_{2})^{\intercal}$ (19)
+As the proportion of the juvenile over the adults is concerned, our objective function is defined as follows:
+$\varphi_{\ell}:\mathbb{R}^{2}\ni x=(x_{1},x_{2})\mapsto\left\{\begin{array}[]{cr}\dfrac{x}{y}&\text{ if }y\neq 0\\ \\ 0&\text{ if }y=0\end{array}\right$.
+Let us introduce $\pi_{1}:\mathbb{R}^{2}\ni x\mapsto x_{1}$ and $\pi_{2}:\mathbb{R}^{2}\ni x\mapsto x_{2}$. Then we obtain the following peak computation problem
 […]
-Proposition 7.
-Let $u\in\Lambda$. Then :
+Then we define:
+$\omega=(\omega_{k})_{k\in\mathbb{N}}\text{ where }\omega_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi_{\ell}(L^{k}x)\text{ and }\omega_{\rm opt}=\sup_{k\in\mathbb{N}}\omega_{k}.$
+3.3 A rapid preliminary analysis
+3.3.1 Asymptotic divergence
+The eigenvalues of $L$ are
+$\lambda_{1}:=\dfrac{a-\sqrt{a^{2}+4bc}}{2}\text{ and }\lambda_{2}:=\dfrac{a+\sqrt{a^{2}+4bc}}{2}.$
+From (18) it is easy to see $\lambda_{2}>1$, $\lambda_{1}<0$ and $\lambda_{1}/\lambda_{2}\in(-1,0)$.
+The system is thus not asymptotically stable and does not admit a $\mathcal{K}\mathcal{L}$ upper bound or a classical Lyapunov function.
+Moreover, for any starting point $x\in X^{\mathrm{in}}_{\ell}$, $(\pi_{1}(L^{k}x))_{k\in\mathbb{N}}$ and $(\pi_{2}(L^{k}x))_{k\in\mathbb{N}}$ are strictly increasing (from $k=2$ for $\pi_{2}$) and both tend to $\infty$ as $k$ goes to $\infty$.
+We give details about this in the Appendix.
+3.3.2 Existence of a stopping integer
+We introduce the following two auxiliary functions:
+$\mathfrak{a}:\mathbb{R}^{2}\ni x\mapsto\lambda_{2}x_{2}-cx_{1}\text{ and }\mathfrak{b}:\mathbb{R}^{2}\ni x\mapsto cx_{1}-\lambda_{1}x_{2}\kern 5.0pt.$ (20)
+Note that for all $x\in(\mathbb{R}_{+}^{*})^{2}$, for all $k\in\mathbb{N}$,
+$\mathfrak{a}(x)\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\mathfrak{b}(x)=cx_{1}\left(1-\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}\right)+\lambda_{1}x_{2}\left(\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k-1}-1\right)$ (21)
+and thus is strictly positive as $\lambda_{1}<0$ and $\lambda_{1}/\lambda_{2}\in(-1,0)$.
+Now, we show that $\omega_{\rm opt}$ is finite and ${\mathrm{R}}_{\omega}^{s}\neq\emptyset$.
+Using an eigendecomposition of $L$, $\mathfrak{a}$ and $\mathfrak{b}$, we have for all $x\in X^{\mathrm{in}}_{\ell}$ and all $k\in\mathbb{N}$
+$\varphi_{\ell}(L^{k}x)=\dfrac{\lambda_{1}^{k+1}\mathfrak{a}(x)+\lambda_{2}^{k+1}\mathfrak{b}(x)}{c\left(\lambda_{1}^{k}\mathfrak{a}(x)+\lambda_{2}^{k}\mathfrak{b}(x)\right)}=\dfrac{\lambda_{2}}{c}\dfrac{\mathfrak{a}(x)\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k+1}+\mathfrak{b}(x)}{\mathfrak{a}(x)\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\mathfrak{b}(x)}.$ (22)
+From (21), both the numerator and denominator are strictly positive.
+As $X^{\mathrm{in}}_{\ell}$ is the Cartesian product of intervals, we can compute the maximum of $\varphi_{\ell}\circ L^{k}$ from the study of the sign of the partial derivatives.
+We conclude that for all $k\in\mathbb{N}$
+$\omega_{2k}=\max_{x\in X^{\mathrm{in}}_{\ell}}\varphi_{\ell}(L^{2k}x)=\varphi_{\ell}(L^{2k}x^{ul})\text{ and }\omega_{2k+1}=\max_{x\in X^{\mathrm{in}}_{\ell}}\varphi_{\ell}(L^{2k+1}x)=\varphi_{\ell}(L^{2k+1}x^{lu}).$ (23)
+From (22), we observe that for all $x\in(\mathbb{R}_{+}^{*})^{2}$, $\lim_{k\to+\infty}\varphi_{\ell}(L^{k}x)=\lambda_{2}/c$.
+Hence, $\limsup_{k\to+\infty}\omega_{k}=\lim_{k\to+\infty}\omega_{k}=\lambda_{2}/c<+\infty$.
+We conclude that $\omega_{\rm opt}<+\infty$.
+To show that ${\mathrm{R}}_{\omega}^{s}\neq\emptyset$, we have to check whether $\omega_{k}>\lambda_{2}/c$ for some $k\in\mathbb{N}$.
+This can be deduced from the properties of $\mathfrak{a}$.
+Let $x^{*}\in X^{\mathrm{in}}_{\ell}$.
+If $\mathfrak{a}(x^{*})<0$, we get $\omega_{0}\geq\varphi_{\ell}(x^{*})>\lambda_{2}/c$ and $0\in{\mathrm{R}}_{\omega}^{s}$.
+Now, we observe that $\mathfrak{a}\circ L=\lambda_{1}\mathfrak{a}$ on $\mathbb{R}\times\mathbb{R}_{+}^{*}$.
+Hence, if $\mathfrak{a}(x^{*})>0$, we have $\mathfrak{a}(Lx^{*})<0$, $\omega_{1}\geq\varphi_{\ell}(Lx^{*})>\lambda_{2}/c$ and thus $1\in{\mathrm{R}}_{\omega}^{s}$.
+If $\mathfrak{a}(x^{*})=0$.
+From the equality $\mathfrak{a}\circ L=\lambda_{1}\mathfrak{a}$, we have for all $k\in\mathbb{N}$, $\mathfrak{a}(L^{k}x^{*})=0$.
+In this case, $\varphi_{\ell}(L^{k}x^{*})=\lambda_{2}/c$ for all $k\in\mathbb{N}$.
+To conclude that ${\mathrm{R}}_{\omega}^{s}\neq\emptyset$, it suffices to show that $X^{\mathrm{in}}_{\ell}$ cannot be included in $\mathfrak{a}^{-1}(\{0\})$.
+By definition, $X^{\mathrm{in}}_{\ell}$ has a nonempty interior, whereas $\mathfrak{a}^{-1}(\{0\})$ has an empty one; hence $\mathfrak{a}(x)\neq 0$ for some $x\in X^{\mathrm{in}}_{\ell}$.
+Finally, to compute $\omega_{\rm opt}$, the initial vectors $x$ are removed such that $\mathfrak{a}(x)=0$ and we introduce
+$X_{\ell}:=X^{\mathrm{in}}_{\ell}\cap\{x\in\mathbb{R}^{2}:x_{1}/x_{2}\neq\lambda_{2}/c\}=X^{\mathrm{in}}\cap\{x\in\mathbb{R}^{2}:\alpha(x)\neq 0\}.$ (24)
+3.3.3 An alternative simpler solution
+We propose a simple alternative approach to solve the peak computation problem ($DPCP_{\ell}$) as follows. We simply use the fact that for all $k\in\mathbb{N}$ and all $x\in X^{\mathrm{in}}_{\ell}$,
+$\varphi_{\ell}(L^{2k+2}x)-\varphi_{\ell}(L^{2k}x)=\dfrac{\lambda_{2}}{c}\dfrac{\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{2k}\mathfrak{a}(x)\mathfrak{b}(x)\left(1+\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{3}-\dfrac{\lambda_{1}}{\lambda_{2}}-\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{2}\right)}{\left(\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{2k+1}\mathfrak{a}(x)+\mathfrak{b}(x)\right)\left(\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{2k}\mathfrak{a}(x)+\mathfrak{b}(x)\right)}.$
+From (21), the denominator of $\varphi_{\ell}(L^{2k+2}x)-\varphi_{\ell}(L^{2k}x)$ is still strictly positive.
+The numerator is strictly negative when $\mathfrak{a}(x)$ is positive.
+When $\mathfrak{a}(x)$ is strictly positive, we consider instead $\varphi_{\ell}(L^{2k+3}x)-\varphi_{\ell}(L^{2k+1}x)=\varphi_{\ell}(L^{2k+2}Lx)-\varphi_{\ell}(L^{2k}Lx)$ which is strictly negative from the equality $\mathfrak{a}\circ L=\lambda_{1}\mathfrak{a}$.
+Then we conclude that, if $\mathfrak{a}(x)<0$, $(\varphi_{\ell}(L^{2k}x))_{k\in\mathbb{N}}$ is strictly decreasing and, if $\mathfrak{a}(x)>0$, $(\varphi_{\ell}(L^{2k+1}x))_{k\in\mathbb{N}}$ is strictly decreasing.
+Finally, this leads to $\omega_{\rm opt}=\max\{\omega_{0},\omega_{1}\}$.
+3.4 The illustration of our work
+For the running example, we will construct
   1. 1.
-    For all $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we have $\limsup_{k\to+\infty}u_{k}\leq h(0)$;
+    a weaker $\mathcal{K}\mathcal{L}$ certificate;
   2. 2.
-    For all $k\in\mathbb{N}^{*}$, $u_{k}<h(1)$ and $u_{0}\leq h(1)$.
-In Prop. 7, the first statement only requires $h$ to be upper semicontinuous. However, a strictly increasing function $f$ from $I$ to $f(I)$ is surjective if and only if it is continuous on $I$.
-Corollary 1.
-Let $u\in\Lambda$. If there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ such that $u_{0}=h(1)$ then $u_{0}=\max_{k\in\mathbb{N}}u_{k}$.
-It is easy to find a pair of a strictly increasing continuous function and a convergent geometric sequence to construct an upper bound of the sequence $u$ when the supremum of $u$ is finite.
-Proposition 8.
-$u\in\Lambda$ if and only if $\mathbb{\Gamma}(u)\neq\emptyset$.
-Proof.
-Suppose $u\in\Lambda$.
-By definition, $\sup_{n\in\mathbb{N}}u_{n}\in\mathbb{R}$ and the function $h:x\mapsto x+\sup_{n\in\mathbb{N}}u_{n}$ belongs to $\Omega([0,1])$ and for all $\beta\in(0,1)$ and all $k\in\mathbb{N}$, $u_{k}\leq\beta^{k}+\sup_{n\in\mathbb{N}}u_{n}$ and $\mathbb{\Gamma}(u)\neq\emptyset$.
-Now, let $u\in\mathbb{R}^{\mathbb{N}}$ and suppose that $\mathbb{\Gamma}(u)\neq\emptyset$.
-Let $(h,\beta)\in\mathbb{\Gamma}(u)$.
-From Prop.
-7 and as $h$ is finite-valued then $\limsup_{k}u_{k}\leq h(0)\in\mathbb{R}$.
-From Prop.
-1, this implies that $u\in\Lambda$.
-∎
-Some upper bounds $h(\beta^{k})$ appearing in $\mathbb{\Gamma}(u)$ cannot be used in our formula. It depends on how $h(0)$ is placed with respect to the terms of $u$.
-Definition 1 (Useful Strictly Increasing Continuous Functions).
-Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. A function $h\in\mathbb{\Gamma}_{\rm fun}(u)$ is said to be useful for $u$ if the set :
-$\mathcal{S}(u,h):=\{k\in\mathbb{N}:u_{k}>h(0)\}$ (3)
-is nonempty.
-By extension, a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ is useful for $u$ if $h$ is useful for $u$.
-From the first statement of Prop.
-7, if $h$ is useful for $u$, there exists $k\in\mathbb{N}$ such that $u_{k}>h(0)\geq\limsup_{n\to+\infty}u_{n}$ i.e.
-${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-We will see that the contrary also holds for i.e.
-if ${\mathrm{R}}_{u}^{s}\neq\emptyset$, then there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$.
-In Prop. 9, we present some theoretical properties on $\mathbb{\Gamma}_{\rm fun}(u)$ and $\mathbb{\Gamma}_{\rm sc}(u,h)$.
-Proposition 9.
-Let $u\in\Lambda$.
-  1. 1.
-    Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$. Let us define :
-    $\underline{\beta}_{u,h}:=\left\{\begin{array}[]{lr}\displaystyle{\sup_{k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}}\exp\left(\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{k}\right)}&\text{ if }\mathcal{S}(u,h)\cap\mathbb{N}^{*}\neq\emptyset\\ \\ 0&\text{ otherwise}\end{array}\right.;$
-    Then $\mathbb{\Gamma}_{\rm sc}(u,h)$ is the interval $(0,1)$ if $\underline{\beta}_{u,h}=0$ and $[\underline{\beta}_{u,h},1)$ if $\underline{\beta}_{u,h}>0$.
-  2. 2.
-    For all finite family families $\{(h_{1},\beta_{1}),(h_{2},\beta_{2}),\ldots,(h_{n},\beta_{n})\}\subset\mathbb{\Gamma}(u)$:
-    $\max_{i=1,\ldots,n}\beta_{i}\in\bigcap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i});$
-  3. 3.
-    The set $\mathbb{\Gamma}_{\rm fun}(u)$ is a convex sublattice (for the classical order, $\min$ and $\max$ on functions) subset of $\Omega([0,1])$ i.e.
-    • (lattice) for all finite families $\{h_{1},h_{2},\ldots,h_{n}\}\subset\mathbb{\Gamma}_{\rm fun}(u)$, $\min_{i=1,\ldots,n}h_{i}\in\mathbb{\Gamma}_{\rm fun}(u)$ and $\max_{i=1,\ldots,n}h_{i}\in\mathbb{\Gamma}_{\rm fun}(u)$;
-    • (convex) for all $t\in[0,1]$ and $h,g\in\mathbb{\Gamma}_{\rm fun}(u)$, $th+(1-t)g\in\mathbb{\Gamma}_{\rm fun}(u)$.
-  4. 4.
-    For all finite families $\{h_{1},h_{2},\ldots,h_{n}\}\subset\mathbb{\Gamma}_{\rm fun}(u)$, we have actually:
-    $\mathbb{\Gamma}_{\rm sc}(u,\min_{i=1,\ldots,n}h_{i})=\bigcap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})\subseteq\mathbb{\Gamma}_{\rm sc}(u,\max_{i=1,\ldots,n}h_{i})$
-Proof.
-1.
-Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$.
-As $h$ is strictly increasing on $[0,1]$, it follows that if $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ it also holds for all $\beta^{\prime}\in[\beta,1)$.
-Now, it suffices to prove that $\underline{\beta}_{u,h}$ is the infimum of $\mathbb{\Gamma}_{\rm sc}(u,h)$ and it belongs to $\mathbb{\Gamma}_{\rm sc}(u,h)$ when it is strictly positive.
-First, we suppose that $\underline{\beta}_{u,h}=0$.
-By definition, this is the same as for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h(0)$.
-As $h$ is strictly increasing, we have for all $\beta\in(0,1)$, for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h(0)<h(\beta^{k})$.
-We have also $u_{0}\leq h(1)=h(\beta^{0})$ for all $\beta\in(0,1)$ as $h\in\mathbb{\Gamma}_{\rm fun}(u)$.
-Then we have $\mathbb{\Gamma}_{\rm sc}(u,h)=(0,1)$ and $\underline{\beta}_{u,h}=0=\inf\mathbb{\Gamma}_{\rm sc}(u,h)$.
-Now, suppose that $\underline{\beta}_{u,h}>0$.
-If $k=0$, $u_{0}\leq h(\underline{\beta}_{u,h}^{0})$ as $h\in\mathbb{\Gamma}_{\rm fun}(u)$.
-If $k\notin\mathcal{S}(u,h)$ and $k\neq 0$ then $u_{k}\leq h(0)<h(\underline{\beta}_{u,h}^{k})$.
-Let $k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}$, $\exp(\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))/k)\leq\underline{\beta}_{u,h}$ and thus ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\leq\underline{\beta}_{u,h}^{k}$ and then $u_{k}=h(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))\leq h(\underline{\beta}_{u,h}^{k})$.
-Let consider $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$.
-We have for all $k\in\mathbb{N}$, $u_{k}\leq h(\beta^{k})$.
-Let $k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}$.
-By simple calculus, we have $\exp(\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))/k)\leq\beta$.
-Taking the supremum over $\mathcal{S}(u,h)\cap\mathbb{N}^{*}$, we obtain $\underline{\beta}_{u,h}\leq\beta<1$.
-We have proved that $\underline{\beta}_{u,h}\in(0,1)$, for all $k\in\mathbb{N}$, $u_{k}\leq h(\underline{\beta}_{u,h}^{k})$ and for all $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$, $\underline{\beta}_{u,h}\leq\beta$ We conclude that $\underline{\beta}_{u,h}\in\mathbb{\Gamma}_{\rm sc}(u,h)$ and $\underline{\beta}_{u,h}=\inf\mathbb{\Gamma}_{\rm sc}(u,h)$.
-2. The result follows from the fact that all elements of $\mathbb{\Gamma}_{\rm fun}(u)$ and for all $k\in\mathbb{N}^{*}$, the functions $x\mapsto x^{k}$ are increasing on the positive reals set.
-3.
-(Lattice) First, the set $\Omega([0,1])$ is a lattice as finitely $\min$ or $\max$ of strictly increasing continuous functions are still strictly increasing and continuous.
-Now, it suffices to prove the results for the case $n=2$.
-Let $h_{1},h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$.
-There exist $\beta_{1},\beta_{2}\in(0,1)$ such that for all $k\in\mathbb{N}$, $u_{k}\leq h_{1}(\beta_{1}^{k})$ and $u_{k}\leq h_{2}(\beta_{2}^{k})$.
-Let us write $\beta=\max\{\beta_{1},\beta_{2}\}$.
-Then for all $k\in\mathbb{N}$, $u_{k}\leq\min\{h_{1}(\beta_{1}^{k}),h_{2}(\beta_{2}^{k})\}\leq\min\{h_{1}(\beta^{k}),h_{2}(\beta^{k}\}=\min\{h_{1},h_{2}\}(\beta^{k})$.
-The same techniques can be used for the maximum.
-(Convexity) Let $t\in[0,1]$, $h_{1},h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$.
-There exist $\beta_{1},\beta_{2}\in(0,1)$ such that for all $k\in\mathbb{N}$, $u_{k}\leq h_{1}(\beta_{1}^{k})$ and $u_{k}\leq h_{2}(\beta_{2}^{k})$.
-Then $u_{k}=tu_{k}+(1-t)u_{k}\leq th_{1}(\beta_{1}^{k})+(1-t)h_{2}(\beta_{2}^{k})\leq th_{1}(\max\{\beta_{1},\beta_{2}\}^{k})+(1-t)h_{2}(\max\{\beta_{1},\beta_{2}\}^{k})=(th_{1}+(1-t)h_{2})(\max\{\beta_{1},\beta_{2}\}^{k})$.
-We conclude that $th_{1}+(1-t)h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$.
-4.
-Let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,\min_{i=1,\ldots,n}h_{i})$.
-Then $u_{k}\leq\min_{i=1,\ldots,n}h_{i}(\beta^{k})\leq h_{i}(\beta^{k})$ for all $k\in\mathbb{N}$.
-Hence, $\beta\in\cap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})$.
-Now, let $\beta\in\cap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})$.
-For all $i=1,\ldots,n$, for all $k\in\mathbb{N}$, $u_{k}\leq h_{i}(\beta^{k})$.
-Taking the minimum over $i$ leads to the conclusion.
-The inclusion on the right-hand side follows readily from the definition of the finite maximum of functions.
-∎
-We present the main result of this section. It asserts that we can find an optimal useful affine function for $u$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-Theorem 1 (Existence of a Useful Optimal Affine Function).
-Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exist $a>0$, $b\in(0,1)$ and $c\in\mathbb{R}$ such that:
-  1. 1.
-    $\{k\in\mathbb{N}:u_{k}>c\}\neq\emptyset$;
-  2. 2.
-    $u_{k}\leq ab^{k}+c$ for all $k\in\mathbb{N}$;
-  3. 3.
-    $u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c$.
-Proof.
-If ${\mathrm{R}}_{u}^{s}$ is nonempty then $u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}>\limsup_{k\to\infty}u_{k}\in\mathbb{R}\cup\{-\infty\}$ and for all $t\in(\limsup_{k\to\infty}u_{k},\max_{k\in\mathbb{N}}u_{k})$ there exists $k\in\mathbb{N}$, such that $u_{k}>t$.
-So let $c$ be in $(\limsup_{k\to\infty}u_{k},\sup_{k\in\mathbb{N}}u_{k})$.
-Now as $c>\limsup_{k\to+\infty}u_{k}$, there exists $N_{c}\in\mathbb{N}$ s.t.
-$\sup_{k\geq N_{c}}x_{k}\leq c$.
-This is the same as for all $k\geq N_{c}$, $u_{k}\leq c$.
-We conclude that for all $k\geq N_{c}$, $u_{k}\leq c+ab^{k}$ for all $a>0$ and all $b\in(0,1)$.
-Note that as $c<\max_{k\in\mathbb{N}}u_{k}$ and $\mathbf{K}^{s}_{u}$ is the greatest element of the argmax of $u$ then $N_{c}$ is strictly greater than $\mathbf{K}^{s}_{u}$.
-Now let introduce the set :
-$S_{c}:=\{(a,b)\in\mathbb{R}_{+}^{*}\times(0,1):u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}=ab^{\mathbf{K}^{s}_{u}}+c\}.$
-It is easy to see that $S_{c}$ is nonempty, indeed, for all $b$ in $(0,1)$, $(\frac{u_{\mathbf{K}^{s}_{u}}-c}{b^{\mathbf{K}^{s}_{u}}},b)\in S_{c}$.
-Moreover, if $\mathbf{K}^{s}_{u}>0$, for all $(a,b)\in S_{c}$, for all $0<k<\mathbf{K}^{s}_{u}$, $u_{k}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c<ab^{k}+c$ as $b\in(0,1)$.
-We also have $u_{0}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c\leq a+c$ as $b\in(0,1)$.
-Hence, the result $u_{k}\leq ab^{k}+c$ holds on the sets $[0,\mathbf{K}^{s}_{u}]\cap\mathbb{N}$ and $[N_{c},+\infty)\cap\mathbb{N}$ with any element $(a,b)\in S_{c}$.
-It suffices to find $(a,b)\in S_{c}$, such that $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]\cap\mathbb{N}$.
-Recall that we have $\mathbf{K}^{s}_{u}+1\leq N_{c}$.
-Let us write $(-\infty,0)\ni\gamma:=\max_{k\in(\mathbf{K}^{s}_{u},N_{c}]}(u_{\mathbf{K}^{s}_{u}}-u_{k})/(\mathbf{K}^{s}_{u}-k)$.
-This cannot be zero since $\mathbf{K}^{s}_{u}$ is the maximal integer $k$ such that $u_{k}=\max_{n\in\mathbb{N}}u_{n}$.
-Now, let $(a,b)\in\mathbb{R}_{+}\times(0,1)$ and define $f_{a,b}:\mathbb{R}\ni y\mapsto ab^{y}+c$.
-The function $f_{a,b}$ is derivable and strictly convex and $f^{\prime}_{a,b}(y)=ab^{y}\ln(b)$ for all $y\in\mathbb{R}$.
-If $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in S_{c}$ then, by convexity of $f_{a,b}$, the tangent of $f$ at $y=\mathbf{K}^{s}_{u}$ satisfies for all $k\in\mathbb{N}$, $f(k)\geq f^{\prime}(\mathbf{K}^{s}_{u})(k-\mathbf{K}^{s}_{u})+f(\mathbf{K}^{s}_{u})$ which is the same as $ab^{k}+c\geq\gamma(k-\mathbf{K}^{s}_{u})+u_{\mathbf{K}^{s}_{u}}$.
-By definition of $\gamma$, we have for all $k\in(\mathbf{K}^{s}_{u},N_{c})$, $(\mathbf{K}^{s}_{u}-k)\gamma\leq u_{\mathbf{K}^{s}_{u}}-u_{k}$ and so $u_{k}\leq u_{\mathbf{K}^{s}_{u}}+(k-\mathbf{K}^{s}_{u})\gamma$.
-We conclude that if $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in S_{c}$, $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]$.
-Finally, we are looking for $(a,b)\in\mathbb{R}_{+}\times(0,1)$ such that :
-$f_{a,b}(\mathbf{K}^{s}_{u})=u_{\mathbf{K}^{s}_{u}}\text{ i.e. }(a,b)\in S_{c}\text{ and }f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$
-From $(a,b)\in S_{c}$, we get $ab^{\mathbf{K}^{s}_{u}}=u_{\mathbf{K}^{s}_{u}}-c$.
-Hence, $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ if and only if $\ln(b)=\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c}<0$.
-So $b=\exp(\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c})\in(0,1)$ and $a=(u_{\mathbf{K}^{s}_{u}}-c)\exp(-\frac{\mathbf{K}^{s}_{u}\gamma}{u_{\mathbf{K}^{s}_{u}}-c})>0$.
-∎
-Th 1 brings deeper details compared with Prop.
-8.
-The existence of a useful and optimal function is important for the construction and the optimality of the proposed formula.
-Moreover, the proof of Th 1 exhibits the parameters of the affine function to choose.
-Unfortunately (but naturally) all the parameters depend on $\mathbf{K}^{s}_{u}$ and $u_{\mathbf{K}^{s}_{u}}$ what is exactly what we want to compute.
-Corollary 2.
-Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-Then, there exists a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ useful for $u$ such that $u_{\mathbf{K}^{s}_{u}}=h(\beta^{\mathbf{K}^{s}_{u}})$.
-Proposition 10.
-Let $u\in\Lambda$. Then ${\mathrm{R}}_{u}^{s}\neq\emptyset$ if and only if there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$.
-Proof.
-Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-From Th 1, there exists a function $h$ of the form $x\mapsto ax+c$ which satisfies $\mathcal{S}(u,h)\neq\emptyset$.
-For the reverse implication, let assume ${\mathrm{R}}_{u}^{s}=\emptyset$.
-From Prop.
-7, for all $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we have $\mathcal{S}(u,h)\subseteq{\mathrm{R}}_{u}^{s}$.
-This proves that the set $\mathcal{S}(u,h)$ is empty for all $h\in\mathbb{\Gamma}_{\rm fun}(u)$.
-∎
-Example 2.
-Let us consider the peaks computation problem mentioned at Subsection 2.2. We now construct an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for the sequence $\nu$ presented at Eq. (2).
-Let $n_{0}$ be the smallest integer $n$ such that $\varphi_{n}(\mu)<0$.
-With simple computations, we get $n_{0}=\lfloor\ln(2p/(3\mu))/\ln(3/2)\rfloor+1$.
-As $\ln(2p/(3\mu))/\ln(3/2)-\ln(3\mu/p)/\ln(2/3)=(\ln(2p)-\ln(p))/\ln(3/2)$, we have $n^{0}\geq\overline{n}$.
-For all $n\geq\overline{n}$, $\nu_{n}=\varphi_{n}(\mu)$ and thus for all $n\geq n_{0}$, $\nu_{n}<0$.
-Following the proof of Th 1, we can impose to $h$ to be linear i.e.
-$h:x\mapsto ax$.
-Indeed, the value $c$ of Th 1 can be chosen to be 0 as we have $\lim_{k\to+\infty}\nu_{k}=-\infty$ and $p^{2}/3>0$.
-In this case, if $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$, $h$ is useful for $\nu$.
-A simple upper bound over $(\nu_{n})_{n}$ is furnished from any couple $(a,\beta)$ such that:
-$a>0,\ \beta\in(0,1)\text{ and }a\beta^{n_{0}}\geq p^{2}/3$
-For example, we can choose $a\beta^{n_{0}}=p^{2}/3$ for which $a=2p^{2}/3$ and $\beta=2^{-1/n_{0}}$ is a solution. ∎
-Th.
-1 provides a theoretical result and to have in hand an optimal function $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$ is very unlikely to happen.
-However, if we have any function $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we are able to produce a safe upper bound of $\mathbf{K}^{s}_{u}$ and then a solution of Problem 2.
-To compute it, we define, for $u\in\mathbb{R}^{\mathbb{N}}$, we introduce the function $\mathfrak{F}_{u}$, from ×$\mathbb{N}\times\Omega([0,1])\times(0,1)\mapsto\mathbb{R}\cup\{+\infty\}$ defined for all $k\in\mathbb{N}$, and $h\in\Omega([0,1])$ and $\beta\in(0,1)$ by:
-$\mathfrak{F}_{u}(k,h,\beta):=\left\{\begin{array}[]{lr}\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{\ln(\beta)}&\text{ if }(h,\beta)\in\mathbb{\Gamma}(u)\text{ and }k\in\mathcal{S}(u,h)\\ +\infty&\text{ otherwise}\end{array}\right.$ (4)
-We will show that $\mathfrak{F}_{u}$ has the desired properties to solve Problem 2.
-Proposition 11.
-Let us take a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ useful for $u$ and let $k\in\mathcal{S}(u,h)$. The following statements hold:
-  1. 1.
-    $\mathfrak{F}_{u}(k,h,\beta)$ is well-defined and strictly positive if $u_{0}<h(1)$ and null if $k=0$ and $u_{0}=h(1)$;
-  2. 2.
-    $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1=\min\{j\in\mathbb{N}:h(\beta^{j})<u_{k}\}=\min\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$;
-  3. 3.
-    $\mathfrak{F}_{u}(k,h,\beta)\geq k$;
-  4. 4.
-    For all $j\in\mathbb{N}$, $j>\mathfrak{F}_{u}(k,h,\beta)$, $u_{j}<u_{k}$;
-  5. 5.
-    Let $j\in\mathcal{S}(u,h)$. If $u_{j}\leq u_{k}$ then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
-Proof.
-1.
-Let $(h,\beta)\in\mathbb{\Gamma}(u)$.
-As $h\in\mathbb{\Gamma}_{\rm fun}(u)$ and we only consider $u_{k}>h(0)$, we have ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})>0$.
-We also have by definition of $h$, if $k\in\mathbb{N}^{*}$, $u_{k}<h(1)$ meaning that ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\in(0,1)$.
-If $u_{0}<h(1)$, the strict inequality is also valid at $k=0$.
-By definition, $\beta\in(0,1)$.
-We can apply the natural logarithm to ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})$ and $\beta$; those values are both strictly negative and then the quotient is positive.
-If we have $u_{0}=h(1)$ and $k=0$, $\mathfrak{F}_{u}(k,h,\beta)=0$.
-2.
-As $h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and $k\in\mathcal{S}(u,h)$, there exists $j\in\mathbb{N}$ such that $h(\beta^{j})<u_{k}$.
-Then the sets $\{j\in\mathbb{N}:h(\beta^{j})<u_{k}\}$ and $\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$ are equal and nonempty.
-Let $j\in\mathbb{N}$ s.t.
-$\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})$.
-Applying the natural logarithm and dividing by $\ln(\beta)<0$ leads to $j>\mathfrak{F}_{u}(k,h,\beta)$ and as $j\in\mathbb{N}$ it implies that $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$.
-By definition of integer parts and as $\beta\in(0,1)$, $\ln(\beta)(\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1)<\mathfrak{F}_{u}(k,h,\beta)\ln(\beta)=\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))$.
-Applying the exponential leads to $\beta^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})$ and $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ belongs to $\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$ and is its minimum integer.
-3.
-The pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(u)$ and thus for all $k\in\mathbb{N}$, $u_{k}\leq h(\beta^{k})$ which is the same as ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\leq\beta^{k}$.
-Applying the natural logarithm and dividing by $\ln(\beta)$ leads to the result.
-4.
-Let $j>\mathfrak{F}_{u}(k,h,\beta)$ then $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ and, as $h$ is increasing, $h(\beta^{j})\leq h(\beta^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1})<u_{k}$ from the second statement.
-Moreover, by definition of $(h,\beta)$, we have $u_{j}\leq h(\beta^{j})$.
-It follows that $u_{j}<u_{k}$.
-[…]
-We relate some structural properties of the sets $\mathbb{\Gamma}_{\rm fun}(u)$ and $\mathbb{\Gamma}_{\rm sc}(u,h)$ depicted at Prop. 9 with the computations of $\mathfrak{F}_{u}(k,h,\beta)$.
-Proposition 12.
-Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. The following properties hold:
-  1. 1.
-    Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$ and $k\in\mathcal{S}(u,h)$.
-Let $\beta,\beta^{\prime}\in\mathbb{\Gamma}_{\rm sc}(u,h)$ such that $\beta\leq\beta^{\prime}$.
-Then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(k,h,\beta^{\prime})$.
-  2. 2.
-    Let $g,h\in\mathbb{\Gamma}_{\rm fun}(u)$.
-Let $k\in\mathcal{S}(u,\min\{g,h\})$.
-Let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,\min\{g,h\})$.
-Then $\mathfrak{F}_{u}(k,\min\{g,h\},\beta)\leq\min\{\mathfrak{F}_{u}(k,g,\beta),\mathfrak{F}_{u}(k,h,\beta)\}$.
-Proof.
-1.
-The function $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is increasing as it is the composition of (×$(\times\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})))\circ{\cdot}^{-1}\circ\ln$ where $\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))$ is negative.
-2.
-First from Prop 9, $f:=\min\{g,h\}\in\mathbb{\Gamma}_{\rm fun}(u)$.
-Second, for all $x\in[0,1]$, $f(x)\leq g(x)$ and we can apply ${{f}_{\left([0,1]\right)}^{-1}}$ which is strictly increasing on $[0,1]$ and $x\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(g(x))$.
-Thus, we write $y=g(x)$ and thus $x=\displaystyle{{g}_{\left([0,1]\right)}^{-1}}(y)$.
-We conclude that ${{g}_{\left([0,1]\right)}^{-1}}(y)\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(y)$.
-Using the same technique, we can prove that ${{h}_{\left([0,1]\right)}^{-1}}(y)\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(y)$.
-The result follows from the fact that $\ln$ is increasing and $\ln(\beta)$ is negative.
-∎
-The first statement of Prop.
-12 has to be read as follows : we have to choose the smallest possible $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ to reduce the approximation gap.
-The second statement has a similar meaning in a functional sense : if we have several functions in $\mathbb{\Gamma}_{\rm fun}(u)$ we have to consider the minimum functions to reduce the overapproximation gap.
-Proposition 13.
-Let $u\in\Lambda$ and $(h,\beta)\in\mathbb{\Gamma}(u)$. Then for all $j\in\mathbb{N}$, $\mathbf{K}^{s}_{u}\leq\lfloor\mathfrak{F}_{u}(j,h,\beta)\rfloor$. Moreover, if $h$ is useful for $u$, we have:
-$\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)\enspace.$
-Proof.
-By definition of $\mathfrak{F}_{u}$, if $u_{j}\leq h(0)$, the inequality holds.
-Now let $j\in\mathcal{S}(u,h)$, we have $u_{j}\leq\sup_{k\in\mathbb{N}}u_{k}=u_{\mathbf{K}^{s}_{u}}$.
-Hence, from the fifth statement of Prop.
-11, $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
-We conclude from the third statement of Prop.
-11 that $\mathbf{K}^{s}_{u}\leq\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor$ and $\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)$ whenever the set $\mathcal{S}(u,h)$ is nonempty.
-Indeed, if $\mathcal{S}(u,h)$ is nonempty, then $u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}$ is also strictly greater than $h(0)$.
-∎
-The result of Prop.
-13 is not surprising but can be directly used in a very particular context.
-Suppose that we have an explicit closed formula for $u_{k}$ but not derivable.
-The function $h$ can regularize and strictly convexify the closed formula.
-We thus obtain with some calculus the exact value of $\mathbf{K}^{s}_{\nu}$ as the integer part (or integer part +1) of the zero of the derivative of $s\to\mathfrak{F}_{u}(s,h,\beta)$ (extended in the reals).
-We can go further than the results of Prop 13 to get a solution of Problem 2. Th. 2 constitutes the next step of Prop 13.
-Theorem 2 (Solution of Problem 2).
-Let $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$ be useful for $u$ and let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ and $k\in\mathcal{S}(u,h)$.
-Then $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor$ is a solution of Problem 2 that is :
-$\max_{n\in\mathbb{N}}u_{n}=\max\{u_{k}:k=0,\ldots,\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor\}.$
-Proof.
-It suffices to combine Prop. 6 and Prop. 13. ∎
-Th.
-2 shows how the function depicted at Eq.
-(4) is used to solve Problem 2.
-It suffices to construct an upper bound of the sequence of interest of the form of $(h(\beta^{k}))_{k\in\mathbb{N}}$ where $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$; moreover, $h(0)$ has to be strictly smaller than some terms of the sequence of interest.
-In theory, when the limsup of the sequence of interest is not the supremum of the sequence, such couples exist and one of them is optimal (Th.
-1) and $h$ can even be chosen as an affine function.
-Example 3.
-To give a practical view of the function provided at Eq.
-(4), we evaluate $\mathfrak{F}_{\nu}(k,h,\beta)$ for the sequence $\nu$ described at Eq (2) of Subsection 2.2.
-The pairs $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ considered are the ones proposed at Example 2.
-We evaluate the function for different values of $p$ and $\mu$ to two choices of pairs $(h,\beta)$.
-In Example 2, we proposed linear functions $h$ characterized by a scalar $a$.
-We recall that the pair are assumed to satisfy the conditions: $a>0$, $\beta\in(0,1)$ and $a\beta^{n_{0}}\geq p^{2}/3$ where $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\mu)<0$.
-First, as the function $\mathfrak{F}_{u}$ depends on the values of the term $\nu_{k}$, we show the seven first $\nu_{0},\ldots,\nu_{6}$ (rounded to the nearest hundredth) for certain values of $p$ and $\mu$ at Table 1.
-$\nu_{k}$
-k 0 1 2 3 4 5 6 $\mathbf{K}^{s}_{\nu}$ : $\nu_{\rm opt}$
-$p=30$ $\mu=\frac{1}{3}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 8 : 300
-$\mu=\frac{1}{10}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 11 : 300
-$\mu=\frac{1}{1000}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 22 : 300
-$p=9$ $\mu=\frac{1}{3}$ 8.25 11.81 16.45 21.83 26.34 27. 25.09 5 : 27
-$\mu=\frac{1}{10}$ 8.25 11.81 16.45 21.83 26.34 27. 27. 8 : 27
-$\mu=\frac{1}{1000}$ 8.25 11.81 16.45 21.83 26.34 27. 27. 19 : 27
-$p=3$ $\mu=\frac{1}{3}$ 2.25 2.81 3 2.95 1.58 -4.03 -20.47 2 : 3
-$\mu=\frac{1}{10}$ 2.25 2.81 3 3. 3. 3. 2.94 5 : 3
-$\mu=\frac{1}{1000}$ 2.25 2.81 3. 3. 3. 3. 3. 17 : 3
-Table 1: Few values of the terms $\nu_{k}$ defined at Eq. (2) depending on the problem parameters $p$ and $\mu$.
-In Table 1, we observe that the terms $\nu_{k}$ strictly increases until the optimal values (the smallest integer maximizer $\mathbf{K}_{\nu}$) and strictly decreases after (after the greatest integer maximizer $\mathbf{K}^{s}_{\nu}$).
-This observation will be formally proved and exploited later on.
-Another point related to this observation is that the supremum of $\nu_{k}$ for $k\notin\operatorname{Argmax}(\nu)$ seems to be either $\nu_{\underline{n}-1}$ or $\nu_{\overline{n}}$.
-For example, for $p=3$ and $\mu=1/3$, $\nu_{3}>\nu_{1}$ whereas for $p=9$ and $\mu=1/3$, $\nu_{4}>\nu_{5}$.
-Actually, we will see that the rank preceding $\underline{n}$ is not the one to consider to compute $\nu_{k}$ for $k\notin\operatorname{Argmax}(\nu)$.
-Now, we evaluate the function $\mathfrak{F}_{\nu}$ for our example.
-To simplify, we write $\mathfrak{F}_{\nu}(k,a,\beta)$ instead of $\mathfrak{F}_{\nu}(k,h,\beta)$.
-We choose two pairs: first, $a=p^{2}$ and $\beta=(2/3)^{1/n_{0}}$ and second $a=2p^{2}/3$ and $\beta=(1/2)^{n_{0}}$.
-The values of $\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ for these two choices are provided at Table 2.
-$\lfloor\mathfrak{F}_{\nu}(k,p^{2},(2/3)^{1/n_{0}})\rfloor$ $\lfloor\mathfrak{F}_{\nu}(k,p^{2}/3,(1/2)^{1/n_{0}})\rfloor$
-k $\underline{n}$ $\overline{n}$ $n_{0}$ 0 1 2 3 4 $\mathbf{K}^{s}_{\nu}$ 0 1 2 3 4 $\mathbf{K}^{s}_{\nu}$
-$p$ $\mu=\frac{1}{3}$ 8 9 10 84 74 65 56 47 27 43 37 32 26 21 9
-= $\mu=\frac{1}{10}$ 8 12 13 97 84 72 61 50 35 56 49 42 35 28 12
-30 $\mu=\frac{1}{1000}$ 8 23 24 202 179 156 131 113 65 104 91 77 64 52 23
-$p$ $\mu=\frac{1}{3}$ 5 6 7 39 33 27 22 19 18 18 15 12 9 7 7
-= $\mu=\frac{1}{10}$ 5 9 10 56 47 39 32 27 27 27 21 17 13 10 9
-9 $\mu=\frac{1}{1000}$ 5 20 21 118 99 82 67 58 56 56 46 36 27 21 20
-$p$ $\mu=\frac{1}{3}$ 2 3 4 13 11 10 $10^{*}$ $17^{*}$ 10 5 4 3 $4^{*}$ $7^{*}$ 3
-= $\mu=\frac{1}{10}$ 2 4 5 23 20 18 18 18 18 9 7 7 7 7 7
-3 $\mu=\frac{1}{1000}$ 2 18 18 61 51 48 48 48 48 25 19 18 18 18 18
-Table 2: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h,\beta)\rfloor$ described at Eq. (4) for the sequence $\nu$ of Eq. (2) and $(h,\beta)\in\mathbb{\Gamma}(\nu)$ proposed at Ex. 2
-The numbers of Table 2 at columns $\lfloor\mathfrak{F}_{\nu}(k,p^{2},(2/3)^{1/n_{0}})\rfloor$ and $\lfloor\mathfrak{F}_{\nu}(k,2p^{2}/3,(1/2)^{1/n_{0}})\rfloor$ represent safe upper bounds on the number of terms of the sequence $\nu$ of Eq.
-(2) to compare to obtain the supremum of the terms $\nu_{k}$.
-From the analysis made at Subsection 2.2, $\mathbf{K}^{s}_{\nu}$ is well identified and is equal to $\overline{n}-1$.
-We observe that all numbers of the concerned columns are greater than $\overline{n}-1$ as stated at Prop.
-13.
-For both choices of $a$ and $\beta$, the best upper bound is furnished by the column $\lfloor\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},a,\beta)$ as asserted at Prop.
-13.
-Note that four integers are followed by an asterisk.
-In practice, those numbers are useless since not computed.
-Indeed, applying the fifth statement of Prop.
-11, it suffices to compute a new stopping integer at $k$ when we have $\nu_{j}<\nu_{k}$ for all $j<k$.
-Otherwise, the stopping integer is greater than at least one previous computed integer.
-Moreover, for the case $a=2p^{2}/3$ and $\beta=(1/2)^{1/n_{0}}$, following Th 2, the terms $\nu_{k}$, for $k>3$, are not computed as $\mathfrak{F}_{\nu}(3,2p^{2}/3,(1/2)^{1/n_{0}})=3$ and we only have to compare $\nu_{0},\nu_{1},\nu_{2}$ and $\nu_{3}$ to obtain the maximal values of the terms $\nu_{k}$.
-Finally, Table 2 highlights the importance of the choice of the pair $(h,\beta)\in\mathbb{\Gamma}(u)$ to compute a less conservative upper bound of $\mathbf{K}^{s}_{u}$.
-For general peaks computation problems of the form Problem 1, those upper bounds serve to fix the number of instances of Problem Pk to solve to compute the optimal value of the peaks computation problem.
-As Problem Pk are constrained optimization problems, a too conservative upper bound can dragstically increase the computational time.
-∎
-Theorem 3 (Formula Optimality).
-For all $u\in\mathbb{R}^{\mathbb{N}}$, we have the following formula:
-$\mathbf{K}^{s}_{u}=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{\ln(\beta)}$
-and the $\inf$ can be replaced by $\min$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-Proof.
-Let us define
-$F:=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta).$
-If $u\notin\Lambda$, then from Prop. 8, $\mathbb{\Gamma}(u)=\emptyset$ and thus $F=+\infty$. Moreover, from Prop. 2, $\mathbf{K}^{s}_{u}=+\infty$.
-Suppose that $u\in\Lambda$.
-From Prop.
-8, $\mathbb{\Gamma}(u)\neq\emptyset$.
-Assume that $\Delta^{s}_{u}$ is empty.
-This is the same as $\mathbf{K}^{s}_{u}=+\infty$.
-From Prop.
-3, ${\mathrm{R}}_{u}^{s}=\emptyset$ and from Prop.
-10, $\mathcal{S}(u,h)=\emptyset$ which leads to $F=+\infty$.
-Assume that $\Delta^{s}_{u}$ is nonempty and thus ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
-From Prop.
-13, $\mathbf{K}^{s}_{u}\leq F$.
-Moreover, from Corollary 2, there exists $(h,\beta)\in\mathbb{\Gamma}(u)$ such that $u_{\mathbf{K}^{s}_{u}}=h(\beta^{\mathbf{K}^{s}_{u}})>h(0)$ and $F<+\infty$.
-Taking the inverse of $h$ on $[0,1]$ and the natural logarithm, we get $\mathbf{K}^{s}_{u}=\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$.
-By definition of $F$, $F\leq\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)=\mathbf{K}^{s}_{u}$.
-We conclude that $\mathbf{K}^{s}_{u}=F$.
-∎
-In general, we cannot rely on pre-analysis of the system as it is done in Subsection 2.2 to solve a peaks computation problem.
-Thanks to Th.
-2, we reduce a peaks computation problem to the construction of a pair $(h,\beta)$ useful for $\nu$.
-In consequence, we are faced to the following problem on the sequence $\nu$ defined at Eq (1).
-Problem 3.
-Construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$.
-Next sections of the paper are focused on classical tools used in stability theory of dynamical systems such as $\mathcal{K}\mathcal{L}$-stability and Lyapunov functions.
-We show how these tools can be adapted to solve Problem 3.
+    a weaker Lyapunov function.
+for the discrete-time peak computation problem ($DPCP_{\ell}$).
+These constructions will be completed independently of the results obtained in Subsubsection 3.3.3.
+To be more explicit, we will apply our techniques to explicit numerical data.
+The following $2\times 2$ Leslie matrix has been found in [RW88].
+$L:=\begin{pmatrix}1&4\\ 0.5&0\end{pmatrix}$. (25)
+From this matrix, we obtain
+$a=:1,\ b:=4,\ c:=0.5,\lambda_{1}:=-1\text{ and }\lambda_{2}:=2.$
+Hence, $\lambda_{2}/c=4$, $(\mathbb{R}_{+}^{*})^{2}\ni x\mapsto\mathfrak{a}(x):=2x_{2}-0.5x_{1}$ and $(\mathbb{R}_{+}^{*})^{2}\ni x\mapsto\mathfrak{b}(x):=0.5x_{1}+x_{2}$.
+In our approach, the choice of the initial condition set is important. We propose a specific initial conditions set for which $\mathfrak{a}$ takes positive and negative values. We thus define:
+$X^{\mathrm{in}}:=[10;40]\times[5;50]$
+that leads to
+$x^{ll}:=(10,5)^{\intercal},\ x^{lu}:=(10,50)^{\intercal},\ x^{ul}:=(40,5)^{\intercal}\text{ and }x^{uu}:=(40,50)^{\intercal}.$
 4 Extended $\mathcal{K}\mathcal{L}$ Theory for Discrete-Time Systems Maximization
-In this section, we come back to the initial problem that is a peaks computation problem i.e.
-the maximization of a function over the reachable values of a discrete-time dynamical system.
-We recall that the data are the following:
+In this paper, we are interested in the resolution of Problem (DPCP), for which the underlying dynamical system $x_{k+1}=T(x_{k})$ is not necessarily asymptotically stable, as shown in the example provided in Section 3.
+However, Inequality (6) holds if and only if the system $x_{k+1}=T(x_{k}),x_{0}\in\mathbb{R}^{d}$ is said to be uniformly global asymptotic stability (UGAS) (see, for example [NT04, Prop.
+1] or [ATP10, Th.
+2]).
+On the other hand, Inequality (10) provides useful information on the asymptotic behavior of the system: the limit superior of $\nu_{k}$ is smaller than $\inf_{k\in\mathbb{N}}\gamma(\sup_{x\in X^{\mathrm{in}}}\psi(x),k)$ and thus finite.
+However, Inequality (10) forces our sequence $\nu$ to have a positive term to obtain the usefulness of the pair proposed in (11).
+Therefore, Inequality (10) constitutes our starting point.
+Thus, we define the class of functions $\mathcal{KL}_{\rm gen}$ that extends the class of functions $\mathcal{K}\mathcal{L}$.
+The class $\mathcal{KL}_{\rm gen}$ permits the construction of a useful element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ and the analog of Inequality (10) is verified by replacing $\gamma$ with a $\mathcal{KL}_{\rm gen}$ function.
+Contrary to the classical $\mathcal{K}\mathcal{L}$ certificate approach, we prove a converse theorem: if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then the analog of Inequality (10) holds for some $\gamma\in\mathcal{KL}_{\rm gen}$.
+We also provide an extension to the famous Sontag’s result [Son98, Proposition 7] recalled above for the $\mathcal{KL}_{\rm gen}$ class.
+It allows to obtain an element $h\in\mathbb{\Gamma}(\nu)$ such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$ when the analogue of (10) is satisfied.
+We recall that the data are as follows:
   • an initial conditions set $X^{\mathrm{in}}\subseteq\mathbb{R}^{d}$;
@@
   • a function $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$ such that $\varphi(0)=0$.
-We recall that we want compute $\nu_{\rm opt}=\sup\{\nu_{k}:k\in\mathbb{N}\}$ where $\nu_{k}=\sup\{\varphi(T^{k}(x):x\in X^{\mathrm{in}}\}$.
-To use the results of Section 3 with the sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$, we suppose that Assumption 1 holds.
-We propose to simplify the notation of the maximal value of an objective function $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ on $X^{\mathrm{in}}$, we introduce the following new notation:
+We recall that we define $\nu_{\rm opt}=\sup\{\nu_{k}:k\in\mathbb{N}\}$ where $\nu_{k}=\sup\{\varphi(T^{k}(x):x\in X^{\mathrm{in}}\}$.
+Naturally, we suppose that Assumption 2 holds.
+To simplify the notation of the maximal value on $X^{\mathrm{in}}$ of a function $f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}$, we introduce the following new notation:
 $\overline{f_{X^{\mathrm{in}}}}:=\sup\{f(x):x\in X^{\mathrm{in}}\}$
-and the set of functions which are bounded from above on $X^{\mathrm{in}}$:
-$\operatorname{Fin}\left(X^{\mathrm{in}}\right):=\{f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}:\overline{f_{X^{\mathrm{in}}}}<+\infty\}$
-With these notations, $\nu_{k}=\overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}$ and by following Assumption 1, we have for all $k\in\mathbb{N}$, $\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$.
-Following Subsection 3.2, we have to study, first, the limsup of the sequence $\nu$ and then construct, when the limsup is finite, an element $(h,\beta)$ of $\mathbb{\Gamma}(\nu)$ i.e.
-$h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and for all $k\in\mathbb{N}$, $h(\beta^{k})$ is greater than $\nu_{k}$.
-We also have to guarantee that if ${\mathrm{R}}_{\nu}^{s}$ is nonempty, $h$ has to be useful for $\nu$ i.e.
-$\mathcal{S}(\nu,h)$ is nonempty.
-Without this latter condition, the function $h$ is so useless as $\mathfrak{F}_{u}(k,h,\beta)=+\infty$.
-In this section, we will revisit $\mathcal{KL}$ class of functions to study the limsup of $\nu$ and to construct our necessary pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$.
-4.1 Using the classical $\mathcal{KL}$ class of functions
-To study global asymptotic stability of discrete-time systems can be done by means of comparison functions (see e.g.
-[GGLW14]).
-For a comprehensive literature about comparison functions, the interested reader can refer to  [Kel14].
-Comparison functions can be used for our problem to study the limsup of the sequence $\nu$.
-We start by recalling some necessary definitions about comparison functions.
-Definition 2 ($\mathcal{K}$/$\mathcal{K}_{\infty}$ functions classes).
-A function $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}$ if $\alpha$ is strictly increasing continuous and $\alpha(0)=0$.
-If moreover, $\lim_{x\to+\infty}\alpha(x)=+\infty$, $\alpha$ is said to belong to $\mathcal{K}_{\infty}$.
-Definition 3 ($\mathcal{L}$ functions class).
-A function $\sigma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to $\mathcal{L}$ if it is continuous, strictly decreasing and $\lim_{x\to+\infty}\sigma(x)=0$.
-Definition 4 ($\mathcal{K}\mathcal{L}$ functions class).
-A function $\gamma:\mathbb{R}_{+}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}\mathcal{L}$ if for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ belongs to $\mathcal{K}$ and if for all $s\in\mathbb{R}_{+}$, $t\mapsto\gamma(s,t)$ belongs to $\mathcal{L}$.
-The class $\mathcal{K}$ is included in $\Omega([0,1])$ but $\mathcal{K}$ functions are strictly positive except at 0 where they are null.
-Hence, $h\in\mathcal{K}\cap\mathbb{\Gamma}_{\rm fun}(\nu)$ is useful for $\nu$ if at least one term $\nu_{k}$ is strictly positive.
-Thus, in this subsection, Assumption 2 must hold.
-Assumption 2.
-There exists $k\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))>0$.
-We recall the result obtained by Sontag.
-Lemma 1 (Proposition 7 of [Son98]).
-For all $\gamma\in\mathcal{K}\mathcal{L}$, there exists $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $s,t\in\mathbb{R}_{+}$, $\gamma(s,t)\leq\theta_{1}(\theta_{2}(s)e^{-t})$.
-Theorem 4.
-Assume that there exist a function $\gamma\in\mathcal{K}\mathcal{L}$ and a function $\psi:\mathbb{R}^{d}\mapsto\mathbb{R}_{+}$ such that $\overline{\psi_{X^{\mathrm{in}}}}>0$.
-We also assume that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$:
-$\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\enspace.$
-Then:
-  1. 1.
-    $\limsup_{n\to+\infty}\nu_{n}\leq 0$ and if, moreover, Assumption 2 holds then ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$;
-  2. 2.
-    there exists $h\in\mathcal{K}_{\infty}$ s.t. $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ and, if, moreover, Assumption 2 holds then $h$ is useful for $\nu$.
-Proof.
-[…]
-2.
-This a direct consequence of the Sontag’s result (Lemma 1 here).
-Indeed, there exist $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\leq\theta_{1}(\theta_{2}(\psi(x))e^{-k})$.
-Then $\nu_{k}\leq h(e^{-1}):=\sup_{x\in X^{\mathrm{in}}}\theta_{1}(\theta_{2}(\psi(x)e^{-k}))$.
-It is easy to see, as $\theta_{1}$ and $\theta_{2}$ are increasing and continuous, that $h$ can be also written $y\mapsto\theta_{1}(\theta_{2}(\overline{\psi_{X^{\mathrm{in}}}}y))$.
-The function $h$ is thus the compositions of $\theta_{1}$, $\theta_{2}$ and $y\mapsto\overline{\psi_{X^{\mathrm{in}}}}y$ which belong to $\mathcal{K}_{\infty}$ (the third as $\overline{\psi_{X^{\mathrm{in}}}}>0$) and thus belongs to $\mathcal{K}_{\infty}$.
-Finally, under Assumption 2, there exists $k\in\mathbb{N}$, such that $\nu_{k}>0=h(0)$.
-∎
-Usually i.e.
-when $\varphi=\|\cdot\|$, the inequality $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)$ (for $\psi=\|\cdot\|$) is true for all $x\in\mathbb{R}^{d}$ and all $k\in\mathbb{N}$ if and only if the system $x_{k+1}=T(x_{k}),x_{0}\in\mathbb{R}^{d}$ is said to be uniformly global asymptotic stability (UGAS) (see for example [NT04, Prop.
-1] or [ATP10, Th.
-2].
-This implies that $x_{k}$ tends to 0 as $k$ tends to +$\infty$+ ∞.
-Thus, with a general function $\varphi$, we obtain a one-sided information on the asymptotic behaviour of the dynamics.
-We could obtain the other side applying the theorem for $-\varphi$.
+and the set of functions that are bounded from above on $X^{\mathrm{in}}$:
+$\operatorname{Fin}\left(X^{\mathrm{in}}\right):=\{f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}:\overline{f_{X^{\mathrm{in}}}}<+\infty\}.$
+With those notations, $\nu_{k}=\overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}$ and by following Assumption 2, we have for all $k\in\mathbb{N}$, $\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$.
 4.1 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
-In this paper, we are not interested in stability of the dynamical system, we can relax the strict increase (or decrease), the continuity and the positivity (on the state variable) conditions required for $\mathcal{KL}$ functions.
+One of the limitations of using classical $\mathcal{K}\mathcal{L}$ functions in (10) is the positivity of $\mathcal{K}\mathcal{L}$ functions.
+To ensure the usefulness of the element $(h,\beta)$ of Equation (11), our sequence $\nu$ must have a positive term.
+Then, we should relax the positivity on the state-variable part of $\mathcal{K}\mathcal{L}$ functions (and not on the time variable).
+We also relax the strict increase (or decrease) and continuity required to be a $\mathcal{KL}$ function.
 Definition 2 (The Class $\mathcal{KL}_{\rm gen}$).
-A function $\gamma:\mathbb{R}\times\mathbb{R}_{+}\mapsto\mathbb{R}$ belongs to $\mathcal{KL}_{\rm gen}$ if for all $s\in\mathbb{R}$, $t\mapsto\gamma(s,t)$ is increasing and for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ is decreasing.
+A function $\gamma:\mathbb{R}\times\mathbb{R}_{+}\mapsto\mathbb{R}$ belongs to $\mathcal{KL}_{\rm gen}$ if
+  • for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ is increasing;
+  • for all $s\in\mathbb{R}$, $t\mapsto\gamma(s,t)$ is decreasing.
+Example 1 (A pertinent $\mathcal{KL}_{\rm gen}$ for Leslie models).
+For the example introduced in Section 3, we compute a $\mathcal{KL}_{\rm gen}$ upper bound for $(X_{\ell},L,\varphi_{\ell})$.
+In this example, we construct a family of $\mathcal{KL}_{\rm gen}$ functions for which some of them will be $\mathcal{KL}_{\rm gen}$ upper bounds for $(X_{\ell},L,\varphi_{\ell})$.
+For this construction, we introduce the following auxiliary function.
+$H:\{(x,y)\in\mathbb{R}^{2}:x+y\neq 0\}\ni(u,v)\mapsto\dfrac{\lambda_{2}}{c}\dfrac{\dfrac{\lambda_{1}}{\lambda_{2}}v+u}{v+u}.$ (26)
+We observe that we have for all $x\in X_{\ell}$ (defined in (24)) and all $k\in\mathbb{N}$
+$\varphi_{\ell}(L^{k}x)=\dfrac{\lambda_{1}^{k+1}\mathfrak{a}(x)+\lambda_{2}^{k+1}\mathfrak{b}(x)}{c\left(\lambda_{1}^{k}\mathfrak{a}(x)+\lambda_{2}^{k}\mathfrak{b}(x)\right)}=\dfrac{\lambda_{2}}{c}\dfrac{\dfrac{\lambda_{1}}{\lambda_{2}}\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)}}{\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)}}=H\left(\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)},\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}\right).$ (27)
+From (27), we propose to construct a $\mathcal{KL}_{\rm gen}$ function $\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\gamma(s,t)$ from $H$.
+Indeed, the final purpose of the constructed $\mathcal{KL}_{\rm gen}$ function is to provide an upper bound of $\varphi_{\ell}\circ L^{k}$ on $X_{\ell}$.
+From the analysis of the sign of the partial derivatives of $H$, we remark that
+  • for all $v>0$ (resp. $v<0$), $u\mapsto H(u,v)$ strictly increases on $(-\infty,-v)$ and on $(-v,+\infty)$ (resp. strictly decreases on $(-\infty,-v)$ and on $(-v,+\infty)$);
+  • for all $u<0$ (resp. $u>0$), $v\mapsto H(u,v)$ strictly increases on $(-\infty,-u)$ and on $(-u,+\infty)$ (resp. strictly decreases on $(-\infty,-u)$ and on $(-u,+\infty)$).
+From these monotonicity properties, $H$ is a candidate to be a $\mathcal{KL}_{\rm gen}$ function on negative values of the state-variable $s$ and with positive second arguments.
+Therefore, we will consider $|\lambda_{1}/\lambda_{2}|^{t}$ rather than $(\lambda_{1}/\lambda_{2})^{t}$ as second argument in $H$.
+For positive values $s$, the $\mathcal{KL}_{\rm gen}$ function $\gamma(s,t)$ cannot be equal to $H(s,|\lambda_{1}/\lambda_{2}|^{t})$ and we will propose a strictly increasing affine function.
+Moreover, $H(s,|\lambda_{1}/\lambda_{2}|^{t})$ is not defined when $s=-|\lambda_{1}/\lambda_{2}|^{t}$ and we have to avoid those points.
+Finally, the main objective of this $\mathcal{KL}_{\rm gen}$ function is to provide an upper bound over $\varphi_{\ell}\circ L^{k}$ on $X_{\ell}$.
+Unfortunately, for all $s\in(-|\lambda_{1}/\lambda_{2}|^{t},0)$, $H(s,|\lambda_{1}/\lambda_{2}|^{t})$ is strictly negative and thus does not provide an upper bound for $\varphi_{\ell}\circ L^{k}$ which is strictly positive on $X_{\ell}$.
+We also remark that $s\mapsto H(s,|\lambda_{1}/\lambda_{2}|^{t})$ is not bounded from above on $(-\infty,-|\lambda_{1}/\lambda_{2}|^{t})$.
+Consequently, we artificially bound it, allowing only values $s$ smaller than $|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon$ for a given strictly positive $\varepsilon$.
+We will put constraints over $\varepsilon$ when we impose the $\mathcal{KL}_{\rm gen}$ function to be an upper bound of $\varphi_{\ell}\circ L^{k}$ on $X_{\ell}$.
+Finally, given $\varepsilon>0$, $H$ is exploited at $(s,|\lambda_{1}/\lambda_{2}|^{t})$ only when $s<-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon$.
+Otherwise, we use an affine function increasing in $s$ and in $|\lambda_{1}/\lambda_{2}|^{t}$ proved to be greater than $H(-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon,|\lambda_{1}/\lambda_{2}|^{t})$.
+Finally, we get a family of $\mathcal{KL}_{\rm gen}$ functions defined for all $\varepsilon>0$ as follows,
+$\gamma_{\varepsilon}:\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\left\{\begin{array}[]{lr}H\left(s,\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}\right)&\text{ if }s\leq-\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}-\varepsilon\\ \left(s+\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}+\varepsilon\right)+H\left(-\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}-\varepsilon,\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}\right)&\text{ otherwise }\end{array}\right$.
+(28)
+We prove in the Appendix that for all $\varepsilon>0$, $\gamma_{\varepsilon}\in\mathcal{KL}_{\rm gen}$.
+Note that we could replace $H(-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon,|\lambda_{1}/\lambda_{2}|^{t})$ by a strictly positive function increasing with respect to $|\lambda_{1}/\lambda_{2}|^{t}$ and greater than $H(-|\lambda_{1}/\lambda_{2}|^{t}-\varepsilon,|\lambda_{1}/\lambda_{2}|^{t}$.
+The current choice of affine function makes $\gamma_{\varepsilon}$ continuous, which is not mandatory to belong to $\mathcal{KL}_{\rm gen}$.
+Applying  (28) to the numerical data of Subsection 3.4, we have
+$\gamma_{\varepsilon}:\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\left\{\begin{array}[]{lr}4\dfrac{0.5^{t+1}+s}{0.5^{t}+s}&\text{ if }s\leq-(0.5)^{t}-\varepsilon\\ \\ s+0.5^{t}(1+6\varepsilon^{-1})+\varepsilon+4&\text{ if }s>-0.5^{t}-\varepsilon\end{array}\right$.
 Definition 3 ($\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds).
-A pair $(\gamma,\theta)\in\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ if and only if :
+A pair $(\gamma,\theta)\in\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ if and only if
 $\forall\,k\in\mathbb{N},\ \forall\,x\in X^{\mathrm{in}},\ \varphi(T^{k}(x))\leq\gamma(\theta(x),k)$
-Moreover, $(\gamma,\theta)$ is said to useful for $\nu$ such that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if the set:
+Moreover, $(\gamma,\theta)$ is said to be useful for $\nu$ such that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if the set
 $\left\{k\in{\mathrm{R}}_{\nu}^{s}:\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}\right\}$
 is nonempty.
-Following the notations introduced earlier, we have the equivalence:
-$\nu\in\Lambda\iff\sup_{k\in\mathbb{N}}\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$
-From Prop. 1 we have the equivalence, we also have:
-$\nu\in\Lambda\iff\limsup_{k\to+\infty}\nu_{k}=\limsup_{k\to+\infty}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))=\limsup_{k\to+\infty}\ \overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}\in\mathbb{R}\cup\{-\infty\}$
-We can establish a first result which states that the existence of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds is equivalent to the fact that the sequence $\nu$ is bounded from above.
+Example 2 (A useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound for $(X_{\ell},L,\varphi_{\ell})$).
+We now propose, as ${\mathrm{R}}_{\omega}^{s}\neq\emptyset$, a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound for the peak computation problem introduced in ($DPCP_{\ell}$).
+We show that for well-chosen $\varepsilon$, $\gamma_{\varepsilon}$ introduced in (28) combined with $\theta_{\ell}:=-|\mathfrak{b}/\mathfrak{a}|$ that belongs to $\operatorname{Fin}\left(X_{\ell}\right)$ provide a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound for $(X_{\ell},L,\varphi_{\ell})$.
+The choice of $\varepsilon$ depends on supremum of $\theta_{\ell}$ over $\mathfrak{a}^{-1}((-\infty,0))$ and over $\mathfrak{a}^{-1}((0,+\infty))$.
+Using the partial derivatives of $|\mathfrak{b}/\mathfrak{a}|$, we conclude that
+[…]
+For all $x\in X_{\ell}$ such that $\mathfrak{a}(x)<0$, we have $\theta_{\ell}(x)<-1$.
+Moreover, the set $\mathfrak{a}^{-1}((-\infty,0))$ is nonempty if and only if $\mathfrak{a}(x^{ul})<0$.
+Thus, the supremum of $\theta_{\ell}$ over $\mathfrak{a}^{-1}((-\infty,0))$ is strictly smaller than -1.
+For all $x\in X_{\ell}$ such that $0<\mathfrak{a}(x)<\mathfrak{b}(x)$, we also have $\theta_{\ell}(x)<-1$.
+However, if the set $\{x\in X_{\ell}:\mathfrak{a}(x)\geq\mathfrak{b}(x)\}$ (included in $\mathfrak{a}^{-1}((0,+\infty))$) is nonempty, for all $\varepsilon>0$, we can find $x\in X_{\ell}$ such that $0<\mathfrak{a}(x)<\mathfrak{b}(x)$ and $1-\varepsilon\leq\theta_{\ell}(x)$.
+Note that $\{x\in X_{\ell}:\mathfrak{a}(x)\geq\mathfrak{b}(x)\}$ is nonempty if and only if $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$ (or equivalently $\theta_{\ell}(x^{lu})\geq-1$).
+Finally, for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)\geq\mathfrak{b}(x)$, we have $\theta_{\ell}(x)\geq-1$.
+From  (21), we also have $|\lambda_{1}/\lambda_{2}|=\lambda_{1}/\lambda_{2}\geq\theta_{\ell}(x)$ for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)\geq\mathfrak{b}(x)$.
+As a corollary, we have
+[…]
+From the computation of $\overline{{\theta_{\ell}}_{X_{\ell}}}$, we define:
+[…]
+and
+$\overline{\varepsilon}=\min\{\varepsilon^{<0},\varepsilon^{>0}\}.$ (30)
+Finally, we have for all $x\in X_{\ell}$ and all $k\in\mathbb{N}$
+$\varphi_{\ell}(L^{k}(x))\leq\gamma_{\overline{\varepsilon}}\left(\theta_{\ell}(x),k\right)\text{ where }\theta_{\ell}=-|\mathfrak{b}/\mathfrak{a}|.$ (31)
+We insist on the fact that $\gamma_{\overline{\varepsilon}}\in\mathcal{KL}_{\rm gen}$ and $\theta_{\ell}\in\operatorname{Fin}\left(X_{\ell}\right)$.
+The details of the proof of (31) is provided in the Appendix.
+From sufficiently large values of $t$, $\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},t)=H(\overline{{\theta_{\ell}}_{X_{\ell}}},|\lambda_{1}/\lambda_{2}|^{t})$ and thus $\inf_{t\geq 0}\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},t)=H\left(\overline{{\theta_{\ell}}_{X_{\ell}}},0\right)=\lambda_{2}/c$.
+It follows that $\{k\in{\mathrm{R}}_{\omega}^{s}:\inf_{t\geq 0}\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},t)<\omega_{k}\}={\mathrm{R}}_{\omega}^{s}$ and finally the pair $(\gamma_{\overline{\varepsilon}},\theta_{\ell})$ is useful.
+Applying our computations to the numerical data of Subsection (3.4), we get
+$\mathfrak{a}(x^{lu})=95>0\text{ and }\mathfrak{a}(x^{ul})=-10<0$
+As $\mathfrak{b}(x^{lu})=55$ and $\mathfrak{b}(x^{ul})=25$, we have
+$\overline{{\theta_{\ell}}_{X_{\ell}}}=\max\{\theta_{\ell}(x^{lu}),\theta_{\ell}(x^{ul})\}=\max\{-11/19,-2.5\}=-11/19$
+and as $\mathfrak{a}(x^{lu})>\mathfrak{b}(x^{lu})$,
+$\overline{\varepsilon}=\min\{0.5(-1+2.5),0.5(-0.5+11/19)\}=3/76$
+Finally, we obtain:
+$\varphi_{\ell}({L_{e}}^{k}x)\leq\gamma_{\overline{\varepsilon}_{e}}(\theta_{\ell}(x),0.5^{t})=\left\{\begin{array}[]{lr}4\dfrac{0.5^{k+1}+\theta_{\ell}(x)}{0.5^{k}+\theta_{\ell}(x)}&\text{ if }\theta_{\ell}(x)\leq-(0.5)^{t}-\dfrac{3}{76}\\ \\ \theta_{\ell}(x)+153\times 0.5^{t}+\dfrac{307}{76}&\text{ if }\theta_{\ell}(x)>-0.5^{t}-\dfrac{3}{76}\end{array}\right.$.
+Following the notations introduced earlier, we have the following equivalence:
+$\nu_{\rm opt}<+\infty\iff\sup_{k\in\mathbb{N}}\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right).$
+We recall that:
+$\nu_{\rm opt}<+\infty\iff\limsup_{k\to+\infty}\nu_{k}=\limsup_{k\to+\infty}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))=\limsup_{k\to+\infty}\ \overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}\in\mathbb{R}\cup\{-\infty\}$
+We can establish the first result, which states that the existence of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds is equivalent to the fact that the sequence $\nu$ is bounded from above.
 This is an adaptation of classical results dealing with stability and $\mathcal{KL}$ upper bounds.
 Theorem 2 ($\mathbb{\Gamma}(\nu)$ to $\mathcal{KL}_{\rm gen}$).
-The sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ belongs to $\Lambda$ if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$.
+The sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ satisfies $\nu_{\rm opt}<+\infty$ if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$.
 If, moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ then $(\gamma,\theta)$ can be chosen to be useful for $\nu$.
@@
 As $s\mapsto\gamma(s,t)$ is increasing for all $t\geq 0$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\gamma(\theta(x),k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ and thus $\nu_{k}=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ for all $k\in\mathbb{N}$.
-Now, taking the limsup on both sides leads to $\limsup_{k\mapsto+\infty}\nu_{k}\leq\inf_{k\in\mathbb{N}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},0)<+\infty$ as $t\mapsto\gamma(s,t)$ is decreasing for all $s\in\mathbb{R}$.
-Suppose that $\nu=(\nu_{k})_{k\in\mathbb{N}}\in\Lambda$.
-From Prop.
-8, $\Gamma(\nu)$ is nonempty and let $(h,\beta)$ be in $\Gamma(\nu)$.
-Then, we define for all $s\in\mathbb{R}$ and $t\in\mathbb{R}_{+}$, $\gamma(s,t):=\min\{s,h(\beta^{t})\}$ and for all $x\in\mathbb{R}^{d}$, $\theta(x):=\sup_{k\in\mathbb{N}}\varphi(T^{k}(x))$.
-It is clear that $\gamma\in\mathcal{KL}_{\rm gen}$ as $\beta\in(0,1)$.
-Now as $\nu\in\Lambda$, $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\nu_{n}\in\mathbb{R}$.
-Moreover, we have $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{x\in X^{\mathrm{in}}}\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))=\overline{\theta_{X^{\mathrm{in}}}}$.
+Now, taking the limsup on both sides leads to
+$\limsup_{k\mapsto+\infty}\nu_{k}\leq\inf_{k\in\mathbb{N}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},0)<+\infty$
+as $t\mapsto\gamma(s,t)$ is decreasing for all $s\in\mathbb{R}$. This proves that $\nu_{\rm opt}<+\infty$.
+Suppose that $\nu=(\nu_{k})_{k\in\mathbb{N}}$ has a finite supremum.
+From the first statement of Th.
+1, $\mathbb{\Gamma}(\nu)$ is nonempty and let $(h,\beta)$ be in $\mathbb{\Gamma}(\nu)$.
+Then, we define for all $s\in\mathbb{R}$ and $t\in\mathbb{R}_{+}$,
+$\gamma(s,t):=\min\{s,h(\beta^{t})\}$
+and for all $x\in\mathbb{R}^{d}$,
+$\theta(x):=\sup_{k\in\mathbb{N}}\varphi(T^{k}(x)).$
+It is clear that $\gamma\in\mathcal{KL}_{\rm gen}$ as $\beta\in(0,1)$. Now by assumption, $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\nu_{n}\in\mathbb{R}$. Moreover, we have
+$\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{x\in X^{\mathrm{in}}}\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))=\overline{\theta_{X^{\mathrm{in}}}}.$
 By definition of $\theta$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\theta(x)$.
@@
 This is the same as for all $x\in X^{\mathrm{in}}$ and for all $k\in X^{\mathrm{in}}$, $\varphi(T^{k}(x))\leq h(\beta^{k})$.
-We conclude that $\varphi(T^{k}(x))\leq\min\{\theta(x),h(\beta^{k})\}$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$.
-Note that for a function $\gamma$ constructed as below, for all $s\in\mathbb{R}$, $\inf_{t\geq 0}\gamma(s,t)=\inf_{t\geq 0}\min\{s,h(\beta^{t})\}=\min\{s,h(0)\}$.
+We conclude that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$
+$\varphi(T^{k}(x))\leq\min\{\theta(x),h(\beta^{k})\}.$
+Note that for the function $\gamma$ constructed as above, for all $s\in\mathbb{R}$
+$\inf_{t\geq 0}\gamma(s,t)=\inf_{t\geq 0}\min\{s,h(\beta^{t})\}=\min\{s,h(0)\}.$
 This follows from the continuity of $h$, the fact $h$ is strictly increasing and $\beta\in(0,1)$.
-If moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then from Lemma 10, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$.
-We could define $\gamma$ using this function $h$.
+If moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then from the second statement of Th 1, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$.
+We can define $\gamma$ using this function $h$.
 Now, let $k\in\mathcal{S}(\nu,h)$.
-As $\nu_{k}\leq\nu_{\rm opt}=\overline{\theta_{X^{\mathrm{in}}}}$, we conclude that $\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)=\min\{\overline{\theta_{X^{\mathrm{in}}}},h(0)\}=h(0)$ which ends the proof.
-∎
-Example 4.
-Let us come back to the analysis of the peaks computation problem of Subsection 2.2.
-We propose to compute a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for the triplet $(X^{\mathrm{in}},T,\varphi)$ defined at Subsection 2.2.
-Following the proof of Th 5, one can define the $\mathcal{KL}_{\rm gen}$ function as $(s,t)\mapsto\min\{s,h(\beta^{t})\}$ where $(h,\beta)$ is a pair proposed at Example 2 and $\theta\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ as $\mathbb{R}^{2}\ni x\mapsto\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))$.
-Another possible pair consists in the $\mathcal{KL}_{\rm gen}$ function: ×$\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\min\{s,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$ and the constant function in $\operatorname{Fin}\left(X^{\mathrm{in}}\right)$: $\mathbb{R}^{2}\ni x\mapsto p^{2}/3$.
-We recall that $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\alpha)<0$.
-We can easily check that $\varphi(T^{k}(x))\leq\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(k+1))^{-1}\}$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
-Moreover, as $\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}=0$, there exists $k\in\mathbb{N}$ such that $\nu_{k}>\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$.
-In the proof of Th.
-5, we exploit Prop.
-8 and Prop.
-10 which link the nonemptiness of the sets $\mathbb{\Gamma}(u)$, $\mathcal{S}(u,h)$ and ${\mathrm{R}}_{u}^{s}$.
-We can reformulate Th 5 as Corollary 3.
+As $\nu_{k}\leq\nu_{\rm opt}=\overline{\theta_{X^{\mathrm{in}}}}$, we conclude that
+$\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)=\min\{\overline{\theta_{X^{\mathrm{in}}}},h(0)\}=h(0)$
+which ends the proof. ∎
+In the proof of Th. 2, we exploit Th. 1 which links the nonemptiness of the sets $\mathbb{\Gamma}(u)$, $\mathcal{S}(u,h)$ and ${\mathrm{R}}_{u}^{s}$. We can reformulate Th 2 as Corollary 1.
 Corollary 1.
@@
 If, moreover, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ then there exists a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$.
-In Th 5 and in Corollary 3, we construct a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ from a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
-Now, the question is to know whether is it possible to construct a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$.
-This is exactly what was done at Th 4 from the Sontag’s result.
-Thus, we provide a positive answer.
-The result is based on an extension of the Sontag’s result for the class $\mathcal{KL}_{\rm gen}$.
+In Th 2 and in Corollary 1, we construct a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ from a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
+Now, we have to check whether it is possible to construct a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$.
+Thus, we provide a positive response.
+The result is based on an extension of Sontag’s result for the class $\mathcal{KL}_{\rm gen}$.
 This permits to fix the value $\beta\in(0,1)$ to only consider problem on functions in $\Omega([0,1])$.
@@
 Lemma 1.
-Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a decreasing function and let $m\geq\inf_{\mathbb{R}_{+}}f$ ($m$ can be equal to $\infty$- ∞ if $\inf_{\mathbb{R}_{+}}f$ is).
-There exists a function $g:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that:
+Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a decreasing function and let $m\geq\inf_{\mathbb{R}_{+}}f$ ($m$ can be equal to -$\infty$ if $\inf_{\mathbb{R}_{+}}f$ is).
+There exists a function $g:\mathbb{R}_{+}\mapsto\mathbb{R}$ that is strictly decreasing and continuous such that:
   1. 1.
@@
 We denote for all couple of pairs $\{(x,y),(z,t)\}$ where $x,y,z,t\in\mathbb{R}$ by $\mathrm{Aff}\{(x,y),(z,t)\}$ the unique affine function $\ell:\mathbb{R}\mapsto\mathbb{R}$ such that $\ell(x)=y$ and $\ell(z)=t$.
-We define the functions $h_{0}$ and $h_{n}$ for all $n\in\mathbb{N}^{*}$ as follows :
+We define the functions $F_{0}$ and $F_{n}$ for all $n\in\mathbb{N}^{*}$ as follows:
 $F_{0}=f(0)\mathbb{1}_{[0,1]}\text{ and }F_{n}=f(n-1)\mathbb{1}_{\left[n;n+\frac{1}{2}\right]}+\mathrm{Aff}\left\{\left(n+\frac{1}{2};f(n-1)\right);\left(n+1;f(n)\right)\right\}\mathbb{1}_{\left[n+\frac{1}{2};n+1\right]}$
 Finally, we define, for all $x\in\mathbb{R}_{+}$, the function $F$:
-$F(x)=\sum_{n\in\mathbb{N}}F_{n}(x)+\dfrac{1}{x+1}\text{ and note that}\sum_{n\in\mathbb{N}}F_{n}(x)=\left\{\begin{array}[]{lr}f(0)&\text{ if }x\in[0,1]\\ f(\lfloor x\rfloor-1)&\text{ if }x\in[\lfloor x\rfloor,\lfloor x\rfloor+1/2]\\ \ell\left(x\right)&\text{ if }x\in(\lfloor x\rfloor+1/2,\lfloor x\rfloor+1)\end{array}\right.$
+[…]
 where $\ell=\mathrm{Aff}\left\{\left(\lfloor x\rfloor+1/2;f(\lfloor x\rfloor-1)\right);\left(\lfloor x\rfloor+1;f(\lfloor x\rfloor)\right)\right\}$.
 The function $\sum_{n\in\mathbb{N}}F_{n}$ is clearly continuous and decreasing (as $f$ does) on $\mathbb{R}_{+}$.
-The function $\sum_{n\in\mathbb{N}}F_{n}$ is not strictly decreasing but adding $x\mapsto(x+1)^{-1}$ makes $F$ to be strictly decreasing and continuous.
-We also have $f\leq\sum_{n\in\mathbb{N}}F_{n}$ as $f$ is decreasing.
-As $x\mapsto(x+1)^{-1}$ is strictly positive we have $f\leq F$.
+The function $\sum_{n\in\mathbb{N}}F_{n}$ is not strictly decreasing, but adding $x\mapsto(x+1)^{-1}$ makes $F$ strictly decreasing and continuous.
+We also have $f\leq\sum_{n\in\mathbb{N}}F_{n}$ because $f$ is decreasing.
+As $x\mapsto(x+1)^{-1}$ is strictly positive, we have $f\leq F$.
 Finally, we have $\inf_{\mathbb{R}_{+}}F=\inf_{\mathbb{R}_{+}}f$.
@@
 The reverse inequality holds as $f\leq F$.
-The function $F$ does not necessarily satisfy $\inf_{\mathbb{R}_{+}}h=m$. Hence, we define for all $x\in\mathbb{R}_{+}$:
+The function $F$ does not necessarily satisfy $\inf_{\mathbb{R}_{+}}F=m$. Hence, we define for all $x\in\mathbb{R}_{+}$ as
 $g(x)=\left\{\begin{array}[]{lr}\max\{F(x),(x+1)^{-1}+m\}&\text{ if }m>\inf_{\mathbb{R}_{+}}f\\ F(x)&\text{ if }m=\inf_{\mathbb{R}_{+}}f\end{array}\right$.
@@
 The function $g$ is clearly strictly decreasing and continuous and satisfies $f\leq g$ as $F$ does.
-As $m>\inf_{\mathbb{R}_{+}}f=\inf_{\mathbb{R}_{+}}F$, there exists $x_{0}$ such that $F(x)<m$ for all $x\geq x_{0}$ and then $g(x)=(x+1)^{-1}+m$ for all $x\geq x_{0}$.
+As $m>\inf_{\mathbb{R}_{+}}f=\inf_{\mathbb{R}_{+}}F$, there exists $x_{0}\in\mathbb{R}_{+}$ such that $F(x)<m$ for all $x\geq x_{0}$ and then $g(x)=(x+1)^{-1}+m$ for all $x\geq x_{0}$.
 This implies that $\inf_{\mathbb{R}_{+}}g=m$.
@@
 Let us define for all $r\in(0,1]$, $h(r)=\sigma(-\ln r)$ and $\lim_{x\to 0^{+}}h(x)=\lim_{y\to+\infty}\sigma(y)=m$.
-Then defining $h(0)=m$ makes $h$ continuous at 0.
+Then, defining $h(0)=m$ makes $h$ continuous at 0.
 Finally, $h$ belongs to $\Omega([0,1])$.
 ∎
-Proof of Th. 6.
-By assumption, there exist a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$.
+Proof of Th. 3.
+By assumption, there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$.
 Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
@@
 This leads to $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$.
-If there exists $k\in\mathbb{N}$ such that $\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}$, we can choose, as it is stated in Lemma 3, $h$ such that $h(0)=m\in(\inf_{t\in\mathbb{R}}\gamma(\overline{\psi},t),\nu_{k})$.
+If there exists $k\in\mathbb{N}$ such that $\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}$, we can choose, as it is stated in Lemma 2, $h$ such that $h(0)=m\in(\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t),\nu_{k})$.
 ∎
-We extend the class of $\mathcal{K}\mathcal{L}$ functions which classically serve for dynamical systems stability.
-In our context, we have a similar use and our $\mathcal{KL}_{\rm gen}$ class constitutes the main part of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds of $\varphi(T^{k}(x))$.
-We prove that such upper bounds exist if and only if $\nu\in\Lambda$.
-Moreover, if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, we are able to construct a function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ and such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
-Unfortunately, in practice, to prove that a pair $(\gamma,\theta)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ seems to be as difficult as proving that a pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(\nu)$.
-We thus propose a new type of Lyapunov functions for which the conditions does not involve the whole trajectories of the analyzed dynamical system.
+Example 3 (An element of $\mathbb{\Gamma}(\omega)$ from $(\gamma_{\overline{\varepsilon}},\theta_{\ell})$).
+From Th 3, we can build an element of $\mathbb{\Gamma}(\omega)$ from the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound obtained in Example 2.
+The situation is simpler in Example 2 as we get directly a scalar in $(0,1)$ from the expression of $\gamma_{\overline{\varepsilon}}$.
+Let
+$h:[0,1]\ni x\mapsto\left\{\begin{array}[]{lr}\displaystyle{\gamma_{\overline{\varepsilon}}\left(\overline{{\theta_{\ell}}_{X_{\ell}}},\dfrac{\ln(x)}{\ln(\beta)}\right)}&\text{ if }x\in(0,1]\\ \dfrac{\lambda_{2}}{c}&x=0\end{array}\right.\text{ and }\beta=\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|$
+where $\gamma_{\varepsilon}$ is defined in (28), $\overline{\varepsilon}$ and $\theta_{\ell}$ are defined in Example 2.
+As $t$ tends to $\infty$, we have $\overline{{\theta_{\ell}}_{X_{\ell}}}<-|\lambda_{1}/\lambda_{2}|^{t}-\overline{\varepsilon}$ for all sufficiently large $t$.
+It follows that $\lim_{x\mapsto 0^{+}}h(x)=\lim_{z\mapsto 0^{+}}H(\overline{{\theta_{\ell}}_{X_{\ell}}},z)=\lambda_{2}/c$.
+Thus, the function $h$ is continuous at 0.
+From (31), we can deduce that $(h,\beta)\in\mathbb{\Gamma}(\omega)$.
+Furthermore, we have, $h([0,1])=[\lambda_{2}/c,\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},0)]$ that contains all terms $\omega_{k}$ greater than the limit superior of $\omega$ and
+$\lambda_{2}/c,\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},0)]\ni y\mapsto\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(y)=\left\{\begin{array}[]{lr}\dfrac{\overline{{\theta_{\ell}}_{X_{\ell}}}(\lambda_{2}-cy)}{cy-\lambda_{1}}&\text{ if }y\leq\dfrac{\overline{\varepsilon}\lambda_{1}-\overline{{\theta_{\ell}}_{X_{\ell}}}(\lambda_{2}-\lambda_{1})}{c\overline{\varepsilon}}\\ \dfrac{\overline{\varepsilon}(cy-c\overline{{\theta_{\ell}}_{X_{\ell}}}-c\overline{\varepsilon}-\lambda_{2})}{c\overline{\varepsilon}+\lambda_{2}-\lambda_{1}}&\text{ if }y>\dfrac{\overline{\varepsilon}\lambda_{1}-\overline{{\theta_{\ell}}_{X_{\ell}}}(\lambda_{2}-\lambda_{1})}{c\overline{\varepsilon}}\end{array}\right.$.
+Then we can use Formula (4) that becomes, if $\omega_{k}>\lambda_{2}/c$:
+$\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\omega_{k}))}{\ln\left(\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|\right)}.$
+Applying those computations to the numerical data of Subsection 3.4, we get, as $\overline{{\theta_{\ell}}_{X_{\ell}}}=-11/19$
+$\left[4,\dfrac{11891}{76}\right]\ni y\mapsto\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(y)=\left\{\begin{array}[]{lr}-\dfrac{11}{19}\dfrac{(2-0.5y)}{0.5y+1}&\text{ if }y\leq 86\\ \\ \dfrac{y}{153}-\dfrac{263}{11628}&\text{ if }y>86\end{array}\right.$.
+Recall that $\omega_{0}=\varphi_{\ell}(x^{ul})=40/5=8$ and then
+$\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(8))}{-\ln(2)}=\dfrac{\ln(22/57)}{-\ln(2)}\approx 1.3735$
+This means that the maximum of terms $(\omega_{k})_{k}$ is either $\omega_{0}$ or $\omega_{1}$.
+Now as $\omega_{1}=\varphi_{\ell}(Lx^{lu})=42$.
+This implies that $\omega_{\rm opt}=42$.
+Moreover, we find that
+$\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(42))}{-\ln(2)}=\dfrac{\ln\left(\dfrac{11\times 19}{19\times 22}\right)}{-\ln(2)}=1.$
+4.2 Summary of the $\mathcal{KL}_{\rm gen}$ extension
+We extend the class of $\mathcal{K}\mathcal{L}$ functions that classically serve for dynamical systems stability.
+In our context, our $\mathcal{KL}_{\rm gen}$ class constitutes the main part of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$.
+We prove that such upper bounds exist if and only if $\nu$ has a finite supremum.
+Moreover, if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, we can construct a function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
+Unfortunately, in practice, proving that a pair $(\gamma,\theta)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ seems to be as difficult as proving that a pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(\nu)$.
+Thus, we propose a new type of Lyapunov function for which the conditions do not involve the entire trajectory of the analyzed dynamical system.
 5 Lyapunov type Functions for Discrete-time Systems Maximization
-Classically, Lyapunov functions are used as certificates of stability (see  [Ela05] or [KP01]).
-They bring a constructive approach to prove the stability of the dynamical system through converse Lyapunov theorems which prove that stability notions are equivalent to the existence of Lyapunov functions [Kha02].
-For some linear, polynomial or piecewise polynomial systems, Lyapunov functions can be computed from numerical optimization solvers based on linear or semidefinite programming (for example see [GH15] and references therein).
-Classical quadratic Lyapunov functions solve our problem when the dynamical system is stable and affine and the objective function is quadratic [Adj21].
-In Subsection 5.1, we generalize this approach theoretically.
-Then in Subsection 5.2, we define a more general class of Lyapunov functions.
-The existence of those functions does not ensure the stability of the system but just the finiteness of the limsup of the sequence $\nu$ defined at Eq.
+Classically, Lyapunov functions (i.e., satisfying conditions (12) and (13)) are used as certificates of stability (see  [Ela05] or [KP01]).
+They present a constructive approach to prove the stability of a dynamical system through converse Lyapunov theorems, which prove that stability notions are equivalent to the existence of Lyapunov functions [Kha02].
+For some linear, polynomial, or piecewise polynomial systems, Lyapunov functions can be computed from numerical optimization solvers based on linear or semidefinite programming (for example, see [GH15] and references therein).
+This numerical opportunity has been exploited for specific peak computation problems.
+Indeed, in [Adj25c] quadratic Lyapunov functions were used to solve our problem when the dynamical system is stable and affine and the objective function is quadratic.
+Next, in [Adj25a], we have proved that more general (local) Lyapunov functions can be used to solve peak computation problems.
+In these previous works, we applied direct-like Lyapunov theorems.
+In Lyapunov stability theory, after obtaining direct theorems, we consider converse theorems.
+The question raised in this case is, which stability notions imply the existence of a Lyapunov function?
+In our context, the converse theorem relies on the following question: which properties of classical Lyapunov functions can be removed to obtain an equivalence between the existence of a useful pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and the existence of a weaker version of the Lyapunov function?
+In this section, we reply to this answer.
+First, in Subsection 5.1, we define a more general class of Lyapunov functions called Opt-Lyapunov functions.
+The existence of these functions does not ensure the stability of the system, but only the finiteness of the limit superior of the sequence $\nu$ defined in Eq.
 (1).
-Last but not least, we etablish an equivalence theorem between the existence of those new Lyapunov functions and the existence of an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
-5.1 Classical Lyapunov Functions Theory
-5.1.1 Definition and Useful Facts
-[…]
-Definition 7 (Classical Lyapunov Functions).
-A function $V:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is said to a Lyapunov function if and only if :
-  1. 1.
-    There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that:
-    $\alpha_{1}\circ\|\cdot\|\leq V\leq\alpha_{2}\circ\|\cdot\|$
-  2. 2.
-    There exists $\lambda\in(0,1)$ such that
-    $V\circ T\leq\lambda V$
-[…]
-  1. 1.
-    there exists $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite111a function $g:X\mapsto\mathbb{R}_{+}$ where $X\subset\mathbb{R}^{d}$ and $0\in X$ is said to be positive definite (on $X$) if $g(0)=0$ and $g(x)>0$ for all $x\in X\backslash\{0\}$ and continuous such that $V\circ T\leq V-\alpha\circ\|\cdot\|$;
-  2. 2.
-    there exists $\rho:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite, continuous and such that $\mathrm{Id}-\rho$ is positive definite satisfying $V\circ T\leq\rho\circ V$.
-We will propose a proof for this equivalence later on the new class of Lyapunov functions used for the maximization problem.
-First, we remark that from the first statement of Def.
-7, $V$ is positive definite.
-If we require $V$ to be continuous and such that $\lim V(x)=+\infty$ as $\|x\|\to+\infty$ (see e.g.
-[Adj24]), then the first statement of Def.
-7 is actually equivalent to $V$ is postive definite.
-Second, if there exists a classical Lyapunov function for $T$, we $T$ must be null at zero [^2] Actually, if $T(e)=e$ for a nonzero vector, we could replace $V$ by $V(\cdot-e)$ and the lower and the upper bounds on $V(\cdot-e)$ by respectively $\alpha_{1}(\|\cdot-e\|)$ and $\alpha_{2}(\|\cdot-e\|)$.
-Finally, from the second statement of Def 7, we can define a kind of norm operator relative to a Lyapunov function.
-We have, for all Lyapunov function $V$ associated to the dynamic $T$, for all $k\in\mathbb{N}$:
-$N_{T^{k}}(V):=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{d}\\ x\neq 0\end{subarray}}\dfrac{V(T^{k}(x))}{V(x)}\in(0,1)$ (5)
-We will simply write $N_{T}(V)$ when $k=1$.
-By definition, we have for all $x\in\mathbb{R}^{d}$, $V(T(x))\leq N_{T}(V)V(x)$.
-Using this and the recurrence principle, we get, for all $k\in\mathbb{N}$, $N_{T^{k}}(V)\leq(N_{T}(V))^{k}$.
-5.1.2 A Direct Classical Lyapunov Theorem
-Let $f$ be a function from $\mathbb{R}^{d}$ to $\mathbb{R}$. We recall that the function from $\mathbb{R}_{+}$ to $\mathbb{R}$:
-$\alpha^{f}:s\mapsto\sup_{\|y\|\leq s}f(y)$ (6)
-is continuous when $f$ is continuous (see e.g.
-[Kha02] or extensions in [Adj24]).
-Obviously, the function $\alpha^{f}$ is increasing.
-Moreover, if $f(0)=0$ then $\alpha^{f}$ is nonnegative and satisfies $\alpha^{f}(0)=0$.
-Last but not least, we have for all $x\in\mathbb{R}^{d}$, $f(x)\leq\alpha^{f}(\|x\|)$.
-Theorem 7 (Classical Lyapunov to $\mathbb{\Gamma}(\nu)$).
-Assume that $\varphi$ is continuous and there exists a Lyapunov function for $T$.
-Then there exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $h(0)=0$.
-Thus, if Assumption 2 holds, $h$ is useful for $\nu$.
-Proof.
-If there exists a Lyapunov function $V$ for $T$, there exists $\alpha_{1}\in\mathcal{K}_{\infty}$ such that $\alpha_{1}\circ\|\cdot\|\leq V$.
-Recall that we have assumed that $\varphi(0)=0$.
-With the continuity assumption of $\varphi$, we conclude that $\alpha^{\varphi}$ (defined at Eq.
-(6)) is nonnegative, increasing, continuous and is equal to 0 at 0.
-Two small issues : $\alpha^{\varphi}$ is not guaranteed to be strictly increasing and surjective from $\mathbb{R}_{+}$ to $\mathbb{R}_{+}$ (the limit of $\alpha^{\varphi}$ at +$\infty$+ ∞ can be finite).
-We thus add to $\alpha^{\varphi}$ a function $\psi\in\mathcal{K}_{\infty}$ and we have for $y\in\mathbb{R}^{d}$, $\varphi(y)\leq(\alpha^{\varphi}+\psi)(\|y\|)$ and $\alpha:=\alpha^{\varphi}+\psi$ belongs to $\mathcal{K}_{\infty}$.
-From $\alpha_{1}\circ\|\cdot\|\leq V$ and $\alpha_{1}\in\mathcal{K}_{\infty}$, we get $\varphi\leq\alpha\circ\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\circ V$ as $\alpha$ is increasing.
-Moreover, we have, for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $V(T^{k}(x))\leq N_{T^{k}}(V)V(x)\leq(N_{T}(V))^{k}V(x)\leq(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}$.
-As $\alpha$ and ${{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}$ are increasing, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\left(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}\right)\right)$.
-We thus define $h:x\mapsto\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}(x\overline{V_{X^{\mathrm{in}}}})\right)$ and $\beta=N_{T}(V)$.
-We have $h(0)=0$ and we conclude that $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$ when Assumption 2 holds.
-∎
-The goal of extending the concept of Lyapunov functions is to weaken the hypothesis of Th 7 by removing the continuity and the positivity assumptions on $\varphi$.
-The extension also allows to prove the equivalence between the existence of useful pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and the existence of this new type of Lyapunov functions.
-5.2 Opt-Lyapunov Functions for Discrete-time Systems Maximization
+Then, in Subsection 5.2, we discuss the notion of usefulness and how it constrains the Opt-Lyapunov functions.
+In the classical case, usefulness constrains the sequence $\nu$ to have some positive terms.
+We replace the functions $\alpha_{1},\alpha_{\varphi}$ in (16) by a certificate of compatibility.
+Finally, in Subsection 5.3, we establish an equivalence theorem between the existence of those Opt-Lyapunov functions and the existence of an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
+5.1 Generalized Lyapunov Functions for Discrete-time Systems Maximization
 5.1.1 Definitions and Useful Facts
@@
 We denote by $\operatorname{dom}(f)$ the domain of a extended real valued function.
+We extend the class of classical Lyapunov functions as follows: An Opt-Lyapunov function is still nonnegative but can be either strictly positive or can have several zeros and can take infinite values.
+The supremum of an Opt-Lyapunov function on the initial conditions set is non-null but smaller than 1.
+Opt-Lyapunov functions keep the decrease condition of classical Lyapunov functions.
 Definition 4 (Opt-Lyapunov functions).
-A function $V$ is said to be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ if and only if:
+A function $V$ is said to be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ if and only if
   1. 1.
@@
     $V\circ T\leq\lambda V$ for some $\lambda\in(0,1)$.
-Example 5.
-Let us consider the dynamical system of the example of Subsection 2.2. Recall that $\mu\in(0,1/3)$ and the set of initial conditions is :
-$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}$
-and the dynamics is the following:
-$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}$
-The function defined, for all $x=(x_{1},x_{2})^{\intercal}\in\mathbb{R}^{2}$ as follows:
-$V(x)=\left\{\begin{array}[]{cr}\dfrac{2\mu^{2}}{x_{1}x_{2}}&\text{ if }x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}\\ \\ 0&\text{ otherwise}\end{array}\right.$
-is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
-Let us prove that it is true.
-Let $x\in\mathbb{R}^{2}$.
-If $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, then $x=(2,1)^{\intercal}y$ with $y\in\mathbb{R}^{*}$.
-First, the coordinates of $x$ has the same sign and so $V$ is nonnegative.
-The fact that $V(x)$ is nonnegative when for $x\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $x=0$ is obvious.
-Second, let $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ and let $y\neq 0$ such that $x=(2,1)^{\intercal}y$.
-Then, $V(T(x))=V((3y,3y/2)^{\intercal})=(4\mu^{2})/(9y^{2})$ and $V(x)=\mu^{2}/y^{2}$ then $V(T(x))/V(x)=(4y^{2})/(9y^{2})=4/9$.
-Any value in $[4/9,1)$ can be chosen.
-Let us take $5/9$.
-As we have $T(z)\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ if and only if $z\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, we conclude that for all $z\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $z=0$, we have $V(T(z))=V(z)=0$.
-We conclude that $V(T(z))\leq 5V(z)/9$ for all $z\in\mathbb{R}^{2}$.
-Finally, since $\mu>0$, we have $\max_{x\in X^{\mathrm{in}}}V(x)=\max_{y\in[\mu,1/2]}(2\mu^{2})/(2y^{2})^{-1}=(2\mu^{2})(2\mu^{2})^{-1}=1$.
+Example 4 (Opt-Lyapunov function for the Leslie Model).
+Returning to the Leslie model example introduced in Section 3.
+As $\lambda_{2}>1$, the system is unstable, that is, $\|x_{k}\|$ goes to $\infty$ as $k$ tends to $\infty$.
+A classical Lyapunov function does not exist for this system.
+However, an Opt-Lyapunov function exists.
+We propose, for $C>0$,
+$\mathbb{R}^{2}\ni x\mapsto V_{\ell}(x)=\left\{\begin{array}[]{cr}\dfrac{C}{x_{1}^{2}+x_{2}^{2}}&\text{ if }x_{1},x_{2}>0\\ +\infty&\text{ otherwise}\end{array}\right.$. (32)
+Let us take $x\in\mathbb{R}^{2}$.
+If $Lx$ has a nonpositive coordinate (and thus $V_{\ell}(Lx)=+\infty$), then either $x_{1}\leq 0$ or $x_{2}\leq 0$ yielding to $V_{\ell}(x)=+\infty$.
+Therefore, $V_{\ell}(Lx)\leq\lambda V_{\ell}(x)$ for any $\lambda\in(0,1)$.
+Now, assume that both coordinates of $Lx$ are strictly positive.
+This implies that $x_{1}>0$.
+We can have $x_{2}\leq 0$.
+In this case, $V_{\ell}(x)=+\infty$ and $V_{\ell}(Lx)\leq\lambda V_{\ell}(x)$ for any $\lambda\in(0,1)$.
+Finally, suppose that $Lx$ and $x$ have strictly positive coordinates.
+We have from  (20)
+$\dfrac{V_{\ell}(Lx)}{V_{\ell}(x)}=\dfrac{x_{1}^{2}+x_{2}^{2}}{(ax_{1}+bx_{2})^{2}+cx_{1}^{2}}=\dfrac{x_{1}^{2}+x_{2}^{2}}{(a^{2}+c^{2})x_{1}^{2}+b^{2}x_{2}^{2}+2abx_{1}x_{2}}<\dfrac{x_{1}^{2}+x_{2}^{2}}{(a^{2}+c^{2})x_{1}^{2}+b^{2}x_{2}^{2}}<1.$
+This inequality is not sufficient to prove condition (3) of Definition 4 as the supremum of $V_{\ell}(Lx)/V_{\ell}(x)$ might be equal to 1.
+We will compute the exact value of the supremum of $V_{\ell}(Lx)/V_{\ell}(x)$ over the open positive orthant of $\mathbb{R}^{2}$.
+This supremum is strictly smaller than 1.
+We can adjust the constant $C$ to have $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}\in(0,1]$ even if we will see that the finiteness of $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}$ is sufficient to obtain the second statement of Definition 4.
+As $V_{\ell}$ is marginally strictly decreasing we have $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}=C/(l_{1}^{2}+l_{2}^{2})$.
+Thus any $C\in(0,l_{1}^{2}+l_{2}^{2}]$ can be chosen.
+Applying those computations to the numerical data of Subsection 3.4, we get for all $x\in(\mathbb{R}_{+}^{*})^{2}$,
+$V_{\ell}(x)=125/x_{1}^{2}+x_{2}^{2}$
+where we choose $C=l_{1}^{2}+l_{2}^{2}$ to obtain $\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}}=1$. ∎
 Proposition 2.
-If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ then :
+If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ then
   1. 1.
@@
 If $V(x)=+\infty$, there is nothing to prove.
-Now, suppose that $V(x)$ is finite.
+Suppose that $V(x)$ is finite.
 From the third statement, $0\leq V(T(x))=V(x)\leq\lambda V(x)$ for some $\lambda\in(0,1)$.
@@
 Corollary 2.
-If $X^{\mathrm{in}}$ is included in the fixed point set of $T$ then a function $V:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that $V\circ T\leq\lambda V$ for all some $\lambda\in(0,1)$ cannot satisfy $\overline{V_{X^{\mathrm{in}}}}>0$.
-Corollary 4 extends the fact that if a classical Lyapunov function (as defined in Def 7) exists for a self-map $T$, $T$ cannot have non-zero fixed point.
+If $X^{\mathrm{in}}$ is included in the fixed point set of $T$ then a function $V:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that $V\circ T\leq\lambda V$ for some $\lambda\in(0,1)$ cannot satisfy $\overline{V_{X^{\mathrm{in}}}}>0$.
+Corollary 2 extends the fact that if a classical Lyapunov function exists for a self-map $T$, $T$ cannot have a nonzero fixed point.
 If an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ exists, $X^{\mathrm{in}}$ contains a vector $x$ such that $x\neq f(x)$.
-Prop. 15 shows that the set of Opt-Lyapunov functions contains the classical Lyapunov functions.
-Proposition 15 (Classical is Opt).
-If $V$ is a classical Lyapunov function for $T$ and $X^{\mathrm{in}}$ is not reduced to $\{0\}$, then $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
-Actually, we can replace equivalently the second statement of Def.
+In fact, we can equivalently replace the second statement of Def.
 4 by $\overline{V_{X^{\mathrm{in}}}}$ is finite and strictly positive.
-Indeed, with this condition, we recover an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ with $\overline{W_{X^{\mathrm{in}}}}\in(0,1]$ by taking $W:=V/\overline{V}$.
+Indeed, under this condition, we recover an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ with $\overline{W_{X^{\mathrm{in}}}}\in(0,1]$ by taking $W:=V/\overline{V_{X^{\mathrm{in}}}}$.
 Proposition 3.
-The existence of a function $W:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that :
+The existence of a function $W:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that
   1. 1.
@@
     • $\rho(x)<x$ for all $x\in\mathbb{R}_{+}^{*}$.
-is equivalent to the existence of a Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
+is equivalent to the existence of an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
 Lemma 3.
@@
 Let $f\in\mathcal{K}_{\infty}$ verifying for all $x>0$, $f(x)>x$.
-Then, there exists a function $g\in\mathcal{K}$ which satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
-Lemma 4 is a slighty modified version of [Kel14, Lemma 25] for which $f-\mathrm{Id}$ is supposed to be in $\mathcal{K}_{\infty}$.
-To be self-contained, we propose in Appendix a proof inspired by the one provided in [Kel14].
+Then, there exists a function $g\in\mathcal{K}$ that satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
+Lemma 3 is a slightly modified version of [Kel14, Lemma 25], for which $f-\mathrm{Id}$ is supposed to be in $\mathcal{K}_{\infty}$.
+To be self-contained, we propose in the Appendix a proof inspired by the one provided in [Kel14].
 Lemma 4.
 Let $\lambda\in(0,1)$. Let $f\in\mathcal{K}$ verifying for all $x>0$, $f(x)<x$. Then, there exists a function $g\in\mathcal{K}$ satisfying $g(f(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
-Again a proof of Lemma 5 is given in [Kel14]. A proof is provided in Appendix for the convenience of the reader.
-Proof of Prop. 16.
-Suppose that such $W$ and $\rho$ exist.
-Let $\lambda\in(0,1)$.
+Again, a proof of Lemma 4 is given in [Kel14]. A proof is provided in the Appendix for the convenience of the reader.
+Proof of Prop. 3.
+Suppose that such $W$ and $\rho$ exist, then let $\lambda\in(0,1)$.
 From Lemma 4, there exists $g$ such that $g(\rho(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
@@
 Then as $g$ is increasing, we have $g(W(T(x)))\leq g(\rho(W(x))=\lambda g(W(x))$ for all $x\in\mathbb{R}^{d}$.
-Now, as $g$ is strictly increasing there exists some $x\in X^{\mathrm{in}}$ such that $g(W(x))$ is strictly positive and $\overline{g\circ W_{X^{\mathrm{in}}}}$ is strictly positive and finite.
+Now, as $g$ is strictly increasing, there exists some $x\in X^{\mathrm{in}}$ such that $g(W(x))$ is strictly positive and $\overline{g\circ W_{X^{\mathrm{in}}}}$ is strictly positive and finite.
 Finally, $(g\circ W)/\overline{g\circ W_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
 The reverse implication is obvious with $\rho(x)=\lambda x$ for all $x\in\mathbb{R}_{+}$. ∎
+Prop. 4 shows that the set of Opt-Lyapunov functions contains the set of classical Lyapunov functions.
+Proposition 4 (Classical is Opt).
+If $V$ is a classical Lyapunov function for $T$ and $X^{\mathrm{in}}$ is bounded and not reduced to $\{0\}$, then $V/\overline{V_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
+Proof.
+Let $V$ be a classical Lyapunov function for $T$.
+If $X^{\mathrm{in}}$ is not reduced to $\{0\}$, as $V$ is positive definite, $V(x)>0$ for some $x\in X^{\mathrm{in}}$.
+From  (12), if $X^{\mathrm{in}}$ is bounded, $\overline{V_{X^{\mathrm{in}}}}$ is finite.
+This implies that $V/\overline{V_{X^{\mathrm{in}}}}$ satisfies the second statement in Definition 4.
+Moreover, it is obvious that $V/\overline{V_{X^{\mathrm{in}}}}$ is positive definite and so verifies the first statement of Definition 4.
+It is easy to see that $V/\overline{V_{X^{\mathrm{in}}}}$ satisfies the third statement of Definition 4 from (13) and $\overline{V_{X^{\mathrm{in}}}}>0$.
+∎
 5.1.2 Operator Ratios From Proper Positive Semidefinite Functions
-We extend the concept of operator ratios (defined at Eq.
-(5)) for which the arguments are positive semidefinite functions rather than positive definite functions.
-The arguments can also be functions which take infinite values.
-So, let us denote by $\mathcal{PSD}$ the set of nonzero proper positive semidefinite functions from $\mathbb{R}^{d}$ to $[0,+\infty]$ i.e.
-:
+We extend the concept of operator ratios developed in [Adj25a] (specialized at Eq.
+(17) for Lyapunov functions) for which the arguments are positive semidefinite functions rather than positive definite functions.
+The arguments can also be functions that take infinite values.
+Therefore, let us denote by $\mathcal{PSD}$ the set of nonzero proper positive semidefinite functions from $\mathbb{R}^{d}$ to $0,+\infty]$, that is,
 $\mathcal{PSD}:=\{f:\mathbb{R}^{d}\mapsto[0,+\infty]:\exists\,x\in\mathbb{R}^{d},\ f(x)\in\mathbb{R}_{+}^{*}\}$ (33)
-We are able to extend the operator ratio function to $\mathcal{PSD}$ functions.
+We can extend the operator ratio function to $\mathcal{PSD}$ functions.
 Definition 5 (Operator Ratios PSD Based).
-Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $k\in\mathbb{N}$. We associate to $U$ and $k$ the operator ratio function $\overline{N_{U^{k}}}:\mathcal{PSD}\mapsto[0,+\infty]$ defined as follows:
-$\overline{N_{U^{k}}}(P):=\sup_{P(x)\in\mathbb{R}_{+}^{*}}\dfrac{P(U^{k}(x))}{P(x)}$
-As for $N_{U^{k}}$, we will simply write $\overline{N_{U}}$ for $k=1$.
-Recall that $\mathrm{Id}$ denotes the identity function $\mathbb{R}^{d}\ni x:\mapsto x$. It is straightforward to see that, if $P\in\mathcal{PSD}$, $\overline{N_{\mathrm{Id}}}(P)=1$.
-Example 6.
-We proved at Example 5 that:
-$\overline{N_{T}}(V)=\sup_{\{x\in\mathbb{R}^{2}:V(x)>0\}}\dfrac{V(T(x))}{V(x)}=\max_{x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}}\dfrac{V(T(x))}{V(x)}=4/9.$
-The operator function is well-defined but the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$ can fail to be true as we can see at Example 7.
+Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. We associate to $U$ the operator ratio function $\overline{N_{U}}:\mathcal{PSD}\mapsto[0,+\infty]$ defined as follows:
+$\mathcal{PSD}\ni P\mapsto\overline{N_{U}}(P):=\sup_{P(x)\in\mathbb{R}_{+}^{*}}\dfrac{P(U(x))}{P(x)}$
+Recall that $\mathrm{Id}$ denotes the identity function $\mathbb{R}^{d}\ni x:\mapsto x$. It is straightforward to see that if $P\in\mathcal{PSD}$, $\overline{N_{\mathrm{Id}}}(P)=1$.
+The operator function is well-defined, but the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$ can fail to be true, as shown in Example 5.
 Example 5 (A counterexample to the power commutation).
-Let us consider the function $U:x\mapsto-x$ and $P:x\mapsto\max\{x,0\}$.
+Let us consider the functions $U:x\mapsto-x$ and $P:x\mapsto\max\{x,0\}$.
 We have $\overline{N_{U}}(P)=\sup_{P(x)>0}P(U(x))/P(x)=\sup_{x>0}P(-x)/x=\sup_{x>0}0/x=0$ whereas $\overline{N_{U^{2}}}(P)=\sup_{x>0}P(x)/x=\sup_{x>0}x/x=1$.
@@
 To obtain the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$, we have to consider a smaller class of operators $P$ and we introduce for a self-map $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, the set $\mathcal{D}\left(U\right)$ of operators defined as follows:
-$\mathcal{D}\left(U\right):=\{P\in\mathcal{PSD}:\forall\,x\in\mathbb{R}^{d},\ P(U(x))=0\implies P(U^{2}(x))=0\}$ (8)
+$\mathcal{D}\left(U\right):=\{P\in\mathcal{PSD}:P(U(x))=0\implies P(U^{2}(x))=0\}$ (34)
 We remark that $P\in\mathcal{D}\left(U\right)$ if and only if $P(U^{2}(x))\in(0,+\infty]\implies P(U(x))\in(0,+\infty]$.
@@
 Example 6 ($P$ not in $\mathcal{D}\left(U\right)$).
-Let consider $U(x)=-x$ and $P(x)=\max\{x,0\}$. Let $x$ be a strictly positive real. Then, $P(U(x))=P(-x)=0$. However, $P(U^{2}(x))=P(x)=x>0$. So $P\notin\mathcal{D}\left(U\right)$.
+The functions $U$ and $P$ are the ones defined in Example 5.
+Let $x>0$ be strictly positive.
+Then, P(U(x))=P(x)=0P(U(x))=P(-x)=0.
+However, $P(U^{2}(x))=P(x)=x>0$.
+So $P\notin\mathcal{D}\left(U\right)$.
 Proposition 5.
@@
 Proof.
-For $k=0$, the result holds as $\overline{N_{\mathrm{Id}}}(P)=1$ if there exists $x\in\mathbb{R}^{d}$ such that $P(x)>0$ and 0 otherwise.
-If $\overline{N_{U}}(P)$ is finite $(\overline{N_{U}}(P))^{0}$ is defined and equal to 1.
+For $k=0$, the result holds as $\overline{N_{\mathrm{Id}}}(P)=1$ and there exists $x\in\mathbb{R}^{d}$ such that $P(x)>0$.
+If $\overline{N_{U}}(P)$ is finite, $(\overline{N_{U}}(P))^{0}$ is defined and equal to 1.
 Now let $k\in\mathbb{N}^{*}$.
-If $\overline{N_{U}}(P)=+\infty$, the inequality is obviously true.
+If $\overline{N_{U}}(P)=+\infty$, then the inequality is obviously true.
 Suppose now that $\overline{N_{U}}(P)<+\infty$.
@@
 Proof.
-First $V^{-1}(\mathbb{R}_{+}^{*})$ is nonempty from the second statement of Def.
-8.
+First, $V^{-1}(\mathbb{R}_{+}^{*})$ is nonempty from the second statement of Def.
+4.
 The third statement implies that $V(T^{2}(x))\leq\lambda V(T(x))$ for all $x\in\mathbb{R}^{d}$ for some $\lambda\in(0,1)$ and thus $V(T^{2}(x))>0$ implies that $V(T(x))>0$.
@@
 ∎
-To prove that a function $V\in\mathcal{PSD}$ is an Opt-Lyapunov for $X^{\mathrm{in}},T)$, we could use any value greater $\overline{N_{T}}(V)$.
-The computation of $\overline{N_{T}}(V)$ ensures to take the best value possible.
-It is important in the context of using the function $\mathfrak{F}_{u}(k,h,\beta)$ with an Opt-Lyapunov function since $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is decreasing (Prop.
-12).
+Example 7 (Operator ratio for $V_{\ell}$).
+We compute the operator ratio of $L$ at the Opt-Lyapunov function $V_{\ell}$ proposed in Equation (32).
+$\overline{N_{L}}(V_{\ell})=\sup_{x\in{V_{\ell}}^{-1}(\mathbb{R}_{+}^{*})}\dfrac{V_{\ell}(Lx)}{V_{\ell}(x)}=\sup_{x\in V_{\ell}^{-1}(\mathbb{R}_{+}^{*})}\dfrac{x^{\intercal}x}{x^{\intercal}L^{\intercal}Lx}=\left(\inf_{x_{1}>0,\ x_{2}>0}\dfrac{x^{\intercal}L^{\intercal}Lx}{x^{\intercal}x}\right)^{-1}$
+Finally, we can use a characterization of copositive matrices (see for example  [HUS10] and references therein) to compute $\overline{N_{L}}{V_{\ell}}$. Indeed, we have:
+$\inf_{\begin{subarray}{c}x_{1}>0\\ x_{2}>0\end{subarray}}\dfrac{x^{\intercal}L^{\intercal}Lx}{x^{\intercal}x}=\sup\left\{\lambda\geq 0:x^{\intercal}L^{\intercal}Lx-\lambda x^{\intercal}x\geq 0\text{ on }(\mathbb{R}_{+}^{*})^{2}\right\}=\sup\left\{\lambda\geq 0:L^{\intercal}L-\lambda\mathrm{Id}\text{ is copositive}\right\}$
+From [HUS10, Proposition 2.1], $L^{\intercal}Lx-\lambda\mathrm{Id}=\begin{pmatrix}a^{2}+c^{2}-\lambda&ab\\ ab&b^{2}-\lambda\end{pmatrix}$ is copositive if and only if $a^{2}+c^{2}-\lambda\geq 0$, $b^{2}-\lambda\geq 0$ and $\sqrt{(a^{2}+c^{2}-\lambda)(b^{2}-\lambda)}+ab\geq 0$.
+Since $a\geq 1$ and $b>1$, the third inequality is a consequence of the first two conditions.
+We conclude that:
+$\overline{N_{L}}(V)=\sup_{x\in V^{-1}(\mathbb{R}_{+}^{*})}\dfrac{x^{\intercal}x}{x^{\intercal}L^{\intercal}Lx}=\dfrac{1}{\min\{a^{2}+c^{2},b^{2}\}}$
+Applying the computations to the numerical data of Subsection 3.4, we have
+$\overline{N_{L}}(V_{\ell})=\dfrac{4}{5}.$
+To prove that a function $V\in\mathcal{PSD}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, we could use any value greater $\overline{N_{T}}(V)$.
+The computation of $\overline{N_{T}}(V)$ ensures that the best value possible is obtained.
+This is important in the context of using the function $\mathfrak{F}_{\nu}(k,h,\beta)$ with an Opt-Lyapunov function because $\beta\mapsto\mathfrak{F}_{\nu}(k,h,\beta)$ is decreasing.
+For classical Lyapunov functions, the extended version (on positive semidefinite functions) of operator ratios (Definition 5) coincides with the positive definite version (recalled in (17)).
 Proposition 7.
@@
 Proof.
-If there exists a Lyapunov function for $U$, then $U(0)=0$.
-From the double inequalities $\alpha\circ\|\cdot\|\leq V\leq\beta\circ\|\cdot\|$ and thus $V(x)=0$ iff $x=0$.
-Thus if $V(U^{2}(x))>0$ then $U^{2}(x)\neq 0$ and thus $U(x)\neq 0$ hence $V(U(x))>0$ and $V\in\mathcal{D}\left(U\right)$.
-The equality between $\overline{N_{U}}(V)$ and $N_{U}(V)$ follows readily from the fact that $\{y\in\mathbb{R}^{d}:V(y)>0\}=\mathbb{R}^{d}\backslash\{0\}$.
-∎
-5.2.3 $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility and $(X^{\mathrm{in}},T,\varphi)$-compatibility
-In the proof of Th.
-7, the construction of $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ from a classical Lyapunov function $V$ is based on the existence of a function $\alpha\in\mathcal{K}_{\infty}$ such that $\varphi\leq\alpha\circ V$.
-The existence of this latter comes from the fact that in Th.
-7, $\varphi$ is continuous and, by definition of $V$, $\alpha^{\prime}\circ\|\cdot\|\leq V$ for some $\alpha^{\prime}\in\mathcal{K}_{\infty}$.
-However, when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$ is not necessarily positive definite and we cannot guarantee the existence of $\alpha\in\mathcal{K}_{\infty}$ such that $\alpha\circ\varphi\leq V$.
-We have to use a different class of function to find a similar relation between $\varphi$ and $V$.
-First, we propose to restrict the interval where a function $\alpha$ should be continuous and strictly increasing.
+[…]
+5.2 $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility and $(X^{\mathrm{in}},T,\varphi)$-compatibility
+The construction of $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ from a classical Lyapunov function $V$ is based on the existence of a function $\alpha\in\mathcal{K}_{\infty}$ such that $\varphi\leq\alpha\circ V$.
+The existence of this latter function $\alpha$ is either assumed (with $\alpha=\mathrm{Id}$) in Equation (16) or in Equation (15) comes from the continuity of $\varphi$ and the definition of $V$ (see Equation (12)).
+However, when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$ is neither positive definite nor continuous.
+Thus, we cannot guarantee the existence of $\alpha\in\mathcal{K}_{\infty}$ such that $\alpha\circ\varphi\leq V$.
+We must use a different class of functions to find a similar relation between $\varphi$ and $V$.
+First, we propose to restrict the interval in which a function $\alpha$ should be continuous and strictly increasing.
 Indeed, the analysis only concerns the values $\{\varphi(T^{k}(x)),k\in\mathbb{N},x\in X^{\mathrm{in}}\}$ and $\{V(x),x\in X^{\mathrm{in}}\}$.
@@
     For all $k\in\mathbb{N}$, $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$.
-The values, if they are finite, $\limsup_{n\in\mathbb{N}}\nu_{n}$ and $\liminf_{n\in\mathbb{N}}\nu_{n}$ also belong to $I_{X^{\mathrm{in}},T}^{\varphi}$.
+If the values are finite, $\limsup_{n\in\mathbb{N}}\nu_{n}$ and $\liminf_{n\in\mathbb{N}}\nu_{n}$ also belong to $I_{X^{\mathrm{in}},T}^{\varphi}$.
   2. 2.
@@
 The values, if they are finite, $\limsup_{k\in\mathbb{N}}\nu_{k}$ and $\liminf_{k\in\mathbb{N}}\nu_{k}$ are limits of subsequences of $(\nu_{k})_{k\in\mathbb{N}}$ and they belongs to $I_{X^{\mathrm{in}},T}^{\varphi}$.
-From Prop.
-1, if $\limsup_{k\in\mathbb{N}}\nu_{k}$ (resp.
+From the definition of $\limsup$ (resp.
+$\liminf$), if $\limsup_{k\in\mathbb{N}}\nu_{k}$ (resp.
 $\liminf_{k\in\mathbb{N}}\nu_{k}$) is infinite, the value $\sup_{k\in\mathbb{N}}\nu_{k}$ (resp.
-$\inf_{k\in\mathbb{N}}\nu_{k}$ is equal to +$\infty$+ ∞ (resp.
-$\infty$- ∞).
-Now, if one of the value $\inf_{k\in\mathbb{N}}\nu_{k}$ or $\sup_{k\in\mathbb{N}}\nu_{k}$ is not finite, the infinity passes to the related bound of $I_{X^{\mathrm{in}},T}^{\varphi}$.
-Now if those values are both finite, we can extract sequences in $I_{X^{\mathrm{in}},T}^{\varphi}$ which converge respectively to $\inf_{m\in\mathbb{N}}\nu_{m}$ and $\sup_{m\in\mathbb{N}}\nu_{m}$.
-For example, we can define $\varphi(T^{k_{n}}(x_{k_{n}}))\subset I_{X^{\mathrm{in}},T}^{\varphi}$ such that for all $n\in\mathbb{N}$, $\inf_{m\in\mathbb{N}}\nu_{m}-1/n\leq\nu_{k_{n}}-1/n\leq\varphi(T^{k_{n}}(x_{k_{n}}))\leq\nu_{k_{n}}\leq\inf_{m\in\mathbb{N}}\nu_{m}+1/n$.
-This sequence converge to $\inf_{m\in\mathbb{N}}\nu_{m}$.
-∎
+$\inf_{k\in\mathbb{N}}\nu_{k}$ is equal to $\infty$ (resp.
+-$\infty$).
+If one of the values $\inf_{k\in\mathbb{N}}\nu_{k}$ or $\sup_{k\in\mathbb{N}}\nu_{k}$ is not finite, the infinity passes to the related bound of $I_{X^{\mathrm{in}},T}^{\varphi}$.
+If those values are both finite, we can extract sequences in $I_{X^{\mathrm{in}},T}^{\varphi}$ which converge respectively to $\inf_{m\in\mathbb{N}}\nu_{m}$ and $\sup_{m\in\mathbb{N}}\nu_{m}$.
+For example, we can define $(\varphi(T^{k_{n}}(x_{k_{n}})))_{n}\subset I_{X^{\mathrm{in}},T}^{\varphi}$ such that for all $n\in\mathbb{N}$,
+$\inf_{m\in\mathbb{N}}\nu_{m}-1/n\leq\nu_{k_{n}}-1/n\leq\varphi(T^{k_{n}}(x_{k_{n}}))\leq\nu_{k_{n}}\leq\inf_{m\in\mathbb{N}}\nu_{m}+1/n.$
+This sequence converges to $\inf_{m\in\mathbb{N}}\nu_{m}$. ∎
 Let $V$ be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
-To pass from $\alpha\circ\varphi\leq V$ to $\varphi\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\alpha\circ V$, $\alpha$ has to be surjective and strictly increasing as a function of $J\to\{V(x),x\in X^{\mathrm{in}}\}\subseteq[0,1]$.
+To pass from $\alpha\circ\varphi\leq V$ to $\varphi\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\alpha\circ V$, $\alpha$ must be surjective and strictly increasing as a function from $J$ to $\{V(x),x\in X^{\mathrm{in}}\}\subseteq[0,1]$.
 From the earlier discussion, $J$ must also contain $I_{X^{\mathrm{in}},T}^{\varphi}$.
-Therefore, we define the following set of functions:
+Therefore, we define the following set of functions
 $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}:=\{\alpha:\mathbb{R}\mapsto\mathbb{R}:\exists\,I\in\mathbb{I}\left(\mathbb{R}\right)\text{ s.t.
@@
 Let us consider $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and $I$ such that $\alpha(I)=[0,1]$.
-As $\alpha$ is continuous and strictly increasing on $I$, ${{\alpha}_{\left(I\right)}^{-1}}$ exists and is also continuous and strictly increasing from $[0,1]\to I$.
+As $\alpha$ is continuous and strictly increasing on $I$, ${{\alpha}_{\left(I\right)}^{-1}}$ exists and is also continuous and strictly increasing from $[0,1]$ to $I$.
 Hence, $I$ is compact.
-Moreover, such interval $I$ is unique.
-Otherwise, consider $I,J\in\mathrm{SIC}(\alpha)$ for which we have $I\neq J$ and $\alpha(I)=\alpha(J)=[0,1]$.
+Moreover, such an interval $I$ is unique.
+Otherwise, consider $I,J\in\mathrm{SIC}(\alpha)$ such that $I\neq J$ and $\alpha(I)=\alpha(J)=[0,1]$.
 Suppose there exists $x\in I$ such that $x\leq\min J$.
@@
 ∎
-We present some useful properties of functions belonging to $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and their relation with $\nu$.
+We present some useful properties of functions belonging to $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and their relation to $\nu$.
 Proposition 10.
-For all $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that there exists an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$, satisfying $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, we have $\alpha(\limsup_{k\to+\infty}\nu_{k})\leq 0$.
+For all $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that there exists an Opt-Lyapunov function $V$ for $(X^{\mathrm{in}},T)$ satisfying $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, we have $\alpha(\limsup_{k\to+\infty}\nu_{k})\leq 0$.
 Proof.
@@
 By assumption, $\alpha(\varphi(T^{k}(x)))\leq V(\varphi(T^{k}(x)))\leq\lambda^{k}V(x)$.
-Taking the sup over $X^{\mathrm{in}}$ and from the increasing condition and the continuity of $\alpha$, $\alpha(\nu_{k})\leq\lambda^{k}\overline{V_{X^{\mathrm{in}}}}$.
+Taking the supremum over $X^{\mathrm{in}}$ and from the increasing condition and the continuity of $\alpha$, $\alpha(\nu_{k})\leq\lambda^{k}\overline{V_{X^{\mathrm{in}}}}$.
 Taking the limsup, we get $\limsup_{k\mapsto+\infty}\alpha(\nu_{k})=\alpha(\limsup_{k\mapsto+\infty}\nu_{k})\leq 0$.
@@
 Corollary 3.
-Suppose there exist $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ $V$ such that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
+Suppose there exist $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and an Opt-Lyapunov function $V$ for $(X^{\mathrm{in}},T)$ such that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
 Then:
   1. 1.
-    $\nu\in\Lambda$;
+    $\nu_{\rm opt}<+\infty$;
   2. 2.
@@
 The first statement follows readily from Prop.
-22.
+10 and the fact that $\limsup_{k\to+\infty}\nu_{k}<+\infty$ iff $\nu_{\rm opt}<+\infty$.
 The second statement follows from Prop.
 1 and Prop.
-22, since we have, if ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$.
+10, since we have, if ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$.
 ∎
 The set $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ defines the class of functions that we must consider to obtain relations $\alpha\circ\varphi\leq V$ when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
-Another important consideration to take into account is usefulness of functions $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ based on $\alpha$.
-An Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is always nonnegative and if $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}$, the inequality is obvious.
-Then $\alpha(\nu_{k})$ must be strictly positive for some $k\in\mathbb{N}$.
+Another important consideration to take into account is the usefulness of functions $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ based on elements $\alpha$ of $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.
+An Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is always nonnegative and if $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}$, $\alpha\circ\varphi\leq V$ obviously holds.
+However, a function $\alpha$ nonpositive on the terms $\nu_{k}$ whereas ${\mathrm{R}}_{\nu}^{s}$ is nonempty is not useful.
+To be useful, $\alpha(\nu_{k})$ must be strictly positive for some $k\in\mathbb{N}$.
 By Prop.
 10 and Corollary 3, we have automatically $k\in{\mathrm{R}}_{\nu}^{s}$.
-We propose a special definition for those particular functions $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.
+We propose a special definition for the particular functions $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.
 Definition 6 ($(X^{\mathrm{in}},T,\varphi)$-Certificate of compatibility).
@@
 A function $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$ is said to be $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$ if and only if there exist $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{k}-\eta$, we have $g(T^{j}(x))>\varepsilon$.
-We will say that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if there exists some $k\in{\mathrm{R}}_{\nu}^{s}$ at which $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
+We say that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if there exists some $k\in{\mathrm{R}}_{\nu}^{s}$ at which $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
+Example 8 ($V_{\ell}$ in (32) is $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-compatible).
+We have to prove that there exists $k\in{\mathrm{R}}_{\omega}^{s}$, $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$, for all $j\in\mathbb{N}$ such that $\varphi_{\ell}(L^{j}x)>\omega_{k}-\eta$, we have $V_{\ell}(L^{j}x)>\varepsilon$.
+Recall that $\overline{\lim}_{n}\omega_{n}=\lambda_{2}/c$ and thus $k\in{\mathrm{R}}_{\omega}^{s}$ iff $\omega_{k}>\lambda_{2}/c$.
+Let $k\in{\mathrm{R}}_{\omega}^{s}$ and let $\eta>0$ any real such that $\omega_{k}-\eta>\lambda_{2}/c$.
+Let $K\in\mathbb{N}$ such that $\omega_{j}\leq\lambda_{2}/c$ for all $j>K$.
+Let $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}_{\ell}$ such that $\varphi_{\ell}(L^{j}(x))>\omega_{k}-\eta$.
+It follows that $j\in\{0,\ldots,K\}$.
+From the discussion in Subsection 3.3.1, $\pi_{2}(L^{j}x)\leq\max\{\pi_{2}(x),\pi_{2}(L^{K}x)\}$.
+Moreover, as for all $x\in\mathbb{R}^{2}$ and all $k\in\mathbb{N}$, $\pi_{2}(L^{k}x)=(\lambda_{2}-\lambda_{1})^{-1}(\lambda_{1}^{k}\mathfrak{a}(x)+\lambda_{2}^{k}\mathfrak{b}(x))$, it is easy to see that $\pi_{2}(L^{j}x)\leq\max\{u_{2},\pi_{2}(L^{K}x^{uu})\}$.
+We conclude that, for all $x\in X^{\mathrm{in}}$ and all $j\in\mathbb{N}$ such that $\varphi_{\ell}(L^{j}x)>\omega_{k}-\eta$, we have
+$\begin{array}[]{ll}V_{\ell}(L^{j}x)=\dfrac{C}{(\pi_{1}(L^{j}x))^{2}+(\pi_{2}(L^{j}x))^{2}}&=\dfrac{C}{(\pi_{2}(L^{j}x))^{2}((\varphi_{\ell}(L^{j}x))^{2}+1)}\\ &\geq\dfrac{C}{(\max\{u_{2},\pi_{2}(L^{K}x^{uu}))\})^{2}(\omega_{\rm opt}^{2}+1)}:=\varepsilon>0\end{array}$
+This proves that $V_{\ell}$ is an Opt-Lyapunov function $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-compatible. ∎
 To be $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$ implies that if $T^{k}(x)$ is $\eta$-optimal for $\varphi$ then $g(T^{k}(x))$ is strictly positive.
@@
 From Prop.
-23, we can suppose that $\nu_{k}>\inf_{n\in\mathbb{N}}\nu_{n}$ and so there exists $\delta>0$ such that $(\nu_{k}-\delta,\nu_{k})\subset I_{X^{\mathrm{in}},T}^{\varphi}$.
-As $\alpha$ is continuous and strictly increasing on $I_{X^{\mathrm{in}},T}^{\varphi}$, for $\varepsilon:=\alpha(\nu_{k})/2$, there exists $\eta\in(0,\delta]$, such that $0<\nu_{k}-s\leq\eta$ implies that $\varepsilon<\alpha(s)<\alpha(\nu_{k})+\varepsilon$.
+11, we can suppose that $\nu_{k}>\inf_{n\in\mathbb{N}}\nu_{n}$ and thus there exists $\zeta>0$ such that $(\nu_{k}-\zeta,\nu_{k})\subset I_{X^{\mathrm{in}},T}^{\varphi}$.
+As $\alpha$ is continuous and strictly increasing on $I_{X^{\mathrm{in}},T}^{\varphi}$, for $\varepsilon:=\alpha(\nu_{k})/2$, there exists $\eta\in(0,\zeta]$, such that $0<\nu_{k}-s\leq\eta$ implies that $\varepsilon<\alpha(s)<\alpha(\nu_{k})+\varepsilon$.
 Now let us consider $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{j}(x))>\nu_{k}-\eta$.
@@
 Now, assume that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at some $k\in{\mathrm{R}}_{\nu}^{s}$.
-Then, from Prop.23, $g$ is also $(X^{\mathrm{in}},T,\varphi)$-compatible at $k=\mathbf{K}_{\nu}$.
+Then, from Prop.11, $g$ is also $(X^{\mathrm{in}},T,\varphi)$-compatible at $k=\mathbf{K}_{\nu}$.
 Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{\rm opt}-\eta$, we have $g(T^{j}(x))>\varepsilon$.
@@
 Finally, as $\alpha$ is strictly increasing, for all $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$, we have $\alpha(\nu_{k})>0$.
-Note that the set $\{k\in\mathbb{N}:\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty as $[\mathbf{K}_{\nu},\mathbf{K}^{s}_{\nu}]\cap\mathbb{N}$ is included in it.
+Note that the set $\{k\in\mathbb{N}:\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty as $\mathbf{K}_{\nu},\mathbf{K}^{s}_{\nu}]\cap\mathbb{N}$ is included in it.
 ∎
-Example 9.
-Let consider the Opt-Lyapunov function $V$ of Example 5. We prove that $V$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
-We define for all $s\in\mathbb{R}$, $\alpha(s):=s-p^{2}/3+\min\{\zeta,\varepsilon\}$ where $0<\zeta<p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ and $\varepsilon:=\min\{\varepsilon_{k},k\in\operatorname{Argmax}{\nu}\}$ where $\varepsilon_{k}=\inf\{V(T^{k}(x)):x\in X^{\mathrm{in}}\}$.
-The scalar $\varepsilon$ is strictly positive as the function $[\mu,1/2]\ni y\mapsto V(T^{k}((2,1)y)$ is continuous and strictly positive for all $k\in\mathbb{N}$ and $\operatorname{Argmax}(\nu)$ is a finite set of integers.
-The function $\alpha$ is clearly strictly increasing and continuous on $\mathbb{R}$.
-We have for all $j\in\mathbb{N}$ and for all $x\in X^{\mathrm{in}}$, $\alpha(\varphi(T^{j}(x)))=\varphi(T^{j}(x))-p^{2}/3+\min\{\zeta,\varepsilon\}<\varphi(T^{j}(x))-p^{2}/3+p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$.
-Then $\alpha(\varphi(T^{j}(x)))$ is strictly negative for all $j\notin\operatorname{Argmax}(\nu)$.
-Now let $j\in\operatorname{Argmax}(\nu)$.
-Suppose that $\varphi(T^{j}(x))>p^{2}/3-\min\{\zeta,\varepsilon\}$.
-As we have $\varphi(T^{j}(x))\leq p^{2}/3$, we get $\alpha(\varphi(T^{j}(x)))\leq\varepsilon\leq V(T^{k}(x))$ for all $k\in\operatorname{Argmax}(\nu)$ and all $x\in X^{\mathrm{in}}$.
-This leads to $\alpha(\varphi(T^{j}(x)))\leq V(T^{j}(x))$.
-On the other hand, $\alpha(\varphi(T^{j}(x)))\leq 0$ for all $j\in\operatorname{Argmax}(\nu)$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))\leq p^{2}/3-\min\{\eta,\varepsilon\}$.
-We have proved that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$.
-Moreover, as the set of $\{x\in X^{\mathrm{in}}:\varphi(T^{j}(x))>p^{2}/3-\min\{\eta,\varepsilon\}\}$ is nonempty for all $j\in\operatorname{Argmax}(\nu)$, we conclude that for all $j\in\operatorname{Argmax}(\nu)$, there exists $x\in X^{\mathrm{in}}$ such that $0<\alpha(\varphi(T^{j}(x)))$ and as $\alpha$ is increasing $0<\alpha(\nu_{j})$ for all $j\in\operatorname{Argmax}(\nu)$ and $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$ making $V$ $(X^{\mathrm{in}},T,\varphi)$-compatible.
-Actually, by simple computations, we have $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$.
-One can compute $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ as follows.
-Let $n\in\mathbb{N}$.
-We have, for all $y\in\mathbb{R}_{+}^{*}$, $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$.
-Thus $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$ is strictly negative if and only if $n\geq n_{y}^{*}:=\max\{\lfloor\ln(4p/(15y))/\ln(3/2)\rfloor+1,0\}$.
-Then for all $y\in\mathbb{R}_{+}^{*}$, $\mathbb{N}\ni n\mapsto\varphi_{n}(y)$ is maximized at $n_{y}^{*}$.
-We deduce that $n_{1/2}^{*}=\lfloor\ln(8p/15)/\ln(3/2)\rfloor+1$ and, as $\mu\in(0,1/3)$, $n_{\mu}^{*}=\lfloor\ln(4p/(15\mu))/\ln(3/2)\rfloor+1$.
-This implies that $n_{1/2}^{*}\leq\underline{n}-1$ and thus we have $\nu_{n}\leq\nu_{n_{1/2}^{*}}$ for all $n\leq\underline{n}-1$.
-This also implies that $n_{\mu}^{*}<\overline{n}$ and so $\nu_{n}<\nu_{\overline{n}}$ for all $n>\overline{n}$.
-Finally, $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}=\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\}$.
-5.2.4 Direct and Converse Opt-Lyapunov Theorems
+Example 9 (A $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$).
+From Th 4 and Example 8, there exists a $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$.
+We will see in Theorem 5 that this certificate is required to construct an element $(h,\beta)\in\mathbb{\Gamma}(\omega)$ from an Opt-Lyapunov function.
+Thus, we construct a certificate of compatibility for $V_{\ell}$ proposed in (32).
+From the definition of $V_{\ell}$, we can easily establish a link between $V_{\ell}$ and $\varphi_{\ell}$.
+Indeed, for all $x\in(\mathbb{R}_{+}^{*})^{2}$:
+$V_{\ell}(x)=\dfrac{C}{x_{1}^{2}+x_{2}^{2}}=\dfrac{C}{x_{1}^{2}(1+(\varphi_{\ell}(x))^{-2})}.$
+The functions of the form $s\mapsto C/(M^{2}(1+s^{-2}))$ are strictly increasing and continuous.
+One way to use these functions as lower bounds of $V_{\ell}$ is to find an upper bound of the set $\{\pi_{1}(L^{k}x):x\in X^{\mathrm{in}}_{\ell},k\in\mathbb{N}\}$.
+However, from the discussion in 3.3.1, $\{\pi_{1}(L^{k}x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}$ is unbounded.
+Moreover, of the form $s\mapsto C/(M^{2}(1+s^{-2}))$ are strictly positive, whereas a certificate of compatibility must take negative values for all $k\in\mathbb{N}$ such that $\omega_{k}<\overline{\lim_{n}}\omega_{n}$.
+Then, the certificate is defined as a piecewise function on $\mathbb{R}$.
+The pieces are one interval and its complement.
+The interval has a lower bound that is strictly greater than the limit superior.
+Hence, the certificate takes the form $s\mapsto C/(M^{2}(1+s^{-2}))$ on the interval.
+The definition of the interval permits us to consider a bounded subset of $\{\pi_{1}(L^{k}x):(x,k)\in X^{\mathrm{in}}_{\ell}\times\mathbb{N}\}$ and thus compute $M$.
+On the complement of the interval, our certificate is an affine form that is proved to be a lower bound of $V_{\ell}$.
+First, from the analysis of $\mathfrak{a}$, we have either $\omega_{0}\in{\mathrm{R}}_{\omega}^{s}$ or $\omega_{1}\in{\mathrm{R}}_{\omega}^{s}$. Then, we define
+$z:=\max\{\omega_{0},\omega_{1}\}.$
+[…]
+$N(x):=\left\lfloor\dfrac{\ln\left(\dfrac{\mathfrak{b}(x)(c(z-\eta)-\lambda_{2})}{|\mathfrak{a}(x)|(c(z-\eta)-\lambda_{1})}\right)}{\ln(\left|\lambda_{1}/\lambda_{2}\right|)}\right\rfloor$
+For a given $x\in X_{\ell}$, the integer $N(x)$, if nonnegative, determines an upper bound of the maximal integer $k$ such that $\varphi_{\ell}(L^{k}x)\geq z-\eta$.
+Hence, the integer $\overline{N}$ defined as follows
+[…]
+is an upper bound for $K$.
+Moreover, it can be proved that $\overline{N}$ is nonnegative.
+Then, we can compute $M=\max_{x\in X^{\mathrm{in}}_{\ell}}\pi_{1}(L^{\overline{N}}x)=\pi_{1}(L^{\overline{N}}x^{uu})$.
+Actually, $\overline{N}$ is an upper bound of the greatest maximizer $\mathbf{K}_{\omega}$.
+However, we continue as the goal is to produce a $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$-certificate of compatibility.
+We are able to define:
+$\alpha_{1}:s\mapsto\dfrac{C}{M^{2}(1+s^{-2})}.$
+Now let us take any $\gamma$ greater than
+$\max\left\{\max_{s\in[z-\eta,z]}\alpha_{1}^{\prime}(s),\dfrac{V_{\ell}(x^{*})}{\eta}\right\}$
+where $x^{*}\in X^{\mathrm{in}}_{\ell}\cup L(X^{\mathrm{in}}_{\ell})$ is the vector such that $z=\varphi_{\ell}(x^{*})$.
+Note that as $\alpha_{1}$ is strictly increasing and $V_{\ell}$ is strictly positive, $\gamma$ is strictly positive.
+It follows that $\alpha_{1}(s)\geq\gamma(s-z)+\alpha_{1}(z)$.
+Moreover, there exists $u\in[z-\eta,z)$ such that $\gamma(u-z)+\alpha_{1}(z)=0$.
+We conclude that
+$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{C}{M(1+s^{-2})}&\text{ if }s\geq z\\ \gamma(s-z)+\dfrac{C}{M^{2}(1+z^{-2})}&\text{ if }s<z\end{array}\right$.
+is a $(X_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$. As for all $s\in[z-\eta,z]$, we have
+$\alpha_{1}^{{}^{\prime}}(s)=\dfrac{2Cs^{-3}}{(M^{2}(1+s^{-2}))^{2}}\leq\dfrac{2C(z-\eta)^{-3}}{(M^{2}(1+z^{-2})},$
+a candidate $\gamma$ could be
+$\max\left\{\dfrac{2C(z-\eta)^{-3}}{(M^{2}(1+z^{-2}))^{2}},\dfrac{V_{\ell}(x^{*})}{\eta}\right\}.$
+Applying those computations to our numerical data of Subsection 3.4, we have $z=\max\{\omega_{0},\omega_{1}\}=42$ and
+$\eta=0.5(42-4)=19$
+We have $\mathfrak{b}(x^{lu})=55$, $\mathfrak{a}(x^{lu})=95$, $\mathfrak{b}(x^{ul})=25$ and $\mathfrak{a}(x^{ul})=-10$. Hence,
+$N(x^{lu})=\left\lfloor\dfrac{\ln\left(\dfrac{319}{665}\right)}{-\ln(2)}\right\rfloor=1\text{ and }N(x^{ul})=\left\lfloor\dfrac{\ln\left(\dfrac{19}{10}\right)}{-\ln(2)}\right\rfloor=-1$
+Thus, the integer $\overline{N}$ is 1.
+We remark that the integer $N(x^{ul})=-1$ which means that to have $\varphi_{\ell}(L^{k}x^{ul})\geq 42-19=23$, $k$ has to be negative.
+Indeed, for all $k\in\mathbb{N}$, $\varphi_{\ell}(L^{k}x^{ul})<23$.
+We could stop here as $\overline{N}$ is an upper bound of the greatest integer maximizer of ($DPCP_{\ell}$).
+We have to compute $M=\pi_{1}(Lx^{uu})^{2}=240^{2}=57600$. We have $x^{*}=Lx^{lu}=(210,5)^{\intercal}$. Then
+$\dfrac{V_{\ell}(x^{*})}{\eta}=\dfrac{125}{(19(210^{2}+5^{2}))}=\dfrac{1}{6707}\text{ and }\dfrac{2C(z-\eta)^{-3}}{M(1+z^{-2})^{2}}=\dfrac{250\times 23^{-3}}{57600(1+42^{-2})^{2}}=\dfrac{21609}{60644708120}$
+and thus $\gamma=1/6707$.
+We conclude that
+$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{125}{57600(1+s^{-2})}&\text{ if }s\geq 42\\ \\ \dfrac{(s-42)}{6707}+\dfrac{49}{22592}&\text{ if }s<42\end{array}\right$.
+5.3 Direct and Converse Opt-Lyapunov Theorems
 We present both direct and converse Opt-Lyapunov theorems using the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion.
-More precisely, the direct version aims to prove that if a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function exists (and such that $\overline{N_{T}}(V)>0$), then we can construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$.
+More precisely, the direct version aims to prove that if a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function exists (and such that $\overline{N_{T}}(V)>0$), then we can construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ that is useful for $\nu$.
 The converse version exhibits a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function when a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$.
-Moreover, we are able to produce a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for this Opt-Lyapunov function.
-We remark that from Prop. 14, if $V$ is an Opt-Lyapunov for $(X^{\mathrm{in}},T)$ then $\{T^{k}(x):k\in\mathbb{N},x\in X^{\mathrm{in}}\}\subseteq\operatorname{dom}(V)$.
+Moreover, we can produce a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for this Opt-Lyapunov function.
+We note that from Prop. 2, if $V$ is an Opt-Lyapunov for $(X^{\mathrm{in}},T)$ then $\{T^{k}(x):k\in\mathbb{N},x\in X^{\mathrm{in}}\}\subseteq\operatorname{dom}(V)$.
 Proposition 12.
@@
 Thus, for all $k\in\mathbb{N}^{*}$, $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))=0$ and as $\alpha$ is increasing and continuous, $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}^{*}$.
-As for some $j\in{\mathrm{R}}_{\nu}^{s}$, $\alpha(\nu_{j})>0$, this leads to $j=0$ and as $\alpha$ is strictly increasing and continuous on $I_{X^{\mathrm{in}},T}^{\varphi}$ then for all $k\in\mathbb{N}^{*}$, $\nu_{0}>\nu_{k}$ and $\mathbf{K}^{s}_{\nu}=0$.
+For some $j\in{\mathrm{R}}_{\nu}^{s}$, $\alpha(\nu_{j})>0$, this leads to $j=0$ and as $\alpha$ is strictly increasing and continuous on $I_{X^{\mathrm{in}},T}^{\varphi}$ then for all $k\in\mathbb{N}^{*}$, $\nu_{0}>\nu_{k}$ and $\mathbf{K}^{s}_{\nu}=0$.
 This also implies that $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$.
@@
 Let $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)$ where $\alpha(I)=[0,1]$.
-We define the function $h:x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$.
+We define the function
+$h:[0,1]\ni x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right).$ (35)
 Then $(h,N_{V}(T))\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$.
 Proof.
-The function ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and continuous.
-Thus, we have, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(V(T^{k}(x))$.
-By Def.
-9 and Prop.
-17, $V(T^{k}(x))\leq(N_{V}(T))^{k}V(x)\leq N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}$.
-As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing, we conclude that $\nu_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}\right)$.
-Note that $x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$ is strictly increasing and continuous on $J\supset[0,1]$ as ${{\alpha}_{\left(J\right)}^{-1}}$ is and $\overline{V_{X^{\mathrm{in}}}}$ is not zero.
+The function ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and continuous on $[0,1]$. Thus, we have, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$,
+$\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(V(T^{k}(x)).$
+By Def. 5 and Prop. 5, $V(T^{k}(x))\leq(N_{V}(T))^{k}V(x)\leq N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing, we conclude that
+$\nu_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}\right).$
+Note that $x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$ is strictly increasing and continuous on $J\supset[0,1]$ as ${{\alpha}_{\left(J\right)}^{-1}}$ is, and $\overline{V_{X^{\mathrm{in}}}}$ is not zero.
 Moreover, by Prop.
@@
 ∎
-Example 10.
-Following Th.
-9, from the Opt-Lyapunov function $V$ defined at Example 5 and the ($X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, $\alpha$ defined at Example 9, the function $h:[0,1]\ni s\mapsto\alpha_{\mathbb{R}}^{-1}(s\overline{V_{X^{\mathrm{in}}}})$ belongs to $\mathbb{\Gamma}_{\rm fun}(\nu)$ and $\overline{N_{T}}(V)=4/9\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
-Then $h_{V}(s)=s+p^{2}/3-\min\{\zeta,\varepsilon\}$ with $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$ and $\zeta\in(0,p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})$ where $n_{1/2}^{*}$ is the maximizer for the function $n\mapsto\varphi_{n}(1/2)$.
-We draw a table as we done at Example 3 to compare the direct (from the function $h$ of Example 3) and the Opt-Lyapunov approaches.
-We have some flexibility about $\zeta$ so we propose two values for a comparison purpose.
-Hence, we compare the computations for $h_{\zeta_{1}}$ associated to $\zeta_{1}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/2$ and $h_{\zeta_{2}}$ associated to $\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$.
-The evaluations for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},4/9)\rfloor$ and for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},4/9)\rfloor$ are given at Table 3.
-$p=30$ $p=9$ $p=3$
-$\mu$
-$\frac{1}{3}$
-$\mu$
-$\frac{1}{10}$
-$\mu$
-$\frac{1}{1000}$
-$\mu$
-$\frac{1}{3}$
-$\mu$
-$\frac{1}{10}$
-$\mu$
-$\frac{1}{1000}$
-$\mu$
-$\frac{1}{3}$
-$\mu$
-$\frac{1}{10}$
-$\mu$
-$\frac{1}{1000}$
-$\mathbf{K}^{s}_{\nu}$ 8 11 22 5 8 19 2 3 17
-$\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ 9 12 23 7 9 20 3 7 18
-$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},\frac{4}{9})\rfloor$ 8 14 36 5 11 33 4 8 31
-$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},\frac{4}{9})\rfloor$ 8 14 36 9 11 33 12 12 31
-Table 3: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},\beta)\rfloor$ described at Eq.
-(4) for the sequence $\nu$ of Eq.
-(2) and two pairs $(h_{\zeta_{1}},4/9)$ and $(h_{\zeta_{2}},4/9)$ of Example 10.
-The evaluations of $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},4/9)\rfloor$ from the Opt-Lyapunov function proposed at Example 5 and associated $(X^{\mathrm{in}},T,\varphi)$-certificates of compatibility provide worst upper bounds when $\mu$ is small (equal to $1/10$ or $1/1000$) but the optimal value for $\mu=1/3$.
-A small value for $\zeta$ ($\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$) makes decrease the precision for $\mu=1/3$.
-We recall that the y-intercept of the affine function $h_{\zeta}$ is $p^{2}/3-\min\{\varepsilon,\eta\}$ where $\varepsilon$ is a very small value when $\mu$ is small.
-Hence, when $\mu=1/10$ or $\mu=1/1000$, $\zeta_{1}$ or $\zeta_{2}$ are still greater than $\varepsilon$.
-Note that the values $h_{\zeta_{i}}^{-1}(\nu_{k})$ are strictly positive if and only if for $k\in\operatorname{Argmax}(\nu)$. This is sufficient to compute an upper bound of $\mathbf{K}^{s}_{\nu}$. ∎
-We establish a ”converse Lyapunov theorem”. The construction of the Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is inspired by the one proposed by Yoshizawa [Yos66].
+We establish a converse Lyapunov theorem. The construction of the Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is inspired by the one proposed by Yoshizawa [Yos66].
 Theorem 6 (Converse Opt-Lyapunov Theorem).
-Let $(h,\beta)\in\mathbb{\Gamma}(\nu)$ be useful for $\nu$.
-Let us define $s\mapsto\omega(s):=\min\left\{h(1),\max\left\{s,h(0)\right\}\right\}$.
-Then, the function $V\in\mathcal{PSD}$ defined as follows:
-$x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega\left(\varphi(T^{k}(x))\right)\right)$
-is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Moreover, the function $\hat{h}$ defined as follows:
+Let $(h,\beta)\in\mathbb{\Gamma}(\nu)$ be useful for $\nu$. Let us define
+$\mathbb{R}\ni s\mapsto\mathfrak{m}(s):=\min\left\{h(1),\max\left\{s,h(0)\right\}\right\}.$
+Then, the function $V\in\mathcal{PSD}$ is defined as follows:
+$\mathbb{R}^{d}\ni x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}\left(\varphi(T^{k}(x))\right)\right)$
+is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Moreover, the function $\hat{h}$ is defined as follows:
 $\hat{h}:\mathbb{R}\mapsto\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in(h(0),h(1)]\end{array}\right$.
@@
 First, we prove that the function $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
-Let us prove that $V$ takes its values in $[0,+\infty]$.
+Let us prove that $V$ takes its values in $0,+\infty]$.
 Let $x\in\mathbb{R}^{d}$.
-As we have $h(1)\geq\omega(s)\geq h(0)$, we have, as ${{h}_{\left([0,1]\right)}^{-1}}$ is increasing, $1\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\omega(\varphi(T^{k}(x))))\geq 0$ for all $k\in\mathbb{N}$.
+As we have for all $s\in\mathbb{R}$ $h(1)\geq\mathfrak{m}(s)\geq h(0)$, we have, as ${{h}_{\left([0,1]\right)}^{-1}}$ is increasing, $1\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\mathfrak{m}(\varphi(T^{k}(x))))\geq 0$ for all $k\in\mathbb{N}$.
 We conclude, as $\beta\in(0,1)$ that $V(x)\geq 0$ and $V(x)=+\infty$ for all $x\in\mathbb{R}^{d}$ such that $\varphi(T^{k}(x))\geq h(1)$ for all $k$ sufficiently large.
-By definition of $\mathbb{\Gamma}(\nu)$, we have for all $x\in X^{\mathrm{in}}$ and all $n\in\mathbb{N}$, $\varphi(T^{n}(x))\leq h(\beta^{n})<h(1)$ and so $V(x)\leq\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(\beta^{k}))=\sup_{k\in\mathbb{N}}\beta^{-k}\beta^{k}=1$.
+By definition of $\mathbb{\Gamma}(\nu)$, we have for all $x\in X^{\mathrm{in}}$ and all $n\in\mathbb{N}$, $\varphi(T^{n}(x))\leq h(\beta^{n})<h(1)$ and so
+$V(x)\leq\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(\beta^{k}))=\sup_{k\in\mathbb{N}}\beta^{-k}\beta^{k}=1.$
 This implies that $\overline{V_{X^{\mathrm{in}}}}\in[0,1]$.
@@
 Let us take $k\in\mathbb{N}$ such that $h(1)>\nu_{k}>h(0)$.
-So there exists $\varepsilon>0$ such that $\nu_{k}-\varepsilon>h(0)$.
+Therefore, there exists $\varepsilon>0$ such that $\nu_{k}-\varepsilon>h(0)$.
 For this $\varepsilon$, there exists $x\in X^{\mathrm{in}}$ such that $\nu_{k}\geq\varphi(T^{k}(x))\geq\nu_{k}-\varepsilon$.
-Finally, for this $x$, $V(x)\geq\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\varphi(T^{k}(x))>\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$ as ${{h}_{\left([0,1]\right)}^{-1}}$ strictly increases.
+Finally, for this $x$,
+$V(x)\geq\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\varphi(T^{k}(x))>\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$
+as ${{h}_{\left([0,1]\right)}^{-1}}$ strictly increases.
 We conclude that $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$.
-Finally, we have to prove that $V\circ T\leq\beta V$.
-Let $y\in\mathbb{R}^{d}$, if $V(y)=+\infty$, the result holds.
-Now suppose that $y\in\operatorname{dom}V$, we have:
-$\begin{array}[]{rl}V(T(y))=&\displaystyle{\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{k+1}(y))\right)}\\ =&\displaystyle{\sup_{j\in\mathbb{N}^{*}}\beta^{-j+1}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ =&\displaystyle{\beta\sup_{j\in\mathbb{N}^{*}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ \leq&\displaystyle{\beta\sup_{j\in\mathbb{N}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)=\beta V(y)}\end{array}$
+Finally, we must prove that $V\circ T\leq\beta V$.
+Let $y\in\mathbb{R}^{d}$.
+If $V(y)=+\infty$, there is nothing to prove.
+Assume now $V(y)<+\infty$.
+We have:
+$\begin{array}[]{rl}V(T(y))=&\displaystyle{\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{k+1}(y)))\right)}\\ =&\displaystyle{\sup_{j\in\mathbb{N}^{*}}\beta^{-j+1}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{j}(y)))\right)}\\ =&\displaystyle{\beta\sup_{j\in\mathbb{N}^{*}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{j}(y)))\right)}\\ \leq&\displaystyle{\beta\sup_{j\in\mathbb{N}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{j}(y)))\right)=\beta V(y)}\end{array}$
 We conclude that $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
 […]
+$\begin{array}[]{l}V(T^{k}(x))=\\ \sup_{n\in\mathbb{N}}\beta^{-n}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\max\left\{\varphi(T^{k+n}(x)),h(0)\right\}\right)\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\max\left\{\varphi(T^{k}(x)),h(0)\right\}\right)\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\varphi(T^{k}(x))\right)\end{array}$
+if $\varphi(T^{k}(x))\geq h(0)$.
+If $\varphi(T^{k}(x))<h(0)$, then $V(T^{k}(x))\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$ and $\hat{h}(\varphi(T^{k}(x)))\leq 0$.
+We conclude that $\hat{h}(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$.
+As $\mathcal{S}(\nu,h)\neq\emptyset$, there exists $k\in\mathbb{N}$, such that $h(1)>\nu_{k}>h(0)$.
+As ${{h}_{\left([0,1]\right)}^{-1}}$ is strictly increasing on $[h(0),h(1)]$, we have $\hat{h}(\nu_{k})=\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\nu_{k})>0$.
+This ends the proof that $\hat{h}$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility of $V$.
+∎
+Example 10.
+As we choose to take $\gamma$ equals
+$\max\left\{\dfrac{2(\omega-\eta)^{-3}}{(M(1+\omega^{-2})},\dfrac{V(x^{*})}{\eta}\right\},$
+the function
+$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{C}{M{}(1+s^{-2})}&\text{ if }s\geq z\\ \gamma(s-z)+\dfrac{C}{M^{2}(1+z^{-2})}&\text{ if }s<z\end{array}\right$.
+is a $(X_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$. Finally, $(h,\beta)\in\mathbb{\Gamma}(\omega)$ where:
+$h=\alpha^{-1}\text{ and }\beta=\dfrac{1}{\min\{a^{2}+c^{2},b^{2}\}}.$
+This leads to the formula for all $k\in\mathbb{N}$ such that $\omega_{k}>\alpha^{-1}(0)$
+$\dfrac{\ln(\alpha(\omega_{k})/\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}})}{-\ln(\min\{a^{2}+c^{2},b^{2}\})}.$
+Applying our computations to the numerical data of Subsection 3.4 leads to $\alpha^{-1}(0)=\dfrac{1757}{64}>8=\omega_{0}$. Then, we cannot use $\alpha$ at $\omega_{0}$. Finally,
+$\dfrac{\ln(\alpha(\omega_{1})/\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}})}{-\ln(\min\{a^{2}+c^{2},b^{2}\})}=\dfrac{\ln\left(\dfrac{49}{22592}\right)}{\ln\left(\dfrac{4}{5}\right)}=27.49$
+This means that from the certificate of compatibility, we assume that the maximal term of $\omega$ is one of the 28th first terms.
+This result is sound but not optimal.
+Indeed, the maximal term is $\omega_{1}$.
+We see that the Opt-Lyapunov approach is less precise than $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ approach.
+For the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ approach, we have to construct an upper bound of $\varphi_{\ell}(L^{k}x)$.
+This bas been done directly from the function $H$ that characterizes $\varphi_{\ell}(L^{k}x)$.
+This seems to explain why the upper bound is tight and thus the function $h\in\mathbb{\Gamma}(\omega)$ obtained provides better results.
+Opt-Lyapunov function $V_{\ell}$ defined in Eq.
+(32) does not depend on $\varphi_{\ell}$.
+The function depends only on the dynamics $L$ and initial conditions set.
+The link between $V_{\ell}$ and $\varphi_{\ell}$ is established in Example 8 and becomes explicit in Example 9.
+More generally, a certificate of compatibility is the function that relates a Lyapunov function to the orbits $\varphi(T^{k}x))$.
+∎
 6 A Technical Summary
-We have proposed a function $\mathfrak{F}_{\nu}$ at Eq.
-(4) whose the integer part of values provide solutions for Problem 2.
-This function is based on couples $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$.
-Without this condition, $(h,\beta)$ is useless as $\mathfrak{F}_{\nu}$ is equal to +$\infty$+ ∞.
-When the supremum of $(\nu_{k})_{k}$ is strictly greater than its limit superior, a useful couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ exists (Th.
-1) and $h$ can be chosen as an affine function.
-Furthermore, we can compute the greatest maximizer of $(\nu_{k})_{k}$ from this couple (Th.
-3) and thus we can solve Problem 1.
-Unfortunately, to directly compute a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is as difficult as solving Problem 1 and we have to construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from alternative methods.
-We propose alternatives coming from classical tools from stability theory of difference equations.
+In this paper, we have proposed a general method to solve optimization problems of the form (DPCP).
+More precisely, focusing on the integer component of the optimal solution, it is necessary to compute an index $k\in\mathbb{N}$ for which the term of the sequence $\nu$ (defined in (1)) is maximal.
+The method is based on a function $\mathfrak{F}_{\nu}$ defined in [Adj25c] and quickly recalled in Eq.
+(4).
+This function $\mathfrak{F}_{\nu}$ requires a pair $(h,\beta)$ where $h$ is a strictly increasing continuous function on $[0,1]$ and $\beta\in(0,1)$ such that $\nu_{k}\leq h(\beta^{k})$ for all $k\in\mathbb{N}$.
+A key notion is the usefulness of $(h,\beta)$, that is, there exists some $k\in\mathbb{N}$ such that $\nu_{k}>h(0)$.
+Without this condition, an upper bound of the greatest index maximizer of $\nu$ cannot be computed from $(h,\beta)$.
+However, the existence theorem [Adj25c, Th.2] (partially recalled in Th 1) is not totally constructive and computing $(h,\beta)\in\mathbb{\Gamma}(\nu)$ has to follow an alternative methodology.
+We propose alternatives based on classical tools from the stability theory of difference equations.
+In [Adj25a], we compute $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from classical $\mathcal{K}\mathcal{L}$ and Lyapunov functions.
+However, this approach suffers from a lack of converse theorems.
+Indeed, some problems of the form (DPCP) (such as Problem ($DPCP_{\ell}$)) cannot be solved using these classes of functions.
 The inequality $\varphi(T^{k}(x))\leq h(\beta^{k})$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ can be viewed as a kind of $\mathcal{KL}$ inequality that appears in asymptotic stability studies.
@@
 We have proved (Th.
-5) that a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ exists if and only if $\nu_{\rm opt}$ is finite.
-Then, we extend Sontag’s result to the $\mathcal{KL}_{\rm gen}$ class (Lemma 2 and Lemma 3) to construct $h\in\Omega([0,1])$, useful for $\nu$, from a $\mathcal{KL}_{\rm gen}$ function and such that $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ (Th.
-6).
-One of the advantage of the Sontag’s result extension is that the scalar part of elements of $\mathbb{\Gamma}(\nu)$ can be fixed and it remains to construct $h$ from the $\mathcal{KL}_{\rm gen}$ certificate.
+2) that a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ exists if and only if $\nu_{\rm opt}$ is finite: Then, we extend Sontag’s result to the $\mathcal{KL}_{\rm gen}$ class (Lemma 1 and Lemma 2) to construct $h\in\Omega([0,1])$, useful for $\nu$, from a $\mathcal{KL}_{\rm gen}$ function and such that $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ (Th.
+3).
+One of the advantages of the extension of Sontag’s result is that the scalar part of the elements of $\mathbb{\Gamma}(\nu)$ can be fixed, and it remains to construct $h$ from the $\mathcal{KL}_{\rm gen}$ certificate.
 In practice, to decide whether a function $(\gamma,\operatorname{Fin}\left(X^{\mathrm{in}}\right))$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$ seems difficult to make.
-One constructive alternative of $\mathcal{K}\mathcal{L}$ certificates consists in Lyapunov functions.
+One constructive alternative to $\mathcal{K}\mathcal{L}$ certificates consists in Lyapunov functions.
 Classical Lyapunov functions satisfy pure functional inequations and are not defined from the orbits of the dynamical system.
-One of the most interesting theoretical result for Lyapunov functions is the necessity of their existence for the stability (namely converse Lyapunov theorems).
-For our specific problem, we propose a new type of Lyapunov function called Opt-Lyapunov functions for which we obtain both direct (Th.
-9) and converse theorem (Th 10).
+One of the most interesting theoretical results for Lyapunov functions is the necessity of their existence for stability (namely, the converse Lyapunov theorems).
+For our specific problem, we propose a new type of Lyapunov function called the Opt-Lyapunov function, for which we obtain both direct (Th.
+5) and the converse theorem (Th 6).
 The main difference between classical and Opt-Lyapunov functions is that Opt-Lyapunov functions are nonnegative but not definite (many zeros) and can take infinite values (outside reachable values taken by the dynamical system).
-The fact that Opt-Lyapunov functions are not definite (and not continuous) makes classical comparison functions impossible to use.
-In the classical case, comparison functions serve as link between our objective function $\varphi$ and the classical Lyapunov function.
+The fact that Opt-Lyapunov functions are not definite (and not continuous) renders classical comparison functions impossible to use.
+In the classical case, comparison functions serve as a link between our objective function $\varphi$ and the classical Lyapunov function.
 This link (and comparison functions) is replaced by the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion (Th.
-8).
-Theorem 11 sums up the equivalences proved in this paper relatively to the peaks computation problem.
+4).
+Theorem 7 summarizes the equivalences proved in this study with respect to the peak computation problem.
 Theorem 7 (Summary).
-The following statements are equivalent:
+The following statements are equivalent.
   1. 1.
@@
   3. 3.
-    There exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)$ is finite and $\nu_{\rm opt}=\max_{k\in\mathbb{N}}\{\nu_{k}:k=0,\ldots,\mathbf{K}^{s}_{\nu}\}$;
+    There exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)$ is well-defined and finite and $\nu_{\rm opt}=\max_{k\in\mathbb{N}}\{\nu_{k}:k=0,\ldots,\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)\}$;
   4. 4.
@@
     There exists an Opt-Lyapunov $(X^{\mathrm{in}},T,\varphi)$ compatible function.
-The proof of Theorem is done by sketching the functions constructed in the proofs of earlier theorems.
-Figure 1 provides a graphical view of different results involved to prove Th.
-11.
-In Figure 1, the notation $\mathcal{OL}_{\rm comp}$ stands for the set of $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov functions.
+The proof of Theorem is obtained by sketching the functions constructed in the proofs of the earlier theorems.
+Figure 2 provides a graphical view of the different results involved in proving Th.
+7.
+In Figure 2, the notation $\mathcal{OL}_{\rm comp}$ denotes the set of $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov functions.
 […]
@@
 As $V\in\mathcal{OL}_{\rm comp}$ is $(X^{\mathrm{in}},T,\varphi)$-compatible, we associate with them a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
-For a such a function $V$, we denote by $\alpha_{V}$ the chosen $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
-We now clarify the circled numbers that appear in Figure 1.
+For such a function $V$, we denote by $\alpha_{V}$ the chosen $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
+We now clarify the circled numbers in Figure 2.
 […]
@@
     $\begin{array}[]{ccc}\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)&\longrightarrow&\mathbb{\Gamma}(\nu)\\ (\gamma,\theta)&\longmapsto&(h,e^{-1})\end{array}$
-    The function $h$ cannot be explicitly formulated and is constructed following the proofs of Lemma 2 and Lemma 3.
+    The function $h$ cannot be explicitly formulated and is constructed following the proofs of Lemmas  1 and  2.
 […]
     $\begin{array}[]{ccccc}\mathcal{OL}_{\rm comp}&\longrightarrow&\mathcal{OL}_{\rm comp}\times\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}&\longrightarrow&\mathbb{\Gamma}(\nu)\\ V&\longmapsto&(V,\alpha_{V})&\longmapsto&(h_{V},\overline{N_{T}}(V))\end{array}$
-    where $h_{V}$ is defined as follows:
-    $[0,1]\ni s\mapsto h_{V}(s)=\displaystyle{{\alpha_{V}}_{\left(J\right)}^{-1}}(s\overline{V_{X^{\mathrm{in}}}})$
+    where $h_{V}$ is defined as
+    $0,1]\ni s\mapsto h_{V}(s)=\displaystyle{{\alpha_{V}}_{\left(J\right)}^{-1}}(s\overline{V_{X^{\mathrm{in}}}})$
     where $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)\in\mathrm{SIC}(\alpha_{V})$ where $\alpha_{V}(I)=[0,1]$.
@@
     • $\bullet$
-      $(-\infty,h(1)]\ni s\mapsto\alpha_{V}(s)=\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in[h(0),h(1)]\end{array}\right.$
-Appendix
-Proof of Lemma 4.
-Lemma 4.
+[…]
+Appendix A
+A.1 Additional proofs for Opt-Lyapunov definition
+We provide details about the proofs of Lemmas  3 and  4.
+Lemma A.3.
 Let $\gamma$ be strictly greater than 1.
@@
 As we have $f(x)>x$ for all $x>0$, we have $f(1)>1$ and thus the interval $[1,f(1))$ is nonempty.
-We define $h:[1,f(1))$ continuous, strictly increasing whose values are strictly positive such that $h(1)=1$ and $\lim_{x\mapsto f(1)^{-}}h(x)=\gamma$.
-We could take for example $h$ the affine map linking $(1,1)$ to $(f(1),\gamma)$.
-Now, as the sequence $(f^{k}(1))_{k}$ is strictly increasing and unbounded (if $M=\sup_{k\in\mathbb{N}}f^{k}(1)>0$ is finite then as $f$ is increasing and continuous $f(M)=M$ but $M<f(M)$ by assumption on $f$).
-As $f$ is invertible on $\mathbb{R}_{+}$ and we have $f^{-1}(x)<x$ for all $x>0$, the sequence $(f^{-k}(1))_{k}$ is strictly decreasing and positive and converges to 0 as $k$ tends to +$\infty$+ ∞.
-Hence $\mathbb{R}_{+}^{*}$ is union of disjoint intervals $[f^{k}(1),f^{k+1}(1))$ where $k\in\mathbb{Z}$.
+We define $h:[1,f(1))$ as continuous and strictly increasing, with strictly positive values such that $h(1)=1$ and $\lim_{x\mapsto f(1)^{-}}h(x)=\gamma$.
+For example, we could take $h$ to be the affine map linking $(1,1)$ to $(f(1),\gamma)$.
+Now, as the sequence $(f^{k}(1))_{k}$ is strictly increasing and unbounded (if $M=\sup_{k\in\mathbb{N}}f^{k}(1)>0$ is finite, then as $f$ is increasing and continuous, $f(M)=M$ but $M<f(M)$ by the assumption on $f$).
+As $f$ is invertible on $\mathbb{R}_{+}$ and we have $f^{-1}(x)<x$ for all $x>0$, the sequence $(f^{-k}(1))_{k}$ is strictly decreasing and positive and converges to 0 as $k$ tends to $\infty$.
+Hence $\mathbb{R}_{+}^{*}$ is union of disjoint intervals $f^{k}(1),f^{k+1}(1))$ where $k\in\mathbb{Z}$.
 Then for all $x>0$, there exists a unique $k\in\mathbb{Z}$ such that $x\in[f^{k}(1),f^{k+1}(1))$.
 We define a function $g$ on $\mathbb{R}_{+}$ as follows $g(0)=0$ and for all $x>0$, $g(x):=\gamma^{k}h(f^{-k}(x))$ where $k$ is the unique integer such that $x\in[f^{k}(1),f^{k+1}(1))$.
-As $g$ is equal to $h$ on $[1,f(1)]$, then $g$ is strictly increasing, continuous and strictly positive on $[1,f(1)]$.
+As $g$ is equal to $h$ on $[1,f(1)]$, then $g$ is strictly increasing, continuous, and strictly positive on $[1,f(1)]$.
 Then, on the open interval $(f^{k}(1),f^{k+1}(1))$ for $k\neq 0$, $g$ is the composition $t\mapsto\gamma^{k}t$, $h$ and $f^{-k}$ and thus is strictly increasing, continuous and strictly positive on $(f^{k}(1),f^{k+1}(1))$ for all $k\in\mathbb{Z}$.
@@
 Now, we prove that $g$ is continuous at 0 (at right).
-Let us consider a sequence $(x_{n})_{n\in\mathbb{N}}$ of positive reals which converges to 0.
+Let us consider a sequence $(x_{n})_{n\in\mathbb{N}}$ of positive reals that converges to 0.
 Let $n\in\mathbb{N}$.
@@
 Finally, $\gamma^{k_{n}}\leq g(x_{n})=\gamma^{k_{n}}h(f^{-k_{n}}(x_{n}))\leq\gamma^{k_{n}+1}$.
-As $(x_{n})_{n\in\mathbb{N}}$ tends to 0, the sequence $(k_{n})_{n\in\mathbb{N}}$ tends to $\infty$- ∞ and we conclude that $(g(x_{n}))_{n\in\mathbb{N}}$ tends to $0=g(0)$ and $g$ is continuous at 0.
-We have already proved that $g$ is strictly increasing on each interval $[f^{k}(1),f^{k+1}(1))$.
-It is easy to see that it is also the case when we pick two reals $x,y$ in two different intervals : $x<y$ implies $g(x)<g(y)$.
-Finally, we have to prove that $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
+As $(x_{n})_{n\in\mathbb{N}}$ tends to 0, the sequence $(k_{n})_{n\in\mathbb{N}}$ tends to -$\infty$ and we conclude that $(g(x_{n}))_{n\in\mathbb{N}}$ tends to $0=g(0)$ and $g$ is continuous at 0.
+We have already proven that $g$ is strictly increasing on each interval $f^{k}(1),f^{k+1}(1))$.
+It is easy to see that this is also the case when we pick two reals $x,y$ in two different intervals: $x<y$ implies $g(x)<g(y)$.
+Finally, we must prove that $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
 As $f(0)=0$, this holds for $x=0$.
-Now let us take $x>0$.
-This real $x$ belongs to some interval $[f^{k}(1),f^{k+1}(1))$.
+Now, let us take $x>0$.
+This real $x$ belongs to some interval $f^{k}(1),f^{k+1}(1))$.
 Thus, $f(x)\in[f^{k+1}(1),f^{k+2}(1))$ and $g(f(x))=\gamma^{k+1}h(f^{-k-1}(f(x))=\gamma\gamma^{k}h(f^{-k}(x))=\gamma g(x)$.
@@
 ∎
-Proof of Lemma 5
-Lemma 5.
+Lemma A.4.
 Let $\lambda\in(0,1)$. Let $f\in\mathcal{K}$ verifying for all $x>0$, $f(x)<x$. Then, there exists a function $g\in\mathcal{K}$ satisfying $g(f(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
 Proof.
-The function $f^{-1}$ is continuous, strictly increasing and definite and satisfies $f^{-1}(x)>x$ for all $x>0$.
+The function $f^{-1}$ is continuous, strictly increasing, and definite, and satisfies $f^{-1}(x)>x$ for all $x>0$.
 Then using Lemma 3, there exists $g\in\mathcal{K}$ such that $g(f^{-1}(x))=\lambda^{-1}g(x)$ for all $x\in\mathbb{R}$.
-Then for all $y\in\mathbb{R}_{+}$, there exists a unique $x\in\mathbb{R}_{+}$ such that $y=f^{-1}(x)$ or equivalently $x=f(y)$.
+Then, for all $y\in\mathbb{R}_{+}$, there exists a unique $x\in\mathbb{R}_{+}$ such that $y=f^{-1}(x)$ or equivalently $x=f(y)$.
 Then, we have $g(y)=\lambda^{-1}g(f(y))$ and we conclude that $g(f(y))=\lambda g(y)$ for all $y\in\mathbb{R}_{+}$.
 ∎
+A.2 Technical details about examples on the application for Leslie models
+We provide technical details about the family of $\mathcal{KL}_{\rm gen}$ functions proposed in Equation (28) and the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound construction proposed in Equation (31).
+Then, we provide details on the construction of the $(X^{\mathrm{in}}_{\ell},L,\varphi_{\ell})$ certificate of compatibility for the Lyapunov function $V_{\ell}$ proposed in Equation (32).
+Recall that passing from $X^{\mathrm{in}}_{\ell}$ to $X_{\ell}$ does not the solutions of Problem ($DPCP_{\ell}$).
+First we give details about the analysis of coordinate functions.
+Lemma A.5.
+For all $x\in(\mathbb{R}_{+}^{*})^{2}$, $(\pi_{1}(L^{k}x))_{k\in\mathbb{N}}$ and $(\pi_{2}(L^{k}x))_{k\in\mathbb{N}}$ are strictly increasing (from $k=2$ for $\pi_{2}$) and both tend to $\infty$ as $k$ goes to $\infty$.
+Proof.
+Let $x\in(\mathbb{R}_{+}^{*})^{2}$, we have $\pi_{1}(Lx)=a\pi_{1}(x)+b\pi_{2}(x)>a\pi_{1}(x)$.
+We conclude that $(\pi_{1}(L^{k}x))_{k}$ is strictly increasing as $(\mathbb{R}_{+}^{*})^{2}$ by $L$.
+Moreover, we have $\pi_{2}(L^{2}x)=c\pi_{1}(Lx)=ca\pi_{1}(x)+cb\pi_{2}(x)>ac\pi_{1}(x)=a\pi_{2}(Lx)$.
+We conclude that $(\pi_{2}(L^{k}x))_{k}$ is strictly increasing from $k\geq 2$.
+Now, we also have
+$\pi_{1}(L^{k}x)=\dfrac{1}{c(\lambda_{2}-\lambda_{1})}\left(\lambda_{1}^{k+1}\mathfrak{a}(x)+\lambda_{2}^{k+1}\mathfrak{b}(x)\right)=\dfrac{\lambda_{2}^{k+1}}{c(\lambda_{2}-\lambda_{1})}\left(\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k+1}\mathfrak{a}(x)+\mathfrak{b}(x)\right)$
+As $\lambda_{2}>1$ and from (21), $\pi_{1}(L^{k}x)$ tends to $\infty$ as $k$ goes to $\infty$. We use the same techniques to prove that $\pi_{1}(L^{k}x)$ tends to $\infty$ as $k$ goes to $\infty$. ∎
+A.2.1 On the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ certificate construction
+First recall that
+$H:\{(x,y)\in\mathbb{R}^{2}:x+y\neq 0\}\ni(u,v)\mapsto\dfrac{\lambda_{2}}{c}\dfrac{\dfrac{\lambda_{1}}{\lambda_{2}}v+u}{v+u}$
+and we have for all $x\in X_{\ell}$ and all $k\in\mathbb{N}$
+$\varphi_{\ell}(L^{k}x)=\dfrac{\lambda_{1}^{k+1}\mathfrak{a}(x)+\lambda_{2}^{k+1}\mathfrak{b}(x)}{c\left(\lambda_{1}^{k}\mathfrak{a}(x)+\lambda_{2}^{k}\mathfrak{b}(x)\right)}=\dfrac{\lambda_{2}}{c}\dfrac{\dfrac{\lambda_{1}}{\lambda_{2}}\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)}}{\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)}}=H\left(\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)},\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}\right).$
+We also have:
+  • for all $v>0$ (resp. $v<0$), $u\mapsto H(u,v)$ is strictly increasing on $(-\infty,-v)$ and on $(-v,+\infty)$ (resp. strictly decreasing on $(-\infty,-v)$ and on $(-v,+\infty)$);
+  • for all $u<0$ (resp. $u>0$), $v\mapsto H(u,v)$ is strictly increasing on $(-\infty,-u)$ and on $(-u,+\infty)$ (resp. strictly decreasing on $(-\infty,-u)$ and on $(-u,+\infty)$).
+Proposition A.1.
+For all $\varepsilon>0$, the function
+$\gamma_{\varepsilon}:\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\left\{\begin{array}[]{lr}H\left(s,\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}\right)&\text{ if }s\leq-\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}-\varepsilon\\ \left(s+\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}+\varepsilon\right)+H\left(-\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}-\varepsilon,\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{t}\right)&\text{ otherwise }\end{array}\right$.
+belongs to $\mathcal{KL}_{\rm gen}$.
+Proof.
+Let $t\in\mathbb{R}_{+}$, for sake of simplicity we write $\omega_{t}=|\lambda_{1}/\lambda_{2}|^{t}$.
+Let $s,s^{\prime}\in\mathbb{R}$ such that $s<s^{\prime}$.
+First suppose that $s<s^{\prime}\leq-\omega_{t}-\varepsilon$.
+As $\omega_{t}>0$, $u\mapsto H(u,\omega_{t})$ is strictly increasing on $(-\infty,-\omega_{t})$.
+Since $s,s^{\prime}<-\omega_{t}$, we get $H(s,\omega_{t})<H(s^{\prime},\omega_{t})$.
+Now, suppose that $s^{\prime}>s>-\omega_{t}-\varepsilon$.
+Then, $(s^{\prime}+\omega_{t}+\varepsilon>(s+\omega_{t}+\varepsilon)$ and thus $\gamma_{\varepsilon}(s^{\prime},t)>\gamma_{\varepsilon}(s,t)$ as $H(-\omega_{t}-\varepsilon,\omega_{t})$ does not depend on $s,s^{\prime}$.
+Finally, suppose that $s<-\omega_{t}-\varepsilon<s^{\prime}$.
+The result follows readily from the fact that $H(\cdot,\omega_{t})$ strictly increases on $(-\infty,-\omega_{t})$ and $(s^{\prime}+\omega_{t}+\varepsilon)$ is strictly positive.
+Now let $s\in\mathbb{R}$.
+Let $t,t^{\prime}\in\mathbb{R}_{+}$ such that $t<t^{\prime}$.
+Keeping the same notation that is $\omega_{t}$ for $|\lambda_{1}/\lambda_{2}|^{t}$, we have $\omega_{t^{\prime}}<\omega_{t}$.
+If $s\leq-\omega_{t}-\varepsilon$, we have $\gamma_{\varepsilon}(s,t)=H(s,\omega_{t})$ and $\gamma_{\varepsilon}(s,t^{\prime})=H(s,\omega_{t^{\prime}})$.
+The conclusion follows from the fact that $v\mapsto H(s,v)$ strictly increases as $s<0$ on $(-\infty,-s)$ and $\omega_{t^{\prime}}<\omega_{t}<\omega_{t}+\varepsilon\leq-s$.
+Now suppose that $s>-\omega_{t^{\prime}}-\varepsilon$.
+Actually, we have for all $u,v$ such that $u+v\neq 0$, $H(u,v)=\lambda_{2}/c(1+(\lambda_{1}/\lambda_{2}-1)v(u+v)^{-1})$ and thus $H(-\omega_{t}-\varepsilon,\omega_{t})=\lambda_{2}/c(1+\varepsilon^{-1}\omega_{t}(1-\lambda_{1}/\lambda_{2}))$ and $H(-\omega_{t^{\prime}}-\varepsilon,\omega_{t^{\prime}})=\lambda_{2}/c(1+\varepsilon^{-1}\omega_{t^{\prime}}(1-\lambda_{1}/\lambda_{2})$.
+We conclude that $v\mapsto(s+v+\varepsilon)+\lambda_{2}/c(1+\varepsilon^{-1}v(1-\lambda_{1}/\lambda_{2})$ strictly increases and $\gamma_{\varepsilon}(s,t^{\prime})<\gamma_{\varepsilon}(s,t)$.
+Finally, suppose that $\omega_{t^{\prime}}-\varepsilon\geq s>-\omega_{t}-\varepsilon$.
+Then $\gamma_{\varepsilon}(s,t^{\prime})=H(s,\omega_{t}^{\prime})$ and $\gamma_{\varepsilon}(s,t)=(s+\omega_{t}+\varepsilon)+H(-\omega_{t}-\varepsilon,\omega_{t})$.
+Then $(s+\omega_{t}+\varepsilon)+H(-\omega_{t}-\varepsilon,\omega_{t})-H(s,\omega_{t}^{\prime})=(s+\omega_{t}+\varepsilon)+\lambda_{2}/c(1+\varepsilon^{-1}\omega_{t}(1-\lambda_{1}/\lambda_{2})-\lambda_{2}/c(1+(\omega_{t^{\prime}}+s)^{-1}\omega_{t^{\prime}}(\lambda_{1}/\lambda_{2}-1)=(s+\omega_{t}+\varepsilon)+\lambda_{2}/c(1-\lambda_{1}/\lambda_{2})(\varepsilon^{-1}\omega_{t}+\omega_{t^{\prime}}(\omega_{t^{\prime}}+s)^{-1})$.
+Now as $\omega_{t^{\prime}}-\varepsilon\geq s$, we have $(\omega_{t^{\prime}}+s)^{-1}\geq-\varepsilon^{-1}$.
+Thus, $(s+\omega_{t}+\varepsilon)+H(-\omega_{t}-\varepsilon,\omega_{t})-H(s,\omega_{t}^{\prime})\geq(s+\omega_{t}+\varepsilon)+\lambda_{2}/c(1-\lambda_{1}/\lambda_{2})\varepsilon^{-1}(\omega_{t}-\omega_{t^{\prime}})>0$ and $\gamma_{\varepsilon}(s,t)>\gamma_{\varepsilon}(s,t^{\prime})$.
+∎
+Recall that
+[…]
+We propose a slightly different approach than that in Example 2.
+We decompose the supremum of $\theta_{\ell}$ over $\mathfrak{a}^{-1}((0,+\infty))$ into two parts.
+Those have an additional constraint that is $(\mathfrak{a}-\mathfrak{b})^{-1}((-\infty,0))$ for the first and $(\mathfrak{a}-\mathfrak{b})^{-1}([0,+\infty))$ for the second.
+This explains the different cases in the definition of $\varepsilon^{>0}$ in (29).
+We recall that $\mathfrak{b}$ is strictly positive on $(\mathbb{R}_{+}^{*})^{2}$. Hence, $\mathfrak{a}(x)\geq\mathfrak{b}(x)$ implies that $\mathfrak{a}(x)>0$.
+Proposition A.2.
+The following equalities are true:
+[…]
+Moreover, if $\mathfrak{a}(x^{lu})<\mathfrak{b}(x^{lu})$, $\{x\in X_{\ell}:\mathfrak{a}(x)\geq\mathfrak{b}(x)\}$ is empty; and if $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$, the set $\{x\in X_{\ell}:\mathfrak{a}(x)>0,\ \mathfrak{a}(x)<\mathfrak{b}(x)\}$, if nonempty, contains elements for which $\theta_{\ell}$ is arbitrarily close to -1.
+Proof.
+Let us consider the supremum of $\theta_{\ell}$ on $C_{1}:=\{x\in X:\mathfrak{a}(x)>0,\ \mathfrak{a}(x)<\mathfrak{b}(x)\}$.
+If $C_{1}$ is empty, the supremum is known to be -$\infty$.
+Suppose that $C_{1}$ is nonempty.
+First, we must have $\mathfrak{a}(x^{lu})>0$.
+Otherwise, $\mathfrak{a}(x^{lu})\leq 0$ which is the same as $\lambda_{2}/c\leq l_{1}/u_{2}$.
+This implies that $\mathfrak{a}(x)\leq 0$ for all $x\in X$ and $C_{1}$ cannot be nonempty.
+Second, $\theta_{\ell}$ is strictly smaller than -1 on $C_{1}$.
+From the monotonicity analysis of $\theta_{\ell}$, we have for all $x$ different from $x^{lu}$ such that $\mathfrak{a}(x)>0$, $\theta_{\ell}(x)<\theta_{\ell}(x^{lu})$.
+If $0<\mathfrak{a}(x^{lu})<\mathfrak{b}(x^{lu})$ then $x^{lu}\in C_{1}$ and it is the unique maximizer of $\theta_{\ell}$ on $C_{1}$.
+Note that $0<\mathfrak{a}(x^{lu})<\mathfrak{b}(x^{lu})$ is equivalent to $\theta_{\ell}(x^{lu})<-1$ and so $\min\{\theta_{\ell}(x^{lu}),-1\}=\theta_{\ell}(x)$.
+Now, suppose that $\mathfrak{a}(x^{ul})\geq\mathfrak{b}(x^{ul})$.
+As $X^{\mathrm{in}}_{\ell}$ is connected and $(\mathfrak{a}-\mathfrak{b})$ continuous, $\mathfrak{a}-\mathfrak{b}(X^{\mathrm{in}}_{\ell})$ is an interval.
+Moreover, $(\mathfrak{a}-\mathfrak{b})(x^{lu})\geq 0$ and as $C_{1}$ is nonempty $(\mathfrak{a}-\mathfrak{b})(y)<0$ for some $y\in X^{\mathrm{in}}_{\ell}$ with $\mathfrak{a}(y)>0$.
+Hence, there exists $z\in X^{\mathrm{in}}_{\ell}$ such that $\mathfrak{a}(z)=\mathfrak{b}(z)$.
+As $\mathfrak{b}$ is strictly positive on $X^{\mathrm{in}}_{\ell}$, $\mathfrak{a}(z)$ is strictly positive.
+Recall that $X^{\mathrm{in}}_{\ell}$ does not contain 0, and thus $y\neq 0$.
+Let $\eta>0$ and write $u_{\eta}=\eta y\|y\|^{-1}+z$.
+We have $\mathfrak{a}(u_{\eta})=\eta\|y\|^{-1}\mathfrak{a}(y)+\mathfrak{a}(z)>0$ and $(\mathfrak{a}-\mathfrak{b})(u_{\eta})=\eta\|y\|^{-1}(\mathfrak{a}-\mathfrak{b})(y)+(\mathfrak{a}-\mathfrak{b})(z)=\eta\|y\|^{-1}(\mathfrak{a}-\mathfrak{b})(y)<0$.
+Then, $u_{\eta}\in C_{1}$ and $\theta_{\ell}(u_{\eta})$ tends to $\theta_{\ell}(z)=-1$ as $\eta$ tends to 0.
+We conclude that $\sup_{C_{1}}\theta_{\ell}=-1$.
+Note that as $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$ we have $1\leq\theta_{\ell}(x^{lu})$ and thus $\min\{-1,\theta_{\ell}(x^{lu})\}=-1$.
+Simple computations show that $0<\mathfrak{a}(x^{lu})<\mathfrak{b}(x^{lu})$ is the same as $a/(2c)<l_{1}/u_{2}$.
+This implies that for all $x\in X_{\ell}$, $a/(2c)<x_{1}/x_{2}$ or equivalently $0<\mathfrak{a}(x)<\mathfrak{b}(x)$ for all $x\in X_{\ell}$.
+We prove that the supremum of $\theta_{\ell}$ on $C_{2}:=\{x\in X:\mathfrak{a}(x)\geq\mathfrak{b}(x)\}$ equals $\theta_{\ell}(x^{lu})$ if $C_{2}$ is nonempty.
+It suffices to prove that $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$ if $C_{2}$ is nonempty.
+Let $x\in C_{2}$.
+This is the same as $a/2c\geq x_{1}/x_{2}$.
+This implies that $a/2c\geq l_{1}/u_{2}$ and so $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$.
+We prove that the supremum of $\theta_{\ell}$ on $C_{3}:=\{x\in X:\mathfrak{a}(x)<0\}$ equals $\theta_{\ell}(x^{ul})$ if $C_{3}$ is nonempty.
+From the monotonicity analysis of $\theta_{\ell}$, it suffices to prove that $\mathfrak{a}(x^{ul})<0$ if $C_{3}$ is nonempty.
+Let $x\in C_{3}$.
+This is the same as $\lambda_{2}/c<x_{1}/x_{2}$.
+This implies that $\lambda_{2}/c<u_{1}/l_{2}$ and so $\mathfrak{a}(x^{ul})<0$.
+∎
+Corollary A.1.
+The maximum of $\theta_{\ell}$ over $X_{\ell}$ can be computed as
+[…]
+Now recall that
+[…]
+and
+$\overline{\varepsilon}=\min\{\varepsilon^{<0},\varepsilon^{>0}\}.$
+Proposition A.3.
+We have for all $x\in X_{\ell}$ and all $k\in\mathbb{N}$:
+$\varphi_{\ell}(L^{k}x)\leq\gamma_{\overline{\varepsilon}}\left(\theta_{\ell}(x),\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right)$
+Proof.
+First, from (21), we have, for all $k\in\mathbb{N}$,
+[…]
+Hence,
+$\forall\,k\in\mathbb{N},\ \forall\,x\text{ s.t.
+}\mathfrak{a}(x)<0,\ \theta_{\ell}(x)<-1\leq\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\text{ and }\forall\,k\in\mathbb{N}^{*},\ \forall\,x\text{ s.t.
+}\mathfrak{a}(x)>0,\ \theta_{\ell}(x)<\dfrac{\lambda_{1}}{\lambda_{2}}\leq\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}.$
+By definition of $\varepsilon^{<0},\varepsilon^{>0}$ and $\overline{\varepsilon}$, we have
+$\forall\,k\in\mathbb{N},\ \forall\,x\text{ s.t. }\mathfrak{a}(x)<0,\ \theta_{\ell}(x)<-1-\overline{\varepsilon}\leq\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}-\overline{\varepsilon}$
+and
+$\forall\,k\in\mathbb{N}^{*},\ \forall\,x\text{ s.t.
+}\mathfrak{a}(x)>0,\ \theta_{\ell}(x)<\dfrac{\lambda_{1}}{\lambda_{2}}-\overline{\varepsilon}\leq\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}-\overline{\varepsilon}.$
+Finally,
+$\forall\,k\in\mathbb{N},\ \forall\,x\text{ s.t.
+}\mathfrak{a}(x)<0,\ \gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),k)=H\left(\theta_{\ell}(x),\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right)$
+and
+$\forall\,k\in\mathbb{N}^{*},\ \forall\,x\text{ s.t.
+}\mathfrak{a}(x)>0,\ \gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),k)=H\left(\theta_{\ell}(x),\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right).$
+Now, from Eq.
+(27) and as for all $x\in X_{\ell}$, $v\mapsto H(\theta_{\ell}(x),v)$ strictly increases on $(-\infty,-\theta_{\ell}(x))$, we conclude for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)<0$ and for all $k\in\mathbb{N}$:
+$\varphi_{\ell}(L^{k}x)=H\left(\theta_{\ell}(x),\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}\right)\leq H\left(\theta_{\ell}(x),\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right)=\gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),k).$
+Recall that $\mathfrak{a}\circ L=\lambda_{1}\mathfrak{a}$ and $\mathfrak{b}\circ L=\lambda_{2}\mathfrak{b}$.
+Thus, for $\mathfrak{a}(x)>0$ we have $\mathfrak{a}(Lx)<0$ and we can apply the previous inequality to $Lx$.
+We have for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)>0$ and for all $k\in\mathbb{N}^{*}$
+$\begin{array}[]{ll}\varphi_{\ell}(L^{k}x)=\varphi_{\ell}(L^{k-1}Lx)&=H\left(\dfrac{\mathfrak{b}(Lx)}{\mathfrak{a}(Lx)},\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k-1}\right)\\ &\leq H\left(\dfrac{\mathfrak{b}(Lx)}{\mathfrak{a}(Lx)},\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k-1}\right)=H\left(-\dfrac{\mathfrak{b}(x)}{\mathfrak{a}(x)},\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right)=H\left(\theta_{\ell}(x),\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right)=\gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),k).\end{array}$
+It remains to prove the inequality at $k=0$ for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)>0$, that is, $\varphi_{\ell}(x)\leq\gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),0)$ for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)>0$.
+Suppose that $0<\mathfrak{a}(x^{lu})<\mathfrak{b}(x^{lu})$.
+It implies that for all $x\in X_{\ell}$, $0<\mathfrak{a}(x)<\mathfrak{b}(x)$ which is the same as $\theta_{\ell}(x)<-1$ and by definition of $\varepsilon^{>0}$ and $\overline{\varepsilon}$, $\theta_{\ell}(x)\leq-1-\overline{\varepsilon}$ for all $x\in X_{\ell}$.
+Then, for all $x\in X_{\ell}$, $\gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),0)=H(\theta_{\ell}(x),1)$.
+Since
+$H\left(\theta_{\ell}(x),1\right)-\varphi(x)=H\left(\theta_{\ell}(x),1\right)-H\left(-\theta_{\ell}(x),1\right)=2\theta_{\ell}(x)\dfrac{\lambda_{2}}{c}\dfrac{1-\frac{\lambda_{1}}{\lambda_{2}}}{\left(1+\theta_{\ell}(x)\right)\left(1-\theta_{\ell}(x)\right)}>0,$
+we conclude that, if $0<\mathfrak{a}(x^{lu})<\mathfrak{b}(x^{lu})$, for all $k\in\mathbb{N}$, for all $x\in X_{\ell}$ such that $\mathfrak{a}(x)>0$:
+$\varphi_{\ell}(L^{k}x)\leq H\left(\theta_{\ell}(x),\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\right)=\gamma_{\overline{\varepsilon}}(\theta_{\ell}(x),k).$
+Now assume that $\mathfrak{a}(x^{lu})\geq\mathfrak{b}(x^{lu})$ and let $x\in X_{\ell}$ such that $\mathfrak{a}(x)>0$.
+If $\{x\in X_{\ell}:0<\mathfrak{a}(x)<\mathfrak{b}(x)\}$ is nonempty, we can have $\theta_{\ell}(y)>-1-\overline{\varepsilon}$ for some $y\in X_{\ell}$ such that $0<\mathfrak{a}(y)<\mathfrak{b}(y)$ as proved in Proposition A.2.
+For the elements $y\in\{x\in X_{\ell}:0<\mathfrak{a}(x)<\mathfrak{b}(x)\}$ for which $\theta_{\ell}(y)\leq-1-\overline{\varepsilon}$, the inequality $\varphi_{\ell}(y)\leq\gamma_{\overline{\varepsilon}}(\theta_{\ell}(y),0)$ has been proved.
+Therefore, suppose that $\theta_{\ell}(x)>-1-\overline{\varepsilon}$.
+We remark that, for all $u>0$ and $\varepsilon>0$
+$H(-u-\varepsilon,u)=\dfrac{\lambda_{2}}{c}\dfrac{u+\varepsilon-\frac{\lambda_{1}}{\lambda_{2}}u}{\varepsilon}=\dfrac{\lambda_{2}}{c}\left(1+\dfrac{u}{\varepsilon}\left(1-\frac{\lambda_{1}}{\lambda_{2}}\right)\right)>\dfrac{\lambda_{2}}{c}.$
+It follows that:
+$\gamma_{\varepsilon}(\theta_{\ell}(x),0)=\theta_{\ell}(x)+1+\overline{\varepsilon}+H(-1-\overline{\varepsilon},1)>\dfrac{\lambda_{2}}{c}>\dfrac{x_{1}}{x_{2}}=\varphi_{\ell}(x).$
+This ends the proof. ∎
+A.2.2 Details on certificate of compatibility
+We bring details about the $(X_{\ell},L,\varphi_{\ell})$ certificate of compatibility proposed in Example 9.
+Lemma A.6.
+Let $z=\max\{\omega_{0},\omega_{1}\}$ and let $\eta>0$ such that $z-\eta>\lambda_{2}/c$.
+[…]
+[…]
+Then,
+[…]
+and
+$M:=\max_{x\in X^{\mathrm{in}}_{\ell}}\pi_{1}(L^{\overline{N}}x)=\pi_{1}(L^{\overline{N}}x^{uu})$
+satisfies $M\geq\pi_{1}(L^{j}x)$ for all $x\in X_{\ell}$, for all $j\in\mathbb{N}$ such that $\varphi_{\ell}(L^{j}x)\geq z-\eta$.
+Proof.
+Let $x\in X_{\ell}$ and $k\in\mathbb{N}$. Then, from (21),
+$\begin{array}[]{ll}\varphi_{\ell}(L^{k}x)\geq z-\eta&\iff\dfrac{\lambda_{2}}{c}\dfrac{\mathfrak{a}(x)\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k+1}+\mathfrak{b}(x)}{\mathfrak{a}(x)\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}+\mathfrak{b}(x)}\geq z-\eta\\ &\iff\left(\dfrac{\lambda_{1}}{\lambda_{2}}\right)^{k}\mathfrak{a}(x)(\lambda_{1}-c(z-\eta))\geq\mathfrak{b}(x)(c(z-\eta)-\lambda_{2})\end{array}$
+Now, if $\mathfrak{a}(x)<0$, we get, applying the natural logarithm,
+$\begin{array}[]{ll}\varphi_{\ell}(L^{k}x)\geq z-\eta&\implies\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|^{k}\geq\dfrac{\mathfrak{b}(x)(c(z-\eta)-\lambda_{2})}{|\mathfrak{a}(x)|(c(z-\eta)-\lambda_{1})}\\ &\implies k\leq N(x)\end{array}$
+[…]
+$\begin{array}[]{ll}N(Lu)=\left\lfloor\dfrac{\ln\left(\dfrac{\mathfrak{b}(Lx)(c(z-\eta)-\lambda_{2})}{|\mathfrak{a}(Lx)|(c(z-\eta)-\lambda_{1})}\right)}{\ln\left(\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|\right)}\right\rfloor&=\left\lfloor\dfrac{\ln\left(\dfrac{\lambda_{2}\mathfrak{b}(x)(c(z-\eta)-\lambda_{2})}{|\lambda_{1}||\mathfrak{a}(x)|(c(z-\eta)-\lambda_{1})}\right)}{\ln\left(\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|\right)}\right\rfloor\\ &=\left\lfloor\dfrac{-\ln\left(\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|\right)+\ln\left(\dfrac{\mathfrak{b}(x)(c(z-\eta)-\lambda_{2})}{|\mathfrak{a}(x)|(c(z-\eta)-\lambda_{1})}\right)}{\ln\left(\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|\right)}\right\rfloor=N(u)-1\end{array}$
+[…]
+Now, we prove the formula for the computation of $\overline{N}$.
+As $\lambda_{1}/\lambda_{2}\in(-1,0)$, the function $\mathbb{R}_{+}^{*}\ni t\mapsto\ln(ct(z-\eta)-\lambda_{2})/(c(z-\eta)-\lambda_{1}))/\ln(|\lambda_{1}/\lambda_{2}|)$ strictly decreases.
+Thus, $N$ is maximal on $X_{\ell}$ when $\mathfrak{b}/|\mathfrak{a}|=-\theta_{\ell}$ is minimal on $X_{\ell}$.
+We have already computed the supremum of $\theta_{\ell}$ over $X_{\ell}$ in Corollary A.1 which corresponds to the opposite of the infimum $\theta_{\ell}$ over $X_{\ell}$.
+Hence, minimizers on $X_{\ell}$ of $\theta_{\ell}$ are the maximizers of $\theta_{\ell}$ over $X_{\ell}$.
+This yields the definition of $\overline{N}$.
+To prove that $\overline{N}$ is nonnegative, it suffices to prove that $\varphi_{\ell}(x^{ul})\geq z-\eta$ when $\mathfrak{a}(x^{lu})\leq 0$, $\varphi_{\ell}(Lx^{lu})\geq z-\eta$ when $\mathfrak{a}(x^{ul})\geq 0$ and $\varphi_{\ell}(x^{ul})\geq z-\eta$ or $\varphi_{\ell}(Lx^{lu})\geq z-\eta$ when $\mathfrak{a}(x^{lu})>0$ and $\mathfrak{a}(x^{ul})<0$.
+We have already proved that $\omega_{2k}=\varphi_{\ell}(L^{2k}x^{ul})\text{ and }\omega_{2k+1}=\varphi_{\ell}(L^{2k+1}x^{lu})$.
+It follows that $\omega_{0}=\varphi_{\ell}(x^{ul})$ and $\omega_{1}=\varphi_{\ell}(Lx^{lu})$.
+First, we assume that $z=\omega_{0}$.
+Then, as $z$ is strictly greater than $\lambda_{2}/c$, it is also the case for $\omega_{0}$.
+Since $\omega_{0}=\varphi_{\ell}(x^{ul})$, we have $\mathfrak{a}(x^{ul})<0$.
+And finally, $N(x^{ul})\geq 0$ if and only $\mathfrak{b}(x^{ul})(c(z-\eta)-\lambda_{2})\leq-\mathfrak{a}(x^{ul})(c(z-\eta)-\lambda_{1})$ that is equivalent to $c(z-\eta)\leq(\lambda_{1}\mathfrak{a}(x^{ul})+\lambda_{2}\mathfrak{b}(x^{ul}))/(\mathfrak{a}(x^{ul})+\mathfrak{b}(x^{ul}))$ that is the same as $z-\eta\leq\varphi(x^{ul})=z$.
+Hence, we conclude that $N(x^{ul})\geq 0$ if $z=\omega_{0}$.
+Now assume that $z=\omega_{1}=\varphi_{\ell}(Lx^{lu})$.
+As $z$ is strictly greater than $\lambda_{2}/c$, $\mathfrak{a}(x^{lu})>0$.
+It follows that $N(x^{lu})\geq 0$ if and only if $c(z-\eta)(\mathfrak{b}(x^{lu})-\mathfrak{a}(x^{lu}))\leq-\lambda_{1}\mathfrak{a}(x^{lu})+\lambda_{2}\mathfrak{b}(x^{lu})$.
+If $\mathfrak{b}(x^{lu})-\mathfrak{a}(x^{lu})\leq 0$, the inequality is valid as $\lambda_{1}<0$ and $\lambda_{2}$, $\mathfrak{a}(x^{lu})$ and $\mathfrak{b}(x^{lu})$ are strictly positive.
+Assume that $\mathfrak{b}(x^{lu})-\mathfrak{a}(x^{lu})>0$.
+Then $N(x^{lu})\geq 0$ if and only if $z-\eta\leq(\lambda_{2}/c)(-\lambda_{1}\mathfrak{a}(x^{lu})/\lambda_{2}+\mathfrak{b}(x^{lu}))/(-\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu}))$.
+Now
+$\begin{array}[]{ll}\dfrac{-\frac{\lambda_{1}}{\lambda_{2}}\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}{-\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}>\dfrac{-\frac{\lambda_{1}}{\lambda_{2}}\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}{\frac{\lambda_{1}}{\lambda_{2}}\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}&=\dfrac{\left(\frac{\lambda_{1}}{\lambda_{2}}\right)^{2}\mathfrak{a}(x^{lu})-\left(\frac{\lambda_{1}}{\lambda_{2}}\right)^{2}\mathfrak{a}(x^{lu})-\frac{\lambda_{1}}{\lambda_{2}}\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}{\frac{\lambda_{1}}{\lambda_{2}}\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}\\ &=\varphi(Lx^{lu})-\dfrac{\frac{\lambda_{1}}{\lambda_{2}}\left(1+\frac{\lambda_{1}}{\lambda_{2}}\right)}{\frac{\lambda_{1}}{\lambda_{2}}\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu})}\end{array}$
+We conclude that $(\lambda_{2}/c)(-\lambda_{1}\mathfrak{a}(x^{lu})/\lambda_{2}+\mathfrak{b}(x^{lu}))/(-\mathfrak{a}(x^{lu})+\mathfrak{b}(x^{lu}))>\varphi(Lx^{lu})=z>z-\eta$ and $N(x^{lu})\geq 0$.
+Then, we can compute $M=\max_{x\in X^{\mathrm{in}}_{\ell}}\pi_{1}(L^{\overline{N}}x)$.
+First, from the discussion of Subsection 3.3.1, for all $k\leq\overline{N}$, $\pi_{1}(L^{k}x)\leq\pi_{1}(L^{\overline{N}}x)\leq M$.
+Now as for all $x\in\mathbb{R}^{2}$ and all $k\in\mathbb{N}$, $\pi_{1}(L^{k}x)=(c(\lambda_{2}-\lambda_{1}))^{-1}(\lambda_{1}^{k+1}\mathfrak{a}(x)+\lambda_{2}^{k+1}\mathfrak{b}(x))$ and thus $\pi_{1}(L^{k}x)$ increases in $x_{1}$ and $x_{2}$ and we conclude that $\pi_{1}(L^{k}x)\leq\pi_{1}(L^{k}x^{uu})$ for all $x\in X^{\mathrm{in}}_{\ell}$ and for all $k\in\mathbb{N}$.
+Finally, $M=\pi_{1}(L^{\overline{N}}x^{uu})$.
+∎
+Proposition A.4.
+Let $z=\max\{\omega_{0},\omega_{1}\}$ and $x^{*}\in X_{\ell}\cup L(X_{\ell})$ such that $\varphi_{\ell}(x^{*})=z$. Let us consider
+$\alpha_{1}:\mathbb{R}\ni s\to\dfrac{C}{M^{2}(1+s^{-2})}.$
+Then for all $\eta>0$ such that $z-\eta>\lambda_{2}/c$ and for all $\gamma$ greater than
+$\max\left\{\max_{s\in[z-\eta,z]}\alpha_{1}^{\prime}(s),\dfrac{V_{\ell}(x^{*})}{\eta}\right\},$
+the function
+$\alpha:\mathbb{R}\ni s\mapsto\left\{\begin{array}[]{lr}\dfrac{C}{M^{2}(1+s^{-2})}&\text{ if }s\geq z\\ \gamma(s-z)+\dfrac{C}{M^{2}(1+z^{-2})}&\text{ if }s<z\end{array}\right$.
+is a $(X_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$.
+Proof.
+From Lemma A.6, for all $x\in X_{\ell}$ and for all $k\in\mathbb{N}$ such that $\varphi(L^{k}x)\geq z-\eta$ we have
+$V_{\ell}(L^{k}x)=\dfrac{C}{\pi_{1}(L^{k}x)^{2}+\pi_{2}(L^{k}x)^{2}}=\dfrac{C}{\pi_{1}(L^{k}x)^{2}(1+\varphi_{\ell}(L^{k}x)^{-2})}\geq\dfrac{C}{M^{2}(1+\varphi_{\ell}(L^{k}x)^{-2})}$
+We define
+$\alpha_{1}:s\mapsto\dfrac{C}{M^{2}(1+s^{-2})}$
+Now let us take any $\gamma$ greater than
+$\max\left\{\max_{s\in[z-\eta,z]}\alpha_{1}^{\prime}(s),\dfrac{V_{\ell}(x^{*})}{\eta}\right\}.$
+Note that as $\alpha_{1}$ is strictly increasing and $V_{\ell}$ is strictly positive, $\gamma$ is strictly positive.
+From the mean value theorem, for all $s\in[z-\eta,z]$, $\alpha_{1}(z)-\alpha_{1}(s)=(z-s)\alpha_{1}^{{}^{\prime}}(t)$ for some $t\in(s,\omega)$.
+Then by definition of $\gamma$, $\alpha_{1}(z)-\alpha_{1}(s)\leq\gamma(z-s)$ or equivalently $\alpha_{1}(z)+\gamma(s-z)\leq\alpha_{1}(z)$.
+By definition of $\alpha_{1}$, we have $\alpha_{1}(z)+\gamma(\varphi_{\ell}(L^{k}x)-z)\leq V(L^{k}z)$ for all $x\in X^{\mathrm{in}}_{\ell}$ and all $k\in\mathbb{N}$ such that $\varphi_{\ell}(L^{k}x)\in[z-\eta,z]$.
+Moreover, there exists $u\in[z-\eta,z)$ such that $\gamma(u-z)+\alpha_{1}(z)=0$.
+Indeed, $V(L^{j}x)\geq\alpha_{1}(\varphi(L^{j}x))$ for all $j\in\mathbb{N}$ and $x\in X_{\ell}$ such that $\varphi(L^{j}x)\geq z-\eta$ and $\varphi(x^{*})=z$ with either $x^{*}\in X_{\ell}$ if $z=\omega_{0}$ or $x^{*}\in L(X_{\ell})$ if $z=\omega_{1}$.
+From this, we have $\gamma(z-\eta-z)+\alpha_{1}(z)\leq-V(x^{*})+\alpha_{1}(z)\leq-\alpha_{1}(z)+\alpha_{1}(z)=0$.
+We also have $\gamma(z-z)+\alpha_{1}(z)>0$.
+Either, $u=z-\eta$ or $\eta\gamma+\alpha_{1}(z)<0$ and $u\in(z-\eta,z)$.
+It follows from the strict positivity of $V_{\ell}$ and the strict increase of $s\mapsto\alpha_{1}(z)+\gamma(s-z)$ that for all $x\in X_{\ell}$ and all $k\in\mathbb{N}$ such that $\varphi_{\ell}(L^{k}x)\leq u$, $\alpha_{1}(z)+\gamma(\varphi_{\ell}(L^{k}x)-z)\leq V_{\ell}(L^{k}x)$.
+We conclude that
+$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{C}{M(1+s^{-2})}&\text{ if }s\geq z\\ \gamma(s-z)+\dfrac{C}{M^{2}(1+z^{-2})}&\text{ if }s<z\end{array}\right$.
+is a $(X_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$. ∎